### v10

# Burnside type results for fusion categories

Abstract.
In this paper, we extend a classical vanishing result of Burnside from the character tables of finite groups to the character tables of commutative fusion rings, or more generally to a certain class of abelian normalizable hypergroups. We also treat the dual vanishing result. We show that any nilpotent fusion categories satisfy both Burnside’s property and its dual. Using Drinfeld’s map, we obtain that the Grothendieck ring of any weakly-integral modular fusion category satisfies both properties. As applications, we prove new identities that hold in the Grothendieck ring of any weakly-integral fusion category satisfying the dual-Burnside’s property, thus providing new categorification criteria. In particular we improve [OY23, Theorem 4.5] as follows: A weakly integral modular fusion category of $\mathrm{FPdim}$ $md$ with $d$ square-free coprime with $m$ and $\mathrm{FPdim}(X)^{2}$ for every simple object $X$, has a pointed modular fusion subcategory of $\mathrm{FPdim}$ $d$. We also prove some new results on the perfect fusion categories.
1991 Mathematics Subject Classification:
18M20; 16T30
The first author is supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS/CCCDI - UEFISCDI, project number PN-III-P4-ID-PCE-2020-0878, within PNCDI III. The second author is supported by BIMSA Start-up Research Fund and Foreign Youth Talent Program from the Ministry of Sciences and Technology of China.
1. Introduction
A classical result of Burnside in the character theory of finite groups states that any irreducible non-linear character of a finite group vanishes on at least one element of the group. This can be stated as follows: in the character table of a finite group, the row of every character of degree $\neq 1$ contains a zero entry.
More recently, the dual version of this result was also intensively studied in the literature. This dual version describes the group elements that vanish on at least one irreducible character, see [INW99] and the references therein. Burnside’s result was extended to all weakly integral fusion categories, initially in the context of modular categories in [GNN09, Appendix], and subsequently in the general framework in [Bur23].
The main goal of this paper is to develop an analogue of Burnside’s result for a certain large class of fusion rings, or even more generally, to a certain class of abelian rational normalizable hypergroups. We also consider the dual version of Burnside’s result in this more general settings.
Recall that in [Bur21] the author introduced the notion of dual of a Grothendieck ring. This was achieved based on work of Harrison [Har79] on dualizable probability groups and on [ZZ21]. More generally, this notion of dual of a fusion ring was extended to arbitrary fusion rings in [BP22]. It was noticed in [Bur21] that the dual of the Grothendieck ring of a pivotal fusion category is isomorphic to the center of the category as defined in [Shi17].
In this paper we show that both Burnside’s property and its dual are related with the ring structure of the dual of the involving Grothendieck/fusion ring. In general, the dual of a fusion ring is no longer a fusion ring, but an abelian normalizable hypergroup, see [Bur21]. The advantage of the class of hypergroups is to be closed under duality, see [Bur21, Proposition 2.3].
For a finite dimensional $\mathbb{C}$-algebra $H$, fix a (finite) $\mathbb{C}$-linear basis $\mathcal{B}$ with the unit $1\in\mathcal{B}$ and write $ab=\sum_{c\in{\mathcal{B}}}N^{c}_{ab}c$ for some scalars $N^{c}_{ab}\in\mathbb{C}$. Note that $H={\mathbb{C}[\mathcal{B}]}$, the complex linear span of $\mathcal{B}$.
Definition 1.1.
A finite dimensional algebra $H$ together with a $\mathbb{C}$-linear basis $\mathcal{B}$ is called a hypergroup if there is an involution map $(-)^{*}:{\mathcal{B}}\rightarrow{\mathcal{B}}$ such that $N^{1}_{aa^{*}}>0$ and $N^{1}_{ab}=0$ if $b\neq a^{*}$. The elements $b\in\mathcal{B}$ are called the standard elements of ${(H,\mathcal{B})}$.
We may refer to ${\mathcal{B}}$ (or $H$) only as a hypergroup, when $H$ (or ${\mathcal{B}}$) is implicitly understood. A hypergroup $\mathcal{B}$ is called:
  • symmetric if $N^{1}_{ab}=N^{1}_{ba}$,
  • normalized if $\sum_{c\in{\mathcal{B}}}N^{c}_{ab}=1$,
  • real if $N_{ab}^{c}\in\mathbb{R}$,
  • rational if $N_{ab}^{c}\in\mathbb{Q}$,
  • real non-negative if $N_{ab}^{c}\geq 0$,
  • abelian if $ab=ba$,
for all $a,b,c\in{\mathcal{B}}$. Throughout this paper we use the abbreviation (A)RN-hypergroup for an (abelian) real non-negative hypergroup.
For a normalized hypergroup, the linear map $\mu_{1}:H\rightarrow\mathbb{C}$ defined on the linear basis ${\mathcal{B}}$ by $\mu_{1}(a)=1$ for all $a\in{\mathcal{B}}$, is a morphism of algebras. It is called the augmentation map of $(H,\mathcal{B})$. The underlying algebra $H$ of a hypergroup $(H,\mathcal{B})$ is a finite dimensional *-algebra, so is semisimple. For a normalized hypergroup ${(H,\mathcal{B})}$ the central primitive idempotent $F_{1}$ corresponding to the character $\mu_{1}$ is denoted by $\lambda_{(H,\mathcal{B})}$ or simply $\lambda_{H}$ if $\mathcal{B}$ is implicitly understood. A formula for this idempotent is described in Corollary 2.8.
A RN-hypergroup is sometimes called a table algebra in the literature. Recall that a fusion ring ${\mathcal{B}}$ is a hypergroup such that $N^{a}_{bc}\in\mathbb{Z}_{\geq 0}$ and $N^{1}_{aa^{*}}=N^{1}_{a^{*}a}=1$ for all $a,b,c\in\mathcal{B}$. The first main result of this paper is the following generalization of Burnside’s result to some weakly-integral fusion rings:
Theorem 1.2.
Let ${(H,\mathcal{B})}$ be a commutative fusion ring with a $h$-integral dual. Let $a$ be a standard element such that $\mathrm{FPdim}(a)>1$. Then there is $\mu\in{\widehat{\mathcal{B}}}$ such that $\mu(a)=0$.
A fusion ring whose dual is h-integral (see Definition 3.6) is always weakly-integral ($\mathrm{FPdim}{(H,\mathcal{B})}\in\mathbb{Z}$). The converse is true for a weakly-integral fusion category (see Remark 3.7). The set $^{\widehat{\mathcal{B}}}$ is the set of characters $\mu_{j}:H\rightarrow\mathbb{C}$ (see §2 for more details). In fact we prove a more general version of Theorem 1.2, for a certain class of abelian rational normalizable hypergroups, see Theorem 4.17.
Definition 1.3.
For any hypergroup ${(H,\mathcal{B})}$, an element $a\in\mathcal{B}$ with
(1.1) $aa^{*}=N_{a,a^{*}}^{1}1\text{ (in other words, }N_{a,a^{*}}^{c}=\delta_{1,c}N_{a,a^{*}}^{1}\text{)}$
is called a grouplike element of ${(H,\mathcal{B})}$.
By Lemma 2.25, the set of grouplike elements form a group (with the multiplication structure) denoted ${G{(H,\mathcal{B})}}$.
Definition 1.4.
A hypergroup ${(H,\mathcal{B})}$ has Burnside’s vanishing property (or shortly, is Burnside) if for all $a\in\mathcal{B}$, the following are equivalent:
  1. (1)
    There is some $\mu\in{\widehat{\mathcal{B}}}$ such that $\mu(a)=0$,
  2. (2)
    the standard element $a$ is not a grouplike element.
Theorem 1.2 states that a commutative fusion ring with h-integral dual is Burnside. Non-Burnside simple integral fusion rings can be found in §12. It is easy to see that a hypergroup is Burnside if and only if the set of invertible standard elements coincides to the set of grouplike elements, see Proposition 4.3.
1.1. The dual hypergroup of a normalized hypergroup
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. By the abelian assumption, $^{\widehat{\mathcal{B}}}$ forms a basis for the dual $H^{*}$, which then is $\mathbb{C}[{\widehat{\mathcal{B}}}]$, and on which one can define a multiplication. On the basis $^{\widehat{\mathcal{B}}}$, the multiplication $\mu_{i}\star\mu_{j}$ is defined by declaring
(1.2) $[\mu_{i}\star\mu_{j}](a):=\mu_{i}(a)\mu_{j}(a),\;\text{for all}\;a\in\mathcal{B}$
and then extending linearly $\mu_{i}\star\mu_{j}$ on the entire ${\mathbb{C}[\mathcal{B}]}$. In this way, one obtains an algebra structure on $H^{*}$.
It follows that there are some non-zero scalars ${\widehat{p}}_{k}(i,j)\in\mathbb{C}$ such that
(1.3) $\mu_{i}\star\mu_{j}=\sum_{k=1}^{m}{\widehat{p}}_{k}(i,j)\mu_{k}.$
By [Bur21, Proposition 2.3], for any abelian normalized hypergroup $(H,{\mathcal{B}})$, the pair $(H^{*},{\widehat{\mathcal{B}}})$ is also an abelian normalized hypergroup. Moreover, the involution on $^{\widehat{\mathcal{B}}}$ is given by $\mu_{j}\mapsto\mu_{j^{\#}}$, where $\mu_{j^{\#}}\in{\widehat{\mathcal{B}}}$ is defined by $\mu_{j^{\#}}(a):=\mu_{j}(a^{*})$ for all $a\in\mathcal{B}$. We denote by $\widehat{{(H,\mathcal{B})}}:=(H^{*},{\widehat{\mathcal{B}}})$ the dual hypergroup of the abelian normalized hypergroup ${(H,\mathcal{B})}$. It is also well known that for a normalized hypergroup ${(H,\mathcal{B})}$, then $\mu_{j}(a^{*})=\overline{\mu_{j}(a)}$ for all $a\in\mathcal{B}$. We also use the notation ${\widehat{h}}_{j}:={\widehat{p}}_{1}(j,j^{\#})^{-1}$, called the order of $\mu_{j}$ in the dual hypergroup ${\widehat{(H,\mathcal{B})}}$, see Lemma 2.6.
For any RN-hypergroup ${(H,\mathcal{B})}$, we denote by ${(H,\mathcal{B})}_{ad}$ its adjoint sub-hypergroup, see §6.2. As for fusion rings, this is defined as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by the element $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}$, where $h_{i}:=1/N^{1}_{i,i^{*}}$.
For any RN-hypergroup ${(H,\mathcal{B})}$, the central primitive idempotent corresponding to the character $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is denoted by $\lambda_{H}$. For $H=K(\mathcal{C})$, the Grothendieck ring of a fusion category $\mathcal{C}$, this primitive idempotent corresponds to the regular element.
Definition 1.5.
A hypergroup ${(H,\mathcal{B})}$ is called normalizable if one obtains a normalized hypergroup after rescaling it. (see §2.2 for more details.)
Based on Lemma 2.2 we denote such a hypergroup by ${(H,\mathcal{B},\mu_{1})}$ where $\mu_{1}\in{\widehat{\mathcal{B}}}$.
Inspired by the notion of dualizable probability groups from [Har79], we propose the following:
Definition 1.6.
An ARN-hypergroup is called dualizable if its dual is also ARN.
1.2. Main results of the paper
We characterize Burnside dualizable ARN-hypergroups as follows:
Theorem 1.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. The following are equivalent (see Definition 2.4 for $n(\_)$):
  1. (1)
    The hypergroup ${(H,\mathcal{B})}$ is Burnside.
  2. (2)
    The following identity holds on ${\widehat{(H,\mathcal{B})}}$:
    (1.4) $\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\frac{1}{n(\widehat{(H,{\mathcal{B}})}_{ad})}\big{(}\sum_{\mu_{j}\in{{\widehat{\mathcal{B}}}}_{ad}}{\widehat{h}}_{j}\mu_{j}\big{)}.$
Note that the right hand side of the above equation coincides to the integral $\lambda_{{}_{{\widehat{(H,\mathcal{B})}}_{ad}}}$ of the adjoint hypergroup ${\widehat{(H,\mathcal{B})}}_{ad}$.
For a fusion category $\mathcal{C}$, its Grothendieck ring is denoted by $K(\mathcal{C})$, and the set of isomorphism class representatives of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$. Let $d_{i}:=\mathrm{FPdim}(X_{i})$ be the Frobenius-Perron dimension of $X_{i}$, $i\in{\mathcal{I}}=\{1,\dots,m\}$, and $x_{i}:=[X_{i}]$ be the class of the simple object $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$.
For any weakly-integral premodular category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is both RN and rational, as demonstrated in [Bur21, Theorem 1.2]. Furthermore, for any unitary fusion category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is also RN, see [LPW21, ENO21], and primary $3$-criterion in [HLPW24]. A normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if its dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside (see examples from group theory in §12). By duality, one obtains the following:
Theorem 1.8.
The following are equivalent for an abelian RN-hypergroup ${(H,\mathcal{B})}$.
  1. (1)
    The hypergroup ${(H,\mathcal{B})}$ is dual-Burnside.
  2. (2)
    The following identity holds:
    (1.5) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{n({{(H,\mathcal{B})}_{ad}})}\big{(}\sum_{x_{i}\in\mathcal{B}_{ad}}d_{i}x_{i}\big{)}$
The dual-Burnside fusion categories can be characterized by:
Theorem 1.9.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then ${K(\mathcal{C})}$ is dual-Burnside if and only if the following holds:
(1.6) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)}.$
About the dual-Burnside hypergroups, we prove the following:
Theorem 1.10.
Let $(H,{\mathcal{B}})$ be an abelian rational RN-hypergroup. If $(H,{\mathcal{B}})$ is dual-Burnside then $\mathrm{FPdim}(H,{\mathcal{B}})\in\mathbb{Q}$.
As a corollary, any commutative dual-Burnside fusion ring is weakly-integral. Similar to the fusion ring settings, [GN08], the universal grading group can be defined for any ARN-hypergroup. Using this grading, we can similarly define the concept of nilpotent ARN-hypergroup, and prove the following:
Theorem 1.11.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. Then ${(H,\mathcal{B})}$ is nilpotent if and only if its dual ${\widehat{(H,\mathcal{B})}}$ is also nilpotent. Moreover, in this case they have the same nilpotency class.
For a nilpotent ARN-hypergroup, we prove the following:
Theorem 1.12.
A nilpotent dualizable ARN-hypergroup is both Burnside and dual-Burnside.
Let $\mathcal{C}$ be a modular fusion category. Then Corollary 10.3 states that ${K(\mathcal{C})}$ is Burnside if and only if it is dual-Burnside. The proof involves the Drinfeld map of $\mathcal{C}$. Then Theorem 1.9 implies that:
Theorem 1.13.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside:
(1.7) $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\frac{1}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{ad}}})}d_{i}x_{i}\big{)}$
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [Bur23, Theorem 2].
Corollary 1.14.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside and with the group of invertible objects $G(\mathcal{C})$ of odd order,
(1.8) $\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{ad}}})}d_{i}x_{i}\big{)}$
For any non-negative integer $n$, let $\mathcal{V}(n)$ be the set of prime divisors of $n$. For a weakly-integral fusion category we also let $\mathcal{V}(\mathcal{C}):=\mathcal{V}(\mathrm{FPdim}(\mathcal{C}))$. From Theorem 1.9, we can also derive the following result:
Theorem 1.15.
Let $\mathcal{C}$ be a fusion category such that ${K(\mathcal{C})}$ is commutative and dual-Burnside. Then
(1.9) $\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
Moreover, if additionally $\mathcal{C}$ is nilpotent then
(1.10) $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2}).$
Note that since any nilpotent fusion category $\mathcal{C}$ is weakly-integral then $d_{i}^{2}\in\mathbb{Z}$ for all $i$, see [GN08, Theorem 3.10]. Note that in the case of an integral fusion category Equation (1.10) can be written as
(1.11) $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}).$
Using the above identities, we prove the following results about the $\mathrm{FPdim}$ of simple objects:
Theorem 1.16.
For any modular fusion category $\mathcal{C}$ such that ${K(\mathcal{C})}$ is dual-Burnside,
(1.12) $\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
For any weakly-integral modular fusion category $\mathcal{C}$:
(1.13) $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})\big{)}$
Again, recall $d_{i}^{2}\in\mathbb{Z}$ for all $i$ by [GN08, Theorem 3.10]. Moreover, if $\mathcal{C}$ is an integral modular fusion category then the above equation becomes:
(1.14) $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i})\big{)}$
Concrete examples on which these results apply can be found in §12.3. Next theorem improves [OY23, Theorem 4.5].
Theorem 1.17.
Let $\mathcal{C}$ be a weakly-integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$ and $d_{i}^{2}$ for all $i$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
$\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
Corollary 1.18.
Let $\mathcal{C}$ be any integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
$\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
In particular, the $\mathrm{FPdim}$ of a perfect integral modular fusion category does not have a powerless prime factor (Corollary 10.11). In particular, a perfect modular fusion category of even $\mathrm{FPdim}$ has its $\mathrm{FPdim}$ divisible by $4$.
Remark 1.19.
By the classification of non-pointed integral modular fusion categories of small rank in [ABPP] and by Corollary 1.18, the $\mathrm{FPdim}$ of a non-pointed integral modular fusion category does not have a powerless prime factor for every rank less than $11\times 3=33$, and every rank less than $17\times 7=119$ in the odd-dimensional case.
Integral modular fusion categories were recently intensively studied in the literature. In [CP22] it was formulated the following conjecture
Conjecture 1.20.
There is no perfect modular fusion category of odd $\mathrm{FPdim}$.
Note that the above result also applies to non-pointed simple integral modular fusion categories (since they are in particular perfect). The importance of the existence of such categories was described in [LPR23, §5], where the following two open statements are proved to be equivalent, whereas Proposition 10.14 states that they imply Conjecture 1.20.
Statement 1.21.
Every simple integral fusion category is weakly group-theoretical.
Statement 1.22.
Every simple integral modular fusion category is pointed.
We finally prove the following result:
Theorem 1.23.
A perfect modular fusion category is (dual-)Burnside if and only if it is integral.
Note that the integral modular fusion category $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if the finite group $G$ is perfect and centerless (more generally, see §11.1).
Shortly, the organization of this paper is the following. In §2 we recall the basics on hypergroups and abelian rational normalizable hypergroups. In §3 we develop few basic facts about Galois symmetries of rational hypergroups. In §4 hypergroups with Burnside property are studied and Theorem 1.7 is proven. In the same section we prove Theorem 1.2 and its more general version Theorem 4.17 for hypergroups. In this section, Burnside’s property for dual hypergroups is also studied and Theorem 1.8 is proven. In §5 we extend the notion of kernels of objects from the settings of fusion categories to the settings of ARN-hypergroups. In §7 we construct the universal grading group for ARN-hypergroups. A connection between this universal grading group and the group of grouplike elements of the dual it is proven in Theorem 6.5. §7 describes the constriction of the universal grading group of an ARN-hypergroup. In §8 we define nilpotent ARN-hypergroup and prove Theorems 1.11 and 1.12. The applications to modular fusion categories are given in §10.1. In §10 we prove Proposition 10.14 showing that any of the two statements from [LPR23, §5] implies Conjecture 1.20.
Next, §11 delves into applications and supplementary results: §11.1 offers a characterization for a Drinfeld center to be perfect; §11.2 establishes some results concerning integrality and Burnside properties, proving Theorems 1.10 and 1.23, particularly highlighting their equivalence in the perfect modular case; §11.3 characterizes the near-group fusion categories which admit a modular structure.
Finally, §12 presents concrete examples and counterexamples: §12.1 introduces non-Burnside simple integral fusion rings; §12.2 classifies certain dual-Burnside $\mathrm{Rep}(G)$; §12.3 outlines the classification of fusion rings lacking modular categorification within certain families, employing skills from this paper.
Throughout of this paper, all vector spaces and linear categories are considered over the ground field $\mathbb{C}$ of complex numbers.
Acknowledgements
We thank Vicenţiu Paşol and Andrew Schopieray for fruitful discussions on Galois groups of fusion rings and hypergroups.
2. Abelian rational normalizable hypergroups
Let ${(H,\mathcal{B})}$ be a hypergroup as in Definition 1.1. A morphism of hypergroups $\psi:(H,{\mathcal{B}})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime})$ is an algebra morphism $\psi:H\rightarrow H^{\prime}$ such that $\psi({\mathcal{B}})={\mathcal{B}}^{\prime}$ and $\psi(b^{*})=\psi(b)^{*}$ for all $b\in\mathcal{B}$. The morphism $\psi$ is called an isomorphism of hypergroups if it is bijective.
Let ${(H,\mathcal{B})}$ be a normalized abelian hypergroup and ${\widehat{(H,\mathcal{B})}}$ its dual as defined in the introduction. There is a formula for the dual fusion coefficients ${\widehat{p}}_{k}(i,j)$ given in [Bur21, Proposition 2.1]. For a normalized abelian hypergroup ${(H,\mathcal{B})}$, this is given by:
(2.1) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{a\in{\mathcal{B}}}h_{a}\mu_{j_{1}}(a)\mu_{j_{2}}(a)\mu_{k}(a^{*})\bigg{)}$
where the scalars $n_{k}$ are defined below (2.2).
Remark 2.1.
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. We usually let :=${\mathcal{B}}:=\{x_{i}\}_{i\in\mathcal{I}}$, with :=$\mathcal{I}:=\{1,\dots m\}$ and $x_{1}=1$, and say that ${(H,\mathcal{B})}$ has rank $m$. We can simplify the notation $N_{x_{i}x_{j}}^{x_{k}}$ by $N_{i,j}^{k}$. Let $\{x_{i}^{\circ}\}_{i\in\mathcal{I}}$be the linear dual basis of the linear basis ${\mathcal{B}}$ of $H$. Therefore $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, for all $i,i^{\prime}\in\mathcal{I}$. It is easy to see that in this case $\widehat{E}_{i}:=x_{i}^{\circ}\in H^{*}$ is a primitive central idempotent of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{x_{i}}:&H^{*}&\to&\mathbb{C}\\ &f&\mapsto&f(x_{i}).\end{array}$
2.1. Function $\tau$ and associative non-degenerate bilinear form
Let $(H,\mathcal{B})$ be any abelian hypergroup. Define $h_{i}:=1/N^{1}_{i,i^{*}}$, for all $i\in\mathcal{I}$. Note that $h_{1}=1$. Then $H$ is a commutative semisimple algebra, and therefore $H\simeq\bigoplus_{j=1}^{m}\mathbb{C}F_{j}$, where $F_{j}$ are the central primitive idempotents of $H$. One can define $\tau\in H^{*}$ with $\tau(x_{i})=\delta_{i,1}$, where as above, $x_{1}=1$. Note that $\tau=x_{1}^{\circ}$ with the above notations. Moreover, in this case
$\begin{array}[]{cccc}(\;,\;)_{\tau}:&H\otimes H&\to&\mathbb{C}\\ &(a,b)&\mapsto&\tau(ab)\end{array}$
is an associative symmetric non-degenerate bilinear form. Thus, one can write
(2.2) $\tau=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}.$
for some non-zero scalars $n_{j}\in\mathbb{C}^{\times}$. These scalars $n_{j}$ are called the formal codegrees of $(H,\mathcal{B})$, see [Ost15]. Note that $\sum_{j=1}^{m}\frac{1}{n_{j}}=\tau(1)=1$.
From the definition of a hypergroup it follows that $\{h_{i}x_{i}\}_{i\in\mathcal{I}}$ and $\{x_{i^{*}}\}_{i\in\mathcal{I}}$ form a pair of dual bases for $(\;,\;)_{\tau}$. Note that in this case $\{n_{j}F_{j}\}_{i\in\mathcal{I}}$ and $\{F_{j}\}_{i\in\mathcal{I}}$ form also a pair of dual basis for the bilinear form $(\;,\;)_{\tau}$. The existence of the two pairs of dual bases implies that:
(2.3) $\sum_{i=1}^{m}{h_{i}}x_{i}\otimes x_{i}^{*}=\sum_{j=1}^{m}n_{j}F_{j}\otimes F_{j}$
From here it follows that
(2.4) $F_{j}:=\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i^{*}})x_{i}\big{)}$
by applying $\operatorname{id}\otimes\mu_{j}$ for the character $\mu_{j}$ corresponding to $F_{j}$. Applying $\mu_{j}\otimes\mu_{k}$ to the above, one obtains the first orthogonality relations
(2.5) $\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i})\mu_{k}(x_{i^{*}})=\delta_{j,k}n_{j}$
Since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$, for $j=k$, it follows from here that $n_{j}>0$. By a classical argument, as for finite groups orthogonality, one obtains the second orthogonality relation:
(2.6) $\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x_{i})\mu_{j}(x_{l}^{*})=\delta_{i,l}h_{i}^{-1}$
2.2. Normalizable abelian hypergroups and their duals
Note that hypergroups can be rescaled as follows. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Let :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Then it is easy to see that $(H,\mathcal{B}^{\prime})$ is also a hypergroup. However note that $\tau^{\prime}(\frac{x_{i}}{\alpha_{{}_{i}}})=\delta_{i,1}$ is the same functional $\tau$ on $H$ and therefore the formal codegrees $n_{j}$ do not change after rescaling.
Recall the notion of a normalizable hypergroup from Definition 1.5. Note that any rescaled normalizable hypergroup is also normalizable.
Lemma 2.2.
Let ${(H,\mathcal{B})}$ be a hypergroup. Then ${(H,\mathcal{B})}$ is normalizable if and only if there is $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$.
Proof.
It is easy to see that if $(H,\{\frac{x_{i}}{\alpha_{{}_{i}}}\})$ is normalized after a rescaling, then $\mu_{1}:H\rightarrow H,\;x_{i}\mapsto\alpha_{{}_{i}}$ is a morphism of algebras on $H$. Thus $\mu_{1}\in{\widehat{\mathcal{B}}}$. Conversely, one can check directly that $(H,\{\frac{x_{i}}{\mu_{1}(x_{i})}\})$ is a normalized hypergroup for any morphism $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $i\in\mathcal{I}$. ∎
We denote by ${(H,\mathcal{B},\mu_{1})}$ the normalizable hypergroup obtained by normalizing ${(H,\mathcal{B})}$ by $\mu_{1}\in{\widehat{\mathcal{B}}}$. We also denote by $\overline{{(H,\mathcal{B},\mu_{1})}}$ the normalized hypergroup obtained by rescaling ${(H,\mathcal{B},\mu_{1})}$ as in Lemma 2.2.
Definition 2.3.
Recall from Subsection 1.1 that any abelian normalized hypergroup $H$ admits a dual hypergroup $\widehat{H}$. In the case of $\overline{{(H,\mathcal{B},\mu_{1})}}$, we denote this dual by ${{\widehat{{(H,\mathcal{B},\mu_{1})}}}}$.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup via $\mu_{1}\in\widehat{\mathcal{B}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$. From Equation (1.2) it follows that the multiplication on the dual of $\widehat{{(H,\mathcal{B},\mu_{1})}}$ can be written as
(2.7) $[f\star g](\frac{x_{i}}{\mu_{1}(x_{i})})=f(\frac{x_{i}}{\mu_{1}(x_{i})})g(\frac{x_{i}}{\mu_{1}(x_{i})})$
for all $f,g\in H^{*}$. Then it is easy to see that the algebra unit of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $\mu_{1}$. Moreover, the involution on $^{\widehat{\mathcal{B}}}$, as given in §1.1, becomes:
(2.8) $\mu_{j}^{\#}(x_{i})=\bigg{(}\frac{\mu_{1}(x_{i})}{\overline{\mu_{1}(x_{i}})}\bigg{)}\mu_{j}(x_{i}^{*}).$
Definition 2.4.
The order of a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is defined by
$n{(H,\mathcal{B},\mu_{1})}:=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}.$
In the abelian case, note that
(2.9) $n{(H,\mathcal{B},\mu_{1})}=n_{1}$
by Equation (2.5).
We shortly write $n(H)$ when $\mathcal{B}$ and $\mu_{1}$ are implicitly understood. In particular, for a normalized hypergroup one may choose $\mu_{1}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n(H)=\sum_{i=1}^{m}h_{i}$. For the rest of this paper we use the notation $d_{i}:=\mu_{1}(x_{i})$ for any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$.
We say that a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is weakly-rational if its order is a rational number, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Q}$. Also, we say that ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral if its order is an integer, i.e. $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Z}$. Note that a fusion ring is weakly-rational if and only if it is weakly-integral since $\mathrm{FPdim}(R)$ is an algebraic number in this case.
Remark 2.5.
Note that the formula from Equation (2.1) can be written for abelian normalizable hypergroups as:
(2.10) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{x_{i}\in{\mathcal{B}}}\frac{h_{i}}{d_{i}}\mu_{j_{1}}(x_{i})\mu_{j_{2}}(x_{i})\mu_{k}(x_{i^{*}})\bigg{)}$
Lemma 2.6.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then with the above notations, $n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
Proof.
By applying (2.10) to ${j_{1}}={j_{2}}^{\#}=j$ and $k=1$, and then (2.5), one obtains (see also [Bur21, Equation (2.10)]),
(2.11) $\widehat{p}_{1}(j,j^{\#})=\frac{n_{j}}{n{(H,\mathcal{B},\mu_{1})}}$
Let ${\widehat{h}}_{j}:=\frac{1}{\widehat{p}_{1}(j,j^{*})}$ be the order of $\mu_{j}$ in ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that $n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}=n{(H,\mathcal{B},\mu_{1})}(\sum_{j=1}^{m}\frac{1}{n_{j}})=n{(H,\mathcal{B},\mu_{1})}$. ∎
Next proposition holds for any normalized hypergroup, not necessarily abelian.
Proposition 2.7.
For any normalized hypergroup ${(H,\mathcal{B})}$,
(2.12) $F_{1}=\frac{1}{n{(H,\mathcal{B})}}\big{(}\sum_{i=1}^{m}h_{i^{*}}x_{i}\big{)}$
where $F_{1}$ is the idempotent corresponding to the augmentation map $\mu_{1}:H\rightarrow\mathbb{C},\;a\mapsto 1$ for all $a\in\mathcal{B}$.
Proof.
Suppose that $F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{i}$ for some $\alpha_{{}_{i}}\in\mathbb{C}$. Recall that $x_{l}F_{i}=\mu_{i}(x_{l})F_{i}$. Then, $x_{l}F_{1}=F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{l}x_{i}$. It follows that $\tau(F_{1})=\alpha_{1}=\tau(x_{l}F_{1})=\alpha_{{}_{l^{*}}}h_{l}^{-1}=\tau(F_{1}x_{l})=\alpha_{l}h_{l^{*}}^{-1}.$ Then $F_{1}=\sum_{i=1}^{m}\alpha_{i}x_{i}={\alpha_{1}}(\sum_{i=1}^{m}{h_{i^{*}}}x_{i})$ and $1=\mu_{1}(F_{1})={\alpha_{1}}(\sum_{i=1}^{m}{h_{i}})$. This implies that $\alpha_{1}=\frac{1}{n(H)}$ and therefore $\alpha_{i}=\frac{h_{i^{*}}}{n(H)}$ for all $i\in\mathcal{I}$. ∎
Corollary 2.8.
In any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$,
(2.13) $F_{1}=\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i=1}^{m}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$
Proposition 2.9.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup. Then ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is also an abelian normalized hypergroup and the formal codegrees of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ are given by
(2.14) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}.$
Proof.
As already mentioned, [Bur21, Proposition 2.3] shows that ${\widehat{(H,\mathcal{B})}}$ is also an abelian normalized hypergroup.
By Remark 2.1, $x_{i}:H^{*}\rightarrow\mathbb{C}$ are the characters of ${\widehat{(H,\mathcal{B})}}$. By definition of $\widehat{\tau}\in(H^{*})^{*}=H$, $\widehat{\tau}(\mu_{j})=\delta_{j,1}$ where $\mu_{1}$ is the unit of $H^{*}$. Thus $\widehat{\tau}=F_{1}$. Equation (2.2) for $H^{*}$ gives that $\widehat{\tau}=\sum_{i=1}^{m}\frac{1}{\widehat{n}_{i}}x_{i}$. Then Equation (2.12) finishes the proof since $n{\widehat{(H,\mathcal{B})}}=n{(H,\mathcal{B})}$, by Lemma 2.6. ∎
Corollary 2.10.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then the primitive central idempotent of the linear character $x_{1}\in\widehat{{\widehat{\mathcal{B}}}}=\mathcal{B}$ is given by
(2.15) $\widehat{F_{1}}=\sum_{j=1}^{m}\frac{\mu_{j}}{n_{j}}\in H^{*}.$
Proof.
Proposition 2.9 shows in its proof that $\widehat{\tau}=F_{1}$. By duality, since $\widehat{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\overline{{(H,\mathcal{B},\mu_{1})}}$, it follows that $\tau=\widehat{F}_{1}$. Equation (2.2) finishes the proof. ∎
A morphism of normalizable hypergroups $\psi:(H,{\mathcal{B}},\mu_{1})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime},\mu_{1}^{\prime})$ is a morphism of hypergroups with the property that $\mu^{\prime}_{1}\circ f=\mu_{1}$.
2.3. Frobenius-Perron theory for a RN-hypergroups
Frobenius-Perron theory can be defined for RN-hypergroups, in the same manner as for fusion rings, see [EGNO15, §3]. We denote by $\mathrm{FPdim}(x_{i})$ the Frobenius-Perron eigenvalue of the left multiplication operator by $x_{i}$ on $H$. As for fusion rings, by a similar argument, it can be shown that $\mathrm{FPdim}(x_{i})=\mathrm{FPdim}(x_{i^{*}})$. Indeed, it easy to see that the matrix $N_{i^{*}}$ corresponding to $x_{i^{*}}$ is the transpose of the matrix $N_{i}$ of $x_{i}$ multiplied by the scalars $\frac{h_{j}}{h_{k}}$ in the $(j,k)$-entry. More precisely, $(N_{i^{*}})_{jk}=\frac{h_{j}}{h_{k}}(N_{i})_{kj}$. Thus, the two matrices have the same set of eigenvalues, in particular the same Frobenius-Perron eigenvalue. Note that any RN-hypergroup ${(H,\mathcal{B})}$ can be normalized by $\mu_{1}=\mathrm{FPdim}$, since $\mu_{1}(x_{i})>0$ for all $1\leq i\leq m$.
Remark 2.11.
Note that any symmetric RN-hypergroup ${(H,\mathcal{B})}$ is transitive in the sense of [EGNO15, Definition 3.3.1]. Indeed the function $\tau$ from §3.1 is symmetric and therefore $\tau(x_{i}x_{j}x_{k})=\tau(x_{j}x_{k}x_{i})=\tau(x_{k}x_{i}x_{j})$. Since $\tau(x_{i}x_{j}x_{k})=h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})$ one has
(2.16) $h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})=h_{j}^{-1}m(x_{j^{*}},x_{k}x_{i})=h_{k}^{-1}m(x_{k^{*}},x_{i}x_{j}).$
which shows that ${(H,\mathcal{B})}$ is transitive.
Then it is easy to see that an analogue [EGNO15, Proposition 3.3.6] also holds for symmetric RN-hypergroups. In particular it follows that $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is the unique which takes positive values on $\mathcal{B}$.
Definition 2.12.
In this case of a RN-hypergroup ${(H,\mathcal{B})}$, one can define the Frobenius-Perron dimension as
$\mathrm{FPdim}{(H,\mathcal{B})}:=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}.$
Lemma 2.13.
For a normalized RN-hypergroup,
$\mathrm{FPdim}{(H,\mathcal{B})}=n{(H,\mathcal{B},\mathrm{FPdim})}.$
Proof.
It is easy to see that [EGNO15, Proposition 3.3.6] also works for normalized RN-hypergroups instead of fusion rings. Therefore $\mu_{1}:H\rightarrow H,x_{i}\mapsto 1$ is the unique morphism $\phi:H\rightarrow\mathbb{C}$ with $\phi(x_{i})>0$ for all $1\leq i\leq m$. This implies that $\mathrm{FPdim}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n{(H,\mathcal{B},\mathrm{FPdim})}=\mathrm{FPdim}{(H,\mathcal{B})}$. ∎
Note that by Lemma 2.2, any RN-hypergroup is normalizable via the character $\mu_{1}=\mathrm{FPdim}$. Lemma 2.13 also shows that
$n(H,\mathcal{B},\mathrm{FPdim})=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}=\mathrm{FPdim}{(H,\mathcal{B})}$
for any RN-hypergroup. For a RN-hypergroup, we sometimes write ${(H,\mathcal{B},\mathrm{FPdim})}$ instead of ${(H,\mathcal{B})}$ to specify that we normalize it by $\mathrm{FPdim}$.
Proposition 2.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be a normalizable hypergroup. Then the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling. If ${(H,\mathcal{B})}$ is RN-hypergroup, then the Frobenius-Perron dimension of ${(H,\mathcal{B})}$ is also invariant under rescaling with real positive numbers.
Proof.
First we show that the order of a normalizable hypergroup is invariant under rescaling. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Consider $(H,\mathcal{B}^{\prime},\mu_{1})$ the rescaled hypergroup with :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Note that the order of $y_{i}$ is given by ${h^{\prime}_{i}}^{-1}=N^{1^{\prime}}_{i^{\prime}i^{{}^{\prime}*}}=\tau(x_{i^{\prime}}x_{i^{{}^{\prime}*}})=\frac{1}{|\alpha_{{}_{i}}|^{2}}h_{i}^{-1}$ and this implies that $h^{\prime}_{i}=|\alpha_{{}_{i}}|^{2}h_{i}>0.$ Therefore
$n(H,\mathcal{B}^{\prime},\mu_{1})=\sum_{i=1}^{m}h^{\prime}_{i}\mid\mu_{1}(y_{i})\mid^{2}=\sum_{i=1}^{m}h_{i}\mid\mu_{1}(x_{i})\mid^{2}=n{(H,\mathcal{B},\mu_{1})}.$
Moreover, if ${(H,\mathcal{B})}$ is a RN-hypergroup and $\alpha_{{}_{i}}=\alpha_{{}_{i^{*}}}>0$, for all $1\leq i\leq m$, it follows that
$\mathrm{FPdim}(H^{\prime},\mathcal{B}^{\prime})$ $=$ $\sum_{i=1}^{m}h^{\prime}_{i}\mathrm{FPdim}(y_{i})^{2}=$
$=$ $\sum_{i=1}^{m}\mid\alpha_{{}_{i}}\mid^{2}h_{i}(\frac{\mathrm{FPdim}(x_{i})^{2}}{\alpha_{{}_{i}}})^{2}$
$=$ $\mathrm{FPdim}{(H,\mathcal{B})}.$
Corollary 2.15.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then with the above notations, the dual formal codegrees are given by
(2.17) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}d_{i}d_{i^{*}}}.$
Proof.
[…]
Remark 2.16.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup. In order to find the central primitive idempotents of $H^{*}$, one can apply Remark 2.1 after normalization. If $x_{i}^{\circ}\in H^{*}$ are defined as above by $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, then ${\widehat{E}}_{i}:=d_{i}x_{i}^{\circ}\in H^{*}$ are the primitive central idempotents of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{{\frac{x_{i}}{d_{i}}}}:&H^{*}&\to&\mathbb{C}\\ &\mu_{j}&\mapsto&\mu_{j}({\frac{x_{i}}{d_{i}}}).\end{array}$
2.4. Grouplike elements in hypergroups
In the rest of this paper all abelian RN-hypergroups are normalized by $\mu_{1}=\mathrm{FPdim}$ and therefore their duals are also considered with respect to $\mu_{1}=\mathrm{FPdim}$.
Lemma 2.17.
For any symmetric RN-hypergroup ${(H,\mathcal{B})}$, $h_{i}d_{i}^{2}\geq 1$. Moreover, $h_{i}d_{i}^{2}=1$ if and only $h_{i}x_{i}x_{i^{*}}=x_{1}.$
Proof.
One has
(2.18) $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}+\sum_{k=1}^{m}N^{k}_{ii^{*}}x_{k}.$
Passing to $\mathrm{FPdim}$, $h_{i}d_{i}^{2}=1+h_{i}\big{(}\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}\big{)}\geq 1.$ If $h_{i}d_{i}^{2}=1$ then $\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}=1$, and therefore $N^{k}_{ii^{*}}=0$ for all $k\neq 1$. Thus $h_{i}x_{i}x_{i^{*}}=x_{1}.$ Conversely, if $h_{i}x_{i}x_{i^{*}}=x_{1}$, applying $\mathrm{FPdim}$, it follows that $h_{i}\mathrm{FPdim}(x_{i})^{2}=1$. ∎
Corollary 2.18.
In any fusion ring ${(H,\mathcal{B})}$, $x_{i}x_{i^{*}}=1$ if and only if $d_{i}=1$.
Recall the set of grouplike elements ${G{(H,\mathcal{B})}}$ from Definition 1.3.
Remark 2.19.
For a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ one has that $x_{i}\in G{(H,\mathcal{B})}$ if and only if
(2.19) ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$
Indeed applying $\mu_{1}$ to Equation 1.1 one has $h_{i}d_{i}d_{i^{*}}=1$ and Equation (2.19) follows. The converse also follows by inspecting the multiplicity of $x_{1}$ in $x_{i}x_{i^{*}}$.
Lemma 2.20.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup. Then $x_{i}\in G{(H,\mathcal{B})}$ if and only if
(2.20) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;j\in\mathcal{J}.$
Proof.
If $x_{i}\in G{(H,\mathcal{B})}$ then applying $\mu_{j}$ to Equation (2.19) one obtains $|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1$ since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$. Conversely, if $\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=1$ for all $\mu_{j}$ then since ${(H,\mathcal{B})}$ is abelian it follows that ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$ and therefore $h_{i}x_{i}x_{i^{*}}=1$. ∎
Dually one has the following:
Lemma 2.21.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if
(2.21) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;i\in\mathcal{I}.$
Proof.
The dual version of Equation (2.19) implies that $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if $\mu_{j}\mu_{j^{\#}}=\mu_{1}.$ Then evaluating at ${\frac{x_{i}}{d_{i}}}$ one gets the desired identity. ∎
Corollary 2.22.
In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ for any $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ one has that $n_{j}=n(H,B,\mu_{1})$.
Proof.
By Lemma 2.21 one has $\mid\mu_{j}(x_{i})|=|d_{i}|$ for all $i\in\mathcal{I}$. By Equation (2.5) one has $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$. ∎
Lemma 2.23.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$,
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then
(2.22) $|\mu_{j}(x_{i})|\leq|d_{i}|\;\;\text{for all}\;i\in\mathcal{I}\;\text{and}\;j\in\mathcal{J}.$
Proof.
Suppose first that ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$. Thus $d_{i}={\mathrm{FPdim}(x_{i})}>0$ in this case. One has $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$ and therefore $\mu_{j}(x_{i})$ is an eigenvalue of $L_{x_{i}}$. It follows that $|\mu_{j}(x_{i})|\leq d_{i}$.
Suppose now that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. It follows that $\mathrm{FPdim}(\mu_{j})=1$, since by Remark 2.11 in this case $\mu_{j}\mapsto 1$ is the only morphism on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ which takes positive values on $^{\widehat{\mathcal{B}}}$. Then, the equation
(2.23) $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$
implies as above that $\big{|}\mu_{j}({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. ∎
Lemma 2.24.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then $n_{j}\leq n{(H,\mathcal{B},\mu_{1})}$ for all $j\in\mathcal{J}$. Moreover, $n_{j}=n{(H,\mathcal{B},\mu_{1})}$ if and only if $\mu_{j}$ is a grouplike element.
Proof.
Lemma 2.23 and Equation (2.5) give:
$n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$
and equality holds if and only if $|\mu_{j}(x_{i})|^{2}=|d_{i}|^{2}$. By Lemma 2.21 this happens if and only if $\mu_{j}\in G{(H,\mathcal{B})}$. ∎
Lemma 2.25.
In any symmetric RN-hypergroup ${(H,\mathcal{B})}$, the set $\{{\frac{x_{i}}{d_{i}}}\}$ of normalized grouplike-elements form a finite group. Moreover, the inverse of ${\frac{x_{i}}{d_{i}}}$ is given by $\frac{x_{i^{*}}}{d_{i^{*}}}$ for all $x_{i}\in{G{(H,\mathcal{B})}}$.
Proof.
If $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}$ and $x_{j}x_{j^{*}}=\frac{1}{h_{j}}x_{1}$ it follows that $(x_{i}x_{j})(x_{i}x_{j})^{*}=\frac{1}{h_{i}h_{j}}x_{1}.$ Thus $(\sum_{k=1}^{m}N^{k}_{ij}x_{k})(\sum_{k=1}^{m}N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1}$ and this implies that $x_{i}x_{j}=\frac{d_{j}d_{i}}{d_{k}}x_{k}$ for a certain $k\in\mathcal{J}$. Moreover, since $h_{i}=\frac{1}{d_{i}^{2}}$ and $h_{j}=\frac{1}{d_{j}^{2}}$ it follows that
$h_{k}^{-1}=m(x_{1},x_{k}x_{k}^{*})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}m(x_{1},x_{i}x_{j}x_{i^{*}}x_{j^{*}})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}\frac{1}{h_{i}h_{j}}=d_{k}^{2}$
which shows that $x_{k}$ is also a grouplike element ($m$ is defined in §5.1). ∎
We denote $\overline{G{(H,\mathcal{B})}}:=\{{\frac{x_{i}}{d_{i}}}\;|\;x_{i}\in G{(H,\mathcal{B})}\}$ the above group of normalized grouplike elements. Clearly $\overline{G{(H,\mathcal{B})}}=G\overline{{(H,\mathcal{B})}}$, the grouplike elements of the normalized hypergroup $\overline{{(H,\mathcal{B})}}$.
By duality one obtains the following:
Lemma 2.26.
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup and the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then ${G{\widehat{(H,\mathcal{B})}}}$ is a group.
Recall that ¯=$\overline{\mathcal{B}}=\{\frac{x_{k}}{d_{k}}\;|\;x_{k}\in{\mathcal{B}}\}$.
Lemma 2.27.
Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup. If $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element then ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}},\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}\in\overline{\mathcal{B}}$, for any $x_{j}\in\mathcal{B}$.
Proof.
Suppose that
$x_{i}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{k},$
where $\mathcal{A}$ is a set such that $N^{k}_{ij}>0$. Multiplying the above equality by $x_{i^{*}}$, one obtains
$\frac{1}{h_{i}}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{i^{*}}x_{k},$
since $h_{i}x_{i^{*}}x_{i}=x_{1}$ as $x_{i}$ is a grouplike element. Since ${(H,\mathcal{B},\mu_{1})}$ is a RN-hypergroup, it follows that $x_{i^{*}}x_{k}=\alpha_{ik}x_{j}$, for some scalar $\alpha_{ik}\in\mathbb{C}$ and all $k\in\mathcal{A}$. Thus $\alpha_{ik}=\frac{d_{i}d_{k}}{d_{j}}$. Moreover, multiplying the last equality by $x_{i}$, one obtains that $\frac{1}{h_{i}}x_{k}=x_{i^{*}}x_{j}$, which shows that $\mathcal{A}$ has a single element. ∎
Lemma 2.28.
Suppose that ${(H,\mathcal{B})}$ is an abelian RN hypergroup and $\mu_{j}\in{\widehat{(H,\mathcal{B})}}$ a character such that $\mu_{j}(x_{m})=d_{m}\omega_{m}$ and $\mu_{j}(x_{n})=d_{n}\omega_{n}$ for some scalars $|\omega_{m}|=|\omega_{n}|=1$. Then $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ for all constituents $x_{p}$ of $x_{m}x_{n}$.
Proof.
Straightforward from the absolute value triangle inequality. ∎
Lemma 2.29.
If ${(H,\mathcal{B})}$ is an abelian RN hypergroup then ${{G{\widehat{(H,\mathcal{B})}}}}$ is a group.
Proof.
Suppose that $\mu_{j},\mu_{k}\in{{G{(H,\mathcal{B})}}}$. By Lemma 2.21 it follows that $\mu_{j}(x_{m})=\omega_{m}\mathrm{FPdim}(x_{m})$ for some roots of unity $\omega_{m}$. Similarly $\mu_{k}(x_{m})=\eta_{m}\mathrm{FPdim}(x_{m})$ for some other root of unity $\eta_{m}$. Moreover, by lemma 2.28, since ${(H,\mathcal{B})}$ is RN it follows that all the constituents $x_{m}x_{n}$ when evaluated at $\mu_{j}$ have the scalar $\omega_{m}\omega_{n}$. On the other hand, evaluated at $\mu_{k}$ these constituents give the scalar $\eta_{m}\eta_{n}$. Thus
$(\mu_{j}\star\mu_{k})(x_{m}x_{n})$ $=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
$=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
$=$ $\omega_{m}\omega_{n}\eta_{m}\eta_{n}\mathrm{FPdim}(x_{m}x_{n})$
On the other hand, note that $\mu_{j}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{j}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\omega_{m}\mathrm{FPdim}(x_{m})$ and similarly $\mu_{k}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{k}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\eta_{m}\mathrm{FPdim}(x_{m}).$ Therefore
$(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})=\omega_{m}\eta_{m}\mathrm{FPdim}(x_{m})\omega_{n}\eta_{n}\mathrm{FPdim}(x_{n})$
This shows that $(\mu_{j}\star\mu_{k})(x_{m}x_{n})=(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})$ which shows that $(\mu_{j}\star\mu_{k})\in{\widehat{\mathcal{B}}}$. Moreover, $(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\omega_{p}\eta_{p}$ and Lemma 2.21 implies that $\mu_{j}\star\mu_{k}\in G{\widehat{(H,\mathcal{B})}}$. ∎
Proposition 2.30.
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup and $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$. It follows that $\mu\star\mu_{k}\in\widehat{B}$ for any $\mu_{k}\in\widehat{B}$.
Proof.
Since $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$ and ${(H,\mathcal{B})}$ is RN, by Lemma 2.21 one has $\mu(x_{i})=\omega_{i}d_{i}$ for some root of unity $\omega_{i}$ and for all $i\in\mathcal{I}$. Here $d_{i}=\mathrm{FPdim}(x_{i})$. Then as above it follows that $\mu(x_{k})=\omega_{i}\omega_{j}d_{k}$ for any constituent $x_{k}$ of $x_{i}x_{j}$.
One can compute directly that $\mu\star\mu_{k}$ and $\mu_{k}\star\mu$ are morphisms of algebras. Indeed, note that $(\mu\star\mu_{k})({\frac{x_{i}}{d_{i}}})=\omega_{i}\mu_{k}({\frac{x_{i}}{d_{i}}})$ and therefore
$(\mu\star\mu_{k})(x_{i}x_{j})=\sum_{l=1}^{m}N^{l}_{ij}(\mu\star\mu_{k})(x_{l})=\omega_{i}\omega_{j}\sum_{l=1}^{m}N^{l}_{ij}\mu_{k}(x_{l})=\omega_{i}\omega_{j}\mu_{k}(x_{i}x_{j}).$
On the other hand
$(\mu\star\mu_{k})(x_{i})(\mu\star\mu_{k})(x_{j})=\omega_{i}\mu_{k}(x_{i})\omega_{j}\mu_{k}(x_{j})=\omega_{i}\omega_{j}\mu_{k}(x_{i}x_{j}).$
Thus $(\mu\star\mu_{k})(x_{i}x_{j})=(\mu\star\mu_{k})(x_{i})(\mu\star\mu_{k})(x_{j})$ which shows that $\mu\star\mu_{k}\in{\widehat{\mathcal{B}}}$. ∎
In the dual situation, Lemma 2.29 and Proposition 2.30 can be written as follows:
Proposition 2.31.
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then $G{(H,\mathcal{B})}$ is a group. Moreover, left multiplication by $L_{{\frac{x_{i}}{d_{i}}}}$ for any $x_{i}\in{G{(H,\mathcal{B})}}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
Note that Lemmata 2.25, 2.26, 2.27, 2.29 and Propositions 2.30, 2.31 imply the following:
Corollary 2.32.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then the following statements hold:
  1. (1)
    Both $\overline{G{(H,\mathcal{B})}}$ and $G{\widehat{(H,\mathcal{B})}}$ are finite groups.
  2. (2)
    The operator given by left multiplication $L_{\frac{x_{i}}{d_{i}}}$ with $x_{i}\in G{(H,\mathcal{B})}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
  3. (3)
    The operator given by left multiplication $L_{\mu_{j}}$ with $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ permutes the standard basis $^{{\widehat{\mathcal{B}}}}$.
2.5. Some results on $P$ and ${\widehat{P}}$.
Lemma 2.33.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $x_{i}\in G{(H,\mathcal{B})}$ be a grouplike element. Then
(2.24) $<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\pm 1.$
Proof.
By the definition of the multiplication in the dual,
$<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}).$
We group together the factors of $\mu_{j}$ and $\mu_{j\#}$ in the above product. Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j^{\#}}({\frac{x_{i}}{d_{i}}})=\mid\mu_{j}({\frac{x_{i}}{d_{i}}})\mid^{2}=1$ by Lemma 2.20. On the other hand, if $\mu_{j}=\mu_{j^{\#}}$, then $\mu_{j}({\frac{x_{i}}{d_{i}}})\in\mathbb{R}$ and therefore $\mu_{j}({\frac{x_{i}}{d_{i}}})=\pm 1$ by the same Lemma 2.20. ∎
Definition 2.34.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. We denote by ${\mathrm{sgn}}(x_{i})$ the above value of the product $\prod_{j=1}^{m}\mu_{j}$ evaluated at ${\frac{x_{i}}{d_{i}}}$.
The dual version of the above result is the following:
Lemma 2.35.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then
(2.25) $<\mu_{j},\;{\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}>=\pm 1.$
By duality we can also define:
Definition 2.36.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. We denote by ${\mathrm{sgn}}(\mu_{j})$ the above value of the product ${\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$ evaluated by $\mu_{j}$.
Remark 2.37.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. Then $L_{{\frac{x_{i}}{d_{i}}}}$, the left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$ induces a permutation on $\overline{\mathcal{B}}$. Because ${\frac{x_{i}}{d_{i}}}=\sum_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})F_{j}$, the determinant $\det(L_{\frac{x_{i}}{d_{i}}})$ of the matrix of the operator $L_{{\frac{x_{i}}{d_{i}}}}$ is $\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})$. On the other hand, $L_{{\frac{x_{i}}{d_{i}}}}$ permutes the normalized basis ¯:=$\overline{\mathcal{B}}:=\{{\frac{x_{i}}{d_{i}}}\}_{i\in\mathcal{I}}$, and therefore, ${\mathrm{sgn}}(x_{i})$ is the signature of the corresponding permutation.
Remark 2.38.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that ${\mathrm{sgn}}(\mu_{j})=\det(L_{\mu_{j}})=\pm 1$ since $\mu_{j}$ determines a permutation of $^{\widehat{\mathcal{B}}}$.
Definition 2.39.
A symmetric hypergroup ${(H,\mathcal{B})}$ is called perfect if it has no nontrivial grouplike elements.
3. Galois symmetries for hypergroups
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any standard element $x_{i}\in\mathcal{B}$, one can write $x_{i}=\sum_{j=1}^{m}{\alpha_{{}_{ij}}}F_{j}$, for some scalars ${\alpha_{{}_{ij}}}\in\mathbb{C}$. Note that ${\alpha_{{}_{ij}}}=\mu_{j}(x_{i})$. Let $\mathbb{K}=\mathbb{Q}({\alpha_{{}_{ij}}})$ be the field obtained by adjoining all ${\alpha_{{}_{ij}}}\in\mathbb{C}$ to $\mathbb{Q}$.
Remark 3.1.
Note that if ${\alpha_{{}_{ij}}}=\alpha_{{}_{ij^{\prime}}}$ for all $i$, then $\mu_{j}(x_{i})=\mu_{j^{\prime}}(x_{i})$, thus $\mu_{j}=\mu_{j^{\prime}}$ and therefore $j=j^{\prime}$.
3.1. Permutation of characters $\mu_{j}$
Let ${(H,\mathcal{B})}$ be any abelian rational hypergroup (i.e. $N^{k}_{ij}\in\mathbb{Q}$). For any character $\mu_{j}:H\rightarrow\mathbb{C}$ and $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, define $\sigma.\mu_{j}\in H^{*}$ as the linear function on $H$ on which the standard basis $\mathcal{B}$ is given by $[\sigma.\mu_{j}](x_{i})=\sigma(\mu_{j}(x_{i}))=\sigma({\alpha_{{}_{ij}}}).$
Lemma 3.2.
For any abelian rational hypergroup ${(H,\mathcal{B})}$, it follows that $\sigma.\mu_{j}:H\rightarrow\mathbb{C}$ is an algebra map. Thus $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$.
Proof.
Indeed, suppose that $x_{i_{1}}x_{i_{2}}=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}$. Since $N^{k}_{{i_{1}},{i_{2}}}\in\mathbb{Q}$,
$[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}](\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}[\sigma.\mu_{j}](x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
On the other hand, note that
$[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$ $=$ $\sigma(\mu_{j}({x_{i_{1}}}))\sigma(\mu_{j}({x_{i_{2}}}))=\sigma(\mu_{j}({x_{i_{1}}})\mu_{j}({x_{i_{2}}}))$
$=$ $\sigma(\mu_{j}({x_{i_{1}}}{x_{i_{2}}}))=\sigma(\mu_{j}(\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}))$
$=$ $\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
Thus $[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$, so that $\sigma.\mu_{j}$ is an algebra map. It is easy to see that, in this case, $\sigma.(\sigma^{\prime}.\mu_{j})=(\sigma\sigma^{\prime}).\mu_{j}.$ It follows that $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$ of all algebra homomorphisms $\mu_{j}:H\rightarrow\mathbb{C}$. ∎
We denote by $\mathcal{J}:=\{1,\dots,m\}$ the set of all indices for the characters $\mu_{j}:H\rightarrow\mathbb{C}$. It follows that for any $j\in\mathcal{J}$, there is an index $\tau(j)=\tau_{\sigma}(j)\in\mathcal{J}$ such that $\sigma.\mu_{j}=\mu_{\tau_{\sigma}(j)}$, and therefore
(3.1) $\sigma({\alpha_{{}_{ij}}})=\alpha_{{}_{i\tau_{\sigma}(j)}}\;\;\text{for all}\;i,j.$
It is easy to see that, in this case, $\tau_{\sigma}$ is a permutation of $\mathcal{J}$.
Corollary 3.3.
Let ${(H,\mathcal{B})}$ be an abelian rational hypergroup. Then $\mathbb{Q}\subseteq\mathbb{K}$ is a Galois extension.
Proof.
Note that Equation (3.1) shows that $\sigma(\mathbb{K})\subseteq\mathbb{K}$, for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Therefore, this is a normal extension and therefore a Galois extension. ∎
Remark 3.4.
Note that the map $\rho:\mathrm{Gal}(\mathbb{K}/\mathbb{Q})\rightarrow\mathbb{S}_{m},\sigma\mapsto\tau_{\sigma}$ is injective. Indeed, if $\tau_{\sigma}=\tau_{\sigma^{\prime}}$, then $\sigma({\alpha_{{}_{ij}}})=\sigma^{\prime}({\alpha_{{}_{ij}}})$, i.e. $\sigma^{-1}\circ\sigma^{\prime}({\alpha_{{}_{ij}}})={\alpha_{{}_{ij}}}$, for all $i,j$. Thus $\sigma=\sigma^{\prime}$ on $\mathbb{K}$.
Proposition 3.5.
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, with the above notations,
(3.2) $\sigma(n_{k})={n_{\tau(k)}}.$
Proof.
Applying $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$ to Equation (2.5), one obtains the following:
$\sum_{i=1}^{m}{h_{i}}\sigma(\alpha_{{}_{il}})\sigma(\alpha_{{}_{i^{*}k}})=\delta_{l,k}\sigma(n_{k}).$
In turn, this equation can be written as
$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{l,k}\sigma(n_{k}).$
On the other hand, by the same orthogonality relation:
$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{\tau(l),\tau(k)}{n_{\tau(k)}}.$
Therefore, for $l=k$, it follows that $\sigma(n_{k})={n_{\tau(k)}}.$
We continue with the following definition.
Definition 3.6.
A hypergroup ${(H,\mathcal{B})}$ is $h$-integral if $h_{i}:=1/N^{1}_{i,i^{*}}\in\mathbb{Z}$, for all $1\leq i\leq m$.
Remark 3.7.
In view of Equation (2.11), the Grothendieck ring of any weakly-integral fusion category $\mathcal{C}$ has a $h$-integral dual ${\widehat{{K(\mathcal{C})}}}$, see [Ost15, Theorem 2.13]. See also Remark 9.1.
Lemma 3.8.
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup with a $h$-integral dual. Then ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral. Moreover, if ${(H,\mathcal{B},\mu_{1})}$ is rational, then $\widehat{h}_{k}=\widehat{h}_{\tau(k)}$, for all $1\leq k\leq m$.
Proof.
By definition,
$n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}\in\mathbb{Z}.$
If ${(H,\mathcal{B},\mu_{1})}$ is rational, then Equation (3.2) can be written as $\sigma\bigg{(}\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{k}}}\bigg{)}=\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{\tau(k)}}}.$ In particular, $\widehat{h}_{k}=\widehat{h}_{\tau(k)}.$, since ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $h$-integral. ∎
4. Burnside’s vanishing property for hypergroups
Let ${(H,\mathcal{B})}$ be any hypergroup. An element $x_{i}\in\mathcal{B}$ is called a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$. Otherwise, $x_{i}$ is called a non-vanishing element.
Let $\mathcal{B}_{0}$ be the set of all vanishing elements of $\mathcal{B}$, and $\mathcal{B}_{1}$ be the set of all non-vanishing elements of $\mathcal{B}$. Thus $\mathcal{B}=\mathcal{B}_{1}\sqcup\mathcal{B}_{0}$. Note that if ${(H,\mathcal{B})}$ is a RN-hypergroup then ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. Note that, in the case of an abelian hypergroup ${(H,\mathcal{B})}$, since $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$, it follows that $x_{i}$ is non-vanishing if and only if it is invertible. Thus, in this case, $\mathcal{B}_{1}$ coincides with the set of all standard elements that are invertible in $H$.
Denote ${\widehat{P}}:=\prod_{j=1}^{m}\mu_{j}$. Note that in any abelian normalizable ${(H,\mathcal{B},\mu_{1})}$ one has:
(4.1) $\prod_{j=1}^{m}\mu_{j}=\sum_{i=1}^{m}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}.$
Proposition 4.1.
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
(4.2) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}}){\widehat{E}}_{i}$
where $\det(L_{{\frac{x_{i}}{d_{i}}}})$ is the determinant of the left multiplication operator by ${\frac{x_{i}}{d_{i}}}$ on $H$.
Proof.
Recall that $<{\widehat{E}}_{i},\frac{x_{j}}{d_{j}}>=\delta_{i,j}$ for all $i,j\in\mathcal{I}$. In order to prove the above equality, it is enough to evaluate both sides at each ${\frac{x_{i}}{d_{i}}}$, and prove that one obtains the same value.
If $x_{i}$ is a vanishing element, then the above left and right hand side both vanish when evaluated at ${\frac{x_{i}}{d_{i}}}$.
If $x_{i}$ is not a vanishing element then
$\big{[}\prod_{j=1}^{m}\mu_{j}\big{]}({\frac{x_{i}}{d_{i}}})=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})=\det(L_{{\frac{x_{i}}{d_{i}}}})$
and the proof is compete. ∎
Dually, applying the above result to ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$, one obtains a decomposition of the basis ^=^${\widehat{\mathcal{B}}}={\widehat{\mathcal{B}}}_{0}\sqcup{\widehat{\mathcal{B}}}_{1}$. The above proposition implies the following:
Proposition 4.2.
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
(4.3) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{\widehat{\mathcal{B}}}_{1}}\det(L_{\mu_{j}})F_{j},$
where $\det(L_{\mu_{j}})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
Recall the definition of an abelian Burnside hypergroup from Definition 1.4. Note that the decomposition $\mathcal{B}=\mathcal{B}_{0}\sqcup\mathcal{B}_{1}$ from above implies the following:
Proposition 4.3.
An abelian hypergroup ${(H,\mathcal{B})}$ is Burnside if and only if one has ${G{(H,\mathcal{B})}}=\mathcal{B}_{1}$.
Corollary 4.4.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if the above identity becomes
(4.4) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
Proof.
It follows from Proposition 4.1. Recall from Definition 2.34 that ${\mathrm{sgn}}(x_{i})$ is the determinant of $L_{\frac{x_{i}}{d_{i}}}$, when $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
Corollary 4.5.
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if
(4.5) $(\prod_{j=1}^{m}\mu_{j})^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
Corollary 4.6.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${G{(H,\mathcal{B})}}$ has odd order. Then ${(H,\mathcal{B})}$ is Burnside if and only if
(4.6) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}$
Proof.
Note that in this case ${\mathrm{sgn}}(x_{i})=1$ for all $x_{i}\in{G{(H,\mathcal{B})}}$, since each ${\frac{x_{i}}{d_{i}}}\in{G{(H,\mathcal{B})}}$ has odd order. ∎
Definition 4.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. We say that the hypergroup ${(H,\mathcal{B})}$ is dual-Burnside if ${\widehat{(H,\mathcal{B})}}$ is Burnside. In other words, for all $\mu_{j}\in{\widehat{\mathcal{B}}}$, the following are equivalent:
  1. (1)
    For all $x_{i}\in{\mathcal{B}}$ then $\mu_{j}(x_{i})\neq 0$,
  2. (2)
    $\mu_{j}\in G\widehat{{(H,\mathcal{B})}}$, i.e $\mu_{j}$ is a grouplike element.
The dual of Corollary 4.4 is the following:
Corollary 4.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if:
(4.7) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}{\mathrm{sgn}}(\mu_{j})F_{j},$
where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
Corollary 4.9.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if :
(4.8) $(\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}})^{2}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}.$
where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
Corollary 4.10.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. If ${{G{\widehat{(H,\mathcal{B})}}}}$ has odd order then ${(H,\mathcal{B})}$ is dual-Burnside if and only if
(4.9) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{{G{\widehat{(H,\mathcal{B})}}}}}F_{j}.$
4.1. Inequalities for $P$ and ${\widehat{P}}$
Proposition 4.11.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $x_{i}\in G{(H,\mathcal{B})}$.
Proof.
First suppose that ${(H,\mathcal{B})}$ is RN. Then $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ by Frobenius-Perron theory. This implies that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|={\prod_{j=1}^{m}}|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$. Then Lemma 2.20 implies that $x_{i}$ is a grouplike element.
Now suppose that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then since $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$ it follows that $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ since $\mathrm{FPdim}(\mu_{j})=1$. This implies again that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$. Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$. Then, again, Lemma 2.20 implies that $x_{i}$ is a grouplike element. ∎
Proposition 4.12.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. One has that ${\widehat{P}}^{2}$ is an idempotent if and only if the non-vanishing elements $x_{i}$ of ${(H,\mathcal{B})}$ satisfy ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$.
Proof.
Note that Equation (4.1) one has ${\widehat{P}}^{4}={\widehat{P}}^{2}$ if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{4}={\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}$ for all $i\in\mathcal{I}$. Therefore if ${\widehat{P}}^{2}$ is idempotent then either ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=0$ or ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ for all $i\in\mathcal{I}$. In the first case, since ${\widehat{P}}({\frac{x_{i}}{d_{i}}})={\prod_{j=1}^{m}}\mu_{j}({\frac{x_{i}}{d_{i}}})$ it follows that $x_{i}$ is a vanishing element. On the other hand if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ implies ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$. ∎
Corollary 4.13.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then ${\widehat{P}}^{2}$ is idempotent if and only if ${(H,\mathcal{B})}$ is Burnside.
Proof.
In this case, by Proposition 4.11, one has $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ for all $j\in\mathcal{J}$ Thus ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$ implies $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=1$ and Lemma 2.20 implies that $x_{i}\in G{(H,\mathcal{B})}$.
Conversely, suppose that ${(H,\mathcal{B})}$ is Burnside. It follows from Equation (4.1) that ${\widehat{P}}=\sum_{x_{i}\in G{(H,\mathcal{B})}}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$. Note that ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\det(L_{\frac{x_{i}}{d_{i}}})$ the operator of left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$. By Corollary 2.32 this is the determinant of a permutation matrix and therefore equal to ${\mathrm{sgn}}(x_{i})=\pm 1$. ∎
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Define also $P={\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$. One has
(4.10) $P=\sum_{j=1}^{m}\mu_{j}(P)F_{j}.$
The dual version of the last three results imply the following:
Proposition 4.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then $|\mu_{j}(P)|\leq 1$ and one has equality if and only if $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$.
Proposition 4.15.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. One has that $P^{2}$ is an idempotent if and only if the non-vanishing elements $\mu_{j}\in{\widehat{\mathcal{B}}}$ satisfy $\mu_{j}(P)=\pm 1$.
Corollary 4.16.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. One has that $P^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside.
4.2. The analogue of Burnside’s theorem
We prove the following more general version of Theorem 1.2. Its proof follows the same lines as the proof of [Bur23, Theorem 2].
Theorem 4.17.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral. Let $x_{i}\in\mathcal{B}$ be an element such that $h_{i}|d_{i}|^{2}>1$ and
(4.11) $h_{i}|\mu_{j}(x_{i})|^{2}\in\mathbb{A}$
for all $j\in\mathcal{J}$. Then there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
Proof.
For any $1\leq i\leq m$, denote $\mathcal{T}_{i}:=\mathcal{T}(x_{i})=\{j\in\mathcal{J}\;|\;\mu_{j}(x_{i})=0\}$ and ${\mathcal{D}}_{i}:=\mathcal{J}\setminus(\mathcal{T}_{i}\cup\{1\})$. One needs to show that $\mathcal{T}_{i}\neq\emptyset$, for any basis element $x_{i}\in\mathcal{B}$ satisfying the hypothesis from the theorem.
Note that the second orthogonality relation can be written as:
(4.12) $\sum_{j=1}^{m}\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}$
Since $\widehat{h}_{1}=1$, for $j=1$, the above equation can be rewritten as:
$\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}=|d_{i}|^{2}+\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}}|\mu_{j}(x_{i})|^{2}$
which gives that
(4.13) $1=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}.$
On the other hand, note that
$n{(H,\mathcal{B},\mu_{1})}=\sum_{j=1}^{m}\widehat{h}_{j}=1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}+\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}.$
Therefore Equation (4.13) can be written as:
(4.14) $1=\frac{1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}-\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}.$
Thus, in order to finish the proof, it is enough to show that
(4.15) $\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}\geq 0,$
since then, it follows that $\frac{1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\geq 1$, i.e. $1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}\geq h_{i}|d_{i}|^{2}$. Since $h_{i}|d_{i}|^{2}>1$, it follows that $\mathcal{T}_{i}\neq\emptyset$.
The inequality from Equation (4.15) can be written as
(4.16) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq 1.$
On the other hand, the weighted AM-GM inequality gives that
(4.17) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq\prod_{j\in{\mathcal{D}}_{i}}\bigg{(}h_{i}\big{(}|\mu_{j}(x_{i})|^{2}\big{)}^{\widehat{h}_{j}}\bigg{)}^{\frac{1}{|{\mathcal{D}}_{i}|}},$
where $|{\mathcal{D}}_{i}|:=\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}$. Note that Equation (3.1) implies that the set ${\mathcal{D}}_{i}$ is stable under the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$. This in turn implies that the product
$P_{i}:=\prod_{j\in{\mathcal{D}}_{i}}(h_{i}|\mu_{j}(x_{i})|^{2})^{\widehat{h}_{j}}$
is fixed by the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, since ${\widehat{h}_{\tau(j)}}=\widehat{h}_{j}$ by Lemma 3.8 . It follows that $P_{i}$ is a rational number. On the other hand, each factor of $P_{i}$ is an algebraic integer (since $\widehat{h}_{j}\in\mathbb{Z}_{{}_{>0}}$), and therefore, the entire product is an integer. Since it is also positive, it must be at least $1$. ∎
Note that Theorem 1.2 follows from the above theorem, since in the case of a fusion ring, $h_{i}=1$ and $\mu_{j}(x_{i})\in\mathbb{A}$, for all $(i,j)\in{\mathcal{I}}\times{\mathcal{J}}$.
Proposition 4.18.
Let ${(H,\mathcal{B})}$ be a rational ARN-hypergroup such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral. Let $x_{i}\in\mathcal{B}$ be an invertible element in $H$ satisfying also Equation (4.11), for all $1\leq j\leq m$. Then $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element.
Proof.
Note that ${(H,\mathcal{B})}$ satisfies the hypothesis of Theorem 4.17. Also, by hypothesis, Equation (4.11) holds for all $1\leq j\leq m$. Suppose that $x_{i}$ is not a grouplike element, i.e. $h_{i}|d_{i}|^{2}>1$, by Lemma 2.17. In this case, by Theorem 4.17, there is $\mu_{j}\in\widehat{B}$ such that $\mu_{j}(x_{i})=0$. Thus, if $x_{i}y=1$ then $10=\mu_{j}(x_{i})\mu_{j}(y)=$ 1, which is a contradiction. ∎
Corollary 4.19.
Any abelian rational normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral and Equation (4.11) is satisfied for all $1\leq i,j\leq m$ is Burnside.
Proof.
By Proposition 4.3, it is enough to show that $\mathcal{B}_{1}={G{(H,\mathcal{B})}}$. Clearly, ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. Suppose that $x_{i}\in\mathcal{B}$ is an element which is invertible in $H$. Since $x_{i}$ is not a vanishing element, it follows from Proposition 4.18 that $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
Corollary 4.20.
Let ${(H,\mathcal{B})}$ be a commutative fusion ring such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral. Then ${(H,\mathcal{B})}$ is Burnside.
The following corollary recovers [Bur23, Theorem 2].
Corollary 4.21.
Any commutative Grothendieck ring of any weakly-integral fusion category is Burnside.
Proof.
It follows from above corollary by Remark 3.7. ∎
Now one can write the corresponding statement of Theorem 4.17 applied to the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ of an abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$. We obtain the following:
Theorem 4.22.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable ${h}$-integral hypergroup such that the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is a rational hypergroup. Let $\mu_{j}\in{\widehat{\mathcal{B}}}$ be an element of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ with $\widehat{h}_{j}>1$, and such that
(4.18) $\frac{\widehat{h}_{j}}{d_{i}^{2}}|\mu_{j}(x_{i})|^{2}\in\mathbb{A},$
for all $i\in\mathcal{I}$. Then, there is $x_{i}\in\mathcal{B}$ such that $\mu_{j}(x_{i})=0$.
Proof.
Note that the condition from Equation (4.11) applied on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ becomes exactly Equation (4.18). ∎
5. Kernels of fusion rings, Brauer’s theorem
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let ${(H,\mathcal{B})}_{+}$ be the set of all elements $x\in H$ such that $x=\sum_{b\in{\mathcal{B}}}x_{b}b$ with $x_{b}\in\mathbb{R}_{\geq 0}$. A standard element $b\in\mathcal{B}$ is called a constituent of $x\in{(H,\mathcal{B})}_{+}$ if $x_{b}>0$.
Let $\psi:H\rightarrow\mathbb{C}$ be a linear character, and $E=E_{\psi}\in H$ be its primitive central idempotent. Then, since $E$ is an eigenvector for any $L_{x}$ with $x\in H_{+}$, it follows that $|\psi(x)|\leq\mathrm{FPdim}(x)$, for any object $x\in{(H,\mathcal{B})}_{+}$.
Definition 5.1.
Let $\mathcal{S}\subseteq\mathcal{B}$ be a subset closed under the involution of ${(H,\mathcal{B})}$, and such that $L:=\mathbb{C}[\mathcal{S}]$ is a subalgebra of $H$. Then ${(L,\;\mathcal{S})}$ is called a sub-hypergroup of ${(H,\mathcal{B})}$.
We mention that the notion of kernels of objects of a fusion category from [Bur16] can be extended to arbitrary ARN-hypergroups.
Definition 5.2.
Let ${(H,\mathcal{B})}$ be any ARN-hypergroup. For any $\psi\in{\widehat{\mathcal{B}}}$, define
$\ker_{{\widehat{(H,\mathcal{B})}}}(\psi):=\{x\in\mathcal{B}\;|\;\psi(x)=\mathrm{FPdim}(x)\}.$
Recall that $^{\widehat{\mathcal{B}}}$ is the set of all algebra morphisms $\mu_{j}:H\rightarrow\mathbb{C}$.
Lemma 5.3.
Let ${(H,\mathcal{B})}$ be any ARN-hypergroup and $\psi\in{\widehat{\mathcal{B}}}$. Then $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is (the basis of) a sub-hypergroup of ${(H,\mathcal{B})}$.
Proof.
Let $x,y\in\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$, i.e. $\psi(x)=\mathrm{FPdim}(x)$ and $\psi(y)=\mathrm{FPdim}(y)$. Suppose that $xy=\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}z$, where $\mathcal{B}^{\prime}$ is the subset of $\mathcal{B}$ consisting of all standard elements for which $N^{z}_{xy}>0$.
It follows that
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)=\psi(x)\psi(y)=\psi(xy)=\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z).$
Using the triangle inequality for the complex absolute-value,
$\mathrm{FPdim}(x)\mathrm{FPdim}(y)$ $=$ $\mid\psi(x)\psi(y)\mid=\mid\psi(xy)\mid=\mid\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z)\mid$
$\leq$ $\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mid\psi(z)\mid\leq\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mathrm{FPdim}(z)$
$=$ $\mathrm{FPdim}(x)\mathrm{FPdim}(y),$
It is easy to see that $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of $H$. ∎
Dually, one can define the following:
Definition 5.4.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. For any $x\in H_{+}$, one can define ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ as the set of all standard elements $\psi\in{\widehat{\mathcal{B}}}$ such that $\psi(x)=\mathrm{FPdim}(x)$.
The dual version of Lemma 5.3 implies that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ is (the basis of) a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$ if ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
5.1. Brauer’s theorem for ARN-hypergroup
For any hypergroup ${(H,\mathcal{B})}$, we define the bilinear function on $H$ such that $m(\sum_{i=1}^{m}\alpha_{i}x_{i},\sum_{i=1}^{m}\beta_{i}x_{i})=\sum_{i=1}^{m}\frac{\alpha_{i}\beta_{i}}{h_{i}}$. In the case of an abelian hypergroup note that Equation (2.2) implies the following:
(5.1) $m(x,y)=\tau(xy^{*})=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x)\mu_{j}(y^{*})\;\;\text{for all}\;\;x,y\in H.$
Lemma 5.5.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. If $x:=\sum_{s\in\mathcal{B}}p_{s}s\in{(H,\mathcal{B})}_{+}$ then
${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s).$
Proof.
Note that $\mathrm{FPdim}(x)=\sum_{\{s\;|\;p_{s}>0\}}p_{s}\mathrm{FPdim}(s)$. If $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
$\mathrm{FPdim}(x)=|\mu_{j}(x)|\leq\sum_{\{s|\;p_{s}>0\}}|\mu_{j}(s)|\leq\sum_{s}p_{s}\mathrm{FPdim}(s)=\mathrm{FPdim}(x),$
and this implies that $\mu_{j}\in\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s)$. The converse is obvious. ∎
Next theorem is a generalization of Brauer’s theorem. Although the proof is the same as the one from [Bur16], we include it here for the completeness of the paper.
For an element $x\in{(H,\mathcal{B})}_{+}$, we denote by $<x>$ the sub-hypergroup generated by $x$. By definition, it is the sub-hypergroup generated by the constituents of all powers $x^{n}$ with $n\geq 1$.
Theorem 5.6.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $x\in{(H,\mathcal{B})}_{+}$. Then $<x>=H$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\{\mu_{1}\}$.
Proof.
Suppose that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\mu_{1}$. Write $x=\sum_{j=1}^{m}\mu_{j}(x)F_{j}$ and then
$x^{n}=\sum_{j=1}^{m}\mu_{j}(x)^{n}F_{j}$
Partition the set $\mathcal{J}=\mathcal{J}_{1}\sqcup\mathcal{J}_{2}\sqcup\cdots\sqcup\mathcal{J}_{s}$ such that the value $\mu_{j}(x)$ is constant on each component $\mathcal{J}_{l}$ of the partition. Denote by $\alpha_{l}:=\mu_{j}(x)$ this constant value of $\mu_{j}(x)$ with $j\in\mathcal{J}_{l}$. Note that by the hypothesis, the partition component containing $\{1\}$ is a singleton.
Also note that for any $s\in\mathcal{B}$ and any $a\in H$, then $m(s,a)=\tau(s^{*}a)$. Thus
${m(s,x^{n})}{}=\tau(s^{*}x^{n})=\sum_{j=1}^{m}\mu_{j}(s^{*})\mu_{j}(x)^{n}\tau(F_{j})=\sum_{l\in\mathcal{P}}\big{(}\sum_{j\in\mathcal{J}_{l}}\tau(F_{j})\mu_{j}(s^{*})\big{)}\alpha_{l}^{n}$
Denote
$a_{l}:=\sum_{j\in\mathcal{A}_{l}}\tau(F_{j})\mu_{j}(s^{*}).$
Note that $a_{1}=n{(H,\mathcal{B})}>0$, by Equation (2.9). If $m(s,x^{n})=0$ for all $n$, then
$\sum_{l\in\mathcal{P}}a_{l}\alpha_{l}^{n}=0$
for all $n\in\mathbb{N}$, and Vandermonde determinant gives $a_{l}=1$. This is a contradiction since $a_{1}\neq 0$.
For the converse, suppose that $<x>=H$. If $\mu_{j}\in\ker_{(H,\mathcal{B})}(x)$, then by Lemma 5.3, $\mu_{j}\in\ker(b)$, for any $b\in\mathcal{B}$. Thus $\mu_{j}=\mu_{1}$. ∎
5.2. The cointegral $\lambda_{(L,S)}$ in terms of kernels
Recall from §1 that for any RN-hypergroup ${(H,\mathcal{B})}$, we denoted by $\lambda_{H}$ the primitive central idempotent corresponding to $\mu_{1}=\mathrm{FPdim}$.
Theorem 5.7.
For any $x\in{(H,\mathcal{B})}_{+}$, define $J_{x}:={\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
$\lambda_{<x>}=\sum_{j\in J_{x}}F_{j}$
Proof.
Suppose that $<x>={(L,\;\mathcal{S})}$. Let $G_{1},\dots,G_{s}$ be the primitive idempotents of ${(L,\;\mathcal{S})}$, and $\psi_{1},\dots,\psi_{s}$ be their associated characters. Moreover, as above one may suppose that $G_{1}=\lambda_{{(L,\;\mathcal{S})}}$, and consequently, $\psi_{1}=\mathrm{FPdim}$.
We look at the restrictions of the characters $\mu_{j}$ at ${(L,\;\mathcal{S})}$. Then, there is a surjective function $f:\{1,\dots,r\}\rightarrow\{1,\dots,s\}$ such that $\mu_{j}\big{|}_{{(L,\;\mathcal{S})}}=\psi_{f(j)}$, for all $1\leq j\leq s$.
With the above notation, it will be shown that $f^{-1}(0)=\mathcal{J}_{x}$ Indeed, since $x$ is a faithful element of ${(L,\;\mathcal{S})}$, it follows that $\ker_{{}_{(L,\;\mathcal{S})}}(x)=\{\psi_{1}\}$. On the other hand, note that $\mu_{j}(x)=\mu_{j}\big{|}_{{}_{(L,\;\mathcal{S})}}(x)=\psi_{f(j)}(x)$. Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\psi_{f(j)}\in\ker_{{}_{(L,\;\mathcal{S})}}(x)$, i.e. $f(j)=0$. ∎
6. The adjoint sub-hypergroup and grouplike elements
6.1. Notion of the center of a standard element
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. We also define the notion of center of a standard element $x_{i}\in\mathcal{B}$:
$Z_{{}_{(H,\mathcal{B})}}(x_{i})=\{\psi\in{\widehat{\mathcal{B}}}\;|\;|\psi(x_{i})|=d_{i}\}.$
If ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup, then as in Lemma 5.3, it is easy to see that $Z_{{}_{(H,\mathcal{B})}}(x_{i})$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$.
Define also the dual notion of the center,
$Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)=\{x_{i}\in{\mathcal{B}}\;|\;|\psi(x_{i})|=d_{i}\}.$
Again, as in Lemma 5.3, it is easy to see that $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of ${(H,\mathcal{B})}$ if ${(H,\mathcal{B})}$ is an ARN-hypergroup..
6.2. Adjoint sub-hypergroup
Proposition 6.1.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}\in H$. Then
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$
Proof.
Note that
$|\psi(I(1))|\leq\sum_{i=1}^{m}|h_{i}\psi(x_{i}x_{i^{*}})|=\sum_{i=1}^{m}h_{i}|\psi(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}d_{i}^{2}=\mathrm{FPdim}{(H,\mathcal{B})}.$
So if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi\in\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$. The converse is immediate. ∎
Definition 6.2.
Denote by ${{(H,\mathcal{B})}_{ad}}$ the sub-hypergroup $<I(1)>$ of $H$ generated by $I(1)$. It is called the adjoint sub-hypergroup of $H$.
6.3. On the support $J_{{(L,\;\mathcal{S})}}$
Definition 6.3.
Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian hypergroup ${(H,\mathcal{B})}$. Its support is the subset $J_{{(L,\;\mathcal{S})}}$ such that
(6.1) ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}=\sum\limits_{j\in{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}}F_{j}.$
The next result is a generalization of [Bur20, Lemma 6.4].
Proposition 6.4.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. With the above notations,
${J_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}\}.$
Proof.
By Theorem 5.7, it is enough to show the following:
${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}({(H,\mathcal{B})})\}.$
Equation (2.5) gives that $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}$, and therefore, $n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}$ if and only if $|\mu_{j}(x_{i})|^{2}=d_{i}^{2}$, for all $i\in\mathcal{I}$. This is equivalent to $\mu_{j}\in Z_{{}_{(H,\mathcal{B})}}(x_{i})$, for all $i\in\mathcal{I}$. The result follows by Proposition 6.1. ∎
Theorem 6.5.
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only $j\in J_{{{(H,\mathcal{B})}_{ad}}}$.
Proof.
It follows from Proposition 6.4 and Lemma 2.24.
Corollary 6.6.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then
(6.2) $\lambda_{{}_{{{(H,\mathcal{B})}_{ad}}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}\widehat{F}_{j}$
Proof.
It follows from Equation (6.1) and Theorem 6.5. ∎
There is a dual version of Theorem 6.5 and its Corollary 6.6:
Proposition 6.7.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $x_{i}\in{G{(H,\mathcal{B})}}$ if and only $i\in\mathcal{I}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}$.
Dually one obtains also the following:
Corollary 6.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
(6.3) $\lambda_{{}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
Proof of Theorem 1.7 and Theorem 1.8.
Proof.
If ${(H,\mathcal{B})}$ is Burnside then by Corollary 4.5,
$\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}.$
Corollary 6.8 implies Equation (1.4). Conversely, if Equation (1.4) holds then
$\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\pm{\widehat{E}}_{i}$
Evaluating both sides at any ${\frac{x_{i}}{d_{i}}}$, one obtains the Burnside property. ∎
By duality one obtains directly Theorem 1.8.
6.4. On the sub-hypergroup generated by $P$
Proposition 6.9.
For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $\mathcal{J}_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\cap{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
Proof.
We prove that $\lambda_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$. Then Equation (6.1) implies the intersection result. Note that for all $s\in\mathcal{S}$ and $t\in\mathcal{T}$, $st{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}d_{t}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$. Lemma 5.3 shows that for any constituent $u$ of $st$ then $u{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{u}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$ also. Thus the same relation holds for any standard element of ${(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}$. ∎
Corollary 6.10.
For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ if and only if ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
Proof.
If ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ then ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}={\lambda_{{}_{{(K,\mathcal{T})}}}}$ and therefore equation (6.1) implies ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$. The converse follows from Proposition 6.9. ∎
Corollary 6.11.
For any two elements $P,Q\in{(H,\mathcal{B})}_{+}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $<P>\subseteq<Q>$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(Q)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$.
Recall $P:=\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}$ for any ${(H,\mathcal{B},\mu_{1})}$.
Proposition 6.12.
In any abelian RN-hypergroup ${(H,\mathcal{B})}$ one has
$<P^{2}>={{(H,\mathcal{B})}_{ad}}.$
Proof.
One can prove first that ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})={\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Indeed $\big{|}\psi(P)\big{|}=\prod_{i=1}^{m}\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$, then $\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}=1$, i.e $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, by Proposition 6.1. Conversely, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$ then $\psi({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}^{2}=1$. Thus, if $i\neq i^{*}$ then grouping together $x_{i}$ and $x_{i^{*}}$ in $P$, one obtains $\psi({\frac{x_{i}}{d_{i}}})\psi(\frac{x_{{i^{*}}}}{{d_{i^{*}}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})^{2}\big{|}=1$. Thus,
$\psi(P)=\prod_{i={i^{*}}}\psi({\frac{x_{i}}{d_{i}}}).$
Note that if $i=i^{*}$ then $\psi({\frac{x_{i}}{d_{i}}})=\pm 1$, since in this case $\psi(x_{i})=\psi(x_{i^{*}})\in\mathbb{R}$. Thus, $\psi(P)=\pm 1$ and $\psi(P^{2})=1$ for all $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
It follows that $\lambda_{<P^{2}>}=\lambda_{<I(1)>}$ by Theorem 5.7, and so $<P^{2}>={{(H,\mathcal{B})}_{ad}}$. ∎
Corollary 6.13.
In any ARN-hypergroup ${(H,\mathcal{B})}$ one has
${{(H,\mathcal{B})}_{ad}}\subseteq<P>.$
Corollary 6.14.
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup. Then the following holds:
  1. (1)
    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is pointed.
  2. (2)
    ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is perfect.
Remark 6.15.
By §7, ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if the universal grading group is trivial, if and only if there is no non-trivial grading (see Corollary 7.7).
Corollary 6.16.
A simple non-pointed abelian RN-hypergroup ${(H,\mathcal{B})}$ has a perfect dual.
Proof.
By simplicity ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$, and therefore ${\widehat{(H,\mathcal{B})}}$ is perfect. ∎
Corollary 6.17.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup with ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ RN. Then the following holds:
  1. (1)
    ${\widehat{(H,\mathcal{B})}}_{ad}=\mathbb{C}$ if and only if ${(H,\mathcal{B})}$ is pointed,
  2. (2)
    ${\widehat{(H,\mathcal{B})}}_{ad}={\widehat{(H,\mathcal{B})}}$ if and only if ${(H,\mathcal{B})}$ is perfect.
7. On the universal grading of hypergroups
We remark that the universal grading construction for fusion rings from [GN08, §2] works word for word in the ARN-hypergroup settings. For the completeness of the paper, we sketch its main steps below.
7.0.1. Based modules are completely reducible
Definition 7.1.
Given an ARN-hypergroup ${(H,\mathcal{B})}$, we define a based ${(H,\mathcal{B})}$-module as a pair $(M,\mathcal{M})$ where $M$ is a $H$-module with $\mathcal{M}$ a finite free $H$-basis for $M$, and if $x_{i}m_{j}=\sum_{k}t^{k}_{ij}m_{k}$, then $t^{k}_{ij}=t^{k^{*}}_{i^{*}j}$, for all $m_{k}\in\mathcal{M}$ and all $x_{i}\in\mathcal{B}$.
Note that the above condition is equivalent to $(L_{x_{i}^{*}})^{t}=L_{x_{i}}$, where $L_{x_{i}}$ is the matrix of the left multiplication operator by $x_{i}$ on $M$.
Then one can define a inner product on $M$ such that for all $m_{i},m_{j}\in\mathcal{M}$
(7.1) $(m_{i},m_{j})=\delta_{i,j}$
This inner product can be used to prove that any based ${(H,\mathcal{B})}$-modules is completely reducible, as in [Ost03a, Lemma 2.1].
Let ${(H,\mathcal{B})}$ be a RN-hypergroup and $x,y\in{(H,\mathcal{B})}_{+}$. We say that $x$ contains $y$ (or $y$ is a constituent of $x$) if the difference $x-y$ is in ${(H,\mathcal{B})}_{+}$.
Definition 7.2.
Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup, $G$ be a finite group. Suppose that $\mathcal{B}=\sqcup_{g\in G}{\mathcal{B}}_{g}$ is a partition of $\mathcal{B}$, and let $H_{g}:=\mathbb{C}[\mathcal{B}_{g}]$. The decomposition ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ is called a grading of ${(H,\mathcal{B})}$ by $G$ if $H_{g}H_{h}\subseteq H_{gh}$, for all $g,h\in G$. The grading is called faithful if $\mathcal{B}_{g}\neq\emptyset$, for all $g\in G$.
Let ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ be a grading of ${(H,\mathcal{B})}$, and define $R_{g}:=\sum_{x_{i}\in\mathcal{B}_{g}}h_{i}d_{i}x_{i}$, for all $g\in G$. The analogue of [ENO05, Proposition 8.20] holds in the settings of ARN-hypergroup:
(7.2) $\mathrm{FPdim}(R_{g})=\frac{\mathrm{FPdim}{(H,\mathcal{B})}}{|G|}.$
Indeed, consider $R:=\sum_{g\in G}R_{g}$, and note that $\frac{1}{\mathrm{FPdim}{(H,\mathcal{B})}}R=F_{1}$, the idempotent corresponding to $\mathrm{FPdim}$ in Corollary 2.8. Thus, $RR_{h}=\mathrm{FPdim}(R_{h})R$. Since $R_{g}R_{h}\subseteq R_{gh}$, one obtains from the previous equality that $R_{g}R_{h}=\mathrm{FPdim}(R_{h})R_{gh}$, for all $g,h\in G$. Passing to $\mathrm{FPdim}$, one get that $\mathrm{FPdim}(R_{gh})=\mathrm{FPdim}(R_{g})$, for all $g,h\in G$.
7.1. Constructing the universal grading
Proposition 7.3.
Let ${(H,\mathcal{B})}$ be a RN-hypergroup. As above, define ${I(1)}:=\sum\limits_{a\in\mathcal{B}}h_{a}aa^{*}$. With the above notations:
  1. (1)
    $I(1)$ is central in $H$.
  2. (2)
    Every based left $H_{ad}$-submodule $M$ of $H$ is also a $H_{ad}$-subbimodule of $H$.
  3. (3)
    A based $H_{ad}$-subbimodule $M$ of $H$ is indecomposable if and only if it is indecomposable as $H_{ad}$-bimodule.
Proof.
  1. (1)
    The proof of this fact is the same as in [GN08].
  2. (2)
    Since $m{I(1)}^{n}={I(1)}^{n}m$, for all $m\in M$, this shows that both the left and the right $H_{ad}$-submodules have the same standard constituents.
  3. (3)
    Clear. ∎
Let $H=\oplus_{a\in A}H_{a}$ be a decomposition of $H$ as a $H_{ad}$-bimodule. This also corresponds to a partition $\mathcal{B}=\sqcup_{a\in A}\mathcal{B}_{a}$. One may suppose that there is an element $1\in A$ such that $H_{1}=H_{ad}$.
7.1.1. Definition of $a^{*}$ for any $a\in A$
Note that if $H_{a}$ is an indecomposable $H_{ad}$-bimodule then, $(H_{a})^{*}$ is also an indecomposable $H_{ad}$-bimodule. Therefore, we may suppose that $(H_{a})^{*}=H_{a^{*}}$, for some $a^{*}\in A$. Next lemma has the same proof as [GN08, Lemma 3.4].
Lemma 7.4.
If $x_{a},y_{a}\in\mathcal{B}_{a}$ then, $x_{a}y_{a}^{*}\in H_{ad}$.
Proof.
Note that
$M:=\{y_{a}\in{\mathcal{B}}_{a}\;|\;m(y_{a},x_{a}{I(1)}^{n})>0,\;\;\text{for some}\;\;n\geq 0\}$
is a $H_{ad}$-submodule of $H_{a}$. Since $H_{a}$ is indecomposable, it follows that $M=H_{a}$. It follows that $m(y_{a}^{*},x_{a}{I(1)}^{n})>0$, which implies that $x_{a}y_{a}^{*}$ has a constituent in $H_{1}$. This gives that $x_{a}y_{a}^{*}$ is contained in $x_{a}x_{a}^{*}{I(1)}^{n}\in R_{1}.$
Theorem 7.5.
There is a well-defined binary operation on $A$ such that $a\star b=c$ if and only if $x_{a}\star y_{b}\in H_{c}$, for all $(x_{a},y_{b})\in\mathcal{B}_{a}\times{\mathcal{B}}_{b}$. It defines a group structure on $A$, with $a^{*}$ the inverse of $a$, and $1$ the unit of $A$.
Proof.
As for the proof of [GN08, Theorem 3.5]. ∎
Definition 7.6.
Above group $A$ is called the universal grading group of ${(H,\mathcal{B})}$. It is denoted $U_{H}$ when the standard basis $\mathcal{B}$ is implicitly understood.
Corollary 7.7.
Every ARN-hypergroup ${(H,\mathcal{B})}$ has a canonical faithful grading by the group $U(H)$. Any other faithful grading of ${(H,\mathcal{B})}$ by a group $G$ is determined by a surjective homomorphism $\pi:U(H)\rightarrow G$.
Proof.
Let ${(H,\mathcal{B})}=\oplus_{g\in G}(S_{g},\mathcal{D}_{g})$ be another grading. It follows that $xx^{*}\in\mathcal{D}_{1}$, for any $x\in\mathcal{D}_{g}$, and therefore, $\mathcal{B}_{ad}\subseteq\mathcal{D}_{1}$. Moreover, each $S_{g}$ is a ${H_{ad}}$-subbimodule of $H$. Since $H_{a}$ is an indecomposable module, then $H_{a}\subseteq S_{\pi(a)}$, for some well-defined $\pi(a)\in G$. It follows that $\pi:U(H)\rightarrow G$ is an epimorphism of groups, since we started with a grading. ∎
7.2. On the dual universal grading group and invertible
In this subsection, we use the notation $H$ for a hypergroup ${(H,\mathcal{B})}$, and $\widehat{H}$ for its dual ${\widehat{(H,\mathcal{B})}}$. The goal of this subsection is to prove the following:
Theorem 7.8.
Let $H$ be an abelian RN-hypergroup. Then,
$U_{H}\simeq{G(\widehat{H})}$
Dually one has the following:
Theorem 7.9.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
$U_{\widehat{H}}\simeq G(H)$
as abelian groups.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying any of the two hypothesis from Lemma 2.23. Recall that this means that al least one of the following holds:
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
By Corollary 2.32, for any such hypergroup, if $\mu_{j}\in{G(\widehat{H})}$ then $\mu_{j}^{n}=\mu_{1}$, for some $n\geq 0$. Thus, for all $i$, $\mu_{j}(x_{i})=\omega_{ij}d_{i}$ for some complex number $|\omega_{ij}|=1$.
Definition 7.10.
Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23. For any standard element $x_{i}\in\mathcal{B}$, we define a morphism $\omega_{i}:{G(\widehat{H})}\rightarrow\mathbb{C},\mu_{j}\mapsto\omega_{ij}=\mu_{j}({\frac{x_{i}}{d_{i}}})$.
Next lemma shows that $\omega_{i}$ is a character of ${G(\widehat{H})}$.
Lemma 7.11.
Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23. For any $x_{i}$, then $\omega_{i}$ is a linear character of ${G(\widehat{H})}$.
Proof.
Since ${\frac{x_{i}}{d_{i}}}:\widehat{H}\rightarrow\mathbb{C}$ are algebra characters, then their restrictions ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}$ are also algebra characters. Note that ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}=\omega_{i}$. ∎
Let ${(H,\mathcal{B})}$ be a satisfying any of the two hypothesis from Lemma 2.23. Given $\psi\in\widehat{{G(\widehat{H})}}$, let
(7.3) $\mathcal{B}_{\psi}:=\{x_{i}\in\mathcal{B}\;\mid\;\omega_{i}=\psi\}=\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j}),\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
Note that the above definition can be rewritten as
$x_{i}\in\mathcal{B}_{\psi}\iff{\frac{x_{i}}{d_{i}}}\big{|}_{G(\widehat{H})}=\psi$
Proposition 7.12.
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then, Equation (7.3) defines a grading on ${(H,\mathcal{B})}$.
Proof.
It reduces to show that if $x_{i}\in\mathcal{B}_{\psi}$ and $x_{k}\in\mathcal{B}_{\phi}$, then $x_{i}x_{k}$ has all constituents in $\mathcal{B}_{\psi\phi}$. Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j})$ and $\mu_{j}(\frac{x_{k}}{d_{k}})=\phi(\mu_{j})$, for all $\mu_{j}\in{G(\widehat{H})}$. It follows that
$\psi\phi(\mu_{j})=\psi(\mu_{j})\phi(\mu_{j})=\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j}(\frac{x_{k}}{d_{k}})=\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}).$
Now, since $|\psi\phi(\mu_{j})|=1$ and ${\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}=\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}\frac{x_{l}}{d_{l}}$, the triangle inequality implies that $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$, for all $l$ with $N^{l}_{ik}>0$. Indeed, note that
$1=|\psi\phi(\mu_{j})|=|\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}|\mu_{j}(\frac{x_{l}}{d_{l}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}=1,$
and this forces $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$. ∎
7.3. Computations of the regular components
For any grading component $\mathcal{B}_{\psi}$, we define (as above) the regular part of the grading:
$R_{\psi}=\sum_{x_{i}\in\mathcal{B}_{\psi}}h_{i}d_{i}x_{i}.$
Lemma 7.13.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. With the above notations, for all $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$,
(7.4) $F_{j}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
Proof.
For any $j$ such that $\mu_{j}\in{G(\widehat{H})}$,
$F_{j}$ $=$ $\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\sum_{x_{i}\in\mathcal{B}_{\psi}}\big{(}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=$
$=$ $\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi^{-1}(\mu_{j})R_{\psi}\big{)}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
If $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ then $n_{j}=n(H)$ by Corrolary 2.22. ∎
Corollary 7.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. With above notations, for any linear character $\psi$ of ${G(\widehat{H})}$, the set $\mathcal{B}_{\psi}$ is not empty.
Proof.
Previous corollary shows that the linear span of $<F_{j}\;|\;\mu_{j}\in{G(\widehat{H})}>$ lies inside the span of $<\mathcal{B}_{\psi}\;|\;\mathcal{B}_{\psi}\neq\emptyset>$. By counting the dimension of these vector spaces, it follows that $\mathcal{B}_{\psi}$ is not empty, for any $\psi$ in the dual. ∎
From Equation (7.4), it follows immediately:
Lemma 7.15.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. Then
$R_{\psi}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}\psi(\mu_{j})F_{j}\big{)}.$
Proof of Theorem 7.8
By Lemma 7.15, if $\mu_{j}=\mu_{1}$, then
$R_{1}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}F_{j}\big{)}=\lambda_{{H_{ad}}}.$
Corollary 6.10 proves that $\mathcal{B}_{1}=\mathcal{B}_{ad}$. Therefore the above grading
$\mathcal{B}=\bigsqcup_{\psi\in\widehat{G(\widehat{H})}}{\mathcal{B}}_{\psi}$
coincides with the universal grading of ${(H,\mathcal{B})}$. Thus $U_{H}={G(\widehat{H})}$. By duality, $U_{\widehat{H}}=G(H)$ if ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is abelian.
7.4. On the perp of the adjoint subring
Recall from Definition 1.6 that an abelian RN-hypergroup is called dualizable if ${\widehat{(H,\mathcal{B})}}$ is also RN.
For any sub-hypergroup ${(L,\;\mathcal{S})}$ of an abelian dualizable hypergroup ${(H,\mathcal{B})}$, define
$\mathcal{S}^{\perp}:=\{\mu_{j}\;|\;\mu_{j}(s)=\mathrm{FPdim}(s)\;\text{for all}\;s\in\mathcal{S}\}.$
With the above notations, $\mathcal{S}^{\perp}=\bigcap_{s\in\mathcal{S}}\ker_{(H,\mathcal{B})}(s)$. Moreover, [Har79, Proposition 2.11] implies that $\mathcal{S}^{\perp}$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$, and $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$.
Corollary 7.16.
Let $H$ be an abelian dualizable RN-hypergroup. With the above notations,
${G(\widehat{H})}^{\perp}=H_{ad}\text{ and }G(H)^{\perp}={(\widehat{H})_{ad}}.$
Proof.
Consider the trivial character $\psi_{1}=\mu_{1}\big{|}_{G(\widehat{H})}$. It follows that
${H_{ad}}=B_{\psi_{1}}$ $=$ $\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=1\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
$=$ $\bigcap_{\mu_{j}\in{G(\widehat{H})}}\ker_{\widehat{H}}(\mu_{j})={G(\widehat{H})}^{\perp}.$
By duality, one obtains the other equality. ∎
7.5. Quotients of hypergroups
In this subsection, we define the quotient hypergroup of a RN-hypergroup ${(H,\mathcal{B})}$ by a sub-hypergroup ${(L,\;\mathcal{S})}$. This is the analogue of the quotient construction from [Har79], in the case of probability groups. Recall that a probability group is a symmetric normalized RN-hypergroup.
One can define the following equivalence relation on $\mathcal{B}$. For two elements $a,b\in\mathcal{B}$, say $a\sim_{\mathcal{S}}b$ if and only if there are $s_{1},s_{2}\in{\mathcal{S}},x\in{\mathcal{B}}$ such that $m(x,as_{1})>0$ and $m(x,s_{2}b)>0$.
Let $[a]_{\mathcal{S}}$ denote the equivalence class of any $a\in\mathcal{B}$, with respect to $\sim_{\mathcal{S}}$. Define $H//{\mathcal{S}}$ as the set of all these equivalence classes of $\sim_{\mathcal{S}}$. Recall that $\lambda_{\mathcal{S}}$ is the primitive central idempotent of ${\mathcal{S}}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
For an element $a\in\mathcal{B}$, one can see that $[a]_{\mathcal{S}}=[b]_{\mathcal{S}}$ if and only if $\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}=\lambda_{\mathcal{S}}b\lambda_{\mathcal{S}}$. Therefore there is a set bijection
$\begin{array}[]{cccc}\phi:&\mathbb{C}[H//{\mathcal{S}}]&\to&\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}\\ &[a]_{\mathcal{S}}&\mapsto&\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}\end{array}$
Then, it easy to verify that $H//{\mathcal{S}}$ becomes a ARN-hypergroup, with the multiplication inherited from $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, via the above isomorphism. We denote by $\overline{m}({[c]},\;[a][b])$ the multiplicity structure of $H//{\mathcal{S}}$. Therefore,
$[a][b]=\sum_{[c]\in H//{\mathcal{S}}}\overline{m}({[c]},\;[a][b])[c].$
We write shortly $[a]$ instead of $[a]_{\mathcal{S}}$ when no confusion is possible. Following [Har79], if ${(H,\mathcal{B})}$ is abelian, one can show that
(7.5) $\overline{m}({[c]},\;[a][b])=\sum_{w\in[c]}m(w,ab).$
It was proven in [Har79, Proposition 2.11] that if ${(H,\mathcal{B})}$ is an abelian dualizable probability group then,
$\begin{array}[]{cccc}\alpha:&{\mathcal{S}}^{\perp}&\to&\widehat{H//{\mathcal{S}}}\\ &\psi&\mapsto&\alpha(\psi)\end{array},$
with $\alpha(\psi)([a]_{\mathcal{S}}):=\psi(a)$, is an isomorphism of probability groups. It is clear that the same results remains true for dualizable ARN-hypergroups.
7.6. Applying Harrison’s results
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. In this subsection, we also use the short notation $H$ (or $\mathcal{B}$) for this hypergroup, if $\mathcal{B}$ (or $H$) are implicitly understood. For any sub-hypergroup ${(L,S)}$ of an ANR-hypergroup ${(H,\mathcal{B})}$, note that [Har79, Proposition 2.11] implies that there is an isomorphism of hypergroups:
$\widehat{H}//\mathcal{S}^{\perp}\simeq\widehat{\mathcal{S}},\;\;[\mu]_{\mathcal{S}^{\perp}}\mapsto\mu\big{|}_{L}.$
In particular, for $\mathcal{S}=H_{ad}$,
(7.6) $\widehat{H}//G(\widehat{H})\simeq\widehat{H_{ad}}$
Dually, for ${\mathcal{S}}={(\widehat{H})_{ad}}$, one obtains that
(7.7) $H//G(H)\simeq\widehat{{(\widehat{H})_{ad}}}$
For $S=G(H)$, then $\widehat{H}//{(\widehat{H})_{ad}}\simeq\widehat{G(H)}$, and dually, for ${\mathcal{S}}={G(\widehat{H})}$, then $H//H_{ad}\simeq\widehat{{G(\widehat{H})}}.$
8. On lower and upper central series of hypergroups
Let $H$ be a RN-hypergroup. Following [GN08], let $H^{(0)}=H$, $H^{(1)}={H_{ad}}$, and $H^{(n)}=H^{(n-1)}_{ad}$, for all $n\geq 1$.
Definition 8.1.
The non-increasing sequence
$H=H^{(0)}\supseteq H^{(1)}\supseteq\dots\supseteq H^{(n)}\supseteq\dots$
will be called the upper central series.
Definition 8.2.
Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian RN-hypergroup ${(H,\mathcal{B})}$. We define the commutator ${(L,\;\mathcal{S})}^{{co}}$ of ${(L,\;\mathcal{S})}$ in ${(H,\mathcal{B})}$ as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by all standard elements $x\in\mathcal{B}$ such that $xx^{*}\in\mathcal{{\mathcal{S}}}$.
As in [GN08, Remark 4.9], it follows that the linear span of ${\mathcal{S}}^{co}$ is already a sub-hypergroup of $H$. Moreover, [GN08, Lemma 4.15] works as well in the settings of ARN-hypergroups. Thus,
(8.1) $(\mathcal{S}^{{co}})_{ad}\subseteq\mathcal{S}\subseteq(\mathcal{S}_{ad})^{{co}}$
for any sub-hypergroup $\mathcal{S}$ of $\mathcal{B}$. By applying $\perp$ to (8.1), one obtains
(8.2) $\big{(}(\mathcal{S}^{{co}})_{ad}\big{)}^{\perp}\supseteq\mathcal{S}^{\perp}\supseteq\big{(}(\mathcal{S}_{ad})^{{co}}\big{)}^{\perp}.$
For an ARN-hypergroup ${(H,\mathcal{B})}$, define $H_{(0)}=\mathbb{C}$ and $H_{(n)}=(H_{(n-1)})^{{co}}$, for all $n\geq 1$. Then, $H_{(1)}=\mathbb{C}^{{co}}=G({(H,\mathcal{B})})=H_{pt}$.
Definition 8.3.
Let $H$ be an ARN-hypergroup. The non-decreasing sequence
$\mathbb{C}=H_{(0)}\subseteq H_{(1)}\subseteq\dots\subseteq H_{(n)}\subseteq\dots$
will be called the lower central series of $H$.
Similarly to [GN08, Theorem 4.16], one can prove that $H^{(n)}=\mathbb{C}$ if and only if $H_{(n)}=H$, where $H$ is an ARN-hypergroup.
Definition 8.4.
An ARN-hypergroup is called nilpotent when $H^{(n)}=\mathbb{C}$, for some $n\in\mathbb{N}$.
Lemma 8.5.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. For any sub-hypergroup ${(L,\;\mathcal{S})}$,
(8.3) $(\mathcal{S}^{\perp})_{ad}\subseteq(\mathcal{S}^{co})^{\perp}.$
Proof.
It reduces to show that $\mu_{j}\mu_{j^{\#}}\in({\mathcal{S}}^{co})^{\perp}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$.
Let $x\in{\mathcal{S}}^{co}$ be a standard element. It follows that $xx^{*}\in{\mathcal{S}}$, and therefore, $\mu_{j}(x){\mu_{j^{\#}}}(x)=\mu_{j}(xx^{*})=\mathrm{FPdim}(x)^{2}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$. This implies
$[\mu_{j}\star\mu_{j^{\#}}]({\frac{x}{\mathrm{FPdim}(x)}})$ $=$ $\mu_{j}({\frac{x}{\mathrm{FPdim}(x)}})\mu_{j^{\#}}({\frac{x}{\mathrm{FPdim}(x)}})$
$=$ $\frac{1}{\mathrm{FPdim}(x)^{2}}\mu_{j}(x)\mu_{j^{\#}}(x)=1,$
which shows that $\mu_{j}\star{\mu_{j^{\#}}}\in({\mathcal{S}}^{co})^{\perp}$. Thus $({\mathcal{S}}^{\perp})_{ad}\subseteq(S^{co})^{\perp}$. ∎
Proposition 8.6.
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. With the above notations, for all $n\geq 0$,
(8.4) $\widehat{H}^{(n)}\subseteq H_{(n)}^{\perp}$
and
(8.5) $H^{(n)}\subseteq(\widehat{H}_{(n)})^{\perp}.$
Proof.
We will prove the first inclusion by induction on $n$. If $n=0$, then
$\widehat{H}^{(0)}=H=\mathbb{C}^{\perp}=H_{(0)}^{\perp}.$
Now, suppose that ${\widehat{H}^{(n)}}\subseteq{H_{(n)}}^{\perp}$. Then
$\widehat{H}^{(n+1)}=({\widehat{H}^{(n)}})_{ad}\subseteq({H_{(n)}}^{\perp})_{ad}\subseteq({H_{(n)}}^{co})^{\perp}=H_{(n+1)}^{\perp}.$
By duality, one obtains the second inclusion. ∎
Observe that the equality holds for $n=1$ also, because
$\widehat{H}^{(1)}={(\widehat{H})_{ad}}=G(H)^{\perp}=H_{(1)}^{\perp}.$
8.1. Proof of Theorem 1.11
Proof.
If ${(H,\mathcal{B})}$ is nilpotent then $H_{(n)}=H$, for some $n$. Then, $\widehat{H}^{(n)}=\mathbb{C}$ by Equation (8.4), and therefore ${\widehat{(H,\mathcal{B})}}$ is nilpotent. Idem for the converse with Equation (8.5).
8.2. Burnside property for nilpotent ARN-hypergroups
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Recall that we call $x_{i}\in\mathcal{B}$ a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
Lemma 8.7.
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup and ${(L,\;\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$. Consider a standard element $x_{i}\in{\mathcal{B}}$ such that $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$. Then, $x_{i}$ is also vanishing element in ${(H,\mathcal{B})}$.
Proof.
Recall that $H//{\mathcal{S}}\simeq\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ as hypergroups, with $[x]\mapsto\lambda_{\mathcal{S}}x\lambda_{\mathcal{S}}$. Moreover, the hypergroup $\widehat{H//{\mathcal{S}}}$ can be identified with a subring of $\widehat{H}$. If $\lambda_{\mathcal{S}}=\sum_{j\in\mathcal{J}_{\mathcal{S}}}F_{j}$, then $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}=\oplus_{j\in\mathcal{J}_{\mathcal{S}}}\mathbb{C}F_{j}$, as a subalgebra of $H$. Thus, the characters $\mu_{j}$, with $j\in\mathcal{J}_{\mathcal{S}}$, have distinct restriction to $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, and cover all the characters of $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$. So if $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$, then $\lambda_{\mathcal{S}}x_{i}\lambda_{\mathcal{S}}$ is a vanishing element in $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, which in turn implies that $x_{i}$ is a vanishing element in $H$. ∎
Lemma 8.8.
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. A grouplike element $g\in{G{(H,\mathcal{B})}}$ is a constituent of $x_{i}x_{i^{*}}$ if and only if $gx_{i}=\mathrm{FPdim}(g)x_{i}$. The set of all grouplike elements that are constituents of $x_{i}x_{i^{*}}$ form a subgroup of ${G{(H,\mathcal{B})}}$.
Proof.
Equation (5.1) gives that
$m(g,x_{i}x_{i}^{*})=\tau(x_{i}x_{i^{*}}g^{*})=\tau(x_{i^{*}}g^{*}x_{i})=m(x_{i^{*}},x_{i^{*}}g^{*})=m(x_{i},gx_{i}).$
Thus, by Lemma 2.27 one has $g^{-1}x_{i}=\mathrm{FPdim}(g^{-1})x_{i}$ if $m(g,x_{i}x_{i}^{*})>0$. The second assertion is a straightforward consequence of this. ∎
Proof of Theorem 1.12 It is enough to prove the following:
Proposition 8.9.
A nilpotent dualizable ARN-hypergroup ${(H,\mathcal{B})}$ is Burnside.
Proof.
We will proceed by induction on the nilpotency class of ${(H,\mathcal{B})}$.
If ${(H,\mathcal{B})}$ is of nilpotency class $1$, then ${(H,\mathcal{B})}$ is pointed and we are done.
Now, suppose that ${(H,\mathcal{B})}$ is nilpotent of class $n$ and that the result holds for the nilpotency classes less than $n$. We will show that any standard element $x_{i}\in\mathcal{B}$ is either a vanishing or a grouplike element. Assume that $x_{i}\notin G(H)$. We will show that $x_{i}$ is a vanishing element.
Consider $[x_{i}]\in H//G(H)$. Since $H//G(H)\simeq{(\widehat{H})_{ad}}$, it is nilpotent, with nilpotency class strictly less that of $H$, by Theorem 1.11. By induction hypothesis, we know that $[x_{i}]$ is either a grouplike element or a vanishing element in $H//G(H)$.
If $[x_{i}]$ is a vanishing element, then Lemma 8.7 shows that $x_{i}$ is a vanishing element.
Now, if $[x_{i}]$ is a grouplike element of $H//G(H)$, then
$\lambda_{G(H)}h_{[x_{i}]}x_{i}x_{i}^{*}\lambda_{G(H)}=\lambda_{G(H)},$
which implies that $h_{[x_{i}]}x_{i}x_{i}^{*}\in G(H)$. Thus, $h_{[x_{i}]}x_{i}x_{i}^{*}=\sum_{g\in G_{1}}g$, for a subgroup $G_{1}\subset G$. Since $x_{i}$ is not a grouplike element in $H$, $G_{1}\neq\{1\}$. Then, there is $\mu_{j}$ such that $\mu_{j}(\Lambda_{G_{1}})=0$ where $\Lambda_{G_{1}}=\frac{1}{|G_{1}|}\big{(}\sum_{x\in G_{1}}x\big{)}$ is the integral of $G_{1}$. This implies that $\mu_{j}(x_{i}x_{i}^{*})=0$. Thus $\mu_{j}(x_{i})=0$, which means that $x_{i}$ is a vanishing element. ∎
Now, since $H$ and $\widehat{H}$ are simultaneously nilpotent, Theorem 1.12 follows.
9. Applications to fusion categories
Let $\mathcal{C}$ be a pivotal fusion category with a commutative Grothendieck ring. As in §1, we denote the set of isomorphism classes of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$ and let :=$\mathcal{I}:=\{1,\dots,m\}$. Let also $d_{i}:=\dim(X_{i})$ be the quantum dimension of $X_{i}$ for all $i\in\mathcal{I}$. For any simple object $X_{i}\in\mathrm{Irr}(\mathcal{C})$ denote by $x_{i}:=[X_{i}]$ the class of $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
Let ${\mathcal{Z}(\mathcal{C})}$ be the Drinfeld center of the fusion category $\mathcal{C}$. The forgetful functor $F:{\mathcal{Z}(\mathcal{C})}\rightarrow\mathcal{C}$ admits a right adjoint functor $R:\mathcal{C}\rightarrow{\mathcal{Z}(\mathcal{C})}$. It is known that $A:=R(\mathbf{1})$ has the structure of a central commutative algebra in $\mathcal{Z}(\mathcal{C})$, where $\mathbf{1}$ is the unit object of $A$. The vector space $\mathrm{{CE}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(\mathbf{1},A)$ is called the set of central elements. The space ${\mathrm{CF}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(A,\mathbf{1})$ is called the space of class functions of $\mathcal{C}$. Recall [Shi17] that ${\mathrm{CF}(\mathcal{C})}\simeq K(\mathcal{C})$, the complexification of the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
For any simple object $X_{i}$ of $\mathcal{C}$ we denote by $\chi_{i}:=\mathrm{ch}(X_{i})\in{\mathrm{CF}(\mathcal{C})}$ its associated character, [Shi17]. The central element space $\mathrm{{CE}(\mathcal{C})}$ has a basis of primitive orthogonal idempotents $E_{i}$ such that $<\chi_{i},E_{j}>=\delta_{i,j}d_{i}$ for all $1\leq i,j\leq m$.
As explained in [Shi17, Theorem 3.8] the adjunction between $F$ and $R$ gives a canonical isomorphism of algebras
(9.1) ${\mathrm{CF}(\mathcal{C})}\xrightarrow{\cong}\mathrm{End}_{{\mathcal{Z}(\mathcal{C})}}(R(\mathbf{1})),$
where ${\mathrm{CF}(\mathcal{C})}$ is the algebra of class functions on $\mathcal{C}$. Since ${\mathcal{Z}(\mathcal{C})}$ is also fusion category we can write $R(\mathbf{1})=\bigoplus_{j=1}^{m}\mathcal{C}^{j}$ as a direct sum of simple objects in ${\mathcal{Z}(\mathcal{C})}$. Recall that $\mathcal{C}^{j}$ are called conjugacy classes for $\mathcal{C}$. The above isomorphism also gives a canonical bijection between the set of central primitive idempotents $\{F_{j}\}_{j=1}^{m}$ and the set of conjugacy classes $\{\mathcal{C}^{j}\}$. We denote as above $\mathcal{J}:=\{1,\dots m\}$.
For any fusion category $\mathcal{C}$, by abuse of notation, the symmetric RN-hypergroup ${(H,\mathcal{B},\mu_{1})}:=(K(\mathcal{C}),\mathrm{Irr}(\mathcal{C}),\mathrm{FPdim})$ is also denoted by $K(\mathcal{C})$. We denote by ${\widehat{{K(\mathcal{C})}}}$ its dual hypergroup.
Recall also from [GN08] the notion of a universal group grading of a fusion category $\mathcal{C}$ and its adjoint subcategory ${\mathcal{C}_{\mathrm{ad}}}$.
Remark 9.1.
By [Bur20, Equation (4.8)], it follows that for any spherical fusion category, $\dim(\mathcal{C}^{j})=\frac{\dim(\mathcal{C})}{n_{j}}$. See also [Ost15, Theorem 2.13] for a related statement. Then Equation (2.11) implies that
(9.2) $\widehat{h}_{j}=\dim(\mathcal{C}^{j})$
in the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$. In particular, for a weakly-integral fusion category $\mathcal{C}$, by [ENO05, Proposition 8.27], $\dim(\mathcal{C}^{j})\in\mathbb{Z}$, and therefore ${\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
Corollary 9.2.
Let $\mathcal{C}$ be a a fusion category with a Burnside commutative Grothendieck ring and let ${\widehat{{K(\mathcal{C})}}}$ be its dual. Then
(9.3) $\prod_{j=1}^{m}\mu_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
where ${\mathrm{sgn}}(x_{i})$ is the determinant of the permutation matrix determined by $L_{{\frac{x_{i}}{d_{i}}}}$ on $K(\mathcal{C})$.
Proof.
Since $K(\mathcal{C})$ is a Burnside hypergroup the result follows from Corollary 4.4. ∎
Note that the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$ is denoted by $\widehat{{\mathrm{CF}(\mathcal{C})}}$ in [Bur21]. Theorem 3.4 from the same paper implies that for any pseudo-unitary fusion category there is a canonical isomorphism of hypergroups
(9.4) $\alpha:\widehat{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})},\;\mu_{j}\mapsto\frac{C_{j}}{\dim(\mathcal{C}^{j})}.$
For a pivotal fusion category, Shimizu has defined $C_{j}:={\mathcal{F}_{\lambda}}^{-1}(F_{j})\in\mathrm{{CE}(\mathcal{C})}$ to be the conjugacy class sums corresponding to the conjugacy class $\mathcal{C}^{j}$. Here $\lambda\in{\mathrm{CF}(\mathcal{C})}$ is a cointegral of $\mathcal{C}$ such that $\langle\lambda,u\rangle=1$, see [Shi17, §5]. Recall also the Fourier transform of $\mathcal{C}$ associated to $\lambda$ is the linear map
(9.5) $\mathcal{F}_{\lambda}:\mathrm{{CE}(\mathcal{C})}\rightarrow{\mathrm{CF}(\mathcal{C})}\;\;\text{given by}\;\;a\mapsto\lambda\leftharpoonup\mathcal{S}(a)$
Since any weakly-integral fusion category is pseudo-unitary, by applying the canonical morphism $\alpha$ one obtains the following:
Corollary 9.3.
For any weakly-internal fusion category $\mathcal{C}$ the following identity holds in $\mathrm{{CE}(\mathcal{C})}$:
(9.6) $\prod_{j=1}^{m}C_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})E_{i}$
where $E_{i}\in\mathrm{{CE}(\mathcal{C})}$ is the primitive central idempotent of $X_{i}\in{\mathcal{C}_{\mathrm{pt}}}$.
For a spherical fusion category note the following:
Proposition 9.4.
Let $\mathcal{C}$ be a spherical fusion category with the Grothendieck ring ${K(\mathcal{C})}$ commutative. Let also $\mu_{j}\in{\widehat{{K(\mathcal{C})}}}$ a character of ${K(\mathcal{C})}$. Suppose that ${\widehat{{K(\mathcal{C})}}}$ is a RN-hypergroup.
Then $\mu_{j}$ is a grouplike element of ${\widehat{{K(\mathcal{C})}}}$ if and only if $\dim(\mathcal{C}^{j})=1$.
Proof.
By Lemma 2.17, $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}d_{j}^{2}=1$. Since ${\widehat{{K(\mathcal{C})}}}$ is normalized it follows that $d_{j}=1$ for all $j\in\mathcal{J}$. Thus $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}=1$. On the other hand, it follows from Equation (9.2) that ${\widehat{h}}_{j}=1$ if and only if $\mathcal{C}^{j}$ is a one dimensional conjugacy class of $\mathcal{C}$. ∎
Suppose now that $\mathcal{C}$ is a pivotal fusion category and ${\mathcal{D}}\subseteq\mathcal{C}$ is a fusion subcategory. There is a subset $\mathcal{J}_{\mathcal{D}}\subseteq\mathcal{J}$ such that
(9.7) $\lambda_{{\mathcal{D}}}=\sum_{j\in\mathcal{J}_{\mathcal{D}}}F_{j}$
since $\lambda_{\mathcal{D}}$ is an idempotent element inside ${\mathrm{CF}(\mathcal{C})}$. Note that this set is denoted by $\mathcal{L}_{\mathcal{D}}$ in [Bur20, §4.2]. Also, this is the corresponding notion of the support from Definition 6.3.
Note that [Bur20, Lemma 4.6 ] implies
(9.8) $J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}=\{j\in\mathcal{J}\;|\;\dim(\mathcal{C}^{j})=1\},$
for any fusion category with a commutative Grothendieck ring ${K(\mathcal{C})}$.
Proposition 9.5.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then ${\widehat{{K(\mathcal{C})}}}$ is Burnside if and only if:
(9.9) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{j\in J_{{\mathcal{C}_{\mathrm{ad}}}}}{\mathrm{sgn}}(\mu_{j})F_{j}$
Proof.
It follows from Corollary 4.8 and Theorem (6.5). ∎
Proof of Theorem 1.9
Proof.
Suppose that the dual ${\widehat{{K(\mathcal{C})}}}$ is Burnside. Squaring Equation (9.9) it follows that
$\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\sum_{j\in J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)}.$
The converse also follows immediately from Proposition 9.5. ∎
Let us provide a dual statement in the group case:
Theorem 9.6.
For any finite nilpotent group $G$ the dual $\widehat{K(G)}$ is Burnside and therefore
(9.10) $\left(\prod_{x_{i}\in\mathrm{Irr}(G)}{\frac{x_{i}}{d_{i}}}\right)^{2}=\frac{|Z(G)|}{|G|}\left(\sum_{x_{i}\in\mathrm{Irr}(G/Z(G))}d_{i}x_{i}\right).$
Proof.
Let $\mathcal{C}=\mathrm{Rep}(G)$. It is easy to verify that $\widehat{K(G)}$ is a real-non negative hypergroup. For, example see [Bur21, Equation (3.12)]. Then Theorem B of [INW99] says that if $G$ is a nilpotent group then ${K(G)}$ is dual-Burnside. ∎
Theorem 1.12 implies the following:
Corollary 9.7.
Suppose that $\mathcal{C}$ is a nilpotent fusion category. Then ${K(\mathcal{C})}$ and ${\widehat{{K(\mathcal{C})}}}$ are both Burnside.
10. Premodular categories
Recall that a premodular category is a braided and spherical fusion category. For the rest of this section suppose that $\mathcal{C}$ is a pseudo-unitary premodular category. By [Shi17, Example 6.14] there is $\mathbb{C}$-algebra map $\mathrm{f}_{Q}:{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})}$ given by the following formula:
(10.1) $\mathrm{f}_{Q}(\chi_{i})=\sum_{i^{\prime}=1}^{m}\frac{s_{ii^{\prime}}}{d_{i^{\prime}}}E_{i^{\prime}}.$
where $S=(s_{ij})$ is the $S$-matrix of $\mathcal{C}$ and $E_{i^{\prime}}$ are the primitive central idempotents of $\mathrm{{CE}(\mathcal{C})}$ as defined in the previous section.
As in [Bur22, §4] there is a braided function $M:\mathcal{I}\rightarrow\mathcal{J}$ such that if
$\mathrm{f}_{Q}(F_{j})=\sum_{i\in\mathcal{B}_{j}}E_{i}$
[…]
Recall from [BB15] the notion of cosets of a fusion category with respect to a fusion subcategory. By [Bur22, Theorem 4.10] two simple objects $X_{i},X_{i^{\prime}}$ of a pseudo-unitary premodular fusion category $\mathcal{C}$ are in the same coset with respect to $\mathcal{Z}_{2}(\mathcal{C})$ if and only if $M(i)=M(i^{\prime})$. For any $j\in\mathcal{J}_{2}$ we denote by $\mathcal{R}_{j}:=\{[X_{i}]\;|\;M(i)=j\}$. The aforementioned theorem implies also that $\mathcal{R}_{j}$ with $j\in\mathcal{J}_{2}$ are exactly the right cosets of $\mathcal{C}$ with respect to $\mathcal{C}^{\prime}:=\mathcal{Z}_{2}(\mathcal{C})$. Denote also ${\mathrm{R}}_{j}:=\sum_{[X_{i}]\in\mathcal{R}_{j}}d_{i}\chi_{i}\in{\mathrm{CF}(\mathcal{C})}$ the regular part of their characters. By [Bur22, Equation (4.20)],
(10.2) $\dim({\mathrm{R}}_{j})=\dim({\mathcal{C}^{\prime}})\dim(\mathcal{C}^{j}),\;\text{for all}\;j\in\mathcal{J}_{2}.$
Proposition 10.1.
Suppose that $\mathcal{C}$ is a weakly-integral premodular category such that $\mathcal{Z}_{2}(\mathcal{C})\subseteq{\mathcal{C}_{\mathrm{pt}}}$ and it acts freely on the set $\mathrm{Irr}(\mathcal{C})$ of isomorphism classes of simple objects of $\mathcal{C}$. Then for any $j\in\mathcal{J}_{2}$ such that $\dim(\mathcal{C}^{j})>1$ there is an $x_{i}$ such that $\mu_{j}(x_{i})\neq 0$.
Proof.
Note that in ${K(\mathcal{C})}$ one has $h_{i}=1$ for all $i$, and therefore ${K(\mathcal{C})}$ is $h$-integral. If $\mathcal{C}$ is weakly-integral, the fact that the dual ${\widehat{{K(\mathcal{C})}}}$ is rational RN follows from [Bur21, Equation (4.5)]. Thus one may try to apply Theorem 4.22 for ${(H,\mathcal{B},\mu_{1})}={K(\mathcal{C})}$. In order to do be able to do this, one needs to verify that the additional condition (4.18) is satisfied for all $1\leq i\leq m$, i.e. $\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}\in\mathbb{A}$ for all $i\in\mathcal{I}$. In [Bur22, Lemma 4.2] it was shown that
(10.3) $\frac{\alpha_{{}_{iM(i^{\prime})}}}{d_{i}}=\frac{s_{{}_{ii^{\prime}}}}{d_{i}d_{i^{\prime}}}=\frac{\alpha_{{}_{i^{\prime}M(i)}}}{d_{i^{\prime}}}.$
for all $1\leq i,i^{\prime}\leq m$. Fix an index $i_{j}\in\mathcal{I}$ such that $M(i_{j})=j$. It follows that
(10.4) ${\alpha_{{}_{ij}}}=\frac{d_{i}}{d_{i_{j}}}\alpha_{i_{j}M(i)}.$
Therefore
[…]
Since $\mathcal{C}$ is spherical one has $d_{i}=\overline{d_{i}}$ and therefore $|d_{i}|^{2}=d_{i}^{2}$. Thus
(10.6) $\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}=\frac{\dim(\mathcal{C}_{j})}{d_{i_{j}}^{2}}\mid\alpha_{i_{j}M(i)}\mid^{2}$
We will show that $\dim(\mathcal{C}^{j})=d_{i_{j}}^{2}$ and then the proof follows since $\alpha_{i_{j}M(i)}\in\mathbb{A}$. Indeed, applying Equation (10.2), then $\dim(\mathcal{C}^{j})=\frac{\dim(R_{j})}{\dim(\mathcal{C}^{\prime})}=d_{i_{j}}^{2}$ since the action of $\mathcal{C}^{\prime}$ is free on $\mathrm{Irr}(\mathcal{C})$. ∎
10.1. Modular category case
In this section let $\mathcal{C}$ be a weakly-integral modular fusion category. We apply the results of previous section to $K(\mathcal{C})$ and prove Theorem 1.13 and Theorem 1.16. Recall the definition of $\mathcal{J}_{{}_{\mathcal{D}}}$ from Equation (9.7).
Composing Drinfeld’s map $\mathrm{f}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\mathrm{{CE}(\mathcal{C})}$ with the inverse of the natural isomorphism $\alpha:\mathrm{{CE}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$ from Equation (9.4) one obtains an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{\mathrm{CF}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$. This can also be written as an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$. It was shown in [Bur21, Theorem 4.1] that in fact this is an isomorphism of normalized hypergroup ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}.$ Recall that $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$.
Remark 10.2.
This isomorphism establishes a canonical bijection between the sets $\mathcal{I}$ and $\mathcal{J}$ which is given by $\mathrm{f}_{Q}(F_{i})=E_{i}$ for all $i\in\mathcal{I}$.
Through the rest of the paper we identify $\mathcal{I}$ and $\mathcal{J}$ via the above bijection.
Corollary 10.3.
Let $\mathcal{C}$ be a modular fusion category. Then $K(\mathcal{C})$ is Burnside if and only if it is dual-Burnside.
Proof.
As mentioned above ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$ is an isomorphism of normalized hypergroup, where $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$. Therefore $\overline{{K(\mathcal{C})}}$ (and thus ${K(\mathcal{C})}$) is Burnside if and only if ${\widehat{{K(\mathcal{C})}}}$ is also Burnside.∎
Lemma 10.4.
Let $\mathcal{C}$ be a modular fusion category. Then with the above notations,
(10.7) $X_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{pt}}})\iff i\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$
Proof.
The result follows since $({\mathcal{C}_{\mathrm{ad}}})^{\prime}={\mathcal{C}_{\mathrm{pt}}}$. Then $X_{i}\in\mathrm{Irr}(({\mathcal{C}_{\mathrm{ad}}})^{\prime})$ if and only if $i\in J_{\mathcal{C}_{\mathrm{ad}}}$ by the second item of [Bur20, Theorem 6.2]. ∎
Theorem 10.5.
In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside the following identity holds:
(10.8) $\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}{\mathrm{sgn}}(x_{j})F_{j}$
Proof.
Since $K(\mathcal{C})$ is Burnside we may apply ${\widehat{\mathrm{f}_{Q}}}^{-1}$ to Equation (9.3). Under the above bijection between $\mathcal{I}$ and $\mathcal{J}$ one has ${\widehat{\mathrm{f}_{Q}}}({\frac{x_{i}}{d_{i}}})=\mu_{i}$, see [Bur20, Theorem 6.1]. Thus the above Equation (9.3) becomes
$\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})F_{i}$
and Lemma 10.4 finishes the proof. ∎
Proof of Theorem 1.13
Proof.
Squaring Equation (10.8) one obtains that $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$. Note that $\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$ by the definition of ${J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$.∎
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside.
Proof of Corollary 1.14
Proof.
In this case every invertible element $x_{j}$ has odd order since this order divides $|{G(\mathcal{C})}|$. Thus every ${\mathrm{sgn}}(x_{j})=1$ for any invertible object $X_{j}\in{\mathcal{C}_{\mathrm{pt}}}$. ∎
Proposition 10.6.
Let $\mathcal{C}$ be a fusion category such that the dual ${\widehat{{K(\mathcal{C})}}}$ is a Burnside hypergroup. Then
$\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
Proof.
For any category $\mathcal{C}$ define $P_{\mathcal{C}}:=\prod_{i=1}^{m}d_{i}$. Suppose that $\big{(}\prod_{i=1}^{m}x_{i}\big{)}^{2}=\sum_{i=1}^{m}N_{i}x_{i}$ with $N_{i}\in\mathbb{Z}_{\geq 0}$. Then equalizing coefficients in Equation (1.7) one obtains the following $\frac{N_{i}}{P_{\mathcal{C}}^{2}}=\frac{d_{i}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. For $i=1$ it follows that $n_{1}\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})=P_{\mathcal{C}}^{2}$ which proves the divisibility. ∎
Proof of Theorem 1.15.
Proof.
By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$ since $\mathcal{C}$ is weakly-integral. Note that since ${\widehat{{K(\mathcal{C})}}}$ is Burnside the first divisibility result follows from Proposition 10.6. Moreover, if $\mathcal{C}$ is nilpotent then [GN08, Corollary 5.3] implies $d_{i}^{2}\mid{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$. Thus $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})$. ∎
Remark 10.7.
In the case of an integral fusion category Equation (1.10) can be written as
$\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}).$
Proof of Theorem 1.16
Proof.
The first divisibility follows from Proposition (10.6). Suppose now that $\mathcal{C}$ is a weakly-integral modular fusion category. By [ENO05, Proposition 8.27] one has that $\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})$ is an integer. By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$. The second item follows since for a modular category one has $\mathrm{FPdim}(\mathcal{C})=\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\cup\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})$. On the other hand, from the first item, $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\subseteq\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})$ which proves the inclusion $\mathcal{V}(\mathcal{C})\subseteq\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})\big{)}$.
Conversely, $\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})\big{)}\subseteq\mathcal{V}(\mathcal{C})$ since $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})\mid\mathrm{FPdim}(\mathcal{C})$ and $d_{i}^{{2}}\mid\mathrm{FPdim}(\mathcal{C})$ for all $1\leq i\leq m$. ∎
Proof of Theorem 1.17
Proof.
Since $g.c.d(d_{i}^{2},\;d)=1$ it follows from Equation (1.14) that $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$. Let now ${\mathcal{D}}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ a pointed fusion subcategory such that $\mathrm{FPdim}({\mathcal{D}})=d$. Since $\mathrm{FPdim}({\mathcal{D}}^{\prime})=N$ and $\mathrm{g.c.d}(d,N)=1$ it follows that ${\mathcal{D}}\cap{\mathcal{D}}^{\prime}=\mathrm{Vec}$ and therefore ${\mathcal{D}}$ is non-degenerate. Thus $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$ and ${\mathcal{D}}^{\prime}$ is non-degenerate as well. ∎
Remark 10.8.
Note that the above Corollary improves [OY23, Theorem 4.5] since the authors also assume the existence of a Tannakian subcategory ${\mathcal{E}}=\mathrm{Rep}(G)\subseteq\mathcal{C}$ such that $\mathcal{C}^{0}_{G}\simeq\mathcal{C}(\mathbb{Z}_{d},q)\boxtimes\mathcal{A}$ for some non-degenerate braided fusion category $\mathcal{A}$.
Proof of Corollary 1.18:
Proof.
Immediate from Theorem 1.17, we just need to check that $g.c.d(d_{i}^{2},\;d)=1$, but it is clear since $d_{i}^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ and $d$ is a square-free factor. ∎
Lemma 10.9.
With the above notations, for an integral perfect modular category, $\mathcal{V}(\mathcal{C})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ and
$d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})\mid P_{\mathcal{C}}^{2}$
Proof.
Since ${\mathcal{C}_{\mathrm{pt}}}$ is trivial then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$, see Remark 11.13. Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ by Theorem 1.16. Since $d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})$ the result follows. ∎
Lemma 10.9 extends trivially to the weakly-integral case because:
Proposition 10.10.
A weakly-integral perfect modular fusion category $\mathcal{C}$ is integral.
Proof.
By the proof of Lemma 10.9 then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$. But [EGNO15, Exercice 9.6.12] states that the adjoint subcategory of a weakly-integral fusion category is integral. ∎
Corollary 10.11.
Let $\mathcal{C}$ be an integral perfect modular fusion category. If $p$ is a prime number dividing $\mathrm{FPdim}(\mathcal{C})$ then $p^{2}\mid\mathrm{FPdim}(\mathcal{C})$.
Corollary 10.12.
Any integral modular category $\mathcal{C}$ such that there is a prime $p$ whose exponent in $\mathrm{FPdim}(\mathcal{C})$ is $1$ (i.e $m\neq 1$) it is not perfect.
Corollary 10.13.
If $\mathcal{C}$ is a perfect, integral, even dimensional modular fusion category then $4\mid\mathrm{FPdim}(\mathcal{C})$.
10.2. On the Statements (1.21) and (1.22) and Conjecture (1.20)
In [LPR23, Theorem 5.6] it was shown that the statements (1.21) and (1.22) are equivalent. Next we will show that any of these statements implies Conjecture 1.20 formulated in [CP22].
Proposition 10.14.
Existence of a perfect odd dimensional modular fusion category implies the existence of a non-pointed simple integral modular fusion category.
Proof.
Suppose that there exists a perfect modular odd-dimensional fusion category $\mathcal{C}$.
Note first that $\mathcal{C}$ has no pointed subcategories since it is perfect. It also has no symmetric subcategories, since then these subcategories are of the type $\mathrm{Rep}(G)$ with $G$ a perfect group. But then $|G|$ is divisible by $4$ and this is a contradiction since the $\mathrm{FPdim}(\mathcal{C})$ is odd.
On the other hand, by a dimension argument, $\mathcal{C}$ has at least one minimal subcategory under the inclusion (and hence a simple subcategory), say ${\mathcal{D}}$. If ${\mathcal{D}}$ is a modular fusion category then ${\mathcal{D}}$ is an example of a non-pointed simple integral modular fusion category.
If ${\mathcal{D}}$ is not modular then ${\mathcal{D}}^{\prime}\cap{\mathcal{D}}$ is a non-trivial symmetric fusion subcategory which as explained above, it is not possible. ∎
Recall that in [CP22] the authors have also shown that Conjecture 1.20 is equivalent to either of the following:
Conjecture 10.15.
Odd-dimensional fusion categories are solvable.
Conjecture 10.16.
Odd-dimensional modular fusion categories are solvable.
11. Applications and extra results
11.1. Perfect Drinfeld center
We will characterize the perfect fusion categories with a perfect center. That will provide a large class of perfect modular fusion categories.
Lemma 11.1.
Let $\mathcal{C}$ be a fusion category. Let $F:\mathcal{Z}(\mathcal{C})\to\mathcal{C}$ be the forgetful functor. Let $F_{1}:{\mathcal{Z}(\mathcal{C})}_{\mathrm{pt}}\rightarrow{\mathcal{C}_{\mathrm{pt}}}$ be the induced group homomorphism. Let $G$ be the universal grading group of $\mathcal{C}$. Then,
  1. (1)
    $\ker F_{1}$ is isomorphic to $\widehat{G_{ab}}$, with $G_{ab}:=G/G^{\prime}$ the abelianization of $G$,
  2. (2)
    $G$ is perfect if and only if $\ker F_{1}$ is trivial (i.e. $F(Z)=1$ implies $Z=1$).
Proof.
By definition of the center $\mathcal{Z}(\mathcal{C})$, see [EGNO15, Definition 7.13.1], $F(Z)=1$ if and only if $Z=(1,\gamma)$ with $\gamma\in\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$. So $\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$ is isomorphic to $\ker F_{1}$. Now, [GN08, Proposition 3.9] states that $\mathrm{Aut}_{\otimes}(\operatorname{id}_{\mathcal{C}})$ is isomorphic to $\widehat{G_{ab}}$. In particular, $\ker F_{1}$ is trivial if and only if $G^{\prime}=G$ (i.e. $G$ is perfect). ∎
Note that Lemma 11.1 covers [DNV15, Lemma 2.1].
Proposition 11.2.
Let $\mathcal{C}$ be a perfect fusion category. Then its universal grading group is perfect if and only if $\mathcal{Z}(\mathcal{C})$ is perfect.
Proof.
Let $Z$ be a simple object of $\mathcal{Z}(\mathcal{C})$ with $\mathrm{FPdim}(Z)=1$. The forgetful functor $F:\mathcal{Z}(\mathcal{C})\to\mathcal{C}$ is a tensor functor, so $\mathrm{FPdim}(F(Z))=1$, but $\mathcal{C}$ is perfect, so $F(Z)=1$. The result follows by Lemma 11.1. ∎
Corollary 11.3.
Let $\mathcal{C}$ be a braided fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if $\mathcal{C}$ is perfect with a perfect universal grading group.
Proof.
Immediate by Proposition 11.2 because $\mathcal{C}$ is isomorphic to a fusion subcategory of $\mathcal{Z}(\mathcal{C})$. ∎
Lemma 11.4.
The universal grading group of $\mathrm{Rep}(G)$ is isomorphic to the center $Z(G)$ of the finite group $G$.
Proof.
Immediate by Theorem 7.8, because a grouplike element in the dual hypergroup of the Grothendieck ring of $\mathrm{Rep}(G)$ corresponds to a column of the character table of $G$ with squared norm $|G|$, so to a conjugacy class of size one, thus to a central element. ∎
Corollary 11.5.
Let $G$ be a finite group. Then $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if $G$ is perfect with a trivial center $Z(G)$.
Proof.
Immediate by Corollary 11.3 and Lemma 11.4, because $\mathrm{Rep}(G)$ is perfect if and only if $G$ is perfect, and (as an abelian group) $Z(G)$ is perfect if and only if it is trivial. ∎
Remark 11.6.
Note that Corollary 11.5 can be proved alternatively using the fact (see [EGNO15, §8.5]) that the simple objects of $\mathcal{Z}(\mathrm{Rep}(G))$ are in bijection with the pairs $(C,V)$ where $C$ is a finite conjugacy class in $G$, and $V$ is an irreducible finite dimensional representation of the centralizer of $g$ in $C$, and the Frobenius-Perron dimension of the object corresponding to a pair $(C,V)$ is $|C|dim_{\mathbb{C}}(V)$.
Corollary 11.7.
Let $G$ be a non-abelian finite simple group. Then $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect.
Proof.
Immediate by Corollary 11.5, because a non-abelian finite simple group is perfect with a trivial center. ∎
By Corollary 11.7, $\mathcal{Z}(\mathrm{Rep}(A_{5}))$ is a perfect integral modular fusion category of $\mathrm{FPdim}$ $360060^{2}=$ 3600. Using Remark 11.6 and GAP, it is of rank $22$ and type
$[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].$
Question 11.8.
Is there a perfect integral modular fusion category of rank less than $22$? of $\mathrm{FPdim}$ less than $3600$?
Recall that a Hopf algebra $A$ is called perfect if $\mathrm{Rep}(A)$ is perfect, i.e. $G(A^{*})$ is a trivial group; that a finite dimensional semisimple Hopf algebra $A$ (over $\mathbb{C}$) is factorizable if and only if $\mathrm{Rep}(A)$ is modular; that $\mathcal{Z}(\mathrm{Rep}(G))=\mathrm{Rep}(D(G))$ where the Hopf algebra $D(G)$ is the Drinfeld double of $G$. So, for any non-abelian finite simple group $G$ then $D(G)$ is a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension $|G|^{2}$.
Question 11.9.
Is there a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension less than $\dim(D(A_{5}))=3600$?
Lemma 11.10.
Let $A$ be a semisimple factorizable Hopf algebra. If $A^{*}$ is perfect then so is $A$.
Proof.
By [Sch01, Theorem 2.3(b)] there is an isomorphism $G(A^{*})\simeq G(A)\cap Z(A)$. If $G(A)=\{1\}$ then clearly $G(A^{*})=\{1\}$. ∎
[…]
11.2. Burnside and integrality properties
Remark 11.11.
Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring $K(\mathcal{C})$. By Theorem 7.8, the dual of $K(\mathcal{C})$ is perfect if and only if its universal grading group is trivial.
Remark 11.12.
The Grothendieck ring of a modular fusion category is perfect if and only if its dual is perfect (as it is isomorphic to its dual).
Remark 11.13.
By Remarks 11.11 and 11.12, a modular fusion category $\mathcal{C}$ is perfect if and only if its universal grading group is trivial (i.e. $\mathcal{C}_{ad}=\mathcal{C}$).
Proof of Theorem 1.10
Proof.
We use the Galois action on $\widehat{\mathcal{B}}$ from Lemma 3.2. Let $\mu_{1}$ be the $\mathrm{FPdim}$ character. By Proposition 3.5, it follows that any Galois conjugate of $\mu_{1}$, say $\mu_{j}=\sigma.\mu_{1}$ has for formal codegree $n_{j}=\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))$. On the other hand, since $\mu_{j}=\sigma.\mu_{1}$ is also a non-vanishing character, it should be a grouplike element, by the definition of a dual-Burnside hypergroup. Thus, its formal codegree should also be $\mathrm{FPdim}(H,{\mathcal{B}})$, which implies that $\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))=\mathrm{FPdim}(H,{\mathcal{B}})$, for all $\sigma$ in the Galois group. Then, we are done since this implies that $\mathrm{FPdim}(H,{\mathcal{B}})$ is a rational number. ∎
The following result is partially due to Andrew Schopieray, see [Sch23].
Corollary 11.14.
Let ${(H,\mathcal{B},\mu_{1})}$ be rational abelian normalizable hypergroup. If the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect and Burnside then ${(H,\mathcal{B},\mu_{1})}$ is integral, i.e $d_{i}\in\mathbb{Q}$.
Proof.
The dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect, so as for the above proof of Theorem 1.10, $\sigma.\mu_{1}=\mu_{1}$ for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Thus $\sigma(d_{i})=d_{i}$, which shows that $d_{i}\in\mathbb{Q}$. ∎
Corollary 11.15.
A commutative dual-Burnside fusion ring with a trivial universal grading group is integral.
Proof.
It follows from Corollary 11.14 and Remark 11.11. ∎
Recall Remark 6.15 about a trivial universal grading group.
Corollary 11.16.
A fusion category with a dual-Burnside commutative Grothendieck ring and a trivial universal grading group is integral.
Proof.
Immediate from Corollary 11.15. ∎
Proof of Theorem 1.23
Proof.
Recall that a modular fusion category is Burnside if and only if it is dual-Burnside (Corollary 10.3). Now, by Remarks 11.11 and 11.12, the universal grading group is trivial, so by Corollary 11.15, it must be integral. The converse follows from [GNN09, Theorem 6.1]. ∎
Remark 11.17.
By Theorem 1.23, a simple non-integral modular fusion category is not (dual-)Burnside. Plenty of modular fusion categories of Lie type (Verlinde categories) are simple non-integral, so is the one from [Sch22, Theorem 1], and the center of Extended-Haagerup, so none is (dual-)Burnside.
11.3. Near-group modular fusion categories
Let $G$ be a finite abelian group and $m$ be a non-negative integer. Let $K(G,m)$ be the fusion ring with basis $G\cup\{\rho\}$ and fusion rules:
$\rho^{2}=\sum_{g\in G}g+m\rho\text{ and }g\rho=\rho g=\rho,$
for all $g\in G$. A fusion category whose Grothendieck ring is $K(G,m)$ is called near-group (or also Tambara-Yamagami when $m=0$), see for example [Izu17].
Proposition 11.18.
There is no modular fusion category whose Grothendieck ring is $K(G,m)$, with $G$ non-trivial and $m>0$.
Proof.
Let $\mathcal{C}$ be such a category and $R=K(\mathcal{C})$ be its Grothendieck ring. Since $\rho^{2}=m\rho+\sum_{g\in G}g$, then $\mathrm{FPdim}(\rho)^{2}=m\mathrm{FPdim}(\rho)+|G|$. Let $x_{1}>0$ and $x_{2}<0$ be the two solutions of the equation $x^{2}-mx-|G|=0$, then $\mathrm{FPdim}(\rho)=x_{1}$.
Any non-trivial character $\eta\in\widehat{G}$ can be extended to a character of $R$ (denoted also by $\eta$) by declaring $\eta(\rho)=0$. On the other hand, the trivial character $\eta=\epsilon$ ($\epsilon(g)=1$, for all $g\in G$) has two extensions on $R$: $\psi_{+}(\rho)=x_{1}$ and $\psi_{-}(\rho)=x_{2}$.
It follows that $\psi_{+}=\mathrm{FPdim}$. Thus
^=${\widehat{\mathcal{B}}}=\{\eta|\;\eta\in\widehat{G},\;\eta\neq\epsilon\}\cup\{\psi_{\pm}\}$
The multiplication on $\widehat{R}$ is given by
$\eta\star\psi_{-}=\eta,\;\eta\star\eta^{\prime}=\eta\eta^{\prime}\in\widehat{G},\;$
Moreover, it is easy to compute that:
$\psi_{-}^{2}=\frac{x_{2}^{2}+|G|}{x_{1}^{2}+|G|}\psi_{+}+\frac{x_{1}^{2}-x_{2}^{2}}{x_{1}^{2}+|G|}\psi_{-}.$
Thus $R$ has RN-dual since $x_{1}^{2}=mx_{1}+|G|>mx_{2}+|G|=x_{2}^{2}$.
Note that $R$ is a self dual hypergroup since $\mathcal{C}$ is modular. In particular, Drinfeld’s map $\mathrm{f}_{Q}:R\rightarrow\widehat{R}$, since an isomorphism of hypergroups, sends grouplike elements of $R$ to grouplike elements of $\widehat{R}$. Since $G(\widehat{R})=1$ and $G(R)=G$ is non-trivial, this shows that $\mathrm{f}_{Q}$ is not bijective, which is a contradiction. ∎
Proposition 11.19.
Let $G$ be a finite group and $\mathcal{C}$ be a modular category whose Grothendieck ring ${K(\mathcal{C})}$ is $K(G,0)$. Then $G\simeq C_{2}$.
Proof.
As above, if $m=0$ then $\psi_{-}^{2}=\mathrm{FPdim}$ and therefore ${\widehat{{K(\mathcal{C})}}}$ has two grouplike elements $\psi_{+}=\mathrm{FPdim}$ and $\psi_{-}$. Note that Drinfeld’s map $\mathrm{f}_{Q}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$ send bijectively grouplike elements of ${K(\mathcal{C})}$ into grouplike elements of ${\widehat{{K(\mathcal{C})}}}$. Since all $g\in G$ are grouplike elements of ${K(\mathcal{C})}$ it follows $|G|=2$. ∎
Remark 11.20.
Ising’s category is the only example of a modular category whose Grothendieck ring is of type $K(C_{2},0)$, see [DGNO07, Appendix B].
Corollary 11.21.
A modular fusion category is near-group if and only if its Grothendieck ring is $K(G,m)$ with $(G,m)=(C_{1},0),(C_{1},1),(C_{2},0)$.
Proof.
Immediate by Propositions 11.18, 11.19, Remark 11.20 and [Ost03b]. ∎
Remark 11.22.
Note that modular (generalized) Tambara-Yamagami categories were classified before, see [Nat13, Lemma 5.3 and Theorem 5.4], [Tho11] and [Sie00].
12. Concrete examples
This section focus on providing some concrete examples of fusion rings/categories to which the main results of the paper apply as categorification criteria, and also some counter-examples showing that some results cannot be generalized.
12.1. Burnside property
Recall that a fusion ring is called Burnside if for every basis element then its fusion matrix has norm $1$ if and only if its determinant is nonzero (in other words, grouplike is equivalent to invertible, see Proposition 4.3). Recall that William Burnside proved that the Grothendieck ring of $\mathrm{Rep}(G)$ is Burnside, for every finite group $G$, and that first author’s [Bur23, Theorem 2] generalizes this result to every weakly-integral fusion category (over $\mathbb{C}$).
A fusion ring $R$ is called $\alpha$-Frobenius if for all basis element $x$ then $\frac{\mathrm{FPdim}(R)^{\alpha}}{\mathrm{FPdim}(x)}$ is an algebraic integer. Kaplansky’s 6th conjecture states that the Grothendieck ring of a complex fusion category is $1$-Frobenius. Under the following bounds:
$\begin{array}[]{c|c|c|c|c|c|c|c|c}\text{rank}(R)&\leq 5&6&7&8&9&10&11&12\\ \hline\cr\mathrm{FPdim}(R)<&1000000&150000&15000&10000&4000&2000&500&200\end{array}$
there are exactly $192$ non-pointed simple integral $1$-Frobenius fusion rings with $\mathrm{FPdim}\neq p^{a}q^{b},pqr$ (the last restriction is motivated by [ENO11]). This is the current state of a classification initiated in [LPW21]. Among them, we found only $4$ non-Burnside ones. They have the type of $\mathrm{Rep}(A_{7})$, but different fusion data.
  • Rank: $9$,
  • FPdim: $2520=2^{3}\times 3^{2}\times 5\times 7$,
  • Type: $[[1,1],[6,1],[10,2],[14,2],[15,1],[21,1],[35,1]]$,
  • Fusion data 1:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&0&1&1&1&0&1&1\\ 1&0&0&0&1&1&1&1&1\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&0&1&1&1&1&2\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&0&0&1&1&1&1&1\\ 0&1&1&0&1&1&0&1&1\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&0&1&1&1&1&1&2\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&1&0&1&1&1&1&2\\ 0&1&0&1&1&1&1&1&2\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&1&1&1&2&1&2&3\\ 0&1&1&1&2&1&2&2&5\\ 0&1&2&2&3&3&3&5&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&1&1&3&2&2&4&4\\ 0&1&1&1&3&2&2&4&4\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&2&2&3&3&3&5&7\\ 0&1&4&4&2&3&5&4&12\\ 1&4&4&4&9&8&7&12&15\end{smallmatrix}$
  • Fusion data 2:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&1&2&1&1&1&1&0\\ 1&0&1&1&1&1&2&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&2&1&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&1&1&1&1&2&1&0\\ 0&1&2&1&1&1&1&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&2&1&1&1&2&1&1\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&2&2&1&2&2&2&2\\ 0&1&1&1&2&1&2&2&5\\ 0&1&1&1&3&3&2&5&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\\ 0&1&0&0&3&2&1&4&5\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&1&1&3&3&2&5&8\\ 0&1&4&4&2&3&5&4&12\\ 1&4&5&5&9&8&8&12&14\end{smallmatrix}$
  • Fusion data 3:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&1&0&0&0&1&1&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&2&0&0&1&1\\ 0&1&0&0&0&3&1&1&0\\ 0&1&1&1&0&1&1&0&1\\ 0&0&0&0&1&1&0&3&1\\ 0&0&1&1&1&0&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 0&1&0&1&0&1&0&0&2\\ 1&0&0&0&0&1&1&0&2\\ 0&0&0&0&1&1&0&2&2\\ 0&0&1&1&1&0&1&1&2\\ 0&1&1&0&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&2&2&3&3&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 1&0&0&0&0&1&1&0&2\\ 0&1&1&0&0&1&0&0&2\\ 0&0&0&0&1&1&0&2&2\\ 0&0&1&1&1&0&1&1&2\\ 0&1&0&1&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&2&2&3&3&4\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&2&0&0&1&1\\ 0&0&0&0&1&1&0&2&2\\ 0&0&0&0&1&1&0&2&2\\ 1&2&1&1&1&1&2&0&3\\ 0&0&1&1&1&0&1&2&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&0&2&3&0&5\\ 0&1&2&2&3&3&3&5&6\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&0&3&1&1&0\\ 0&0&1&1&1&0&1&1&2\\ 0&0&1&1&1&0&1&1&2\\ 0&0&1&1&1&0&1&2&3\\ 1&3&0&0&0&6&2&3&0\\ 0&1&1&1&1&2&2&2&2\\ 0&1&1&1&2&3&2&3&3\\ 0&0&2&2&3&0&2&3&9\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&1&1&1&0&1&1&0&1\\ 0&1&1&0&0&1&1&0&3\\ 0&1&0&1&0&1&1&0&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&1&1&1&2&2&2&2\\ 1&1&1&1&0&2&2&0&4\\ 0&0&0&0&3&2&0&5&4\\ 0&1&3&3&3&2&4&4&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&0&0&0&1&1&0&3&1\\ 0&0&0&0&2&1&0&3&3\\ 0&0&0&0&2&1&0&3&3\\ 0&1&2&2&0&2&3&0&5\\ 0&1&1&1&2&3&2&3&3\\ 0&0&0&0&3&2&0&5&4\\ 1&3&3&3&0&3&5&0&7\\ 0&1&3&3&5&3&4&7&10\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&1&0&1&1&4\\ 0&1&2&2&2&2&3&3&4\\ 0&1&2&2&2&2&3&3&4\\ 0&1&2&2&3&3&3&5&6\\ 0&0&2&2&3&0&2&3&9\\ 0&1&3&3&3&2&4&4&7\\ 0&1&3&3&5&3&4&7&10\\ 1&4&4&4&6&9&7&10&17\end{smallmatrix}$
  • Fusion data 4:
    $\begin{smallmatrix}1&0&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&0&0&0\\ 0&0&0&0&1&0&0&0&0\\ 0&0&0&0&0&1&0&0&0\\ 0&0&0&0&0&0&1&0&0\\ 0&0&0&0&0&0&0&1&0\\ 0&0&0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0&0&0\\ 1&0&0&0&0&1&0&1&0\\ 0&0&0&1&0&0&1&0&1\\ 0&0&1&0&0&0&1&0&1\\ 0&0&0&0&1&2&0&2&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&0&1&1\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0&0&0\\ 0&0&0&1&0&0&1&0&1\\ 1&0&0&2&1&1&2&1&0\\ 0&1&2&1&1&1&1&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&2&1&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0&0&0\\ 0&0&1&0&0&0&1&0&1\\ 0&1&2&1&1&1&1&1&0\\ 1&0&1&1&1&1&2&1&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&1&1&1&2\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0&0&0\\ 0&0&0&0&1&2&0&2&0\\ 0&0&1&1&0&0&1&0&3\\ 0&0&1&1&0&0&1&0&3\\ 1&1&0&0&3&3&0&5&0\\ 0&2&0&0&3&0&2&2&2\\ 0&0&1&1&0&2&1&2&3\\ 0&2&0&0&5&2&2&4&2\\ 0&0&3&3&0&2&3&2&9\end{smallmatrix},$
    $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&2&1&1&1&2&1&1\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&2&2&1&2&2&2&2\\ 0&1&1&1&2&1&2&2&5\\ 0&1&1&1&3&3&2&5&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\\ 0&1&0&0&3&2&1&4&5\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&1&1&3&3&2&5&8\\ 0&1&4&4&2&3&5&4&12\\ 1&4&5&5&9&8&8&12&14\end{smallmatrix}$
The determinant of the second fusion matrix (of each fusion data above) is ±$\pm 36$ whereas its norm is $6$, proving that these integral fusion ring are non-Burnside, and so cannot be categorified.
Similar classification were done in the non-1-Frobenius case, and unexpectedly, most (not all) of the ones we found are non-Burnside (see one example below), so without categorification, which can be seen as an evidence for Kaplansky 6th conjecture.
  • Rank: $7$,
  • FPdim: $798=2\times 3\times 7\times 19$,
  • Type: $[[1,1],[7,1],[8,1],[9,3],[21,1]]$,
  • Fusion data:
    $\begin{smallmatrix}1&0&0&0&0&0&0\\ 0&1&0&0&0&0&0\\ 0&0&1&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&1&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0\\ 1&0&0&1&1&1&1\\ 0&0&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&5\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0\\ 0&0&1&1&1&1&1\\ 1&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&1&1&1&6\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 1&1&2&1&1&2&1\\ 0&1&1&1&2&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&2&2&1\\ 1&1&2&2&1&1&1\\ 0&1&1&2&1&2&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&2&1&2&1\\ 1&1&2&1&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1\\ 0&1&1&1&1&1&5\\ 0&1&1&1&1&1&6\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 1&5&6&7&7&7&8\end{smallmatrix}$
The determinant of the third fusion matrix is $16$ whereas its norm is $8$, proving that this integral fusion ring is non-Burnside, and so cannot be categorified.
12.2. Dual-Burnside property
We made the abbreviation (A)RN for (abelian) real non-negative. Note that [LPW21, Corollary 8.5] means that the Grothendieck ring, if commutative, of a unitary fusion category, has a RN dual (hypergroup).
Question 12.1.
Is the dual of the Grothendieck ring of a complex fusion category always RN?
Recall that a commutative fusion ring $F$ with RN dual is called dual-Burnside when a column of its character table has a zero entry if and only if the squared norm of the column (i.e. the corresponding formal codegree) has value strictly less than $\mathrm{FPdim}(F)$, or alternatively, when $P^{2}$ is an idempotent, where
$P=\prod_{i\in I}\frac{x_{i}}{\mathrm{FPdim}(x_{i})}$
and $\{x_{i}\}_{i\in I}$ is the set of basis elements (Corollary 4.16). We warn the reader that for a non-abelian group $G$, the product of all its elements depends on the ordering, in fact, the set of all possible products forms a $G^{\prime}$-coset, see [DH82].
Recall that [INW99, Theorem B] states that for all finite nilpotent group $G$ then the Grothendieck ring of $\mathrm{Rep}(G)$ is dual-Burnside. It is extended to every nilpotent dualizable ARN-hypergroup by Theorem 1.12. So every commutative nilpotent fusion ring with RN dual is both Burnside and dual-Burnside. Note that all the commutative nilpotent fusion rings we checked have RN dual.
Question 12.2.
Is there a commutative nilpotent fusion ring without RN dual?
Let us call a finite group $G$ dual-Burnside if $K(\mathrm{Rep}(G))$ is dual-Burnside. So a nilpotent finite group is dual-Burnside. There are exactly $144$ finite groups of order less than $32$, and among them, exactly $30$ ones are non-nilpotent, and among them, only $\mathrm{SL}(2,3)$ is dual-Burnside (so the $29$ other ones, e.g. $S_{3}$ and $D_{5}$, are not dual-Burnside).
Observe that a finite group $G$ is centerless if and only if $K(\mathrm{Rep}(G))$ has a perfect dual (see Definition 2.39, Theorem 7.8 and Lemma 11.4). So a centerless finite group is dual-Burnside if and only if every non-$\mathrm{FPdim}$ column of its character table has a zero entry. We checked that every centerless and dual-Burnside finite group $G$ of order $|G|\leq 1000$ is almost simple (i.e. $S\subseteq G\subseteq\mathrm{Aut}(S)$ with $S$ non-abelian simple); but $G=A_{5}\times A_{5}$, of order $3600$, is not almost simple. A finite non-abelian simple group is centerless, so by [Pal] and its answers, all the non-alternating finite simple groups are dual-Burnside, except the Mathieu groups $M_{22}$ and $M_{24}$. The alternating group $A_{n}$, $5\leq n\leq 19$, is dual-Burnside if and only if $n\not\in\{7,11,13,15,16,18,19\}$. Recall that a finite group $G$ is simple if and only if $K(\mathrm{Rep}(G))$ is simple.
12.3. Modular categories
Let us provide some applications of Theorem 1.16. There are exactly $43$ different types for a $\frac{1}{2}$-Frobenius integral fusion ring up to rank $11$ (see [ABPP]), and among them, exactly $7$ ones (below) are excluded from modular categorification by Theorem 1.18.
$[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],$
$[1,1,1,1,1,1,1,1,1,3],[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4].$
For example, let us provide all the fusion data for the first type above:
  • Rank: $6$,
  • FPdim: $12=2^{2}\times 3$,
  • Type: $[1,1,1,1,2,2]$,
  • Fusion data 1:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 1&0&0&1&0&1\\ 0&1&1&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&1&1&0&1&0\\ 1&0&0&1&0&1\end{smallmatrix}$
  • Fusion data 2:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&0&0&0&1\\ 0&0&1&1&1&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&1&1&1&0\\ 1&1&0&0&0&1\end{smallmatrix}$
  • Fusion data 3:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&2\\ 1&1&1&1&0&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&0&0\\ 0&0&0&0&2&0\end{smallmatrix}$
  • Fusion data 4:
    $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&2\\ 1&1&1&1&0&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&0&0\\ 0&0&0&0&2&0\end{smallmatrix}$
They are excluded by Theorem 1.18 from modular categorification, because $3$ is a powerless factor of the $\mathrm{FPdim}$, whereas $3$ does not divide $4$, the number of grouplike elements. Note that the above fusion data $3$ and $4$ can be generalized as follows. Let $G$ be a group of order $n^{2}$, and let $K$ be a group of order $m+1$. Consider the following $\frac{1}{2}$-Frobenius integral fusion ring:
  • Rank: $n^{2}+m$,
  • FPdim: $n^{2}(m+1)$,
  • Type: $[[1,n^{2}],[n,m]]$,
  • Basis: =$\mathcal{B}=\{x_{g}\}_{g\in G}\cup\{\rho_{k}\}_{k\in K\setminus\{e\}}$,
  • Fusion data:
    • $\rho_{e}:=\frac{1}{n}\sum_{g\in G}x_{g}$,
    • $x_{g}x_{h}=x_{gh}$, for all $g,h\in G$,
    • $\rho_{k}\rho_{l}=n\rho_{kl}$, for all $k,l\in K$,
    • $x_{g}\rho_{k}=\rho_{k}x_{g}=\rho_{k}$, for all $(g,k)\in G\times K$,
It is excluded by Theorem 1.16 (1.13) from modular categorification if $n>1$ and $m+1$ has a prime factor not dividing $n$.
Availability of data and materials
The datasets generated during and/or analyzed during the current study are available from the second author on reasonable request.
Declaration
Conflict of interests: The authors declare that they have no conflict of interest.

### v11

@@
 We also treat the dual vanishing result.
-We show that any nilpotent fusion categories satisfy both Burnside’s property and its dual.
+We show that any nilpotent unitary fusion categories satisfy both Burnside’s property and its dual.
 Using Drinfeld’s map, we obtain that the Grothendieck ring of any weakly-integral modular fusion category satisfies both properties.
@@
 In particular we improve [OY23, Theorem 4.5] as follows: A weakly integral modular fusion category of $\mathrm{FPdim}$ $md$ with $d$ square-free coprime with $m$ and $\mathrm{FPdim}(X)^{2}$ for every simple object $X$, has a pointed modular fusion subcategory of $\mathrm{FPdim}$ $d$.
-We also prove some new results on the perfect fusion categories.
+We also present new results on perfect modular fusion categories, including a Cauchy-type theorem.
 1991 Mathematics Subject Classification:
-18M20; 16T30
+18M20; 20N20; 20C15; 18N25; 16T20; 16T30
 The first author is supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS/CCCDI - UEFISCDI, project number PN-III-P4-ID-PCE-2020-0878, within PNCDI III.
-The second author is supported by BIMSA Start-up Research Fund and Foreign Youth Talent Program from the Ministry of Sciences and Technology of China.
+The second author is supported by BIMSA Start-up Research Fund, Foreign Youth Talent Program from the Ministry of Sciences and Technology of China and National Natural Science Foundation of China (NSFC, Grant no.
+12471031)
 1. Introduction
 A classical result of Burnside in the character theory of finite groups states that any irreducible non-linear character of a finite group vanishes on at least one element of the group.
-This can be stated as follows: in the character table of a finite group, the row of every character of degree $\neq 1$ contains a zero entry.
+This can be stated as follows: in the character table of a finite group, the row of every character of degree 1$\neq$ 1≠ 1 contains a zero entry.
 More recently, the dual version of this result was also intensively studied in the literature.
 This dual version describes the group elements that vanish on at least one irreducible character, see [INW99] and the references therein.
-Burnside’s result was extended to all weakly integral fusion categories, initially in the context of modular categories in [GNN09, Appendix], and subsequently in the general framework in [Bur23].
+Burnside’s result was extended to every weakly integral fusion category with a commutative Grothendieck ring, initially in the context of modular categories in [GNN09, Appendix], and subsequently in the general framework in [Bur23].
 The main goal of this paper is to develop an analogue of Burnside’s result for a certain large class of fusion rings, or even more generally, to a certain class of abelian rational normalizable hypergroups.
@@
 This was achieved based on work of Harrison [Har79] on dualizable probability groups and on [ZZ21].
-More generally, this notion of dual of a fusion ring was extended to arbitrary fusion rings in [BP22].
+More generally, this notion of dual of a fusion ring was extended to arbitrary fusion rings in [BP25].
 It was noticed in [Bur21] that the dual of the Grothendieck ring of a pivotal fusion category is isomorphic to the center of the category as defined in [Shi17].
@@
 For any hypergroup ${(H,\mathcal{B})}$, an element $a\in\mathcal{B}$ with
-(1.1) $aa^{*}=N_{a,a^{*}}^{1}1\text{ (in other words, }N_{a,a^{*}}^{c}=\delta_{1,c}N_{a,a^{*}}^{1}\text{)}$
-is called a grouplike element of ${(H,\mathcal{B})}$.
+(1.1) $aa^{*}=N_{a,a^{*}}^{1}1\text{ (or equivalently, }h_{a}aa^{*}=1\text{, with }h_{a}:=1/N_{a,a^{*}}^{1})$
+is called a grouplike element of ${(H,\mathcal{B})}$. In other words, $N_{a,a^{*}}^{c}=\delta_{1,c}N_{a,a^{*}}^{1}$.
 By Lemma 2.32, the set of grouplike elements form a group (with the multiplication structure) denoted ${G{(H,\mathcal{B})}}$.
@@
 Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup.
-By the abelian assumption, $^{\widehat{\mathcal{B}}}$ forms a basis for the dual $H^{*}$, which then is $\mathbb{C}[{\widehat{\mathcal{B}}}]$, and on which one can define a multiplication.
+By the abelian assumption, $^{\widehat{\mathcal{B}}}$ forms a basis for the dual $H^{*}$, which then is $\mathbb{C}[{\widehat{\mathcal{B}}}]$, and on which we can define a multiplication.
 On the basis $^{\widehat{\mathcal{B}}}$, the multiplication $\mu_{i}\star\mu_{j}$ is defined by declaring
 (1.2) $[\mu_{i}\star\mu_{j}](a):=\mu_{i}(a)\mu_{j}(a),\;\text{for all}\;a\in\mathcal{B}$
-and then extending linearly $\mu_{i}\star\mu_{j}$ on the entire ${\mathbb{C}[\mathcal{B}]}$. In this way, one obtains an algebra structure on $H^{*}$.
+and then extending linearly $\mu_{i}\star\mu_{j}$ on the entire ${\mathbb{C}[\mathcal{B}]}$. In this way, we obtain an algebra structure on $H^{*}$.
 It follows that there are some non-zero scalars ${\widehat{p}}_{k}(i,j)\in\mathbb{C}$ such that
@@
 It is also well known that for a normalized hypergroup ${(H,\mathcal{B})}$, then $\mu_{j}(a^{*})=\overline{\mu_{j}(a)}$ for all $a\in\mathcal{B}$.
-We also use the notation ${\widehat{h}}_{j}:={\widehat{p}}_{1}(j,j^{\#})^{-1}$, called the order of $\mu_{j}$ in the dual hypergroup ${\widehat{(H,\mathcal{B})}}$, see Lemma 2.6.
+We also use the notation ${\widehat{h}}_{j}:={\widehat{p}}_{1}(j,j^{\#})^{-1}$, called the order of $\mu_{j}$ in the dual hypergroup $^{\widehat{(H,\mathcal{B})}}$, see Lemma 2.8.
 For any RN-hypergroup ${(H,\mathcal{B})}$, we denote by ${(H,\mathcal{B})}_{ad}$ its adjoint sub-hypergroup, see §6.2.
-As for fusion rings, this is defined as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by the element $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}$, where $h_{i}:=1/N^{1}_{i,i^{*}}$.
+As for fusion rings, this is defined as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by the element $I(1):=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}$, where $h_{i}:=1/N^{1}_{i,i^{*}}$.
 For any RN-hypergroup ${(H,\mathcal{B})}$, the central primitive idempotent corresponding to the character $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is denoted by $\lambda_{H}$.
@@
 Definition 1.5.
-A hypergroup ${(H,\mathcal{B})}$ is called normalizable if one obtains a normalized hypergroup after rescaling it. (see §2.2 for more details.)
+A hypergroup ${(H,\mathcal{B})}$ is called normalizable if we obtain a normalized hypergroup after rescaling it. (see §2.2 for more details.)
 Based on Lemma 2.3 we denote such a hypergroup by ${(H,\mathcal{B},\mu_{1})}$ where $\mu_{1}\in{\widehat{\mathcal{B}}}$.
@@
 1.2. Main results of the paper
-We characterize Burnside dualizable ARN-hypergroups as follows:
 Theorem 1.7.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. The following are equivalent (see Definition 2.4 for $n(\_)$):
-  1. (1)
-    The hypergroup ${(H,\mathcal{B})}$ is Burnside.
-  2. (2)
-    The following identity holds on ${\widehat{(H,\mathcal{B})}}$:
+Let $(H,B,\mu_{1})$ be an abelian normalizable hypergroup such that $G{(H,\mathcal{B})}$ is a finite group. Then $(H,{\mathcal{B}})$ is Burnside if and only if
+$(\prod_{j=1}^{m}\mu_{j})^{2}=\frac{|G{(H,\mathcal{B})}|}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{j\in\mathcal{I}_{G{(H,\mathcal{B})}}}{\widehat{h}}_{j}\mu_{j}\big{)}$
+See Definition 2.5 for $n(\_)$). We characterize Burnside dualizable ARN-hypergroups as follows:
+Theorem 1.8.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
+Then $(H,{\mathcal{B}})$ is Burnside if and only if the following identity holds on $^{\widehat{(H,\mathcal{B})}}$:
 (1.4) $\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\frac{1}{n(\widehat{(H,{\mathcal{B}})}_{ad})}\big{(}\sum_{\mu_{j}\in{{\widehat{\mathcal{B}}}}_{ad}}{\widehat{h}}_{j}\mu_{j}\big{)}.$
-Note that the right hand side of the above equation coincides to the integral $\lambda_{{}_{{\widehat{(H,\mathcal{B})}}_{ad}}}$ of the adjoint hypergroup ${\widehat{(H,\mathcal{B})}}_{ad}$.
+Note that the RHS of the equation above corresponds to the integral $\lambda_{{}_{{\widehat{(H,\mathcal{B})}}_{ad}}}$ of the adjoint hypergroup ${\widehat{(H,\mathcal{B})}}_{ad}$.
+Remark 1.9.
+If the abelian normalizable hypergroup ${(H,\mathcal{B})}$ is either RN or dual RN, then according to Lemmas 2.32 and 2.36, the group $G{(H,\mathcal{B})}$ must be finite.
+As a result, we can apply Theorem 1.7.
+Specifically, by comparing this with Theorem 1.8, we can infer that in the dual RN case, $n(\widehat{(H,{\mathcal{B}})}_{ad})|G{(H,\mathcal{B})}|=n{(H,\mathcal{B})}$.
+By duality, we also have $n((H,{\mathcal{B}})_{ad})|G{\widehat{(H,\mathcal{B})}}|=n{(H,\mathcal{B})}$ in the RN case.
+These equalities can also be obtained from Harrison’s Equations (7.7) and (7.6), although in the more restrictive dualizable ARN case.
 For a fusion category $\mathcal{C}$, its Grothendieck ring is denoted by $K(\mathcal{C})$, and the set of isomorphism class representatives of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$.
 Let $d_{i}:=\mathrm{FPdim}(X_{i})$ be the Frobenius-Perron dimension of $X_{i}$, $i\in{\mathcal{I}}=\{1,\dots,m\}$, and $x_{i}:=[X_{i}]$ be the class of the simple object $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$.
-For any weakly-integral premodular category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is both RN and rational, as demonstrated in [Bur21, Theorem 1.2].
-Furthermore, for any unitary fusion category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is also RN, see [LPW21, ENO21], and primary $3$-criterion in [HLPW24].
-A normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if its dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside (see examples from group theory in §12).
-By duality, one obtains the following:
+For any weakly-integral premodular category $\mathcal{C}$, $^{\widehat{{K(\mathcal{C})}}}$ is both RN and rational, as demonstrated in [Bur21, Theorem 1.2].
+Furthermore, for any unitary fusion category $\mathcal{C}$, $^{\widehat{{K(\mathcal{C})}}}$ is also RN, see [LPW21, ENO25], and primary 333-criterion in [HLPW24].
+A normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if its dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside (see examples from group theory in §12).
+The dual version of Theorem 1.7, is the following:
 Theorem 1.10.
-The following are equivalent for an abelian RN-hypergroup ${(H,\mathcal{B})}$.
+Let $(H,B,\mu_{1})$ be an abelian normalizable hypergroup such that $G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is a finite group. Then $(H,{\mathcal{B}})$ is dual-Burnside if and only if
+$(\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}})^{2}=\frac{|G{\widehat{{(H,\mathcal{B},\mu_{1})}}}|}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}_{G{\widehat{{(H,\mathcal{B},\mu_{1})}}}}}h_{i}d_{i^{*}}x_{i}\big{)}$
+On the other hand, the dual version of Theorem 1.8 is:
+Theorem 1.11.
+The following are equivalent for an ARN-hypergroup ${(H,\mathcal{B})}$.
   1. (1)
@@
     The following identity holds:
-    (1.5) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{n({{(H,\mathcal{B})}_{ad}})}\big{(}\sum_{x_{i}\in\mathcal{B}_{ad}}d_{i}x_{i}\big{)}$
+    (1.5) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{n({{(H,\mathcal{B})}_{ad}})}\big{(}\sum_{x_{i}\in\mathcal{B}_{ad}}h_{i}d_{i^{*}}x_{i}\big{)}$
 The dual-Burnside fusion categories can be characterized by:
@@
 Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup.
-Then ${(H,\mathcal{B})}$ is nilpotent if and only if its dual ${\widehat{(H,\mathcal{B})}}$ is also nilpotent.
+Then ${(H,\mathcal{B})}$ is nilpotent if and only if its dual $^{\widehat{(H,\mathcal{B})}}$ is also nilpotent.
 Moreover, in this case they have the same nilpotency class.
@@
 (1.7) $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\frac{1}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{ad}}})}d_{i}x_{i}\big{)}$
-Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [Bur23, Theorem 2].
+Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [GNN09, Appendix] or [Bur23, Theorem 2].
 Corollary 1.17.
@@
 (1.14) $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i})\big{)}$
-Concrete examples on which these results apply can be found in §12.3. Next theorem improves [OY23, Theorem 4.5].
+In particular,
+Remark 1.20 (Cauchy-type theorem).
+For every perfect integral modular fusion category,
+(1.15) $\mathcal{V}(\mathcal{C})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}),$
+thus, for all prime $p$ dividing $\mathrm{FPdim}(\mathcal{C})$, there is a simple object $X$ in $\mathcal{C}$ such that $p$ divides $\mathrm{FPdim}(X)$; which can be interpreted as a Cauchy-type theorem.
+Consequently, $p^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ since $\mathrm{FPdim}(X)^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ by [EGNO15, Proposition 8.14.6].
+Concrete examples where these results apply can be found in §12.3. Next theorem improves [OY23, Theorem 4.5].
 Theorem 1.21.
-Let $\mathcal{C}$ be a weakly-integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$ and $d_{i}^{2}$ for all $i$.
+Let $\mathcal{C}$ be a weakly-integral modular fusion category of $\mathrm{FPdim}=dm$ with $d$ square-free coprime with $m$ and $d_{i}^{2}$ for all $i$.
 Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
 $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
-where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
+where ${\mathcal{D}}$ is a pointed modular fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a modular fusion category of dimension $m$.
 Corollary 1.22.
-Let $\mathcal{C}$ be any integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$.
+Let $\mathcal{C}$ be any integral modular fusion category of $\mathrm{FPdim}=dm$ with $d$ square-free coprime with $m$.
 Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
 $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$
-where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
+where ${\mathcal{D}}$ is a pointed modular fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a modular fusion category of dimension $m$.
 In particular, the $\mathrm{FPdim}$ of a perfect integral modular fusion category does not have a powerless prime factor (Corollary 10.10).
-In particular, a perfect modular fusion category of even $\mathrm{FPdim}$ has its $\mathrm{FPdim}$ divisible by $4$.
+In particular, a perfect modular fusion category of even $\mathrm{FPdim}$ has its $\mathrm{FPdim}$ divisible by 444.
 Remark 1.23.
-By the classification of non-pointed integral modular fusion categories of small rank in [ABPP] and by Corollary 1.18, the $\mathrm{FPdim}$ of a non-pointed integral modular fusion category does not have a powerless prime factor for every rank less than $11\times 3=33$, and every rank less than $17\times 7=119$ in the odd-dimensional case.
-Integral modular fusion categories were recently intensively studied in the literature. In [CP22] it was formulated the following conjecture
+By the classification of non-pointed integral modular fusion categories of small rank in [ABPP] and by Corollary 1.22, the $\mathrm{FPdim}$ of a non-pointed integral modular fusion category does not have a powerless prime factor for every rank less than 11×3=3311$\times$ 3=3311 × 3 = 33, and every rank less than 17×7=11917$\times$ 7=11917 × 7 = 119 in the odd-dimensional case.
+Integral modular fusion categories were recently intensively studied in the literature. The following conjecture is formulated in [CP22]:
 Conjecture 1.24.
-There is no perfect modular fusion category of odd $\mathrm{FPdim}$.
+There is no non-trivial perfect modular fusion category of odd $\mathrm{FPdim}$.
+This conjecture was checked for ranks less than 252525 in [ABPP] or [CGP23].
 Note that the above result also applies to non-pointed simple integral modular fusion categories (since they are in particular perfect).
 The importance of the existence of such categories was described in [LPR23, §5], where the following two open statements are proved to be equivalent, whereas Proposition 10.13 states that they imply Conjecture 1.24.
-Statement 1.21.
+Statement 1.25.
 Every simple integral fusion category is weakly group-theoretical.
-Statement 1.22.
+Statement 1.26.
 Every simple integral modular fusion category is pointed.
@@
 In §3 we develop few basic facts about Galois symmetries of rational hypergroups.
-In §4 hypergroups with Burnside property are studied and Theorem 1.7 is proven.
-In the same section we prove Theorem 1.2 and its more general version Theorem 4.17 for hypergroups.
-In this section, Burnside’s property for dual hypergroups is also studied and Theorem 1.8 is proven.
+In §4 hypergroups with Burnside property are studied.
+In the same section, we prove Theorem 1.2 and its more general version Theorem 4.17 for hypergroups.
+In this section, Burnside’s property for dual hypergroups is also studied.
 In §5 we extend the notion of kernels of objects from the settings of fusion categories to the settings of ARN-hypergroups.
-In §7 we construct the universal grading group for ARN-hypergroups.
-A connection between this universal grading group and the group of grouplike elements of the dual it is proven in Theorem 6.5.
-§7 describes the constriction of the universal grading group of an ARN-hypergroup.
-In §8 we define nilpotent ARN-hypergroup and prove Theorems 1.11 and 1.12.
-The applications to modular fusion categories are given in §10.1.
-In §10 we prove Proposition 10.14 showing that any of the two statements from [LPR23, §5] implies Conjecture 1.20.
-Next, §11 delves into applications and supplementary results: §11.1 offers a characterization for a Drinfeld center to be perfect; §11.2 establishes some results concerning integrality and Burnside properties, proving Theorems 1.10 and 1.23, particularly highlighting their equivalence in the perfect modular case; §11.3 characterizes the near-group fusion categories which admit a modular structure.
-Finally, §12 presents concrete examples and counterexamples: §12.1 introduces non-Burnside simple integral fusion rings; §12.2 classifies certain dual-Burnside $\mathrm{Rep}(G)$; §12.3 outlines the classification of fusion rings lacking modular categorification within certain families, employing skills from this paper.
+In §6 and §7, we construct the universal grading group for ARN-hypergroups.
+A connection between this universal grading group and the group of grouplike elements of the dual is proven in Theorem 6.5.
+Theorems 1.8 and 1.11 are proven.
+In §8 we define nilpotent ARN-hypergroup and prove Theorems 1.14 and 1.15.
+In §9 we prove Theorem 1.12.
+In §10, the applications to modular fusion categories are given (e.g.
+the proofs from Theorem 1.16 to Corollary 1.22), and we prove Proposition 10.13 showing that any of the two statements from [LPR23, §5] implies Conjecture 1.24.
+Next, §11 delves into applications and supplementary results: §11.1 offers a characterization for a Drinfeld center to be perfect; §11.2 establishes some results concerning integrality and Burnside properties, proving Theorems 1.13 and 1.27, particularly highlighting their equivalence in the perfect modular case; §11.3 characterizes the near-group fusion categories which admit a modular structure.
+Finally, §12 presents concrete examples and counterexamples: §12.1 introduces non-Burnside simple integral fusion rings; §12.2 classifies certain dual-Burnside $\mathrm{Rep}(G)$; §12.3 provides certain families of fusion rings lacking modular categorification by Theorems 1.19 or 1.22.
 Throughout of this paper, all vector spaces and linear categories are considered over the ground field $\mathbb{C}$ of complex numbers.
@@
 The morphism $\psi$ is called an isomorphism of hypergroups if it is bijective.
-Let ${(H,\mathcal{B})}$ be a normalized abelian hypergroup and ${\widehat{(H,\mathcal{B})}}$ its dual as defined in the introduction.
-There is a formula for the dual fusion coefficients ${\widehat{p}}_{k}(i,j)$ given in [Bur21, Proposition 2.1].
-For a normalized abelian hypergroup ${(H,\mathcal{B})}$, this is given by:
-(2.1) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{a\in{\mathcal{B}}}h_{a}\mu_{j_{1}}(a)\mu_{j_{2}}(a)\mu_{k}(a^{*})\bigg{)}$
-where the scalars $n_{k}$ are defined below (2.2).
+Let ${(H,\mathcal{B})}$ be a normalized abelian hypergroup and $^{\widehat{(H,\mathcal{B})}}$ its dual as defined in the introduction.
+There is a formula for the dual fusion coefficients ${\widehat{p}}_{k}(i,j)$ given in [Bur21, Proposition 2.1] by:
+(2.1) $\widehat{p}_{k}(i,j)=\frac{1}{n_{k}}\bigg{(}\sum_{a\in{\mathcal{B}}}h_{a}\mu_{i}(a)\mu_{j}(a)\mu_{k}(a^{*})\bigg{)}$
+where the scalars $n_{k}$ are defined below (2.2), and $h_{a}:=1/N^{1}_{a,a^{*}}$ (then $h_{1}=1$).
 Remark 2.1.
@@
 We usually let :=${\mathcal{B}}:=\{x_{i}\}_{i\in\mathcal{I}}$, with :=$\mathcal{I}:=\{1,\dots m\}$ and $x_{1}=1$, and say that ${(H,\mathcal{B})}$ has rank $m$.
-We can simplify the notation $N_{x_{i}x_{j}}^{x_{k}}$ by $N_{i,j}^{k}$.
+We can simplify the notation $N_{x_{i}x_{j}}^{x_{k}}$ by $N_{i,j}^{k}$, and $h_{x_{i}}$ by $h_{i}$.
+Remark 2.2.
 Let $\{x_{i}^{\circ}\}_{i\in\mathcal{I}}$ be the linear dual basis of the linear basis ${\mathcal{B}}$ of $H$.
-Therefore $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, for all $i,i^{\prime}\in\mathcal{I}$.
-It is easy to see that in this case $\widehat{E}_{i}:=x_{i}^{\circ}\in H^{*}$ is a primitive central idempotent of $H^{*}$ corresponding to the character
+Therefore $x_{i}^{\circ}(x_{j})=\delta_{i,j}$, for all $i,j\in\mathcal{I}$.
+It is easy to see that in this case $\widehat{F}_{i}:=x_{i}^{\circ}\in H^{*}$ is a primitive central idempotent of $H^{*}$ corresponding to the character
 $\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{x_{i}}:&H^{*}&\to&\mathbb{C}\\ &f&\mapsto&f(x_{i}).\end{array}$
@@
 Let $(H,\mathcal{B})$ be any abelian hypergroup.
-Define $h_{i}:=1/N^{1}_{i,i^{*}}$, for all $i\in\mathcal{I}$.
-Note that $h_{1}=1$.
-Then $H$ is a commutative semisimple algebra, and therefore $H\simeq\bigoplus_{j=1}^{m}\mathbb{C}F_{j}$, where $F_{j}$ are the central primitive idempotents of $H$.
-One can define $\tau\in H^{*}$ with $\tau(x_{i})=\delta_{i,1}$, where as above, $x_{1}=1$.
+The finite dimensional *-algebra $H$ (so semismple) is commutative, therefore $H\simeq\bigoplus_{i=1}^{m}\mathbb{C}F_{i}$, where $(F_{i})$ are the central primitive idempotents of $H$.
+We can define $\tau\in H^{*}$ with $\tau(x_{i})=\delta_{i,1}$, where as above, $x_{1}=1$.
 Note that $\tau=x_{1}^{\circ}$ with the above notations.
 Moreover, in this case
-$\begin{array}[]{cccc}(\;,\;)_{\tau}:&H\otimes H&\to&\mathbb{C}\\ &(a,b)&\mapsto&\tau(ab)\end{array}$
-is an associative symmetric non-degenerate bilinear form. Thus, one can write
-(2.2) $\tau=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}.$
-for some non-zero scalars $n_{j}\in\mathbb{C}^{\times}$. These scalars $n_{j}$ are called the formal codegrees of $(H,\mathcal{B})$, see [Ost15]. Note that $\sum_{j=1}^{m}\frac{1}{n_{j}}=\tau(1)=1$.
+$\begin{array}[]{cccc}(\;,\;)_{\tau}:&H\times H&\to&\mathbb{C}\\ &(a,b)&\mapsto&\tau(ab)\end{array}$
+is an associative symmetric non-degenerate bilinear form.
+Thus, $\tau(F_{i})$ must be nonzero for all $i$.
+The nonzero scalars $n_{i}:=\tau(F_{i})^{-1}$ are called the formal codegrees of $(H,\mathcal{B})$, see [Ost15].
+We can write
+(2.2) $\tau=\sum_{i=1}^{m}\tau(F_{i})\mu_{i}=\sum_{i=1}^{m}\frac{1}{n_{i}}\mu_{i}.$
+Note that $\sum_{i\in\mathcal{I}}\frac{1}{n_{i}}=\tau(1)=1$.
 From the definition of a hypergroup it follows that $\{h_{i}x_{i}\}_{i\in\mathcal{I}}$ and $\{x_{i^{*}}\}_{i\in\mathcal{I}}$ form a pair of dual bases for $(\;,\;)_{\tau}$.
-Note that in this case $\{n_{j}F_{j}\}_{i\in\mathcal{I}}$ and $\{F_{j}\}_{i\in\mathcal{I}}$ form also a pair of dual basis for the bilinear form $(\;,\;)_{\tau}$.
+Note that in this case $\{n_{j}F_{j}\}_{j\in\mathcal{I}}$ and $\{F_{j}\}_{j\in\mathcal{I}}$ form also a pair of dual basis for the bilinear form $(\;,\;)_{\tau}$.
 The existence of the two pairs of dual bases implies that:
-(2.3) $\sum_{i=1}^{m}{h_{i}}x_{i}\otimes x_{i}^{*}=\sum_{j=1}^{m}n_{j}F_{j}\otimes F_{j}$
-From here it follows that
-(2.4) $F_{j}:=\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i^{*}})x_{i}\big{)}$
-by applying $\operatorname{id}\otimes\mu_{j}$ for the character $\mu_{j}$ corresponding to $F_{j}$. Applying $\mu_{j}\otimes\mu_{k}$ to the above, one obtains the first orthogonality relations
-(2.5) $\sum_{i=1}^{m}{h_{i}}\mu_{j}(x_{i})\mu_{k}(x_{i^{*}})=\delta_{j,k}n_{j}$
+(2.3) $\sum_{i\in\mathcal{I}}{h_{i}}x_{i}\otimes x_{i}^{*}=\sum_{j\in\mathcal{I}}n_{j}F_{j}\otimes F_{j}$
+By applying $\operatorname{id}\otimes\mu_{j}$, for the character $\mu_{j}$ corresponding to $F_{j}$, it follows that
+(2.4) $F_{j}=\frac{1}{n_{j}}\sum_{i\in\mathcal{I}}{h_{i}}\mu_{j}(x_{i^{*}})x_{i}$
+Applying $\mu_{j}\otimes\mu_{k}$ to the above, we obtain the first orthogonality relation
+(2.5) $\sum_{i\in\mathcal{I}}{h_{i}}\mu_{j}(x_{i})\mu_{k}(x_{i^{*}})=\delta_{j,k}n_{j}.$
 Since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$, for $j=k$, it follows from here that $n_{j}>0$.
-By a classical argument, as for finite groups orthogonality, one obtains the second orthogonality relation:
-(2.6) $\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x_{i})\mu_{j}(x_{l}^{*})=\delta_{i,l}h_{i}^{-1}$
+By a classical argument (see, for example, the proof of [LPR23, Theorem 2.4]), as for finite groups orthogonality, we obtain the second orthogonality relation:
+(2.6) $\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x_{i})\mu_{j}(x_{l^{*}})=\delta_{i,l}h_{i}^{-1}.$
 2.2. Normalizable abelian hypergroups and their duals
@@
 Let :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$.
-Then it is easy to see that $(H,\mathcal{B}^{\prime})$ is also a hypergroup.
-However note that $\tau^{\prime}(\frac{x_{i}}{\alpha_{{}_{i}}})=\delta_{i,1}$ is the same functional $\tau$ on $H$ and therefore the formal codegrees $n_{j}$ do not change after rescaling.
+Then, $(H,\mathcal{B}^{\prime})$ is also a hypergroup, with coefficients rescaled as $\frac{\alpha_{k}}{\alpha_{i}\alpha_{j}}N_{i,j}^{k}$; and then $h_{i}:=(N_{i,i^{*}}^{1})^{-1}$ rescaled as $|\alpha_{i}|^{2}h_{i}$.
+Consider $\tau^{\prime}$ defined by $\tau^{\prime}(y_{i}):=\delta_{i,1}$, then $\tau^{\prime}(x_{i})=\alpha_{i}\delta_{i,1}=\delta_{i,1}$, because $\alpha_{1}=1$.
+Therefore $\tau^{\prime}=\tau$ and the formal codegrees $(n_{j})$ are invariant by rescaling.
 Recall the notion of a normalizable hypergroup from Definition 1.5. Note that any rescaled normalizable hypergroup is also normalizable.
 Lemma 2.3.
-Let ${(H,\mathcal{B})}$ be a hypergroup. Then ${(H,\mathcal{B})}$ is normalizable if and only if there is $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$.
+An hypergroup ${(H,\mathcal{B})}$ is normalizable if and only if there is $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})$ nonzero, for all $i\in\mathcal{I}$.
 Proof.
-It is easy to see that if $(H,\{\frac{x_{i}}{\alpha_{{}_{i}}}\})$ is normalized after a rescaling, then $\mu_{1}:H\rightarrow H,\;x_{i}\mapsto\alpha_{{}_{i}}$ is a morphism of algebras on $H$.
-Thus $\mu_{1}\in{\widehat{\mathcal{B}}}$.
-Conversely, one can check directly that $(H,\{\frac{x_{i}}{\mu_{1}(x_{i})}\})$ is a normalized hypergroup for any morphism $\mu_{1}\in{\widehat{\mathcal{B}}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $i\in\mathcal{I}$.
+The rescaling $(H,\{\frac{x_{i}}{\alpha_{{}_{i}}}\})$ is normalized if and only if $\sum_{i\in\mathcal{I}}\frac{\alpha_{k}}{\alpha_{i}\alpha_{j}}N_{i,j}^{k}=1$, if and only if $\alpha_{i}\alpha_{j}=\sum_{i\in\mathcal{I}}N_{i,j}^{k}\alpha_{k}$, if and only if $\mu_{1}:x_{i}\mapsto\alpha_{i}$ is an algebra morphism (i.e.
+an element of $^{\widehat{\mathcal{B}}}$) with $\mu_{1}(x_{i})$ nonzero, for all $i\in\mathcal{I}$.
 ∎
-We denote by ${(H,\mathcal{B},\mu_{1})}$ the normalizable hypergroup obtained by normalizing ${(H,\mathcal{B})}$ by $\mu_{1}\in{\widehat{\mathcal{B}}}$.
-We also denote by $\overline{{(H,\mathcal{B},\mu_{1})}}$ the normalized hypergroup obtained by rescaling ${(H,\mathcal{B},\mu_{1})}$ as in Lemma 2.2.
+The normalizable hypergroup ${(H,\mathcal{B})}$ together with $\mu_{1}$ as in Lemma 2.3 is denoted ${(H,\mathcal{B},\mu_{1})}$.
+The normalized hypergroup $(H,\{\frac{x_{i}}{\mu_{1}(x_{i})}\})$ is denoted $\overline{{(H,\mathcal{B},\mu_{1})}}$.
 Definition 2.4.
-Recall from Subsection 1.1 that any abelian normalized hypergroup $H$ admits a dual hypergroup $\widehat{H}$.
-In the case of $\overline{{(H,\mathcal{B},\mu_{1})}}$, we denote this dual by ${{\widehat{{(H,\mathcal{B},\mu_{1})}}}}$.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup via $\mu_{1}\in\widehat{\mathcal{B}}$ such that $\mu_{1}(x_{i})\neq 0$ for all $1\leq i\leq m$.
-From Equation (1.2) it follows that the multiplication on the dual of $\widehat{{(H,\mathcal{B},\mu_{1})}}$ can be written as
-(2.7) $[f\star g](\frac{x_{i}}{\mu_{1}(x_{i})})=f(\frac{x_{i}}{\mu_{1}(x_{i})})g(\frac{x_{i}}{\mu_{1}(x_{i})})$
+Recall from §1.1 that any abelian normalized hypergroup $H$ admits a dual hypergroup $H^{*}$.
+In the case of $\overline{{(H,\mathcal{B},\mu_{1})}}$, we denote this dual by $^{{\widehat{{(H,\mathcal{B},\mu_{1})}}}}$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. From Equation (1.2), it follows that the multiplication on the dual $\widehat{{(H,\mathcal{B},\mu_{1})}}$ can be written as
+(2.7) $[f\star g]\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)=f\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)g\left(\frac{x_{i}}{\mu_{1}(x_{i})}\right)$
 for all $f,g\in H^{*}$.
-Then it is easy to see that the algebra unit of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $\mu_{1}$.
+Then it is easy to see that the algebra unit of $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $\mu_{1}$.
 Moreover, the involution on $^{\widehat{\mathcal{B}}}$, as given in §1.1, becomes:
-(2.8) $\mu_{j}^{\#}(x_{i})=\bigg{(}\frac{\mu_{1}(x_{i})}{\overline{\mu_{1}(x_{i}})}\bigg{)}\mu_{j}(x_{i}^{*}).$
+(2.8) $\mu_{j^{\#}}(\frac{x_{i}}{\mu_{1}(x_{i})})=\mu_{j}(\frac{x_{i^{*}}}{\overline{\mu_{1}(x_{i}})}).$
 Definition 2.5.
 The order of a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is defined by
-$n{(H,\mathcal{B},\mu_{1})}:=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}.$
-In the abelian case, note that
+$n{(H,\mathcal{B},\mu_{1})}:=\sum_{i\in\mathcal{I}}h_{i}|\mu_{1}(x_{i})|^{2}.$
+In the abelian case, the following holds by Equation (2.5):
 (2.9) $n{(H,\mathcal{B},\mu_{1})}=n_{1}$
-by Equation (2.5).
+Proposition 2.6.
+Let ${(H,\mathcal{B},\mu_{1})}$ be a normalizable hypergroup. Then the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling.
+Proof.
+Let $x^{\prime}_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$.
+Consider $(H,\mathcal{B}^{\prime},\mu_{1})$ the rescaled hypergroup with :=$\mathcal{B}^{\prime}:=\{x^{\prime}_{i}\}_{i\in\mathcal{I}}$.
+Recall that the order of $x^{\prime}_{i}$ is $h^{\prime}_{i}=|\alpha_{{}_{i}}|^{2}h_{i}>0.$ Therefore
+$n(H,\mathcal{B}^{\prime},\mu_{1})=\sum_{i\in\mathcal{I}}h^{\prime}_{i}\mid\mu_{1}(x^{\prime}_{i})\mid^{2}=\sum_{i\in\mathcal{I}}h_{i}\mid\mu_{1}(x_{i})\mid^{2}=n{(H,\mathcal{B},\mu_{1})}.\qed$
 We shortly write $n(H)$ when $\mathcal{B}$ and $\mu_{1}$ are implicitly understood.
-In particular, for a normalized hypergroup one may choose $\mu_{1}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n(H)=\sum_{i=1}^{m}h_{i}$.
+In particular, for a normalized hypergroup we may choose $\mu_{1}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n(H)=\sum_{i\in\mathcal{I}}h_{i}$.
 For the rest of this paper we use the notation $d_{i}:=\mu_{1}(x_{i})$ for any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$.
@@
 $n{(H,\mathcal{B},\mu_{1})}\in\mathbb{Z}$.
-Note that a fusion ring is weakly-rational if and only if it is weakly-integral since $\mathrm{FPdim}(R)$ is an algebraic number in this case.
+Note that a fusion ring is weakly-rational if and only if it is weakly-integral since $\mathrm{FPdim}(R)$ is an algebraic integer in this case.
 Remark 2.7.
-Note that the formula from Equation (2.1) can be written for abelian normalizable hypergroups as:
-(2.10) $\widehat{p}_{k}({j_{1}},{j_{2}})=\frac{1}{n_{k}}\bigg{(}\sum_{x_{i}\in{\mathcal{B}}}\frac{h_{i}}{d_{i}}\mu_{j_{1}}(x_{i})\mu_{j_{2}}(x_{i})\mu_{k}(x_{i^{*}})\bigg{)}$
+Note that the formula from Equation (2.1) can be written for abelian normalizable hypergroups as follows (using the fact that $h_{l}$ rescales as $|d_{l}|^{2}h_{l}$):
+(2.10) $\widehat{p}_{k}(i,j)=\frac{1}{n_{k}}\bigg{(}\sum_{x_{l}\in{\mathcal{B}}}\frac{h_{l}}{d_{l}}\mu_{i}(x_{l})\mu_{j}(x_{l})\mu_{k}(x_{l^{*}})\bigg{)}$
 Lemma 2.8.
@@
 Proof.
-By applying (2.10) to ${j_{1}}={j_{2}}^{\#}=j$ and $k=1$, and then (2.5), one obtains (see also [Bur21, Equation (2.10)]),
-(2.11) $\widehat{p}_{1}(j,j^{\#})=\frac{n_{j}}{n{(H,\mathcal{B},\mu_{1})}}$
-Let ${\widehat{h}}_{j}:=\frac{1}{\widehat{p}_{1}(j,j^{*})}$ be the order of $\mu_{j}$ in ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
-It follows that $n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}=n{(H,\mathcal{B},\mu_{1})}(\sum_{j=1}^{m}\frac{1}{n_{j}})=n{(H,\mathcal{B},\mu_{1})}$.
+By applying (2.10) to $i=j^{\#}$ and $k=1$, and then (2.5), we obtain (see also [Bur21, Equation (2.10)]),
+(2.11) $\widehat{p}_{1}(i,i^{\#})=\frac{n_{i}}{n{(H,\mathcal{B},\mu_{1})}}$
+Let ${\widehat{h}}_{i}:=\frac{1}{\widehat{p}_{1}(i,i^{\#})}$ be the order of $\mu_{i}$ in $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
+It follows that $n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{i\in\mathcal{I}}{\widehat{h}}_{i}=n{(H,\mathcal{B},\mu_{1})}(\sum_{i\in\mathcal{I}}\frac{1}{n_{i}})=n{(H,\mathcal{B},\mu_{1})}$.
 ∎
-Next proposition holds for any normalized hypergroup, not necessarily abelian.
+The next proposition holds for any normalized hypergroup, not necessarily abelian.
 Proposition 2.9.
 For any normalized hypergroup ${(H,\mathcal{B})}$,
-(2.12) $F_{1}=\frac{1}{n{(H,\mathcal{B})}}\big{(}\sum_{i=1}^{m}h_{i^{*}}x_{i}\big{)}$
+(2.12) $F_{1}=\frac{1}{n{(H,\mathcal{B})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}x_{i}\big{)}$
 where $F_{1}$ is the idempotent corresponding to the augmentation map $\mu_{1}:H\rightarrow\mathbb{C},\;a\mapsto 1$ for all $a\in\mathcal{B}$.
 Proof.
-Suppose that $F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{i}$ for some $\alpha_{{}_{i}}\in\mathbb{C}$.
+Suppose that $F_{1}=\sum_{i\in\mathcal{I}}\alpha_{{}_{i}}x_{i}$ for some $\alpha_{{}_{i}}\in\mathbb{C}$.
 Recall that $x_{l}F_{i}=\mu_{i}(x_{l})F_{i}$.
-Then, $x_{l}F_{1}=F_{1}=\sum_{i=1}^{m}\alpha_{{}_{i}}x_{l}x_{i}$.
-It follows that $\tau(F_{1})=\alpha_{1}=\tau(x_{l}F_{1})=\alpha_{{}_{l^{*}}}h_{l}^{-1}=\tau(F_{1}x_{l})=\alpha_{l}h_{l^{*}}^{-1}.$ Then $F_{1}=\sum_{i=1}^{m}\alpha_{i}x_{i}={\alpha_{1}}(\sum_{i=1}^{m}{h_{i^{*}}}x_{i})$ and $1=\mu_{1}(F_{1})={\alpha_{1}}(\sum_{i=1}^{m}{h_{i}})$.
+Then, $x_{l}F_{1}=F_{1}=\sum_{i\in\mathcal{I}}\alpha_{{}_{i}}x_{l}x_{i}$.
+It follows that $\tau(F_{1})=\alpha_{1}=\tau(x_{l}F_{1})=\alpha_{l^{*}}h_{l}^{-1}.$ Thus $\alpha_{l^{*}}=h_{l}\alpha_{1}$, so $\alpha_{l}=h_{l^{*}}\alpha_{1}$, $F_{1}=\sum_{i\in\mathcal{I}}\alpha_{i}x_{i}={\alpha_{1}}(\sum_{i\in\mathcal{I}}{h_{i^{*}}}x_{i})$ and $1=\mu_{1}(F_{1})={\alpha_{1}}(\sum_{i\in\mathcal{I}}{h_{i}})$.
 This implies that $\alpha_{1}=\frac{1}{n(H)}$ and therefore $\alpha_{i}=\frac{h_{i^{*}}}{n(H)}$ for all $i\in\mathcal{I}$.
@@
 In any normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$,
-(2.13) $F_{1}=\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i=1}^{m}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$
+(2.13) $F_{1}=\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$
+Remark 2.11.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian (rational) normalizable hypergroup.
+In order to find the central primitive idempotents of $H^{*}$, we can apply Remark 2.2 after normalization.
+If $x_{i}^{\circ}\in H^{*}$ are defined as above by $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, then $\widehat{F}_{i}:=d_{i}x_{i}^{\circ}\in H^{*}$ are the primitive central idempotents of $H^{*}$ corresponding to the character
+$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{\overline{x}_{i}}:&H^{*}&\to&\mathbb{C}\\ &\mu_{j}&\mapsto&\mu_{j}(\overline{x}_{i}),\end{array}$
+where $\overline{x}_{i}=\frac{x_{i}}{d_{i}}$. By identifying $H^{**}$ with $H$, we get that $\operatorname{ev}_{\overline{x}_{i}}=\overline{x}_{i}$.
+Recall that $\overline{x}_{1}=x_{1}$.
 Proposition 2.12.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup.
-Then ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is also an abelian normalized hypergroup and the formal codegrees of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ are given by
-(2.14) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}.$
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is an abelian normalized hypergroup whose formal codegrees are
+(2.14) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}|d_{i}|^{2}}.$
 Proof.
-As already mentioned, [Bur21, Proposition 2.3] shows that ${\widehat{(H,\mathcal{B})}}$ is also an abelian normalized hypergroup.
-By Remark 2.1, $x_{i}:H^{*}\rightarrow\mathbb{C}$ are the characters of ${\widehat{(H,\mathcal{B})}}$.
-By definition of $\widehat{\tau}\in(H^{*})^{*}=H$, $\widehat{\tau}(\mu_{j})=\delta_{j,1}$ where $\mu_{1}$ is the unit of $H^{*}$.
+As already mentioned, [Bur21, Proposition 2.3] shows that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is an abelian normalized hypergroup.
+Let $\overline{x}_{i}=\frac{x_{i}}{d_{i}}$ be the normalization of $x_{i}$.
+By Remark 2.11, $\overline{x}_{i}:H^{*}\rightarrow\mathbb{C}$ are the characters of $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
+By definition of $\widehat{\tau}\in(H^{*})^{*}=H$, $\widehat{\tau}(\mu_{j})=\delta_{j,1}$ as $\mu_{1}$ is the unit of $H^{*}$.
 Thus $\widehat{\tau}=F_{1}$.
-Equation (2.2) for $H^{*}$ gives that $\widehat{\tau}=\sum_{i=1}^{m}\frac{1}{\widehat{n}_{i}}x_{i}$.
-Then Equation (2.12) finishes the proof since $n{\widehat{(H,\mathcal{B})}}=n{(H,\mathcal{B})}$, by Lemma 2.6.
-∎
+Equation (2.2) for $H^{*}$ gives that $\widehat{\tau}=\sum_{i\in\mathcal{I}}\frac{1}{\widehat{n}_{i}}\overline{x}_{i}$.
+So, by Equation (2.13),
+$\frac{1}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}=F_{1}=\widehat{\tau}=\sum_{i\in\mathcal{I}}\frac{1}{\widehat{n}_{i}}\frac{x_{i}}{d_{i}}.$
+The result follows by $h_{i^{*}}=(N_{i^{*},i}^{1})^{-1}=(N_{i,i^{*}}^{1})^{-1}=h_{i}$, in the abelian case. ∎
 Corollary 2.13.
 Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then the primitive central idempotent of the linear character $x_{1}\in\widehat{\widehat{\mathcal{B}}}=\mathcal{B}$ is given by
-(2.15) $\widehat{F_{1}}=\sum_{j=1}^{m}\frac{\mu_{j}}{n_{j}}\in H^{*}.$
+(2.15) $\widehat{F_{1}}=\sum_{j\in\mathcal{I}}\frac{\mu_{j}}{n_{j}}\in H^{*}.$
 Proof.
@@
 ∎
-A morphism of normalizable hypergroups $\psi:(H,{\mathcal{B}},\mu_{1})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime},\mu_{1}^{\prime})$ is a morphism of hypergroups with the property that $\mu^{\prime}_{1}\circ f=\mu_{1}$.
+A morphism of normalizable hypergroups $\psi:(H,{\mathcal{B}},\mu_{1})\rightarrow(H^{\prime},{\mathcal{B}}^{\prime},\mu_{1}^{\prime})$ is a morphism of hypergroups with the property that $\mu^{\prime}_{1}\circ\psi=\mu_{1}$.
 2.3. Frobenius-Perron theory for a RN-hypergroups
 Frobenius-Perron theory can be defined for RN-hypergroups, in the same manner as for fusion rings, see [EGNO15, §3].
-We denote by $\mathrm{FPdim}(x_{i})$ the Frobenius-Perron eigenvalue of the left multiplication operator by $x_{i}$ on $H$.
-As for fusion rings, by a similar argument, it can be shown that $\mathrm{FPdim}(x_{i})=\mathrm{FPdim}(x_{i^{*}})$.
-Indeed, it easy to see that the matrix $N_{i^{*}}$ corresponding to $x_{i^{*}}$ is the transpose of the matrix $N_{i}$ of $x_{i}$ multiplied by the scalars $\frac{h_{j}}{h_{k}}$ in the $(j,k)$-entry.
-More precisely, $(N_{i^{*}})_{jk}=\frac{h_{j}}{h_{k}}(N_{i})_{kj}$.
-Thus, the two matrices have the same set of eigenvalues, in particular the same Frobenius-Perron eigenvalue.
-Note that any RN-hypergroup ${(H,\mathcal{B})}$ can be normalized by $\mu_{1}=\mathrm{FPdim}$, since $\mu_{1}(x_{i})>0$ for all $1\leq i\leq m$.
-Remark 2.11.
-Note that any symmetric RN-hypergroup ${(H,\mathcal{B})}$ is transitive in the sense of [EGNO15, Definition 3.3.1].
-Indeed the function $\tau$ from §3.1 is symmetric and therefore $\tau(x_{i}x_{j}x_{k})=\tau(x_{j}x_{k}x_{i})=\tau(x_{k}x_{i}x_{j})$.
-Since $\tau(x_{i}x_{j}x_{k})=h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})$ one has
-(2.16) $h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})=h_{j}^{-1}m(x_{j^{*}},x_{k}x_{i})=h_{k}^{-1}m(x_{k^{*}},x_{i}x_{j}).$
-which shows that ${(H,\mathcal{B})}$ is transitive.
-Then it is easy to see that an analogue [EGNO15, Proposition 3.3.6] also holds for symmetric RN-hypergroups.
-In particular it follows that $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is the unique which takes positive values on $\mathcal{B}$.
-Definition 2.12.
-In this case of a RN-hypergroup ${(H,\mathcal{B})}$, one can define the Frobenius-Perron dimension as
-$\mathrm{FPdim}{(H,\mathcal{B})}:=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}.$
+We denote by $\mathrm{FPdim}(x_{i})$ the Frobenius-Perron eigenvalue of the left multiplication operator by $x_{i}$ on $H$, i.e.
+the ($\ell^{2}$) matrix norm of $N_{i}=(N_{i,j}^{k})_{j,k}$.
+Recall that a hypergroup is called symmetric when $N_{a,b}^{1}=N_{b,a}^{1}$, for all $a,b\in\mathcal{B}$, but $N_{a,b}^{1}=\delta_{a^{*},b}h_{a}^{-1}$, so symmetric means that $h_{a}=h_{a^{*}}$, for all $a\in\mathcal{B}$
 Lemma 2.14.
-For a normalized RN-hypergroup,
-$\mathrm{FPdim}{(H,\mathcal{B})}=n{(H,\mathcal{B},\mathrm{FPdim})}.$
+An abelian hypergroup is symmetric.
 Proof.
-It is easy to see that [EGNO15, Proposition 3.3.6] also works for normalized RN-hypergroups instead of fusion rings.
-Therefore $\mu_{1}:H\rightarrow H,x_{i}\mapsto 1$ is the unique morphism $\phi:H\rightarrow\mathbb{C}$ with $\phi(x_{i})>0$ for all $1\leq i\leq m$.
-This implies that $\mathrm{FPdim}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n{(H,\mathcal{B},\mathrm{FPdim})}=\mathrm{FPdim}{(H,\mathcal{B})}$.
+By definition and abelian assumption, $h_{i}^{-1}:=N_{i,i^{*}}^{1}=N_{i^{*},i}^{1}=h_{i^{*}}^{-1}$. ∎
+Lemma 2.15.
+If $N_{i,j}^{k}$ is nonzero then
+(2.16) $\frac{h_{i^{*}}h_{j^{*}}h_{k}}{h_{i}h_{j}h_{k^{*}}}=1.$
+Proof.
+The associativity of $H$ reformulates as $\sum_{s}N_{i,j}^{s}N_{s,k}^{t}=\sum_{s}N_{j,k}^{s}N_{i,s}^{t}$, for all $i,j,k,t$.
+If $t=1$, we get that $\sum_{s}N_{i,j}^{s}N_{s,k}^{1}=\sum_{s}N_{j,k}^{s}N_{i,s}^{1}$.
+But recall that $N_{a,b}^{1}=\delta_{a^{*},b}h_{a}^{-1}$.
+It follows that $N_{i,j}^{k^{*}}h_{k^{*}}^{-1}=N_{j,k}^{i^{*}}h_{i}^{-1}$, in other words, $N_{i,j}^{k}=\frac{h_{k}}{h_{i}}N_{j,k^{*}}^{i^{*}}$.
+By applying this last equality, we get that $N_{j,k^{*}}^{i^{*}}=\frac{h_{i^{*}}}{h_{j}}N_{k^{*},i}^{j^{*}}$, and $N_{k^{*},i}^{j^{*}}=\frac{h_{j^{*}}}{h_{k^{*}}}N_{i,j}^{k}$.
+Thus
+(2.17) $N_{i,j}^{k}=\frac{h_{k}}{h_{i}}N_{j,k^{*}}^{i^{*}}=\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}N_{k^{*},i}^{j^{*}}=\frac{h_{i^{*}}h_{j^{*}}h_{k}}{h_{i}h_{j}h_{k^{*}}}N_{i,j}^{k}$
+The result follows. ∎
+Theorem 2.16.
+A (finite-dimensional) hypergroup is symmetric, i.e. $h_{i^{*}}=h_{i}$, for all $i\in\mathcal{I}$. In particular, the function $\tau$ from §2.1 is symmetric.
+Proof.
+Introduce the notation $r_{i}:=\frac{h_{i^{*}}}{h_{i}}$. Lemma 2.15 reformulates as:
+$N_{i,j}^{k}\neq 0\Rightarrow r_{i}r_{j}=r_{k}.$
+We aim to demonstrate that $r_{i}$ is a root of unity for every $i\in\mathcal{I}$.
+First, select $k_{2}$ such that $N_{i,i}^{k_{2}}$ is nonzero, which implies $r_{i}^{2}=r_{k_{2}}$.
+Next, choose $k_{3}$ such that $N_{i,k_{2}}^{k_{3}}$ is nonzero, leading to $r_{i}r_{k_{2}}=r_{k_{3}}$ and thus $r_{i}^{3}=r_{k_{3}}$.
+Continuing this process, we find that the sequence $\{r_{i}^{n}\}$ for $n\in\mathbb{N}$ matches the set $\{r_{j}\}$ for $j\in\mathcal{J}$, with $\mathcal{J}$ being a subset of $\mathcal{I}$.
+Due to the finite dimensionality of the hypergroup, $\mathcal{I}$ is finite, which implies that $\mathcal{J}$ is also finite.
+Consequently, there exist integers $m$ and $n$ with $m>n$ such that $r_{i}^{m}=r_{i}^{n}$.
+Thus, $r_{i}^{s}=1$ where $s=m-n>0$, confirming that $r_{i}$ is a root of unity.
+Additionally, the axioms of the hypergroup state that $N_{a,a^{*}}^{1}$ is positive for all $a\in\mathcal{B}$; hence $h_{i}$ is positive for all $i\in\mathcal{I}$.
+Therefore, $r_{i}=\frac{h_{i^{*}}}{h_{i}}$ is a positive root of unity, which must be 111.
 ∎
-Note that by Lemma 2.2, any RN-hypergroup is normalizable via the character $\mu_{1}=\mathrm{FPdim}$. Lemma 2.13 also shows that
-$n(H,\mathcal{B},\mathrm{FPdim})=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}=\mathrm{FPdim}{(H,\mathcal{B})}$
-for any RN-hypergroup. For a RN-hypergroup, we sometimes write ${(H,\mathcal{B},\mathrm{FPdim})}$ instead of ${(H,\mathcal{B})}$ to specify that we normalize it by $\mathrm{FPdim}$.
-Proposition 2.14.
-Let ${(H,\mathcal{B},\mu_{1})}$ be a normalizable hypergroup.
-Then the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling.
-If ${(H,\mathcal{B})}$ is RN-hypergroup, then the Frobenius-Perron dimension of ${(H,\mathcal{B})}$ is also invariant under rescaling with real positive numbers.
+In the rest of the paper, we assume that the involution on $\mathcal{B}$ extends into an anti-involution on $H$, i.e. $N_{i,j}^{k}=N_{j^{*},i^{*}}^{k^{*}}$, for all $i,j,k$.
+Proposition 2.17 (Frobenius Reciprocity).
+The following equalities hold:
+(2.18) $h_{k}^{-1}N_{i,j}^{k}=h_{j}^{-1}N_{i^{*},k}^{j}=h_{i}^{-1}N_{j,k^{*}}^{i^{*}}=h_{k}^{-1}N_{j^{*},i^{*}}^{k^{*}}=h_{j}^{-1}N_{k^{*},i}^{j^{*}}=h_{i}^{-1}N_{k,j^{*}}^{i}.$
 Proof.
-First we show that the order of a normalizable hypergroup is invariant under rescaling.
-Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$.
-Consider $(H,\mathcal{B}^{\prime},\mu_{1})$ the rescaled hypergroup with :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$.
-Note that the order of $y_{i}$ is given by ${h^{\prime}_{i}}^{-1}=N^{1^{\prime}}_{i^{\prime}i^{{}^{\prime}*}}=\tau(x_{i^{\prime}}x_{i^{{}^{\prime}*}})=\frac{1}{|\alpha_{{}_{i}}|^{2}}h_{i}^{-1}$ and this implies that $h^{\prime}_{i}=|\alpha_{{}_{i}}|^{2}h_{i}>0.$ Therefore
-$n(H,\mathcal{B}^{\prime},\mu_{1})=\sum_{i=1}^{m}h^{\prime}_{i}\mid\mu_{1}(y_{i})\mid^{2}=\sum_{i=1}^{m}h_{i}\mid\mu_{1}(x_{i})\mid^{2}=n{(H,\mathcal{B},\mu_{1})}.$
-Moreover, if ${(H,\mathcal{B})}$ is a RN-hypergroup and $\alpha_{{}_{i}}=\alpha_{{}_{i^{*}}}>0$, for all $1\leq i\leq m$, it follows that
-$\mathrm{FPdim}(H^{\prime},\mathcal{B}^{\prime})$ $=$ $\sum_{i=1}^{m}h^{\prime}_{i}\mathrm{FPdim}(y_{i})^{2}=$
-$=$ $\sum_{i=1}^{m}\mid\alpha_{{}_{i}}\mid^{2}h_{i}(\frac{\mathrm{FPdim}(x_{i})^{2}}{\alpha_{{}_{i}}})^{2}$
-$=$ $\mathrm{FPdim}{(H,\mathcal{B})}.$
-Corollary 2.15.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then with the above notations, the dual formal codegrees are given by
-(2.17) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}d_{i}d_{i^{*}}}.$
-Proof.
-[…]
-Remark 2.16.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup.
-In order to find the central primitive idempotents of $H^{*}$, one can apply Remark 2.1 after normalization.
-If $x_{i}^{\circ}\in H^{*}$ are defined as above by $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, then ${\widehat{E}}_{i}:=d_{i}x_{i}^{\circ}\in H^{*}$ are the primitive central idempotents of $H^{*}$ corresponding to the character
-$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{{\frac{x_{i}}{d_{i}}}}:&H^{*}&\to&\mathbb{C}\\ &\mu_{j}&\mapsto&\mu_{j}({\frac{x_{i}}{d_{i}}}).\end{array}$
-2.4. Grouplike elements in hypergroups
-In the rest of this paper all abelian RN-hypergroups are normalized by $\mu_{1}=\mathrm{FPdim}$ and therefore their duals are also considered with respect to $\mu_{1}=\mathrm{FPdim}$.
-Lemma 2.17.
-For any symmetric RN-hypergroup ${(H,\mathcal{B})}$, $h_{i}d_{i}^{2}\geq 1$. Moreover, $h_{i}d_{i}^{2}=1$ if and only $h_{i}x_{i}x_{i^{*}}=x_{1}.$
-Proof.
-One has
-(2.18) $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}+\sum_{k=1}^{m}N^{k}_{ii^{*}}x_{k}.$
-Passing to $\mathrm{FPdim}$, $h_{i}d_{i}^{2}=1+h_{i}\big{(}\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}\big{)}\geq 1.$ If $h_{i}d_{i}^{2}=1$ then $\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}=1$, and therefore $N^{k}_{ii^{*}}=0$ for all $k\neq 1$.
-Thus $h_{i}x_{i}x_{i^{*}}=x_{1}.$ Conversely, if $h_{i}x_{i}x_{i^{*}}=x_{1}$, applying $\mathrm{FPdim}$, it follows that $h_{i}\mathrm{FPdim}(x_{i})^{2}=1$.
+By (2.17), $N_{i,j}^{k}=\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}N_{k^{*},i}^{j^{*}}$, then by anti-involution, $N_{k^{*},i}^{j^{*}}=N_{i^{*},k}^{j}$, and by (2.16), $\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}=\frac{h_{k^{*}}}{h_{j^{*}}}$.
+But $h_{k^{*}}=h_{k}$ and $h_{j^{*}}=h_{j}$ by Theorem 2.16.
+The first equality follows.
+The rest is similar.
 ∎
 Corollary 2.18.
-In any fusion ring ${(H,\mathcal{B})}$, $x_{i}x_{i^{*}}=1$ if and only if $d_{i}=1$.
+For every $x_{i}\in\mathcal{B}$ then $\mathrm{FPdim}(x_{i})=\mathrm{FPdim}(x_{i^{*}})$.
+Proof.
+Let $N_{i}$ be the matrix $(N_{i,j}^{k})_{j,k}$.
+By the first equality of (2.18), $N_{i,j}^{k}=\frac{h_{k}}{h_{j}}N_{i^{*},k}^{j}$, so $N_{i}=DN_{i^{*}}^{T}D^{-1}$, with $D={\rm diag}(h_{i})$ and $()^{T}$ the matrix transpose.
+Thus
+$\mathrm{FPdim}(x_{i^{*}})=\|N_{i^{*}}\|=\|N_{i^{*}}^{T}\|=\|D^{-1}N_{i}D\|=\|N_{i}\|=\mathrm{FPdim}(x_{i}).$
+The second last equality hold because $D^{-1}N_{i}D$ have the same eigenvalues than $N_{i}$, because if $N_{i}v=\lambda v$ then $D^{-1}N_{i}Dw=\lambda w$, with $w=D^{-1}v$. ∎
+Since we are dealing with the finite-dimensional case, by Theorem 2.16, every hypergroup is symmetric.
+Therefore, we can omit this assumption in the rest of the paper, as it is automatically satisfied.
+Remark 2.19.
+Note that any hypergroup ${(H,\mathcal{B})}$ is transitive in the sense of [EGNO15, Definition 3.3.1], i.e.
+$\forall i,j\in\mathcal{I}$ there are $k_{1},k_{2}\in\mathcal{I}$ such that $N_{i,k_{1}}^{j}$ and $N_{k_{2},i}^{j}$ are nonzero.
+Indeed, take $k_{1},k_{2}$ such that $N_{i^{*},j}^{k_{1}}$ and $N_{j,i^{*}}^{k_{2}}$ are nonzero, the result follows by Frobenius reciprocity (2.18).
+Observe that an analogue of [EGNO15, Proposition 3.3.6] holds for RN-hypergroups as it requires Frobenius-Perron [EGNO15, Theorem 3.2.1] involving a RN-matrix (i.e.
+with real nonnegative entries).
+In particular, $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is the unique algebra morphism which takes positive values on $\mathcal{B}$.
+Definition 2.20.
+The Frobenius-Perron dimension of a RN-hypergroup ${(H,\mathcal{B})}$ is
+$\mathrm{FPdim}{(H,\mathcal{B})}:=\sum_{i\in\mathcal{I}}h_{i}\mathrm{FPdim}(x_{i})^{2}.$
+By Lemma 2.3 and Remark 2.19, any RN-hypergroup is normalizable via the linear character $\mu_{1}=\mathrm{FPdim}$. By Definition 2.5 and positivity of $\mathrm{FPdim}$,
+$\mathrm{FPdim}{(H,\mathcal{B})}=n{(H,\mathcal{B},\mathrm{FPdim})}.$
+Moreover, in the normalized case, by uniqueness in Remark 2.19, $\mathrm{FPdim}$ must be the augmentation map, i.e. $\mathrm{FPdim}(x_{i})=1$ for all $i\in\mathcal{I}$.
+Recall from Proposition 2.6 that for a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling.
+Proposition 2.21.
+If ${(H,\mathcal{B})}$ is a RN-hypergroup, then the Frobenius-Perron dimension of ${(H,\mathcal{B})}$ is invariant under rescaling with real positive numbers.
+Proof.
+If ${(H,\mathcal{B})}$ is a RN-hypergroup and $\alpha_{{}_{i}}=\alpha_{{}_{i^{*}}}>0$, for all $i\in\mathcal{I}$, then $(H^{\prime},\mathcal{B}^{\prime})=(H,\{\frac{x_{i}}{\alpha_{i}}\})$ is RN and
+$\mathrm{FPdim}(H^{\prime},\mathcal{B}^{\prime})$ $=$ $\sum_{i\in\mathcal{I}}h^{\prime}_{i}\mathrm{FPdim}(\frac{x_{i}}{\alpha_{i}})^{2}=$
+$=$ $\sum_{i\in\mathcal{I}}|\alpha_{{}_{i}}|^{2}h_{i}\frac{\mathrm{FPdim}(x_{i})^{2}}{\alpha_{{}_{i}}^{2}}$
+$=$ $\mathrm{FPdim}{(H,\mathcal{B})}.\qed$
+2.4. Fourier transform
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Define
+(2.19) $\mathcal{F}:{(H,\mathcal{B})}\rightarrow{\widehat{(H,\mathcal{B})}},\;x_{i}\mapsto\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i^{*}}}x_{i^{*}}^{\circ}.$
+It is clear that $\mathcal{F}$ is a linear isomorphism. Additionally, as noted in Remark 2.11, we have $\widehat{F}_{i}:=d_{i}x_{i}^{\circ}$. Therefore:
+(2.20) $\mathcal{F}(x_{i})=\frac{n{(H,\mathcal{B},\mu_{1})}}{d_{i^{*}}h_{i^{*}}}\widehat{F}_{i^{*}}.$
+For the sake of brevity, we define $|H|:=n{(H,\mathcal{B},\mu_{1})}$.
+Lemma 2.22.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. With the above notations, for all $x,y\in H$:
+$\langle\mathcal{F}(y),x\rangle=|H|\tau(xy).$
+Proof.
+It suffices to verify the above identity for $(x,y)=(x_{i},x_{j})$, for all $i,j$. We have:
+[…]
+since $h_{i}=h_{i^{*}}$ (see Lemma 2.14), and $\delta_{i,j^{*}}h_{i}^{-1}=\tau(x_{i}x_{j})$, because $h_{i}^{-1}=N_{i,i^{*}}^{1}$ and $\tau=x_{1}^{\circ}$. ∎
+Proposition 2.23.
+Using the above notations, we have that
+(2.21) $\mathcal{F}(F_{k})=\frac{|H|}{n_{k}}\mu_{k}.$
+Proof.
+By taking $(x,y)=(F_{j},F_{k})$ in Lemma 2.22, we obtain:
+$\langle\mathcal{F}(F_{k}),F_{j}\rangle=|H|\tau(F_{j}F_{k})=\delta_{j,k}|H|\tau(F_{j}).$
+From Equation (2.4), we have:
+$\tau(F_{j})=\frac{1}{n_{j}}\big{(}\sum_{i\in\mathcal{I}}h_{i}\mu_{j}(x_{i^{*}})\tau(x_{i})\big{)}=\frac{1}{n_{j}}.$
+Thus
+$\langle\mathcal{F}(F_{k}),F_{j}\rangle=\delta_{j,k}\frac{|H|}{n_{j}},$
+which shows that $\mathcal{F}(F_{k})=\frac{|H|}{n_{k}}\mu_{k}$, since $\mu_{i}(F_{j})=\delta_{i,j}$. ∎
+2.5. Grouplike elements in hypergroups
+In the rest of this paper, all abelian RN-hypergroups are normalized by $\mu_{1}=\mathrm{FPdim}$, and therefore their duals are also considered with respect to $\mu_{1}=\mathrm{FPdim}$.
+Recall that $d_{i}:=\mu_{1}(x_{i})=\mathrm{FPdim}(x_{i})$.
+Lemma 2.24.
+For any RN-hypergroup ${(H,\mathcal{B})}$, $h_{i}d_{i}^{2}\geq 1$. Moreover, $h_{i}d_{i}^{2}=1$ if and only $h_{i}x_{i}x_{i^{*}}=x_{1}.$
+Proof.
+The following equality holds
+(2.22) $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}+\sum_{k=2}^{m}N^{k}_{ii^{*}}x_{k}.$
+Passing to $\mathrm{FPdim}$, $h_{i}d_{i}^{2}=1+h_{i}\big{(}\sum_{k=2}^{m}N^{k}_{ii^{*}}d_{k}\big{)}\geq 1.$ If $h_{i}d_{i}^{2}=1$ then $\sum_{k=2}^{m}N^{k}_{ii^{*}}d_{k}=0$, and therefore $N^{k}_{ii^{*}}=0$ for all $k\neq 1$.
+Thus $h_{i}x_{i}x_{i^{*}}=x_{1}.$ Conversely, if $h_{i}x_{i}x_{i^{*}}=x_{1}$, applying $\mathrm{FPdim}$, it follows that $h_{i}d_{i}^{2}=1$.
+∎
+Corollary 2.25.
+In any fusion ring ${(H,\mathcal{B})}$, $x_{i}x_{i^{*}}=x_{1}$ if and only if $d_{i}=1$.
 Recall the set of grouplike elements ${G{(H,\mathcal{B})}}$ from Definition 1.3.
 Remark 2.26.
-For a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ one has that $x_{i}\in G{(H,\mathcal{B})}$ if and only if
-(2.19) ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$
-Indeed applying $\mu_{1}$ to Equation 1.1 one has $h_{i}d_{i}d_{i^{*}}=1$ and Equation (2.19) follows. The converse also follows by inspecting the multiplicity of $x_{1}$ in $x_{i}x_{i^{*}}$.
+For a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, we establish that $x_{i}\in G{(H,\mathcal{B})}$ if and only if:
+(2.23) $\frac{x_{i}x_{i^{*}}}{d_{i}d_{i^{*}}}=x_{1}.$
+Applying $\mu_{1}$ to Equation (1.1), we find that $h_{i}d_{i}d_{i^{*}}=1$, from which Equation (2.23) follows. The converse is similar.
 Lemma 2.27.
 Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup. Then $x_{i}\in G{(H,\mathcal{B})}$ if and only if
-(2.20) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;j\in\mathcal{J}.$
+(2.24) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;j\in\mathcal{I}.$
 Proof.
-If $x_{i}\in G{(H,\mathcal{B})}$ then applying $\mu_{j}$ to Equation (2.19) one obtains $|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1$ since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$.
-Conversely, if $\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=1$ for all $\mu_{j}$ then since ${(H,\mathcal{B})}$ is abelian it follows that ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$ and therefore $h_{i}x_{i}x_{i^{*}}=1$.
+If $x_{i}\in G{(H,\mathcal{B})}$ then applying $\mu_{j}$ to Equation (2.23) results in $|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1$, since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$.
+Conversely, if $\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=1$ for all $\mu_{j}$, then ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=x_{1}$, since ${(H,\mathcal{B})}$ is abelian, and therefore $x_{i}\in G{(H,\mathcal{B})}$ by Remark 2.26.
 ∎
-Dually one has the following:
+Dually we have the following:
 Lemma 2.28.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ if and only if
 (2.25) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;i\in\mathcal{I}.$
 Proof.
-The dual version of Equation (2.19) implies that $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if $\mu_{j}\mu_{j^{\#}}=\mu_{1}.$ Then evaluating at ${\frac{x_{i}}{d_{i}}}$ one gets the desired identity.
+The dual version of Equation (2.23) implies that $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ if and only if $\mu_{j}\mu_{j^{\#}}=\mu_{1}.$ Evaluating at ${\frac{x_{i}}{d_{i}}}$ results in the desired identity.
 ∎
 Corollary 2.29.
-In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ for any $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ one has that $n_{j}=n(H,B,\mu_{1})$.
+In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, for any $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, it holds that $n_{j}=n(H,B,\mu_{1})$.
 Proof.
-By Lemma 2.21 one has $\mid\mu_{j}(x_{i})|=|d_{i}|$ for all $i\in\mathcal{I}$.
-By Equation (2.5) one has $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$.
+By Lemma 2.28, we find that $\mid\mu_{j}(x_{i})|=|d_{i}|$ for all $i\in\mathcal{I}$.
+From Equation (2.5), we deduce that $n_{j}=\sum_{i\in\mathcal{I}}h_{i}|\mu_{j}(x_{i})|^{2}=\sum_{i\in\mathcal{I}}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$.
 ∎
@@
   2. (2)
-    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
+    $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
 Then
-(2.22) $|\mu_{j}(x_{i})|\leq|d_{i}|\;\;\text{for all}\;i\in\mathcal{I}\;\text{and}\;j\in\mathcal{J}.$
+(2.26) $|\mu_{j}(x_{i})|\leq|\mu_{1}(x_{i})|=|d_{i}|\;\;\text{, for all}\;i,j\in\mathcal{I}.$
 Proof.
 Suppose first that ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$.
-Thus $d_{i}={\mathrm{FPdim}(x_{i})}>0$ in this case.
-One has $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$ and therefore $\mu_{j}(x_{i})$ is an eigenvalue of $L_{x_{i}}$.
-It follows that $|\mu_{j}(x_{i})|\leq d_{i}$.
-Suppose now that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
-It follows that $\mathrm{FPdim}(\mu_{j})=1$, since by Remark 2.11 in this case $\mu_{j}\mapsto 1$ is the only morphism on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ which takes positive values on $^{\widehat{\mathcal{B}}}$.
+Thus $d_{i}={\mathrm{FPdim}(x_{i})}>0$.
+But $x_{i}=\sum_{j\in\mathcal{I}}\mu_{j}(x_{i})F_{j}$, and therefore $x_{i}F_{j}=\mu_{j}(x_{i})F_{j}$ meaning that $\mu_{j}(x_{i})$ is an eigenvalue of $L_{x_{i}}$.
+It follows that $|\mu_{j}(x_{i})|\leq d_{i}$, by definition of $\mathrm{FPdim}$.
+Now, suppose that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
+It follows that $\mathrm{FPdim}(\mu_{j})=1$, since by Remark 2.19, in this case, $\mu_{j}\mapsto 1$ is the only algebra morphism on $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ which takes positive values on $^{\widehat{\mathcal{B}}}$.
 Then, the equation
-(2.23) $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$
+(2.27) $\mu_{j}=\sum_{i\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}$
 implies as above that $\big{|}\mu_{j}({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$. ∎
 Lemma 2.31.
-Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that one of the following holds:
-  1. (1)
-    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
-  2. (2)
-    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
-Then $n_{j}\leq n{(H,\mathcal{B},\mu_{1})}$ for all $j\in\mathcal{J}$. Moreover, $n_{j}=n{(H,\mathcal{B},\mu_{1})}$ if and only if $\mu_{j}$ is a grouplike element.
+Under the same assumptions as Lemma 2.30, we have $n_{j}\leq n{(H,\mathcal{B},\mu_{1})}$, for all $j\in\mathcal{I}$.
+Moreover, $n_{j}=n{(H,\mathcal{B},\mu_{1})}$ if and only if $\mu_{j}$ is a grouplike element.
 Proof.
 Lemma 2.30 and Equation (2.5) give:
-$n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$
-and equality holds if and only if $|\mu_{j}(x_{i})|^{2}=|d_{i}|^{2}$. By Lemma 2.21 this happens if and only if $\mu_{j}\in G{(H,\mathcal{B})}$. ∎
+$n_{j}=\sum_{i\in\mathcal{I}}h_{i}|\mu_{j}(x_{i})|^{2}\leq\sum_{i\in\mathcal{I}}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1}),$
+and the equality holds if and only if $|\mu_{j}(x_{i})|^{2}=|d_{i}|^{2}$, if and only if $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, by Lemma 2.28. ∎
 Lemma 2.32.
-In any symmetric RN-hypergroup ${(H,\mathcal{B})}$, the set $\{{\frac{x_{i}}{d_{i}}}\}$ of normalized grouplike-elements form a finite group.
+In any RN-hypergroup ${(H,\mathcal{B})}$, the set $\{{\frac{x_{i}}{d_{i}}}\}$ of normalized grouplike-elements forms a finite group.
 Moreover, the inverse of ${\frac{x_{i}}{d_{i}}}$ is given by $\frac{x_{i^{*}}}{d_{i^{*}}}$ for all $x_{i}\in{G{(H,\mathcal{B})}}$.
 Proof.
-If $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}$ and $x_{j}x_{j^{*}}=\frac{1}{h_{j}}x_{1}$ it follows that $(x_{i}x_{j})(x_{i}x_{j})^{*}=\frac{1}{h_{i}h_{j}}x_{1}.$ Thus $(\sum_{k=1}^{m}N^{k}_{ij}x_{k})(\sum_{k=1}^{m}N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1}$ and this implies that $x_{i}x_{j}=\frac{d_{j}d_{i}}{d_{k}}x_{k}$ for a certain $k\in\mathcal{J}$.
-Moreover, since $h_{i}=\frac{1}{d_{i}^{2}}$ and $h_{j}=\frac{1}{d_{j}^{2}}$ it follows that
-$h_{k}^{-1}=m(x_{1},x_{k}x_{k}^{*})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}m(x_{1},x_{i}x_{j}x_{i^{*}}x_{j^{*}})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}\frac{1}{h_{i}h_{j}}=d_{k}^{2}$
-which shows that $x_{k}$ is also a grouplike element ($m$ is defined in §5.1). ∎
+If $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}$ and $x_{j}x_{j^{*}}=\frac{1}{h_{j}}x_{1}$ then $(x_{i}x_{j})(x_{i}x_{j})^{*}=\frac{1}{h_{i}h_{j}}x_{1}.$ Thus $(\sum_{k=1}^{m}N^{k}_{ij}x_{k})(\sum_{k=1}^{m}N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1},$ so that there is a unique $k\in\mathcal{I}$ such that $N_{i,j}^{k}$ is nonzero (if it were not the case, then by the axioms of hypergroups coupled with the RN assumption, the left-hand side of the aforementioned equation would decompose into several components).
+Consequently, we have $(N^{k}_{ij}x_{k})(N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1}$, and furthermore, $x_{i}x_{j}=N^{k}_{ij}x_{k}$.
+However, applying $\mathrm{FPdim}$ yields $N^{k}_{ij}=(d_{i}d_{j})/d_{k}$, so $\frac{x_{i}}{d_{i}}\frac{x_{j}}{d_{j}}=\frac{x_{k}}{d_{k}}$, and by the definition of a grouplike element, we have $h_{i}^{-1}=d_{i}^{2}$.
+Synthesizing all these equalities, we deduce that $((d_{i}d_{j})/d_{k})^{2}x_{k}x_{k^{*}}=(d_{i}d_{j})^{2}x_{1}$, leading to the conclusion that $x_{k}x_{k^{*}}=d_{k}^{2}x_{1}=h_{k}^{-1}x_{1}$.
+There remain to prove that if $x_{i}\in{G{(H,\mathcal{B})}}$ then so is $x_{i^{*}}$.
+By Lemma 2.24, $x_{i^{*}}$ is grouplike if and only if $h_{i^{*}}d_{i^{*}}^{2}=1$.
+But $h_{i^{*}}=h_{i}$ by Theorem 2.16, and $d_{i^{*}}=d_{i}$ by Corollary 2.18, so $h_{i^{*}}d_{i^{*}}^{2}=h_{i}d_{i}^{2}=1$, because $x_{i}$ is grouplike.
+Finally, by above, there is $k$ such that $\frac{x_{i}}{d_{i}}\frac{x_{i^{*}}}{d_{i^{*}}}=\frac{x_{k}}{d_{k}}$, and by the hypergroup axioms, $k$ must be 111.
+∎
 We denote $\overline{G{(H,\mathcal{B})}}:=\{{\frac{x_{i}}{d_{i}}}\;|\;x_{i}\in G{(H,\mathcal{B})}\}$ the above group of normalized grouplike elements.
 Clearly $\overline{G{(H,\mathcal{B})}}=G\overline{{(H,\mathcal{B})}}$, the grouplike elements of the normalized hypergroup $\overline{{(H,\mathcal{B})}}$.
-By duality one obtains the following:
+By duality we obtain the following:
 Lemma 2.33.
-If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup and the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then ${G{\widehat{(H,\mathcal{B})}}}$ is a group.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that its dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $^{G{\widehat{(H,\mathcal{B})}}}$ is a group.
 Recall that ¯=$\overline{\mathcal{B}}=\{\frac{x_{k}}{d_{k}}\;|\;x_{k}\in{\mathcal{B}}\}$.
 Lemma 2.34.
-Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup.
-If $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element then ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}},\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}\in\overline{\mathcal{B}}$, for any $x_{j}\in\mathcal{B}$.
+Let ${(H,\mathcal{B})}$ be a RN-hypergroup.
+Let $x_{i}\in{G{(H,\mathcal{B})}}$, then ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}},\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}\in\overline{\mathcal{B}}$, for any $x_{j}\in\mathcal{B}$.
 Proof.
@@
 $x_{i}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{k},$
-where $\mathcal{A}$ is a set such that $N^{k}_{ij}>0$. Multiplying the above equality by $x_{i^{*}}$, one obtains
+where $\mathcal{A}$ is a set such that $N^{k}_{ij}>0$. Multiplying the above equality by $x_{i^{*}}$, we obtain
 $\frac{1}{h_{i}}x_{j}=\sum_{k\in\mathcal{A}}N^{k}_{ij}x_{i^{*}}x_{k},$
-since $h_{i}x_{i^{*}}x_{i}=x_{1}$ as $x_{i}$ is a grouplike element.
-Since ${(H,\mathcal{B},\mu_{1})}$ is a RN-hypergroup, it follows that $x_{i^{*}}x_{k}=\alpha_{ik}x_{j}$, for some scalar $\alpha_{ik}\in\mathbb{C}$ and all $k\in\mathcal{A}$.
+since $h_{i^{*}}x_{i^{*}}x_{i}=x_{1}$ as $x_{i^{*}}$ is also grouplike by Lemma 2.32, and $h_{i^{*}}=h_{i}$ by Theorem 2.16.
+By RN assumption, $x_{i^{*}}x_{k}=\alpha_{ik}x_{j}$, for some scalar $\alpha_{ik}>0$ and all $k\in\mathcal{A}$.
 Thus $\alpha_{ik}=\frac{d_{i}d_{k}}{d_{j}}$.
-Moreover, multiplying the last equality by $x_{i}$, one obtains that $\frac{1}{h_{i}}x_{k}=x_{i^{*}}x_{j}$, which shows that $\mathcal{A}$ has a single element.
+Multiplying the second last equality by $x_{i}$, we get that $h_{i}^{-1}x_{k}=\alpha_{ik}x_{i}x_{j}$, which reformulates as ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}}=\frac{x_{k}}{d_{k}}\in\overline{\mathcal{B}}$.
+Idem for $\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}$ using right multiplications.
 ∎
 Lemma 2.35.
-Suppose that ${(H,\mathcal{B})}$ is an abelian RN hypergroup and $\mu_{j}\in{\widehat{(H,\mathcal{B})}}$ a character such that $\mu_{j}(x_{m})=d_{m}\omega_{m}$ and $\mu_{j}(x_{n})=d_{n}\omega_{n}$ for some scalars $|\omega_{m}|=|\omega_{n}|=1$.
+Suppose that ${(H,\mathcal{B})}$ is an abelian RN hypergroup and $\mu_{j}\in\widehat{\mathcal{B}}$ a character such that $\mu_{j}(x_{m})=d_{m}\omega_{m}$ and $\mu_{j}(x_{n})=d_{n}\omega_{n}$ for some scalars $|\omega_{m}|=|\omega_{n}|=1$.
 Then $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ for all constituents $x_{p}$ of $x_{m}x_{n}$.
 Proof.
-Straightforward from the absolute value triangle inequality. ∎
+Apply the absolute value triangle inequality and Lemma 2.30. ∎
 Lemma 2.36.
-If ${(H,\mathcal{B})}$ is an abelian RN hypergroup then ${{G{\widehat{(H,\mathcal{B})}}}}$ is a group.
+If ${(H,\mathcal{B})}$ is an abelian RN hypergroup then $^{{G{\widehat{(H,\mathcal{B})}}}}$ is a group.
 Proof.
-Suppose that $\mu_{j},\mu_{k}\in{{G{(H,\mathcal{B})}}}$.
-By Lemma 2.21 it follows that $\mu_{j}(x_{m})=\omega_{m}\mathrm{FPdim}(x_{m})$ for some roots of unity $\omega_{m}$.
-Similarly $\mu_{k}(x_{m})=\eta_{m}\mathrm{FPdim}(x_{m})$ for some other root of unity $\eta_{m}$.
-Moreover, by lemma 2.28, since ${(H,\mathcal{B})}$ is RN it follows that all the constituents $x_{m}x_{n}$ when evaluated at $\mu_{j}$ have the scalar $\omega_{m}\omega_{n}$.
-On the other hand, evaluated at $\mu_{k}$ these constituents give the scalar $\eta_{m}\eta_{n}$.
-Thus
-$(\mu_{j}\star\mu_{k})(x_{m}x_{n})$ $=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
-$=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
-$=$ $\omega_{m}\omega_{n}\eta_{m}\eta_{n}\mathrm{FPdim}(x_{m}x_{n})$
-On the other hand, note that $\mu_{j}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{j}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\omega_{m}\mathrm{FPdim}(x_{m})$ and similarly $\mu_{k}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{k}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\eta_{m}\mathrm{FPdim}(x_{m}).$ Therefore
-$(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})=\omega_{m}\eta_{m}\mathrm{FPdim}(x_{m})\omega_{n}\eta_{n}\mathrm{FPdim}(x_{n})$
-This shows that $(\mu_{j}\star\mu_{k})(x_{m}x_{n})=(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})$ which shows that $(\mu_{j}\star\mu_{k})\in{\widehat{\mathcal{B}}}$.
-Moreover, $(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\omega_{p}\eta_{p}$ and Lemma 2.21 implies that $\mu_{j}\star\mu_{k}\in G{\widehat{(H,\mathcal{B})}}$.
+Suppose that $\mu_{j},\mu_{k}\in{{G{\widehat{(H,\mathcal{B})}}}}$.
+By Lemma 2.28, $\mu_{j}(x_{m})=\omega_{m}d_{m}$ and $\mu_{k}(x_{m})=\eta_{m}d_{m}$, for some roots of unity $\omega_{m}$ and $\eta_{m}$.
+Now,
+$(\mu_{j}\star\mu_{k})(x_{m}x_{n})=\sum_{p}N^{p}_{m,n}d_{p}(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}}),$
+but by Equation (2.7),
+$(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}})=\mu_{j}({\frac{x_{p}}{d_{p}}})\mu_{k}({\frac{x_{p}}{d_{p}}}),$
+and by Lemma 2.35, $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ and $\mu_{k}(x_{p})=\eta_{m}\eta_{n}d_{p}$, when $N^{p}_{m,n}$ is nonzero. Thus
+$(\mu_{j}\star\mu_{k})(x_{m}x_{n})=\sum_{p}N^{p}_{m,n}d_{p}\omega_{m}\omega_{n}\eta_{m}\eta_{n}=\omega_{m}\omega_{n}\eta_{m}\eta_{n}d_{m}d_{n}$
+On the other hand,
+$(\mu_{j}\star\mu_{k})(x_{m})=d_{m}(\mu_{j}\star\mu_{k})({\frac{x_{m}}{d_{m}}})=d_{m}\mu_{j}({\frac{x_{m}}{d_{m}}})\mu_{k}({\frac{x_{m}}{d_{m}}})=d_{m}\omega_{m}\eta_{m}.$
+Idem, $(\mu_{j}\star\mu_{k})(x_{n})=d_{n}\omega_{n}\eta_{n}$. So
+$(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})=d_{m}\omega_{m}\eta_{m}d_{n}\omega_{n}\eta_{n}=(\mu_{j}\star\mu_{k})(x_{m}x_{n}),$
+by above, which shows that $(\mu_{j}\star\mu_{k})\in{\widehat{\mathcal{B}}}$.
+Moreover, $(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}})=\mu_{j}({\frac{x_{p}}{d_{p}}})\mu_{k}({\frac{x_{p}}{d_{p}}})=\omega_{p}\eta_{p}$, which by Lemma 2.28, implies that $\mu_{j}\star\mu_{k}\in G{\widehat{(H,\mathcal{B})}}$.
 ∎
 Proposition 2.37.
-Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup and $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$. It follows that $\mu\star\mu_{k}\in\widehat{B}$ for any $\mu_{k}\in\widehat{B}$.
+Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup and $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$. Then, $\mu\star\mu_{k}\in\widehat{B}$, for any $\mu_{k}\in\widehat{B}$.
 Proof.
-Since $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$ and ${(H,\mathcal{B})}$ is RN, by Lemma 2.21 one has $\mu(x_{i})=\omega_{i}d_{i}$ for some root of unity $\omega_{i}$ and for all $i\in\mathcal{I}$.
-Here $d_{i}=\mathrm{FPdim}(x_{i})$.
-Then as above it follows that $\mu(x_{k})=\omega_{i}\omega_{j}d_{k}$ for any constituent $x_{k}$ of $x_{i}x_{j}$.
-One can compute directly that $\mu\star\mu_{k}$ and $\mu_{k}\star\mu$ are morphisms of algebras.
-Indeed, note that $(\mu\star\mu_{k})({\frac{x_{i}}{d_{i}}})=\omega_{i}\mu_{k}({\frac{x_{i}}{d_{i}}})$ and therefore
+Since $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$ and ${(H,\mathcal{B})}$ is RN, by Lemma 2.28, $\mu(x_{i})=\omega_{i}d_{i}$ for some root of unity $\omega_{i}$ and for all $i\in\mathcal{I}$.
+By Lemma 2.35, $\mu(x_{l})=\omega_{i}\omega_{j}d_{l}$ for any constituent $x_{l}$ of $x_{i}x_{j}$, therefore
 $(\mu\star\mu_{k})(x_{i}x_{j})=\sum_{l=1}^{m}N^{l}_{ij}(\mu\star\mu_{k})(x_{l})=\omega_{i}\omega_{j}\sum_{l=1}^{m}N^{l}_{ij}\mu_{k}(x_{l})=\omega_{i}\omega_{j}\mu_{k}(x_{i}x_{j}).$
@@
 Thus $(\mu\star\mu_{k})(x_{i}x_{j})=(\mu\star\mu_{k})(x_{i})(\mu\star\mu_{k})(x_{j})$ which shows that $\mu\star\mu_{k}\in{\widehat{\mathcal{B}}}$. ∎
-In the dual situation, Lemma 2.29 and Proposition 2.30 can be written as follows:
+By duality, Lemma 2.36 and Proposition 2.37 reformulate as follows:
 Proposition 2.38.
-If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then $G{(H,\mathcal{B})}$ is a group.
-Moreover, left multiplication by $L_{{\frac{x_{i}}{d_{i}}}}$ for any $x_{i}\in{G{(H,\mathcal{B})}}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
-Note that Lemmata 2.25, 2.26, 2.27, 2.29 and Propositions 2.30, 2.31 imply the following:
+If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN, then $\overline{G{(H,\mathcal{B})}}$ is a group.
+Moreover, the left multiplication $L_{{\frac{x_{i}}{d_{i}}}}$, for any $x_{i}\in{G{(H,\mathcal{B})}}$, permutes the normalized basis $\overline{\mathcal{B}}$.
+Lemmas 2.32, 2.33, 2.34, 2.36 and Propositions 2.37, 2.38 imply the following:
 Corollary 2.39.
@@
   2. (2)
-    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
-Then the following statements hold:
+    $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
+Then, the following statements hold:
   1. (1)
@@
   2. (2)
-    The operator given by left multiplication $L_{\frac{x_{i}}{d_{i}}}$ with $x_{i}\in G{(H,\mathcal{B})}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
+    The left multiplication operator $L_{\frac{x_{i}}{d_{i}}}$, with $x_{i}\in G{(H,\mathcal{B})}$, permutes the normalized basis $\overline{\mathcal{B}}$.
   3. (3)
-    The operator given by left multiplication $L_{\mu_{j}}$ with $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ permutes the standard basis $^{{\widehat{\mathcal{B}}}}$.
-2.5. Some results on $P$ and ${\widehat{P}}$.
+    The left multiplication operator $L_{\mu_{i}}$, with $\mu_{i}\in G{\widehat{(H,\mathcal{B})}}$, permutes the basis $^{{\widehat{\mathcal{B}}}}$.
+2.6. Some results on $P$ and $^{\widehat{P}}$.
+They denote $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$ and $\prod_{j\in\mathcal{I}}\mu_{j}$, respectively.
 Lemma 2.40.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $x_{i}\in G{(H,\mathcal{B})}$ be a grouplike element. Then
-(2.24) $<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\pm 1.$
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $x_{i}\in G{(H,\mathcal{B})}$ be a grouplike element. Then
+(2.28) $(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})=\pm 1.$
 Proof.
-By the definition of the multiplication in the dual,
-$<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}).$
-We group together the factors of $\mu_{j}$ and $\mu_{j\#}$ in the above product.
-Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j^{\#}}({\frac{x_{i}}{d_{i}}})=\mid\mu_{j}({\frac{x_{i}}{d_{i}}})\mid^{2}=1$ by Lemma 2.20.
-On the other hand, if $\mu_{j}=\mu_{j^{\#}}$, then $\mu_{j}({\frac{x_{i}}{d_{i}}})\in\mathbb{R}$ and therefore $\mu_{j}({\frac{x_{i}}{d_{i}}})=\pm 1$ by the same Lemma 2.20.
+By Equation (2.7), defining the multiplication in the dual,
+$(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}}).$
+By Equation (2.8), defining the involution in the dual, and then Lemma 2.27,
+$\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j^{\#}}({\frac{x_{i}}{d_{i}}})=|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1.$
+We group together the factors $\mu_{j}$ and $\mu_{j\#}$ in the product above.
+Now, if $\mu_{j}=\mu_{j^{\#}}$, then $\mu_{j}({\frac{x_{i}}{d_{i}}})\in\mathbb{R}$, and therefore, $\mu_{j}({\frac{x_{i}}{d_{i}}})=\pm 1$ by Lemma 2.27.
 ∎
 Definition 2.41.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element.
-We denote by ${\mathrm{sgn}}(x_{i})$ the above value of the product $\prod_{j=1}^{m}\mu_{j}$ evaluated at ${\frac{x_{i}}{d_{i}}}$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element.
+Then $(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})$ will be denoted ${\mathrm{sgn}}(x_{i})$.
 The dual version of the above result is the following:
 Lemma 2.42.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then
-(2.25) $<\mu_{j},\;{\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}>=\pm 1.$
-By duality we can also define:
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then
+(2.29) $\mu_{j}(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})=\pm 1.$
+By duality, we can also define:
 Definition 2.43.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element.
-We denote by ${\mathrm{sgn}}(\mu_{j})$ the above value of the product ${\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$ evaluated by $\mu_{j}$.
-Remark 2.37.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element.
-Then $L_{{\frac{x_{i}}{d_{i}}}}$, the left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$ induces a permutation on $\overline{\mathcal{B}}$.
-Because ${\frac{x_{i}}{d_{i}}}=\sum_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})F_{j}$, the determinant $\det(L_{\frac{x_{i}}{d_{i}}})$ of the matrix of the operator $L_{{\frac{x_{i}}{d_{i}}}}$ is $\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})$.
-On the other hand, $L_{{\frac{x_{i}}{d_{i}}}}$ permutes the normalized basis ¯:=$\overline{\mathcal{B}}:=\{{\frac{x_{i}}{d_{i}}}\}_{i\in\mathcal{I}}$, and therefore, ${\mathrm{sgn}}(x_{i})$ is the signature of the corresponding permutation.
-Remark 2.38.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
-It follows that ${\mathrm{sgn}}(\mu_{j})=\det(L_{\mu_{j}})=\pm 1$ since $\mu_{j}$ determines a permutation of $^{\widehat{\mathcal{B}}}$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element.
+Then $\mu_{j}(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})$ will be denoted ${\mathrm{sgn}}(\mu_{j})$.
+Proposition 2.44.
+Following the assumptions of Corollary 2.39, let $x_{i}\in{G{(H,\mathcal{B})}}$ and $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ be grouplike elements. Then
+${\mathrm{sgn}}(x_{i})=\det(L_{\frac{x_{i}}{d_{i}}})\text{ and }{\mathrm{sgn}}(\mu_{j})=\det(L_{\mu_{j}})$
+are the signature of the permutations of $\overline{\mathcal{B}}$ by $L_{{\frac{x_{i}}{d_{i}}}}$, and of $\widehat{\mathcal{B}}$ by $L_{\mu_{j}}$, respectively.
+Proof.
+By Corollary 2.39, the left multiplication operator $L_{{\frac{x_{i}}{d_{i}}}}$ permutes $\overline{\mathcal{B}}$, but the signature of this permutation is the determinant $\det(L_{\frac{x_{i}}{d_{i}}})$.
+Now ${\frac{x_{i}}{d_{i}}}=\sum_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})F_{j}$, so $\det(L_{\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})$.
+Idem for ${\mathrm{sgn}}(\mu_{j})$.
+∎
 Definition 2.45.
-A symmetric hypergroup ${(H,\mathcal{B})}$ is called perfect if it has no nontrivial grouplike elements.
+A hypergroup ${(H,\mathcal{B})}$ is called perfect if it has no nontrivial grouplike elements.
 3. Galois symmetries for hypergroups
-Let $(H,\mathcal{B})$ be an abelian rational hypergroup.
-For any standard element $x_{i}\in\mathcal{B}$, one can write $x_{i}=\sum_{j=1}^{m}{\alpha_{{}_{ij}}}F_{j}$, for some scalars ${\alpha_{{}_{ij}}}\in\mathbb{C}$.
-Note that ${\alpha_{{}_{ij}}}=\mu_{j}(x_{i})$.
-Let $\mathbb{K}=\mathbb{Q}({\alpha_{{}_{ij}}})$ be the field obtained by adjoining all ${\alpha_{{}_{ij}}}\in\mathbb{C}$ to $\mathbb{Q}$.
+Let $(H,\mathcal{B})$ be an abelian rational hypergroup (i.e.
+$N^{k}_{ij}\in\mathbb{Q}$).
+For any standard element $x_{i}\in\mathcal{B}$, we can write $x_{i}=\sum_{j\in\mathcal{I}}{\alpha_{{}_{ij}}}F_{j}$, with ${\alpha_{{}_{ij}}}=\mu_{j}(x_{i})\in\mathbb{C}$.
+Let $\mathbb{K}=\mathbb{Q}({\alpha_{{}_{ij}}})$ be the field obtained by adjoining all ${\alpha_{{}_{ij}}}$ to $\mathbb{Q}$.
 Remark 3.1.
@@
 3.1. Permutation of characters $\mu_{j}$
-Let ${(H,\mathcal{B})}$ be any abelian rational hypergroup (i.e.
-$N^{k}_{ij}\in\mathbb{Q}$).
-For any character $\mu_{j}:H\rightarrow\mathbb{C}$ and $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, define $\sigma.\mu_{j}\in H^{*}$ as the linear function on $H$ on which the standard basis $\mathcal{B}$ is given by $[\sigma.\mu_{j}](x_{i})=\sigma(\mu_{j}(x_{i}))=\sigma({\alpha_{{}_{ij}}}).$
+Let ${(H,\mathcal{B})}$ be any abelian rational hypergroup.
+For any character $\mu_{j}:H\rightarrow\mathbb{C}$ and $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, define $\sigma.\mu_{j}\in H^{*}$ as the linear function on $H$ such that $[\sigma.\mu_{j}](x_{i})=\sigma(\mu_{j}(x_{i}))=\sigma({\alpha_{{}_{ij}}})$, for all $x_{i}\in{\mathcal{B}}$.
 Lemma 3.2.
-For any abelian rational hypergroup ${(H,\mathcal{B})}$, it follows that $\sigma.\mu_{j}:H\rightarrow\mathbb{C}$ is an algebra map.
+For any abelian rational hypergroup ${(H,\mathcal{B})}$, the function $\sigma.\mu_{j}$ is an algebra map.
 Thus $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$.
 Proof.
-Indeed, suppose that $x_{i_{1}}x_{i_{2}}=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}$. Since $N^{k}_{{i_{1}},{i_{2}}}\in\mathbb{Q}$,
-$[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}](\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}[\sigma.\mu_{j}](x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
-On the other hand, note that
+Suppose that $x_{i_{1}}x_{i_{2}}=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k}$. Since $N^{k}_{{i_{1}},{i_{2}}}\in\mathbb{Q}$,
+$[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}](\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k})=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}[\sigma.\mu_{j}](x_{k})=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
+On the other hand,
 $[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$ $=$ $\sigma(\mu_{j}({x_{i_{1}}}))\sigma(\mu_{j}({x_{i_{2}}}))=\sigma(\mu_{j}({x_{i_{1}}})\mu_{j}({x_{i_{2}}}))$
-$=$ $\sigma(\mu_{j}({x_{i_{1}}}{x_{i_{2}}}))=\sigma(\mu_{j}(\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}))$
-$=$ $\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
-Thus $[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$, so that $\sigma.\mu_{j}$ is an algebra map.
-It is easy to see that, in this case, $\sigma.(\sigma^{\prime}.\mu_{j})=(\sigma\sigma^{\prime}).\mu_{j}.$ It follows that $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$ of all algebra homomorphisms $\mu_{j}:H\rightarrow\mathbb{C}$.
-∎
-We denote by $\mathcal{J}:=\{1,\dots,m\}$ the set of all indices for the characters $\mu_{j}:H\rightarrow\mathbb{C}$.
-It follows that for any $j\in\mathcal{J}$, there is an index $\tau(j)=\tau_{\sigma}(j)\in\mathcal{J}$ such that $\sigma.\mu_{j}=\mu_{\tau_{\sigma}(j)}$, and therefore
-(3.1) $\sigma({\alpha_{{}_{ij}}})=\alpha_{{}_{i\tau_{\sigma}(j)}}\;\;\text{for all}\;i,j.$
-It is easy to see that, in this case, $\tau_{\sigma}$ is a permutation of $\mathcal{J}$.
+$=$ $\sigma(\mu_{j}({x_{i_{1}}}{x_{i_{2}}}))=\sigma(\mu_{j}(\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k}))$
+$=$ $\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
+Thus $[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$, so that $\sigma.\mu_{j}$ is an algebra map. Now,
+$[\sigma.(\sigma^{\prime}.\mu_{j})](x_{i})=\sigma([\sigma^{\prime}.\mu_{j}](x_{i}))=\sigma(\sigma^{\prime}(\mu_{j}(x_{i})))=(\sigma\sigma^{\prime})(\mu_{j}(x_{i}))=[\sigma\sigma^{\prime}.\mu_{j}](x_{i})$
+It follows that $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$. ∎
+Thus, there is a permutation $\tau_{\sigma}$ of ${\mathcal{I}}$ such that $\sigma.\mu_{j}=\mu_{\tau_{\sigma}(j)}$, and therefore
+(3.1) $\sigma({\alpha_{{}_{ij}}})=\alpha_{{}_{i\tau_{\sigma}(j)}}\;\;\text{for all}\;i,j\in{\mathcal{I}}.$
 Corollary 3.3.
@@
 Proof.
-Note that Equation (3.1) shows that $\sigma(\mathbb{K})\subseteq\mathbb{K}$, for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$.
-Therefore, this is a normal extension and therefore a Galois extension.
+Equation (3.1) shows that $\sigma(\mathbb{K})\subseteq\mathbb{K}$, for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$.
+Therefore, this is a normal extension and consequently a Galois extension (as ${\mathrm{c}har}(\mathbb{Q})=0$).
 ∎
 Remark 3.4.
-Note that the map $\rho:\mathrm{Gal}(\mathbb{K}/\mathbb{Q})\rightarrow\mathbb{S}_{m},\sigma\mapsto\tau_{\sigma}$ is injective.
+The map $\rho:\mathrm{Gal}(\mathbb{K}/\mathbb{Q})\rightarrow\mathbb{S}_{\mathcal{I}},\sigma\mapsto\tau_{\sigma}$ is injective.
 Indeed, if $\tau_{\sigma}=\tau_{\sigma^{\prime}}$, then $\sigma({\alpha_{{}_{ij}}})=\sigma^{\prime}({\alpha_{{}_{ij}}})$, i.e.
 $\sigma^{-1}\circ\sigma^{\prime}({\alpha_{{}_{ij}}})={\alpha_{{}_{ij}}}$, for all $i,j$.
-Thus $\sigma=\sigma^{\prime}$ on $\mathbb{K}$.
+Thus $\sigma=\sigma^{\prime}$.
 Proposition 3.5.
-Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, with the above notations,
-(3.2) $\sigma(n_{k})={n_{\tau(k)}}.$
+Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, let $\tau=\tau_{\sigma}$, with the above notations,
+(3.2) $\sigma(n_{j})={n_{\tau(j)}}.$
 Proof.
-Applying $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$ to Equation (2.5), one obtains the following:
-$\sum_{i=1}^{m}{h_{i}}\sigma(\alpha_{{}_{il}})\sigma(\alpha_{{}_{i^{*}k}})=\delta_{l,k}\sigma(n_{k}).$
-In turn, this equation can be written as
-$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{l,k}\sigma(n_{k}).$
+Applying $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$ to the orthogonality relation (2.5), we obtain:
+$\sum_{i\in\mathcal{I}}h_{i}\mu_{\tau(j)}(x_{i})\mu_{\tau(k)}(x_{i^{*}})=\delta_{j,k}\sigma(n_{j}),$
 On the other hand, by the same orthogonality relation:
-$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{\tau(l),\tau(k)}{n_{\tau(k)}}.$
-Therefore, for $l=k$, it follows that $\sigma(n_{k})={n_{\tau(k)}}.$
-We continue with the following definition.
+$\sum_{i\in\mathcal{I}}h_{i}\mu_{\tau(j)}(x_{i})\mu_{\tau(k)}(x_{i^{*}})=\delta_{\tau(j),\tau(k)}{n_{\tau(j)}}.$
+Therefore, with $j=k$, we get that $\sigma(n_{j})={n_{\tau(j)}}$. ∎
 Definition 3.6.
-A hypergroup ${(H,\mathcal{B})}$ is $h$-integral if $h_{i}:=1/N^{1}_{i,i^{*}}\in\mathbb{Z}$, for all $1\leq i\leq m$.
+A hypergroup ${(H,\mathcal{B})}$ is $h$-integral if $h_{i}:=1/N^{1}_{i,i^{*}}\in\mathbb{Z}$, for all $i\in\mathcal{I}$.
 Remark 3.7.
-In view of Equation (2.11), the Grothendieck ring of any weakly-integral fusion category $\mathcal{C}$ has a $h$-integral dual ${\widehat{{K(\mathcal{C})}}}$, see [Ost15, Theorem 2.13].
+Let ${(H,\mathcal{B},\mu_{1})}$ be the Grothendieck ring of any weakly-integral fusion category $\mathcal{C}$.
+By [Ost15, Theorem 2.13] and Equation (2.11), $\widehat{h}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{n_{i}}$ is an integer, for all $i\in\mathcal{I}$.
+Thus, the dual $^{\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
 See also Remark 9.1.
@@
 Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup with a $h$-integral dual.
-Then ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral.
-Moreover, if ${(H,\mathcal{B},\mu_{1})}$ is rational, then $\widehat{h}_{k}=\widehat{h}_{\tau(k)}$, for all $1\leq k\leq m$.
+Then it is weakly-integral.
+Moreover, if it is rational, then $\widehat{h}_{k}=\widehat{h}_{\tau(k)}$, for all $k\in\mathcal{I}$.
 Proof.
-By definition,
-$n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}\in\mathbb{Z}.$
-If ${(H,\mathcal{B},\mu_{1})}$ is rational, then Equation (3.2) can be written as $\sigma\bigg{(}\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{k}}}\bigg{)}=\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{\tau(k)}}}.$ In particular, $\widehat{h}_{k}=\widehat{h}_{\tau(k)}.$, since ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $h$-integral.
+By Lemma 2.8 and Definition 2.5,
+$n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{i\in\mathcal{I}}{\widehat{h}}_{i}\in\mathbb{Z}.$
+If ${(H,\mathcal{B},\mu_{1})}$ is rational, then Equation (3.2) can be written as $\sigma\big{(}\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h}_{i}}\big{)}=\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h}_{\tau(i)}}.$ In particular, $\widehat{h}_{k}=\widehat{h}_{\tau(i)}$, since $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $h$-integral.
 ∎
@@
 Thus $\mathcal{B}=\mathcal{B}_{1}\sqcup\mathcal{B}_{0}$.
-Note that if ${(H,\mathcal{B})}$ is a RN-hypergroup then ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$.
-Note that, in the case of an abelian hypergroup ${(H,\mathcal{B})}$, since $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$, it follows that $x_{i}$ is non-vanishing if and only if it is invertible.
+By the proof of Lemma 2.27, if ${(H,\mathcal{B})}$ is a RN-hypergroup then ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$.
+In the abelian case, since $x_{i}=\sum_{j\in\mathcal{I}}\mu_{j}(x_{i})F_{j}$, it follows that $x_{i}$ is non-vanishing if and only if it is invertible in $H$ (i.e.
+$\det(L_{x_{i}})$ nonzero).
 Thus, in this case, $\mathcal{B}_{1}$ coincides with the set of all standard elements that are invertible in $H$.
-Denote ${\widehat{P}}:=\prod_{j=1}^{m}\mu_{j}$. Note that in any abelian normalizable ${(H,\mathcal{B},\mu_{1})}$ one has:
-(4.1) $\prod_{j=1}^{m}\mu_{j}=\sum_{i=1}^{m}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}.$
+In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$:
+(4.1) ${\widehat{P}}:=\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{i\in\mathcal{I}}{\widehat{P}}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}.$
 Proposition 4.1.
-Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
-(4.2) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}}){\widehat{E}}_{i}$
+Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then:
+(4.2) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}$
 where $\det(L_{{\frac{x_{i}}{d_{i}}}})$ is the determinant of the left multiplication operator by ${\frac{x_{i}}{d_{i}}}$ on $H$.
 Proof.
-Recall that $<{\widehat{E}}_{i},\frac{x_{j}}{d_{j}}>=\delta_{i,j}$ for all $i,j\in\mathcal{I}$.
-In order to prove the above equality, it is enough to evaluate both sides at each ${\frac{x_{i}}{d_{i}}}$, and prove that one obtains the same value.
-If $x_{i}$ is a vanishing element, then the above left and right hand side both vanish when evaluated at ${\frac{x_{i}}{d_{i}}}$.
-If $x_{i}$ is not a vanishing element then
-$\big{[}\prod_{j=1}^{m}\mu_{j}\big{]}({\frac{x_{i}}{d_{i}}})=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})=\det(L_{{\frac{x_{i}}{d_{i}}}})$
-and the proof is compete. ∎
-Dually, applying the above result to ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$, one obtains a decomposition of the basis ^=^${\widehat{\mathcal{B}}}={\widehat{\mathcal{B}}}_{0}\sqcup{\widehat{\mathcal{B}}}_{1}$.
-The above proposition implies the following:
+Recall that $\widehat{F}_{i}(\frac{x_{j}}{d_{j}})=\delta_{i,j}$, and $\frac{x_{j}}{d_{j}}=\sum_{i\in\mathcal{I}}\mu_{i}(\frac{x_{j}}{d_{j}})F_{i}$. Thus, for all $j\in\mathcal{I}$,
+$\left[\sum_{i\in\mathcal{I}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}\right](\frac{x_{j}}{d_{j}})=\det(L_{\frac{x_{j}}{d_{j}}})=\prod_{i\in\mathcal{I}}\mu_{i}(\frac{x_{j}}{d_{j}})=\left[\prod_{i\in\mathcal{I}}\mu_{i}\right](\frac{x_{j}}{d_{j}})$
+It follows that
+$\prod_{i\in\mathcal{I}}\mu_{i}=\sum_{i\in\mathcal{I}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}.$
+Indeed, $x_{i}\in{\mathcal{B}_{1}}$ if and only if $\det(L_{{\frac{x_{i}}{d_{i}}}})$ is nonzero, justifying the last equality.∎
+Dually, applying the above result to $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, we obtain a decomposition of the basis ^=^${\widehat{\mathcal{B}}}={\widehat{\mathcal{B}}}_{0}\sqcup{\widehat{\mathcal{B}}}_{1}$.
+Proposition 4.1 implies the following:
 Proposition 4.2.
-Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
-(4.3) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{\widehat{\mathcal{B}}}_{1}}\det(L_{\mu_{j}})F_{j},$
+Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then:
+(4.3) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{\widehat{\mathcal{B}}}_{1}}\det(L_{\mu_{j}})F_{j},$
 where $\det(L_{\mu_{j}})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
@@
 Proposition 4.3.
-An abelian hypergroup ${(H,\mathcal{B})}$ is Burnside if and only if one has ${G{(H,\mathcal{B})}}=\mathcal{B}_{1}$.
+An abelian hypergroup ${(H,\mathcal{B})}$ is Burnside if and only if we have ${G{(H,\mathcal{B})}}=\mathcal{B}_{1}$.
 Corollary 4.4.
-An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if the above identity becomes
-(4.4) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
+An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if the following equality holds:
+(4.4) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\mathrm{sgn}}(x_{i})\widehat{F}_{i},$
+where the notation ${\mathrm{sgn}}(x_{i})$ was introduced in Definition 2.41.
 Proof.
-It follows from Proposition 4.1. Recall from Definition 2.34 that ${\mathrm{sgn}}(x_{i})$ is the determinant of $L_{\frac{x_{i}}{d_{i}}}$, when $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
+Immediate from Propositions 4.1 and 4.3. ∎
 Corollary 4.5.
 An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if
-(4.5) $(\prod_{j=1}^{m}\mu_{j})^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
+(4.5) $(\prod_{j\in\mathcal{I}}\mu_{j})^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
 Corollary 4.6.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${G{(H,\mathcal{B})}}$ has odd order. Then ${(H,\mathcal{B})}$ is Burnside if and only if
-(4.6) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}$
+An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, such that ${G{(H,\mathcal{B})}}$ has odd order, is Burnside if and only if
+(4.6) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in G(H,B)}\widehat{F}_{i}.$
 Proof.
-Note that in this case ${\mathrm{sgn}}(x_{i})=1$ for all $x_{i}\in{G{(H,\mathcal{B})}}$, since each ${\frac{x_{i}}{d_{i}}}\in{G{(H,\mathcal{B})}}$ has odd order. ∎
+Every ${\frac{x_{i}}{d_{i}}}\in{G{(H,\mathcal{B})}}$ has odd order, so ${\mathrm{sgn}}(x_{i})=1$. ∎
 Definition 4.7.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup.
-We say that the hypergroup ${(H,\mathcal{B})}$ is dual-Burnside if ${\widehat{(H,\mathcal{B})}}$ is Burnside.
-In other words, for all $\mu_{j}\in{\widehat{\mathcal{B}}}$, the following are equivalent:
+An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside; in other words, if for all $\mu_{j}\in{\widehat{\mathcal{B}}}$, the following are equivalent:
   1. (1)
@@
     $\mu_{j}\in G\widehat{{(H,\mathcal{B})}}$, i.e $\mu_{j}$ is a grouplike element.
-The dual of Corollary 4.4 is the following:
+This can be shortened to $G\widehat{{(H,\mathcal{B})}}={\widehat{\mathcal{B}}}_{1}$.
+Here are the dual of Corollaries 4.4, 4.5 and 4.6:
 Corollary 4.8.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if:
-(4.7) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}{\mathrm{sgn}}(\mu_{j})F_{j},$
-where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
+An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if:
+(4.7) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}{\mathrm{sgn}}(\mu_{j})F_{j}.$
 Corollary 4.9.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if :
-(4.8) $(\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}})^{2}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}.$
-where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
+An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if :
+(4.8) $(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})^{2}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}.$
 Corollary 4.10.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. If ${{G{\widehat{(H,\mathcal{B})}}}}$ has odd order then ${(H,\mathcal{B})}$ is dual-Burnside if and only if
-(4.9) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{{G{\widehat{(H,\mathcal{B})}}}}}F_{j}.$
-4.1. Inequalities for $P$ and ${\widehat{P}}$
+An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, with $^{{G{\widehat{(H,\mathcal{B})}}}}$ of odd order, is dual-Burnside if and only if
+(4.9) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{{G{\widehat{(H,\mathcal{B})}}}}}F_{j}.$
+4.1. Inequalities for $P$ and $^{\widehat{P}}$
+Recall that $P:=\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}$ and ${\widehat{P}}:=\prod_{j\in\mathcal{I}}\mu_{j}$.
 Proposition 4.11.
@@
 Then $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$.
-Moreover one has equality if and only if $x_{i}\in G{(H,\mathcal{B})}$.
+Moreover, the equality holds if and only if $x_{i}\in G{(H,\mathcal{B})}$.
 Proof.
-First suppose that ${(H,\mathcal{B})}$ is RN.
+First, suppose that ${(H,\mathcal{B})}$ is RN.
 Then $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ by Frobenius-Perron theory.
-This implies that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|={\prod_{j=1}^{m}}|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$.
-Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$.
-Then Lemma 2.20 implies that $x_{i}$ is a grouplike element.
-Now suppose that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
-Then since $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$ it follows that $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ since $\mathrm{FPdim}(\mu_{j})=1$.
-This implies again that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$.
-Moreover one has equality if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{J}$.
-Then, again, Lemma 2.20 implies that $x_{i}$ is a grouplike element.
+This implies that $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=\prod_{j\in\mathcal{I}}|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$.
+Moreover, the equality holds if and only if $|{\frac{x_{i}}{d_{i}}}(\mu_{j})|=1$ for all $j\in\mathcal{I}$, if and only if $x_{i}$ is grouplike, by Lemma 2.27.
+Next, suppose that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
+Since $\mu_{j}=\sum_{i\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}$, then $\{\mu_{j}({\frac{x_{i}}{d_{i}}})\}_{i\in\mathcal{I}}$ are the eigenvalues of $L_{\mu_{j}}$.
+By Frobenius-Perron theory, $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq\mathrm{FPdim}(\mu_{j})=1$, as $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is normalized.
+Thus, $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|\leq 1$, again.
+The rest is as before.
 ∎
@@
 Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup.
-One has that ${\widehat{P}}^{2}$ is an idempotent if and only if the non-vanishing elements $x_{i}$ of ${(H,\mathcal{B})}$ satisfy ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$.
+Then, ${\widehat{P}}^{2}$ is an idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$, for all $x_{i}\in\mathcal{B}_{1}$.
 Proof.
-Note that Equation (4.1) one has ${\widehat{P}}^{4}={\widehat{P}}^{2}$ if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{4}={\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}$ for all $i\in\mathcal{I}$.
-Therefore if ${\widehat{P}}^{2}$ is idempotent then either ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=0$ or ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ for all $i\in\mathcal{I}$.
-In the first case, since ${\widehat{P}}({\frac{x_{i}}{d_{i}}})={\prod_{j=1}^{m}}\mu_{j}({\frac{x_{i}}{d_{i}}})$ it follows that $x_{i}$ is a vanishing element.
-On the other hand if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$ implies ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$.
+By Equation (4.1), ${\widehat{P}}^{4}={\widehat{P}}^{2}$ if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{4}={\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}$ for all $i\in\mathcal{I}$.
+Therefore, ${\widehat{P}}^{2}$ is idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=0$ or ${\widehat{P}}({\frac{x_{i}}{d_{i}}})^{2}=1$, for all $i\in\mathcal{I}$.
+Since ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})$, the second case occurs if and only if $x_{i}\in\mathcal{B}_{1}$, if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$.
 ∎
 Corollary 4.13.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.23. Then ${\widehat{P}}^{2}$ is idempotent if and only if ${(H,\mathcal{B})}$ is Burnside.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30.
+Then, ${\widehat{P}}^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is Burnside.
 Proof.
-In this case, by Proposition 4.11, one has $|\mu_{j}({\frac{x_{i}}{d_{i}}})|\leq 1$ for all $j\in\mathcal{J}$ Thus ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$ implies $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=1$ and Lemma 2.20 implies that $x_{i}\in G{(H,\mathcal{B})}$.
-Conversely, suppose that ${(H,\mathcal{B})}$ is Burnside.
-It follows from Equation (4.1) that ${\widehat{P}}=\sum_{x_{i}\in G{(H,\mathcal{B})}}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$.
-Note that ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\det(L_{\frac{x_{i}}{d_{i}}})$ the operator of left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$.
-By Corollary 2.32 this is the determinant of a permutation matrix and therefore equal to ${\mathrm{sgn}}(x_{i})=\pm 1$.
-∎
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Define also $P={\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$. One has
-(4.10) $P=\sum_{j=1}^{m}\mu_{j}(P)F_{j}.$
-The dual version of the last three results imply the following:
+By Proposition 4.12, ${\widehat{P}}^{2}$ is an idempotent if and only if ${\widehat{P}}({\frac{x_{i}}{d_{i}}})=\pm 1$, for all $x_{i}\in\mathcal{B}_{1}$.
+But, by assumption and Proposition 4.11, $|{\widehat{P}}({\frac{x_{i}}{d_{i}}})|=1$ if and only if $x_{i}\in G{(H,\mathcal{B})}$.
+Thus, ${\widehat{P}}^{2}$ idempotent implies that ${\mathcal{B}_{1}}\subset G{(H,\mathcal{B})}$.
+But $G{(H,\mathcal{B})}\subset\mathcal{B}_{1}$, so the equality holds, i.e.
+${(H,\mathcal{B})}$ is Burnside.
+Conversely, if ${(H,\mathcal{B})}$ is Burnside, then ${\widehat{P}}^{2}=\sum_{x_{i}\in G{(H,\mathcal{B})}}\widehat{F}_{i}$, by Corollary 4.5, which is an idempotent. ∎
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and $P:=\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$. Then
+(4.10) $P=\sum_{j\in\mathcal{I}}\mu_{j}(P)F_{j}.$
+Here are the dual version of the last three results:
 Proposition 4.14.
 Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30.
-Then $|\mu_{j}(P)|\leq 1$ and one has equality if and only if $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$.
+Then $|\mu_{j}(P)|\leq 1$.
+Moreover, the equality holds if and only if $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$.
 Proposition 4.15.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup.
-One has that $P^{2}$ is an idempotent if and only if the non-vanishing elements $\mu_{j}\in{\widehat{\mathcal{B}}}$ satisfy $\mu_{j}(P)=\pm 1$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then, $P^{2}$ is an idempotent if and only if $\mu_{j}(P)=\pm 1$, for all $\mu_{j}\in\widehat{\mathcal{B}}_{1}$.
 Corollary 4.16.
 Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying the hypothesis of Lemma 2.30.
-One has that $P^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside.
+Then, $P^{2}$ is an idempotent if and only if ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside.
 4.2. The analogue of Burnside’s theorem
-We prove the following more general version of Theorem 1.2. Its proof follows the same lines as the proof of [Bur23, Theorem 2].
+We will prove the following generalization of Theorem 1.2, using the same approach as in [Bur23, Theorem 2].
 Theorem 4.17.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral.
-Let $x_{i}\in\mathcal{B}$ be an element such that $h_{i}|d_{i}|^{2}>1$ and
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup with ${h}$-integral dual. Let $x_{i}\in\mathcal{B}$. If $h_{i}|d_{i}|^{2}>1$, and
 (4.11) $h_{i}|\mu_{j}(x_{i})|^{2}\in\mathbb{A}$
-for all $j\in\mathcal{J}$. Then there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
+for all $j\in\mathcal{I}$. Then $x_{i}\in\mathcal{B}_{0}$
 Proof.
-For any $1\leq i\leq m$, denote $\mathcal{T}_{i}:=\mathcal{T}(x_{i})=\{j\in\mathcal{J}\;|\;\mu_{j}(x_{i})=0\}$ and ${\mathcal{D}}_{i}:=\mathcal{J}\setminus(\mathcal{T}_{i}\cup\{1\})$.
-One needs to show that $\mathcal{T}_{i}\neq\emptyset$, for any basis element $x_{i}\in\mathcal{B}$ satisfying the hypothesis from the theorem.
-Note that the second orthogonality relation can be written as:
-(4.12) $\sum_{j=1}^{m}\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}$
-Since $\widehat{h}_{1}=1$, for $j=1$, the above equation can be rewritten as:
+For any $i\in\mathcal{I}$, denote $\mathcal{T}_{i}:=\{j\in\mathcal{I}\;|\;\mu_{j}(x_{i})=0\}$ and ${\mathcal{D}}_{i}:=\mathcal{I}\setminus(\mathcal{T}_{i}\cup\{1\})$.
+We need to show that $\mathcal{T}_{i}\neq\emptyset$, for any $x_{i}\in\mathcal{B}$ satisfying the theorem’s hypothesis.
+The second orthogonality relation (2.6) and Equation (2.11) implies
+(4.12) $\sum_{j\in\mathcal{I}}\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}.$
+Since $\widehat{h}_{1}=1$, the above equation can be written as:
 $\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}}=|d_{i}|^{2}+\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}}|\mu_{j}(x_{i})|^{2}$
@@
 (4.13) $1=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}.$
-On the other hand, note that
-$n{(H,\mathcal{B},\mu_{1})}=\sum_{j=1}^{m}\widehat{h}_{j}=1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}+\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}.$
-Therefore Equation (4.13) can be written as:
+On the other hand,
+$n{(H,\mathcal{B},\mu_{1})}=\sum_{j\in\mathcal{I}}\widehat{h}_{j}=1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}+\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}.$
+Therefore, Equation (4.13) can be written as
 (4.14) $1=\frac{1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}-\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}.$
@@
 (4.15) $\big{(}\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}|\mu_{j}(x_{i})|^{2}}{|d_{i}|^{2}}-\sum_{j\in{\mathcal{D}}_{i}}\frac{\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\big{)}\geq 0,$
-since then, it follows that $\frac{1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\geq 1$, i.e.
-$1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}\geq h_{i}|d_{i}|^{2}$.
-Since $h_{i}|d_{i}|^{2}>1$, it follows that $\mathcal{T}_{i}\neq\emptyset$.
-The inequality from Equation (4.15) can be written as
-(4.16) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq 1.$
+since then, $\frac{1+\sum_{j\in\mathcal{T}_{i}}\widehat{h}_{j}}{h_{i}|d_{i}|^{2}}\geq 1$, i.e.
+$1+\sum\limits_{j\in\mathcal{T}_{i}}\widehat{h}_{j}\geq h_{i}|d_{i}|^{2}$; and $h_{i}|d_{i}|^{2}>1$ implies $\mathcal{T}_{i}\neq\emptyset$.
+The inequality (4.15) can be written as
+(4.16) $(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq 1.$
 On the other hand, the weighted AM-GM inequality gives that
-(4.17) $\frac{1}{\sum\limits_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq\prod_{j\in{\mathcal{D}}_{i}}\bigg{(}h_{i}\big{(}|\mu_{j}(x_{i})|^{2}\big{)}^{\widehat{h}_{j}}\bigg{)}^{\frac{1}{|{\mathcal{D}}_{i}|}},$
-where $|{\mathcal{D}}_{i}|:=\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j}$.
-Note that Equation (3.1) implies that the set ${\mathcal{D}}_{i}$ is stable under the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$.
-This in turn implies that the product
+(4.17) $(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}(\sum_{j\in{\mathcal{D}}_{i}}{\widehat{h}_{j}h_{i}|\mu_{j}(x_{i})|^{2}})\geq\bigg{(}\prod_{j\in{\mathcal{D}}_{i}}\big{(}h_{i}|\mu_{j}(x_{i})|^{2}\big{)}^{\widehat{h}_{j}}\bigg{)}^{(\sum_{j\in{\mathcal{D}}_{i}}\widehat{h}_{j})^{-1}}.$
+By rational assumption, $h_{i}$ is rational and $\mu_{j}(x_{i})$ is an algebraic number.
+Equation (3.1) implies that the set ${\mathcal{D}}_{i}$ is stable under the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$.
+It follows that the product
 $P_{i}:=\prod_{j\in{\mathcal{D}}_{i}}(h_{i}|\mu_{j}(x_{i})|^{2})^{\widehat{h}_{j}}$
-is fixed by the Galois group $\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, since ${\widehat{h}_{\tau(j)}}=\widehat{h}_{j}$ by Lemma 3.8 .
-It follows that $P_{i}$ is a rational number.
+is fixed by the action of $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, since ${\widehat{h}_{\tau(j)}}=\widehat{h}_{j}$ by Lemma 3.8.
+Thus $P_{i}$ is a rational number.
 On the other hand, each factor of $P_{i}$ is an algebraic integer (since $\widehat{h}_{j}\in\mathbb{Z}_{{}_{>0}}$), and therefore, the entire product is an integer.
-Since it is also positive, it must be at least $1$.
+Since it is also positive, it must be at least 111.
 ∎
-Note that Theorem 1.2 follows from the above theorem, since in the case of a fusion ring, $h_{i}=1$ and $\mu_{j}(x_{i})\in\mathbb{A}$, for all $(i,j)\in{\mathcal{I}}\times{\mathcal{J}}$.
+Note that Theorem 1.2 follows from the above theorem, since in the case of a fusion ring, $h_{i}=1$, and $\mu_{j}(x_{i})\in\mathbb{A}$, for all $i,j\in{\mathcal{I}}$.
 Proposition 4.18.
-Let ${(H,\mathcal{B})}$ be a rational ARN-hypergroup such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral.
-Let $x_{i}\in\mathcal{B}$ be an invertible element in $H$ satisfying also Equation (4.11), for all $1\leq j\leq m$.
-Then $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element.
+Let ${(H,\mathcal{B})}$ be a rational ARN-hypergroup with $h$-integral dual.
+Let $x_{i}\in\mathcal{B}$ be an invertible element in $H$ satisfying also Equation (4.11), for all $j\in\mathcal{I}$.
+Then $x_{i}\in{G{(H,\mathcal{B})}}$.
 Proof.
-Note that ${(H,\mathcal{B})}$ satisfies the hypothesis of Theorem 4.17.
-Also, by hypothesis, Equation (4.11) holds for all $1\leq j\leq m$.
 Suppose that $x_{i}$ is not a grouplike element, i.e.
 $h_{i}|d_{i}|^{2}>1$, by Lemma 2.24.
-In this case, by Theorem 4.17, there is $\mu_{j}\in\widehat{B}$ such that $\mu_{j}(x_{i})=0$.
-Thus, if $x_{i}y=1$ then $10=\mu_{j}(x_{i})\mu_{j}(y)=$ 1, which is a contradiction.
+Then, all the assumptions of Theorem 4.17 are satisfied, so that, $x_{i}\in\mathcal{B}_{0}$, contradiction with $x_{i}$ invertible.∎
+Corollary 4.19.
+Any rational ARN-hypergroup ${(H,\mathcal{B})}$ with $h$-integral dual, such that Equation (4.11) holds for all $i,j\in\mathcal{I}$, is Burnside.
+Proof.
+By Proposition 4.3, it is enough to show that ${\mathcal{B}_{1}}\subseteq{G{(H,\mathcal{B})}}$.
+Let $x_{i}\in\mathcal{B}_{1}$, thus $x_{i}$ is invertible, so by Proposition 4.18, $x_{i}\in{G{(H,\mathcal{B})}}$.
 ∎
 Corollary 4.20.
-Any abelian rational normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is ${h}$-integral and Equation (4.11) is satisfied for all $1\leq i,j\leq m$ is Burnside.
-Proof.
-By Proposition 4.3, it is enough to show that $\mathcal{B}_{1}={G{(H,\mathcal{B})}}$.
-Clearly, ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$.
-Suppose that $x_{i}\in\mathcal{B}$ is an element which is invertible in $H$.
-Since $x_{i}$ is not a vanishing element, it follows from Proposition 4.18 that $x_{i}\in{G{(H,\mathcal{B})}}$.
-∎
-Corollary 4.20.
-Let ${(H,\mathcal{B})}$ be a commutative fusion ring such that $\widehat{(H,\mathcal{B})}$ is ${h}$-integral. Then ${(H,\mathcal{B})}$ is Burnside.
+Any commutative fusion ring with $h$-integral dual is Burnside.
 The following corollary recovers [Bur23, Theorem 2].
 Corollary 4.21.
-Any commutative Grothendieck ring of any weakly-integral fusion category is Burnside.
+A weakly-integral fusion category with a commutative Grothendieck ring is Burnside.
 Proof.
-It follows from above corollary by Remark 3.7. ∎
-Now one can write the corresponding statement of Theorem 4.17 applied to the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ of an abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$.
-We obtain the following:
+It follows from Corollary 4.20 by Remark 3.7. ∎
+We can write a kind of dual version of Theorem 4.17:
 Theorem 4.22.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable ${h}$-integral hypergroup such that the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is a rational hypergroup.
-Let $\mu_{j}\in{\widehat{\mathcal{B}}}$ be an element of ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ with $\widehat{h}_{j}>1$, and such that
-(4.18) $\frac{\widehat{h}_{j}}{d_{i}^{2}}|\mu_{j}(x_{i})|^{2}\in\mathbb{A},$
-for all $i\in\mathcal{I}$. Then, there is $x_{i}\in\mathcal{B}$ such that $\mu_{j}(x_{i})=0$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable $h$-integral hypergroup with a rational dual. Let $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\widehat{h}_{j}>1$ and
+(4.18) $\widehat{h}_{j}\left|\mu_{j}(\frac{x_{i}}{d_{i}})\right|^{2}\in\mathbb{A},$
+for all $i\in\mathcal{I}$. Then, $\mathcal{B}_{0}$ is non-empty.
 Proof.
-Note that the condition from Equation (4.11) applied on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ becomes exactly Equation (4.18). ∎
+The condition from Equation (4.11) applied on $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ becomes exactly Equation (4.18).
+So we can apply Theorem 4.17 to $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$.
+Thus $\mu_{j}\in\widehat{\mathcal{B}}_{0}$, meaning the existence of $x_{i}\in\mathcal{B}$ such that $\mu_{j}(\frac{x_{i}}{d_{i}})=0$, so $x_{i}\in\mathcal{B}_{0}$.
+∎
 5. Kernels of fusion rings, Brauer’s theorem
@@
 A standard element $b\in\mathcal{B}$ is called a constituent of $x\in{(H,\mathcal{B})}_{+}$ if $x_{b}>0$.
-Let $\psi:H\rightarrow\mathbb{C}$ be a linear character, and $E=E_{\psi}\in H$ be its primitive central idempotent.
-Then, since $E$ is an eigenvector for any $L_{x}$ with $x\in H_{+}$, it follows that $|\psi(x)|\leq\mathrm{FPdim}(x)$, for any object $x\in{(H,\mathcal{B})}_{+}$.
+Recall that $^{\widehat{\mathcal{B}}}$ is the set of all algebra morphisms $\mu_{j}:H\rightarrow\mathbb{C}$.
+Let $\psi\in{\widehat{\mathcal{B}}}$.
+For all $x\in{(H,\mathcal{B})}_{+}$,
+$|\psi(x)|=|\psi(\sum_{b\in{\mathcal{B}}}x_{b}b)|\leq\sum_{b\in{\mathcal{B}}}x_{b}|\psi(b)|\leq\sum_{b\in{\mathcal{B}}}x_{b}\mathrm{FPdim}(b)=\mathrm{FPdim}(x),$
+by Frobenius-Perron theory.
 Definition 5.1.
-Let $\mathcal{S}\subseteq\mathcal{B}$ be a subset closed under the involution of ${(H,\mathcal{B})}$, and such that $L:=\mathbb{C}[\mathcal{S}]$ is a subalgebra of $H$.
+Let $\mathcal{S}\subseteq\mathcal{B}$ be a subset such that
+  • $\mathcal{S}$ is closed under the involution of ${(H,\mathcal{B})}$,
+  • $L:=\mathbb{C}[\mathcal{S}]$ is a subalgebra of $H$.
 Then ${(L,\mathcal{S})}$ is called a sub-hypergroup of ${(H,\mathcal{B})}$.
-We mention that the notion of kernels of objects of a fusion category from [Bur16] can be extended to arbitrary ARN-hypergroups.
+The notion of kernel of object of a fusion category from [Bur16] can be extended to arbitrary ARN-hypergroups.
 Definition 5.2.
@@
 $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi):=\{x\in\mathcal{B}\;|\;\psi(x)=\mathrm{FPdim}(x)\}.$
-Recall that $^{\widehat{\mathcal{B}}}$ is the set of all algebra morphisms $\mu_{j}:H\rightarrow\mathbb{C}$.
 Lemma 5.3.
@@
 Using the triangle inequality for the complex absolute-value,
-$\mathrm{FPdim}(x)\mathrm{FPdim}(y)$ $=$ $\mid\psi(x)\psi(y)\mid=\mid\psi(xy)\mid=\mid\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z)\mid$
-$\leq$ $\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mid\psi(z)\mid\leq\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mathrm{FPdim}(z)$
-$=$ $\mathrm{FPdim}(x)\mathrm{FPdim}(y),$
-It is easy to see that $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of $H$. ∎
-Dually, one can define the following:
+$\mathrm{FPdim}(x)\mathrm{FPdim}(y)$ $=$ $\left|\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\psi(z)\right|$
+$\leq$ $\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}|\psi(z)|\leq\sum_{z\in\mathcal{B}^{\prime}}N^{z}_{xy}\mathrm{FPdim}(z)$
+$=$ $\mathrm{FPdim}(x)\mathrm{FPdim}(y).$
+It is easy to deduce that $\ker_{{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of $H$. ∎
+Dually, we can define the following:
 Definition 5.4.
-Let ${(H,\mathcal{B})}$ be an ARN-hypergroup.
-For any $x\in H_{+}$, one can define ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ as the set of all standard elements $\psi\in{\widehat{\mathcal{B}}}$ such that $\psi(x)=\mathrm{FPdim}(x)$.
-The dual version of Lemma 5.3 implies that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ is (the basis of) a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$ if ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Let $x\in H_{+}$. Define
+${\ker_{{}_{{(H,\mathcal{B})}}}}(x):=\{\psi\in{\widehat{\mathcal{B}}}\;|\;\psi(x)=\mathrm{FPdim}(x)\}.$
+The dual version of Lemma 5.3 implies that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ is (the basis of) a sub-hypergroup of $^{\widehat{(H,\mathcal{B})}}$ if $^{\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
 5.1. Brauer’s theorem for ARN-hypergroup
-For any hypergroup ${(H,\mathcal{B})}$, we define the bilinear function on $H$ such that $m(\sum_{i=1}^{m}\alpha_{i}x_{i},\sum_{i=1}^{m}\beta_{i}x_{i})=\sum_{i=1}^{m}\frac{\alpha_{i}\beta_{i}}{h_{i}}$.
-In the case of an abelian hypergroup note that Equation (2.2) implies the following:
-(5.1) $m(x,y)=\tau(xy^{*})=\sum_{j=1}^{m}\frac{1}{n_{j}}\mu_{j}(x)\mu_{j}(y^{*})\;\;\text{for all}\;\;x,y\in H.$
+For any hypergroup ${(H,\mathcal{B})}$, we define the bilinear function $m$ on $H$ by
+$m(\sum_{i\in\mathcal{I}}\alpha_{i}x_{i},\sum_{i\in\mathcal{I}}\beta_{i}x_{i})=\sum_{i\in\mathcal{I}}\frac{\alpha_{i}\beta_{i}}{h_{i}}.$
+In the abelian case, Equation (2.2) implies
+(5.1) $m(x,y)=\tau(xy^{*})=\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x)\mu_{j}(y^{*}),\;\;\text{for all}\;\;x,y\in H.$
 Lemma 5.5.
 Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. If $x:=\sum_{s\in\mathcal{B}}p_{s}s\in{(H,\mathcal{B})}_{+}$ then
-${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s).$
+${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\bigcap_{{}_{\{s\in\mathcal{B}|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s).$
 Proof.
-Note that $\mathrm{FPdim}(x)=\sum_{\{s\;|\;p_{s}>0\}}p_{s}\mathrm{FPdim}(s)$. If $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
-$\mathrm{FPdim}(x)=|\mu_{j}(x)|\leq\sum_{\{s|\;p_{s}>0\}}|\mu_{j}(s)|\leq\sum_{s}p_{s}\mathrm{FPdim}(s)=\mathrm{FPdim}(x),$
-and this implies that $\mu_{j}\in\bigcap_{{}_{\{s|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s)$. The converse is obvious. ∎
-Next theorem is a generalization of Brauer’s theorem. Although the proof is the same as the one from [Bur16], we include it here for the completeness of the paper.
-For an element $x\in{(H,\mathcal{B})}_{+}$, we denote by $<x>$ the sub-hypergroup generated by $x$.
-By definition, it is the sub-hypergroup generated by the constituents of all powers $x^{n}$ with $n\geq 1$.
+Note that $\mathrm{FPdim}(x)=\sum_{\{s\in\mathcal{B}\;|\;p_{s}>0\}}p_{s}\mathrm{FPdim}(s)$. Let $\mu_{j}\in\widehat{B}$ then
+$|\mu_{j}(x)|\leq\sum_{\{s\in\mathcal{B}|\;p_{s}>0\}}p_{s}|\mu_{j}(s)|\leq\sum_{s}p_{s}\mathrm{FPdim}(s)=\mathrm{FPdim}(x).$
+Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\mu_{j}\in\bigcap_{{}_{\{s\in\mathcal{B}|\;p_{s}>0\}}}{\ker_{{}_{{(H,\mathcal{B})}}}}(s)$. ∎
+The following theorem is a generalization of Brauer’s theorem. Although the proof is the same as that in [Bur16], we include it here for the sake of completeness.
+For $x\in{(H,\mathcal{B})}_{+}$, we denote by $\langle x\rangle$ the sub-hypergroup generated by $x$, i.e. whose basis is the set of constituents of all powers $x^{n}$, with $n\geq 1$.
 Theorem 5.6.
-Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $x\in{(H,\mathcal{B})}_{+}$. Then $<x>=H$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\{\mu_{1}\}$.
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup, and $x\in{(H,\mathcal{B})}_{+}$. Then $\langle x\rangle=H$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\{\mu_{1}\}$.
 Proof.
-Suppose that ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)=\mu_{1}$. Write $x=\sum_{j=1}^{m}\mu_{j}(x)F_{j}$ and then
-$x^{n}=\sum_{j=1}^{m}\mu_{j}(x)^{n}F_{j}$
-Partition the set $\mathcal{J}=\mathcal{J}_{1}\sqcup\mathcal{J}_{2}\sqcup\cdots\sqcup\mathcal{J}_{s}$ such that the value $\mu_{j}(x)$ is constant on each component $\mathcal{J}_{l}$ of the partition.
-Denote by $\alpha_{l}:=\mu_{j}(x)$ this constant value of $\mu_{j}(x)$ with $j\in\mathcal{J}_{l}$.
-Note that by the hypothesis, the partition component containing $\{1\}$ is a singleton.
-Also note that for any $s\in\mathcal{B}$ and any $a\in H$, then $m(s,a)=\tau(s^{*}a)$. Thus
-${m(s,x^{n})}{}=\tau(s^{*}x^{n})=\sum_{j=1}^{m}\mu_{j}(s^{*})\mu_{j}(x)^{n}\tau(F_{j})=\sum_{l\in\mathcal{P}}\big{(}\sum_{j\in\mathcal{J}_{l}}\tau(F_{j})\mu_{j}(s^{*})\big{)}\alpha_{l}^{n}$
-Denote
-$a_{l}:=\sum_{j\in\mathcal{A}_{l}}\tau(F_{j})\mu_{j}(s^{*}).$
-Note that $a_{1}=n{(H,\mathcal{B})}>0$, by Equation (2.9). If $m(s,x^{n})=0$ for all $n$, then
-$\sum_{l\in\mathcal{P}}a_{l}\alpha_{l}^{n}=0$
-for all $n\in\mathbb{N}$, and Vandermonde determinant gives $a_{l}=1$. This is a contradiction since $a_{1}\neq 0$.
-For the converse, suppose that $<x>=H$. If $\mu_{j}\in\ker_{(H,\mathcal{B})}(x)$, then by Lemma 5.3, $\mu_{j}\in\ker(b)$, for any $b\in\mathcal{B}$. Thus $\mu_{j}=\mu_{1}$. ∎
-5.2. The cointegral $\lambda_{(L,S)}$ in terms of kernels
-Recall from §1 that for any RN-hypergroup ${(H,\mathcal{B})}$, we denoted by $\lambda_{H}$ the primitive central idempotent corresponding to $\mu_{1}=\mathrm{FPdim}$.
+Partition the set $\mathcal{I}=\bigsqcup_{l\in\mathcal{P}}\mathcal{I}_{l}$ such that the values $\mu_{j}(x)$ are constant on each component $\mathcal{I}_{l}$ of the partition, but distinct otherwise (i.e.
+for all $l,l^{\prime}\in\mathcal{P}$, $j\in\mathcal{I}_{l}$, $j^{\prime}\in\mathcal{I}_{l^{\prime}}$, $\mu_{j}(x)=\mu_{j^{\prime}}(x)$ if and only if $l=l^{\prime}$).
+Let $s\in\mathcal{B}$.
+By Equation (5.1),
+${m(x^{n},s)}{}=\sum_{j\in\mathcal{I}}\frac{1}{n_{j}}\mu_{j}(x)^{n}\mu_{j}(s^{*})=\sum_{l\in\mathcal{P}}\left(\sum_{j\in\mathcal{I}_{l}}\frac{1}{n_{j}}\mu_{j}(s^{*})\right)\alpha_{l}^{n}=\sum_{l\in\mathcal{P}}a_{l}\alpha_{l}^{n}$
+with
+$a_{l}:=\sum_{j\in\mathcal{I}_{l}}\frac{1}{n_{j}}\mu_{j}(s^{*}),\text{ and }\alpha_{l}:=\mu_{j}(x)\text{ when }j\in\mathcal{I}_{l}.$
+Note that $s\not\in\langle x\rangle$ if and only if $m(x^{n},s)=0$ for all $n$, if and only if $Mv=0$, with $M$ be the matrix $(\alpha_{l}^{n})$ and $v$ be the vector $(a_{l})$.
+But $M$ is invertible because its determinant is nonzero, because all $\alpha_{l}$ are distinct (by Vandermonde determinant formula).
+Thus, $s\in\langle x\rangle$ if and only if there is $l\in\mathcal{P}$ such that $a_{l}$ is nonzero.
+[…]
+Finally, $\mu_{j}\in\ker_{(H,\mathcal{B})}(x)$ if and only if $x\in\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})$, if and only if $\langle x\rangle\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})$ by Lemma 5.3.
+So if $\langle x\rangle=H$ then $\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})=H$, i.e $\mu_{j}=\mu_{1}$.
+∎
+5.2. The cointegral $\lambda_{(L,\mathcal{S})}$ in terms of kernels
+Recall from §1 that for any ARN-hypergroup ${(H,\mathcal{B})}$, we denoted by $\lambda_{H}$ the primitive central idempotent $F_{1}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
+Let ${(L,\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$, and let $\lambda_{(L,\mathcal{S})}$ be the idempotent in $L\subset H$ corresponding to $\mu_{1}\big{|}_{(L,\mathcal{S})}$.
+By Equation (2.4), we get that:
+(5.2) $\lambda_{(L,\mathcal{S})}=\frac{1}{n{(L,\mathcal{S})}}\sum_{x_{i}\in\mathcal{S}}h_{i}d_{i}x_{i}.$
+In particular, $\lambda_{(L,\mathcal{S})}=\lambda_{(K,\mathcal{T})}$ if and only if ${(L,\mathcal{S})}={(K,\mathcal{T})}$.
 Theorem 5.7.
-For any $x\in{(H,\mathcal{B})}_{+}$, define $J_{x}:={\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ then
-$\lambda_{<x>}=\sum_{j\in J_{x}}F_{j}$
+For any $x\in{(H,\mathcal{B})}_{+}$, let $\mathcal{I}_{x}$ be the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$. Then
+$\lambda_{\langle x\rangle}=\sum_{j\in\mathcal{I}_{x}}F_{j}$
 Proof.
-Suppose that $<x>={(L,\;\mathcal{S})}$.
-Let $G_{1},\dots,G_{s}$ be the primitive idempotents of ${(L,\;\mathcal{S})}$, and $\psi_{1},\dots,\psi_{s}$ be their associated characters.
-Moreover, as above one may suppose that $G_{1}=\lambda_{{(L,\;\mathcal{S})}}$, and consequently, $\psi_{1}=\mathrm{FPdim}$.
+Suppose that $\langle x\rangle={(L,\mathcal{S})}$.
+Let $G_{1},\dots,G_{s}$ be the primitive central idempotents of ${(L,\mathcal{S})}$, and $\psi_{1},\dots,\psi_{s}$ be their associated characters.
+As above, let us take $G_{1}=\lambda_{{(L,\mathcal{S})}}$, and consequently, $\psi_{1}=\mathrm{FPdim}$.
 We look at the restrictions of the characters $\mu_{j}$ at ${(L,\mathcal{S})}$.
-Then, there is a surjective function $f:\{1,\dots,r\}\rightarrow\{1,\dots,s\}$ such that $\mu_{j}\big{|}_{{(L,\;\mathcal{S})}}=\psi_{f(j)}$, for all $1\leq j\leq s$.
-With the above notation, it will be shown that $f^{-1}(0)=\mathcal{J}_{x}$ Indeed, since $x$ is a faithful element of ${(L,\;\mathcal{S})}$, it follows that $\ker_{{}_{(L,\;\mathcal{S})}}(x)=\{\psi_{1}\}$.
-On the other hand, note that $\mu_{j}(x)=\mu_{j}\big{|}_{{}_{(L,\;\mathcal{S})}}(x)=\psi_{f(j)}(x)$.
-Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\psi_{f(j)}\in\ker_{{}_{(L,\;\mathcal{S})}}(x)$, i.e.
-$f(j)=0$.
-∎
+Then, there is a surjective function $f:\mathcal{I}\rightarrow\{1,\dots,s\}$ such that $\mu_{j}\big{|}_{{(L,\mathcal{S})}}=\psi_{f(j)}$, for all $j\in\mathcal{I}$.
+With the above notation, we are reduced to show that $f^{-1}(\{1\})=\mathcal{I}_{x}$ By Theorem 5.6, since $\langle x\rangle={(L,\mathcal{S})}$, then $\ker_{{}_{(L,\mathcal{S})}}(x)=\{\psi_{1}\}$.
+On the other hand,
+$\mu_{j}(x)=\mu_{j}\big{|}_{{}_{(L,\mathcal{S})}}(x)=\psi_{f(j)}(x).$
+Thus, $\mu_{j}\in{\ker_{{}_{{(H,\mathcal{B})}}}}(x)$ if and only if $\psi_{f(j)}\in\ker_{{}_{(L,\mathcal{S})}}(x)$, i.e. $f(j)=1$. ∎
 6. The adjoint sub-hypergroup and grouplike elements
@@
 $Z_{{}_{(H,\mathcal{B})}}(x_{i})=\{\psi\in{\widehat{\mathcal{B}}}\;|\;|\psi(x_{i})|=d_{i}\}.$
-If ${\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup, then as in Lemma 5.3, it is easy to see that $Z_{{}_{(H,\mathcal{B})}}(x_{i})$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$.
 Define also the dual notion of the center,
 $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)=\{x_{i}\in{\mathcal{B}}\;|\;|\psi(x_{i})|=d_{i}\}.$
-Again, as in Lemma 5.3, it is easy to see that $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of ${(H,\mathcal{B})}$ if ${(H,\mathcal{B})}$ is an ARN-hypergroup..
+As in Lemma 5.3, it is easy to see that $Z_{{}_{\widehat{(H,\mathcal{B})}}}(\psi)$ is a sub-hypergroup of ${(H,\mathcal{B})}$.
+Idem for $Z_{{}_{(H,\mathcal{B})}}(x_{i})$, if $^{\widehat{(H,\mathcal{B})}}$ is a RN-hypergroup.
 6.2. Adjoint sub-hypergroup
 Proposition 6.1.
-Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}\in H$. Then
-${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and $I(1):=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}\in H$. Then
+${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\bigcap_{i\in\mathcal{I}}Z_{{}_{(H,\mathcal{B})}}(x_{i})$
 Proof.
 Note that
-$|\psi(I(1))|\leq\sum_{i=1}^{m}|h_{i}\psi(x_{i}x_{i^{*}})|=\sum_{i=1}^{m}h_{i}|\psi(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}d_{i}^{2}=\mathrm{FPdim}{(H,\mathcal{B})}.$
-So if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi\in\bigcap_{i=1}^{m}Z_{{}_{(H,\mathcal{B})}}(x_{i})$. The converse is immediate. ∎
+$\psi(I(1))=\sum_{i\in\mathcal{I}}h_{i}\psi(x_{i}x_{i^{*}})=\sum_{i\in\mathcal{I}}h_{i}|\psi(x_{i})|^{2}\leq\sum_{i\in\mathcal{I}}h_{i}d_{i}^{2}=\mathrm{FPdim}{(H,\mathcal{B})}.$
+So if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi\in\bigcap_{i\in\mathcal{I}}Z_{{}_{(H,\mathcal{B})}}(x_{i})$. The converse is immediate. ∎
 Definition 6.2.
-Denote by ${{(H,\mathcal{B})}_{ad}}$ the sub-hypergroup $<I(1)>$ of $H$ generated by $I(1)$. It is called the adjoint sub-hypergroup of $H$.
-6.3. On the support $J_{{(L,\;\mathcal{S})}}$
+Denote by ${{(H,\mathcal{B})}_{ad}}$ the sub-hypergroup $\langle I(1)\rangle$ of $H$ generated by $I(1)$. It is called the adjoint sub-hypergroup of $H$.
+6.3. On the support ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$
 Definition 6.3.
-Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian hypergroup ${(H,\mathcal{B})}$. Its support is the subset $J_{{(L,\;\mathcal{S})}}$ such that
-(6.1) ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}=\sum\limits_{j\in{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}}F_{j}.$
-The next result is a generalization of [Bur20, Lemma 6.4].
+Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an abelian hypergroup ${(H,\mathcal{B})}$. Its support is the subset ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}\subset\mathcal{I}$ such that
+(6.1) ${\lambda_{{}_{{(L,\mathcal{S})}}}}=\sum_{j\in{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}}F_{j}.$
+In particular, by Equation (5.2), ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}={\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}$ if and only if ${(L,\mathcal{S})}={(K,\mathcal{T})}$.
+The next result generalizes [Bur20, Lemma 6.4].
 Proposition 6.4.
 Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. With the above notations,
-${J_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}\}.$
+${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{I}\;|\;n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}\}.$
 Proof.
-By Theorem 5.7, it is enough to show the following:
-${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))=\{j\in\mathcal{J}\;|\;n_{j}=\mathrm{FPdim}({(H,\mathcal{B})})\}.$
-Equation (2.5) gives that $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}$, and therefore, $n_{j}=\mathrm{FPdim}{(H,\mathcal{B})}$ if and only if $|\mu_{j}(x_{i})|^{2}=d_{i}^{2}$, for all $i\in\mathcal{I}$.
-This is equivalent to $\mu_{j}\in Z_{{}_{(H,\mathcal{B})}}(x_{i})$, for all $i\in\mathcal{I}$.
-The result follows by Proposition 6.1.
-∎
+By Theorem 5.7 and Definition 6.2, ${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}={\mathcal{I}}_{I(1)}$, the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$. Thus,
+${{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}=\{j\in\mathcal{I}\;|\;\mu_{j}(I(1))=\mathrm{FPdim}(I(1))\}.$
+But $\mathrm{FPdim}(I(1))=\mathrm{FPdim}{(H,\mathcal{B})}$, and $\mu_{j}(I(1))=n_{j}$ by Equation (2.5). ∎
 Theorem 6.5.
-Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only $j\in J_{{{(H,\mathcal{B})}_{ad}}}$.
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only $j\in{{\mathcal{I}}_{{{(H,\mathcal{B})}_{ad}}}}$.
 Proof.
-It follows from Proposition 6.4 and Lemma 2.24.
+It follows from Proposition 6.4 and Lemma 2.31. ∎
 Corollary 6.6.
 Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then
-(6.2) $\lambda_{{}_{{{(H,\mathcal{B})}_{ad}}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}\widehat{F}_{j}$
+(6.2) $\lambda_{{}_{{{(H,\mathcal{B})}_{ad}}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}$
 Proof.
 It follows from Equation (6.1) and Theorem 6.5. ∎
-There is a dual version of Theorem 6.5 and its Corollary 6.6:
+Here are dual versions of Theorem 6.5 and Corollary 6.6:
+Theorem 6.7.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
+Then $x_{i}\in{G{(H,\mathcal{B})}}$ if and only $i\in\mathcal{I}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}$.
+Corollary 6.8.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
+(6.3) $\lambda_{{}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
+6.4. Proof of Theorems 1.7 and 1.10
+Recall Definition 6.3, which defines the support $\mathcal{I}_{{(L,\mathcal{S})}}\subset\mathcal{I}$ of a sub-hypergroup ${(L,\mathcal{S})}$ within an abelian hypergroup ${(H,\mathcal{B})}$.
+For the sake of brevity, we define $|L|:=n(L,\mathcal{S},\mu_{1})$.
+Theorem 6.9.
+Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$. With the above notations, we have:
+(6.4) $\sum\limits_{x_{i}\in\mathcal{S}}\widehat{F}_{i}=\frac{|L|}{|H|}\big{(}\sum\limits_{j\in\mathcal{I}_{{(L,\mathcal{S})}}}{\widehat{h}}_{j}\mu_{j}\big{)}.$
+Proof.
+Applying $\mathcal{F}$ to Equation (6.1) we have that
+[…]
+On the other hand, from Equation (2.13) we have ${\lambda_{{}_{{(L,\mathcal{S})}}}}=\frac{1}{|L|}\big{(}\sum\limits_{x_{i}\in\mathcal{S}}h_{i^{*}}d_{i^{*}}x_{i}\big{)}$ and therefore:
+[…]
+Comparing (6.5) and (6.6), we obtain:
+$\sum\limits_{x_{i}\in\mathcal{S}}\widehat{F}_{i}=|L|\big{(}\sum\limits_{j\in\mathcal{I}_{(L,\mathcal{S})}}\frac{\mu_{j}}{n_{j}}\big{)}.$
+According to Equation (2.11), we have ${\widehat{h}}_{j}=\frac{|H|}{n_{j}}$, thus concluding the proof. ∎
+Proof of Theorem 1.7: By applying Corollary 4.5, we get that
+$\big{(}\prod_{j\in\mathcal{I}}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
+On the other hand, by utilizing Theorem 6.9 with $\mathcal{S}=G{(H,\mathcal{B})}$, we arrive at the desired conclusion.
+Proof of Theorem 1.10: To derive the result through duality from Theorem 1.7, we need to consider the normalized basis $(\frac{x_{i}}{d_{i}})_{i\in\mathcal{I}}$.
+Recall that by definition, $\langle x_{i}x_{i^{*}},x_{1}\rangle^{-1}=h_{i}$.
+Therefore, $\langle\frac{x_{i}}{d_{i}}\frac{x_{i^{*}}}{d_{i^{*}}},x_{1}\rangle^{-1}=d_{i}d_{i^{*}}h_{i}$.
+Consequently, $d_{i}d_{i^{*}}h_{i}\frac{x_{i}}{d_{i}}=d_{i^{*}}h_{i}x_{i}$, which precisely matches the summand on the RHS.
+6.5. Proof of Theorems 1.8 and 1.11.
+Proof.
+By Corollary 4.5, ${(H,\mathcal{B})}$ is Burnside if and only if
+$\big{(}\prod_{j\in\mathcal{I}}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in G(H,B)}\widehat{F}_{i},$
+if and only if Equation (1.4) holds, by Corollary 6.8, (the dual version of) Equation (2.12) and Lemma 2.14.∎
+By duality, we obtain directly Theorem 1.11.
+6.6. On the sub-hypergroup generated by $P$
+Lemma 6.10.
+For any sub-hypergroup ${(L,\mathcal{S})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${\mathcal{I}}_{{(L,\mathcal{S})}}=\bigcap_{s\in\mathcal{S}}{\mathcal{I}}_{s}$.
+Thus, $s\in\mathcal{S}$ if and only if $s\lambda_{{(L,\mathcal{S})}}=d_{s}\lambda_{{(L,\mathcal{S})}}$.
+Proof.
+Note that ${(L,\mathcal{S})}=\langle x\rangle$, with $x=\sum_{s\in\mathcal{S}}s\in{(H,\mathcal{B})}_{+}$.
+Recall that ${\mathcal{I}}_{x}$ is the index set of ${\ker_{{}_{{(H,\mathcal{B})}}}}(x)$.
+By Theorem 5.7, ${\mathcal{I}_{x}}={\mathcal{I}_{\langle x\rangle}}$, and by Lemma 5.5, ${\mathcal{I}_{x}}=\bigcap_{s\in\mathcal{S}}{\mathcal{I}_{s}}$.
+The first sentence follows.
+Recall that $x=\sum_{j\in\mathcal{I}}\mu_{j}(x)F_{j}$, thus $x\lambda_{{(L,\mathcal{S})}}=\sum_{j\in\mathcal{I}_{{(L,\mathcal{S})}}}\mu_{j}(x)F_{j}$.
+But the first sentence means that $j\in{\mathcal{I}}_{{(L,\mathcal{S})}}$ if and only if $\mu_{j}(s)=\mathrm{FPdim}(s)$, for all $s\in\mathcal{S}$.
+So if $s\in\mathcal{S}$ then $s\lambda_{{(L,\mathcal{S})}}=d_{s}\lambda_{{(L,\mathcal{S})}}$.
+Regarding the converse, we deduce by positivity in Equation (5.2) that $sS\subseteq S$.
+Specifically, since $s=sx_{1}\in sS$, it follows that $s\in S$.
+∎
+In particular, for an ARN-hypergroup with $\mu_{1}=\mathrm{FPdim}$, then ${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}=\{j\in{\mathcal{I}}\text{ such that }\mu_{j}|_{(L,\mathcal{S})}=\mathrm{FPdim}\}$.
+Let ${(L,\mathcal{S})},{(K,\mathcal{T})}$ be two sub-hypergroups of an ARN-hypergroup ${(H,\mathcal{B})}$.
+Let ${(L,\mathcal{S})}\vee{(K,\mathcal{T})}$ be the sub-hypergroup generated by ${(L,\mathcal{S})}$ and ${(K,\mathcal{T})}$.
 Proposition 6.11.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
-Then $x_{i}\in{G{(H,\mathcal{B})}}$ if and only $i\in\mathcal{I}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}$.
-Dually one obtains also the following:
+With above notations, $\mathcal{I}_{{}_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}}={\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\cap{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$.
+In other words, $\lambda_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}=\lambda_{{(L,\mathcal{S})}}\lambda_{{(K,\mathcal{T})}}$.
+Proof.
+Let ${\mathcal{S}}\vee{\mathcal{T}}$ denotes the basis of ${(L,\mathcal{S})}\vee{(K,\mathcal{T})}$. By Lemmas 6.10 and 5.3,
+${\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}\cap{\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}=\bigcap_{a\in{\mathcal{S}}\cup{\mathcal{T}}}{\mathcal{I}}_{a}$ =$=\{j\in{\mathcal{I}}\ |\ \mu_{j}\in\ker_{{(H,\mathcal{B})}}(a),\ \forall a\in{\mathcal{S}}\cup{\mathcal{T}}\}$
+=$=\{j\in{\mathcal{I}}\ |\ {\mathcal{S}}\cup{\mathcal{T}}\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})\}$
+=$=\{j\in{\mathcal{I}}\ |\ {\mathcal{S}}\vee{\mathcal{T}}\subset\ker_{{\widehat{(H,\mathcal{B})}}}(\mu_{j})\}$
+=$=\{j\in{\mathcal{I}}\ |\ \mu_{j}\in\ker_{{(H,\mathcal{B})}}(a),\ \forall a\in{\mathcal{S}}\vee{\mathcal{T}}\}$
+$=\bigcap_{a\in{\mathcal{S}}\vee{\mathcal{T}}}{\mathcal{I}}_{a}=\mathcal{I}_{{}_{{(L,\mathcal{S})}\vee{(K,\mathcal{T})}}}.\qed$
+Proposition 6.12.
+For any two sub-hypergroups ${(L,\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${(L,\mathcal{S})}\subseteq{(K,\mathcal{T})}$ if and only if ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$.
+Proof.
+By Lemma 6.10, if ${(L,\mathcal{S})}\subseteq{(K,\mathcal{T})}$ then ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$.
+By Proposition 6.11, if ${\mathcal{I}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{I}_{{}_{{(L,\mathcal{S})}}}}$, then ${\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}={\lambda_{{}_{{(K,\mathcal{T})}}}}$.
+Pick $s\in\mathcal{S}$, then
+$s{\lambda_{{}_{{(K,\mathcal{T})}}}}=s{\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}{\lambda_{{}_{{(L,\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}{\lambda_{{}_{{(K,\mathcal{T})}}}},$
+thus $s\in\mathcal{T}$, by Lemma 6.10. ∎
 Corollary 6.13.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
-(6.3) $\lambda_{{}_{{{\widehat{(H,\mathcal{B})}}}_{ad}}}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
-Proof of Theorem 1.7 and Theorem 1.8.
+For any two elements $P,Q\in{(H,\mathcal{B})}_{+}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $\langle P\rangle\subseteq\langle Q\rangle$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(Q)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$.
+For any ${(H,\mathcal{B},\mu_{1})}$, recall that $P:=\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$.
+Proposition 6.14.
+For any abelian RN-hypergroup ${(H,\mathcal{B})}$ then
+$\langle P^{2}\rangle={{(H,\mathcal{B})}_{ad}}.$
 Proof.
-If ${(H,\mathcal{B})}$ is Burnside then by Corollary 4.5,
-$\big{(}\prod_{j=1}^{m}\mu_{j}\big{)}^{2}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}.$
-Corollary 6.8 implies Equation (1.4). Conversely, if Equation (1.4) holds then
-$\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\pm{\widehat{E}}_{i}$
-Evaluating both sides at any ${\frac{x_{i}}{d_{i}}}$, one obtains the Burnside property. ∎
-By duality one obtains directly Theorem 1.8.
-6.4. On the sub-hypergroup generated by $P$
-Proposition 6.9.
-For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $\mathcal{J}_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\cap{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
-Proof.
-We prove that $\lambda_{{}_{{(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}}}={\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$.
-Then Equation (6.1) implies the intersection result.
-Note that for all $s\in\mathcal{S}$ and $t\in\mathcal{T}$, $st{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{s}d_{t}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$.
-Lemma 5.3 shows that for any constituent $u$ of $st$ then $u{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}=d_{u}{\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}$ also.
-Thus the same relation holds for any standard element of ${(L,\;\mathcal{S})}\vee{(K,\mathcal{T})}$.
+By Definition 6.2, ${{(H,\mathcal{B})}_{ad}}:=\langle I(1)\rangle$, so we are reduced to show that $\mathcal{I}_{\langle P^{2}\rangle}=\mathcal{I}_{\langle I(1)\rangle}$, so (by Theorem 5.7) that $\mathcal{I}_{P^{2}}=\mathcal{I}_{I(1)}$ i.e.
+${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})={\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
+Note that $\big{|}\psi(P)\big{|}=\prod_{i\in\mathcal{I}}\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$.
+Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$, then $\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}=1$, i.e $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, by Proposition 6.1.
+Thus, ${\ker_{{}_{{(H,\mathcal{B})}}}}(P)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, and the same argument shows that ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
+Conversely, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$ then $\psi({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}^{2}=1$.
+Thus, if $i\neq i^{*}$ then, grouping together $x_{i}$ and $x_{i^{*}}$ in $P$, we obtain $\psi({\frac{x_{i}}{d_{i}}})\psi(\frac{x_{{i^{*}}}}{{d_{i^{*}}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})^{2}\big{|}=1$.
+Thus,
+$\psi(P)=\prod_{i={i^{*}}}\psi({\frac{x_{i}}{d_{i}}}).$
+Note that if $i=i^{*}$, then $\psi({\frac{x_{i}}{d_{i}}})=\pm 1$, since in this case $\psi(x_{i})=\psi(x_{i^{*}})\in\mathbb{R}$.
+Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, then $\psi(P)=\pm 1$ and $\psi(P^{2})=1$, i.e.
+$\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})$.
 ∎
 Corollary 6.15.
-For any two sub-hypergroups ${(L,\;\mathcal{S})},{(K,\mathcal{T})}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ if and only if ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
+For any ARN-hypergroup ${(H,\mathcal{B})}$, then ${{(H,\mathcal{B})}_{ad}}\subseteq\langle P\rangle$.
+Corollary 6.16.
+Let ${(H,\mathcal{B})}$ be an ARN hypergroup. The following holds:
+  1. (1)
+    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if ${(H,\mathcal{B})}$ is pointed,
+  2. (2)
+    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is pointed,
+  3. (3)
+    ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is perfect.
 Proof.
-If ${(L,\;\mathcal{S})}\subseteq{(K,\mathcal{T})}$ then ${\lambda_{{}_{{(L,\;\mathcal{S})}}}}{\lambda_{{}_{{(K,\mathcal{T})}}}}={\lambda_{{}_{{(K,\mathcal{T})}}}}$ and therefore equation (6.1) implies ${\mathcal{J}_{{}_{{(K,\mathcal{T})}}}}\subseteq{\mathcal{J}_{{}_{{(L,\;\mathcal{S})}}}}$.
-The converse follows from Proposition 6.9.
+First, recall that ${{(H,\mathcal{B})}_{ad}}=\langle I(1)\rangle$, so ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $I(1)=n{(H,\mathcal{B})}x_{1}$, if and only if ${(H,\mathcal{B})}$ is pointed, as $I(1)=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}$.
+Next, by Equation (5.2), ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if $\lambda_{{{(H,\mathcal{B})}_{ad}}}=\lambda_{\mathbb{C}}=x_{1}=\operatorname{id}=\sum_{i\in\mathcal{I}}F_{i},$ if and only if $G{\widehat{(H,\mathcal{B})}}=\widehat{\mathcal{B}}$, by Equation 6.2, meaning that $^{\widehat{(H,\mathcal{B})}}$ is pointed.
+Similary, ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if $\lambda_{{{(H,\mathcal{B})}_{ad}}}=F_{1}$, if and only if $G{\widehat{(H,\mathcal{B})}}=\{\mu_{1}\}$, meaning that $^{\widehat{(H,\mathcal{B})}}$ is perfect.
 ∎
-Corollary 6.11.
-For any two elements $P,Q\in{(H,\mathcal{B})}_{+}$ of an ARN-hypergroup ${(H,\mathcal{B})}$, then $<P>\subseteq<Q>$ if and only if ${\ker_{{}_{{(H,\mathcal{B})}}}}(Q)\subseteq{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$.
-Recall $P:=\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}$ for any ${(H,\mathcal{B},\mu_{1})}$.
-Proposition 6.12.
-In any abelian RN-hypergroup ${(H,\mathcal{B})}$ one has
-$<P^{2}>={{(H,\mathcal{B})}_{ad}}.$
-Proof.
-One can prove first that ${\ker_{{}_{{(H,\mathcal{B})}}}}(P^{2})={\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
-Indeed $\big{|}\psi(P)\big{|}=\prod_{i=1}^{m}\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}\leq 1$.
-Thus, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(P)$, then $\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}=1$, i.e $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$, by Proposition 6.1.
-Conversely, if $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$ then $\psi({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})\big{|}^{2}=1$.
-Thus, if $i\neq i^{*}$ then grouping together $x_{i}$ and $x_{i^{*}}$ in $P$, one obtains $\psi({\frac{x_{i}}{d_{i}}})\psi(\frac{x_{{i^{*}}}}{{d_{i^{*}}}})=\big{|}\psi({\frac{x_{i}}{d_{i}}})^{2}\big{|}=1$.
-Thus,
-$\psi(P)=\prod_{i={i^{*}}}\psi({\frac{x_{i}}{d_{i}}}).$
-Note that if $i=i^{*}$ then $\psi({\frac{x_{i}}{d_{i}}})=\pm 1$, since in this case $\psi(x_{i})=\psi(x_{i^{*}})\in\mathbb{R}$.
-Thus, $\psi(P)=\pm 1$ and $\psi(P^{2})=1$ for all $\psi\in{\ker_{{}_{{(H,\mathcal{B})}}}}(I(1))$.
-It follows that $\lambda_{<P^{2}>}=\lambda_{<I(1)>}$ by Theorem 5.7, and so $<P^{2}>={{(H,\mathcal{B})}_{ad}}$. ∎
-Corollary 6.13.
-In any ARN-hypergroup ${(H,\mathcal{B})}$ one has
-${{(H,\mathcal{B})}_{ad}}\subseteq<P>.$
-Corollary 6.14.
-Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup. Then the following holds:
-  1. (1)
-    ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is pointed.
-  2. (2)
-    ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$ if and only if ${\widehat{(H,\mathcal{B})}}$ is perfect.
 Remark 6.17.
@@
 Corollary 6.18.
-A simple non-pointed abelian RN-hypergroup ${(H,\mathcal{B})}$ has a perfect dual.
+A simple non-pointed ARN-hypergroup ${(H,\mathcal{B})}$ has a perfect dual.
 Proof.
-By simplicity ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$, and therefore ${\widehat{(H,\mathcal{B})}}$ is perfect. ∎
+By simplicity ${{(H,\mathcal{B})}_{ad}}=\mathbb{C}$ or ${(H,\mathcal{B})}$.
+By non-pointed assumption and Corollary 6.16, ${{(H,\mathcal{B})}_{ad}}\neq\mathbb{C}$, therefore ${{(H,\mathcal{B})}_{ad}}={(H,\mathcal{B})}$, and so $^{\widehat{(H,\mathcal{B})}}$ is perfect by Corollary 6.16.
+∎
 Corollary 6.19.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup with ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ RN. Then the following holds:
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup with $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ RN. Then the following holds:
+  1. (1)
+    ${\widehat{(H,\mathcal{B})}}_{ad}=\mathbb{C}$ if and only if $^{\widehat{(H,\mathcal{B})}}$ is pointed,
   2. (2)
@@
   3. (3)
-    ${\widehat{(H,\mathcal{B})}}_{ad}={\widehat{(H,\mathcal{B})}}$ if and only if ${(H,\mathcal{B})}$ is perfect.
+    $^{\widehat{(H,\mathcal{B})}}_{ad}={\widehat{(H,\mathcal{B})}}$ if and only if ${(H,\mathcal{B})}$ is perfect.
 7. On the universal grading of hypergroups
-We remark that the universal grading construction for fusion rings from [GN08, §2] works word for word in the ARN-hypergroup settings.
-For the completeness of the paper, we sketch its main steps below.
-7.0.1. Based modules are completely reducible
+The universal grading construction for fusion rings from [GN08, §2] works word for word in the ARN-hypergroup settings. For the completeness of the paper, we sketch its main steps below.
+7.1. Based modules are completely reducible
 Definition 7.1.
-Given an ARN-hypergroup ${(H,\mathcal{B})}$, we define a based ${(H,\mathcal{B})}$-module as a pair $(M,\mathcal{M})$ where $M$ is a $H$-module with $\mathcal{M}$ a finite free $H$-basis for $M$, and if $x_{i}m_{j}=\sum_{k}t^{k}_{ij}m_{k}$, then $t^{k}_{ij}=t^{k^{*}}_{i^{*}j}$, for all $m_{k}\in\mathcal{M}$ and all $x_{i}\in\mathcal{B}$.
-Note that the above condition is equivalent to $(L_{x_{i}^{*}})^{t}=L_{x_{i}}$, where $L_{x_{i}}$ is the matrix of the left multiplication operator by $x_{i}$ on $M$.
-Then one can define a inner product on $M$ such that for all $m_{i},m_{j}\in\mathcal{M}$
-(7.1) $(m_{i},m_{j})=\delta_{i,j}$
-This inner product can be used to prove that any based ${(H,\mathcal{B})}$-modules is completely reducible, as in [Ost03a, Lemma 2.1].
+Given an ARN-hypergroup ${(H,\mathcal{B})}$, we define a based left ${(H,\mathcal{B})}$-module as a pair $(M,\mathcal{M})$ where $M$ is a $H$-module with $\mathcal{M}$ a finite free $H$-basis for $M$, i.e.
+for all $x_{i}\in\mathcal{B}$ and all $m_{j}\in\mathcal{M}$, then $x_{i}m_{j}=\sum_{k}t^{k}_{ij}m_{k}$; and moreover, for all $i,j,k$ then $t^{k}_{ij}$ is zero if and only if $t^{j}_{i^{*}k}$ is zero.
+The last condition above means that $(x_{i}m_{j},m_{k})$ vanishes if and only if $(m_{j},x_{i^{*}}m_{k})$ vanishes, for the inner product on $M$ such that for all $m_{i},m_{j}\in\mathcal{M}$
+(7.1) $(m_{i},m_{j})=\delta_{i,j}.$
+It is easy to deduce that any based ${(H,\mathcal{B})}$-modules is completely reducible, as in [Ost03a, Lemma 2.1].
+Indeed, if $N\subset M$ and $HN\subset N$ then $(N,N^{\perp})=\{0\}=(HN,N^{\perp})=(N,HN^{\perp})$.
+Thus, $HN^{\perp}\subset N^{\perp}$.
 Let ${(H,\mathcal{B})}$ be a RN-hypergroup and $x,y\in{(H,\mathcal{B})}_{+}$. We say that $x$ contains $y$ (or $y$ is a constituent of $x$) if the difference $x-y$ is in ${(H,\mathcal{B})}_{+}$.
 Definition 7.2.
-Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup, $G$ be a finite group.
+Let ${(H,\mathcal{B})}$ be a RN-hypergroup, $G$ be a finite group.
 Suppose that $\mathcal{B}=\sqcup_{g\in G}{\mathcal{B}}_{g}$ is a partition of $\mathcal{B}$, and let $H_{g}:=\mathbb{C}[\mathcal{B}_{g}]$.
@@
 The grading is called faithful if $\mathcal{B}_{g}\neq\emptyset$, for all $g\in G$.
-Let ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ be a grading of ${(H,\mathcal{B})}$, and define $R_{g}:=\sum_{x_{i}\in\mathcal{B}_{g}}h_{i}d_{i}x_{i}$, for all $g\in G$.
+Please note that a faithful grading group of an abelian hypergroup must also be abelian.
+Let ${(H,\mathcal{B})}=\bigoplus_{g\in G}(H_{g},\mathcal{B}_{g})$ be a faithful grading of ${(H,\mathcal{B})}$, and define $R_{g}:=\sum_{x_{i}\in\mathcal{B}_{g}}h_{i}d_{i}x_{i}$, for all $g\in G$.
 The analogue of [ENO05, Proposition 8.20] holds in the settings of ARN-hypergroup:
@@
 Thus, $RR_{h}=\mathrm{FPdim}(R_{h})R$.
-Since $R_{g}R_{h}\subseteq R_{gh}$, one obtains from the previous equality that $R_{g}R_{h}=\mathrm{FPdim}(R_{h})R_{gh}$, for all $g,h\in G$.
-Passing to $\mathrm{FPdim}$, one get that $\mathrm{FPdim}(R_{gh})=\mathrm{FPdim}(R_{g})$, for all $g,h\in G$.
+Since $H_{g}H_{h}\subseteq H_{gh}$, we obtain from the previous equality that $R_{g}R_{h}=\mathrm{FPdim}(R_{h})R_{gh}$, for all $g,h\in G$.
+Passing to $\mathrm{FPdim}$, we get that $\mathrm{FPdim}(R_{g})\mathrm{FPdim}(R_{h})=\mathrm{FPdim}(R_{h})\mathrm{FPdim}(R_{gh})$; but $\mathrm{FPdim}(R_{h})\neq 0$ by faithfulness, so $\mathrm{FPdim}(R_{gh})=\mathrm{FPdim}(R_{g})$, for all $g,h\in G$.
 7.2. Constructing the universal grading
 Proposition 7.3.
-Let ${(H,\mathcal{B})}$ be a RN-hypergroup. As above, define ${I(1)}:=\sum\limits_{a\in\mathcal{B}}h_{a}aa^{*}$. With the above notations:
+Let ${(H,\mathcal{B})}$ be a RN-hypergroup. For all $x\in H$, define $I(x):=\sum_{j\in\mathcal{I}}h_{j}x_{j}xx_{j^{*}}$. With the above notations:
   1. (1)
-    $I(1)$ is central in $H$.
+    $I(x)$ is central in $H$.
   2. (2)
@@
   3. (3)
-    A based $H_{ad}$-subbimodule $M$ of $H$ is indecomposable if and only if it is indecomposable as $H_{ad}$-bimodule.
+    A based $H_{ad}$-subbimodule $M$ of $H$ is indecomposable if and only if it is indecomposable as left $H_{ad}$-module.
 Proof.
   1. (1)
-    The proof of this fact is the same as in [GN08].
+    For all $i\in\mathcal{I}$, by Frobenius reciprocity (2.18):
+    $x_{i}I(x)=\sum_{j\in\mathcal{I}}h_{j}x_{i}x_{j}xx_{j^{*}}=\sum_{j,k\in\mathcal{I}}h_{j}N_{i,j}^{k}x_{k}xx_{j^{*}}=\sum_{j,k\in\mathcal{I}}h_{k}N_{k^{*},i}^{j^{*}}x_{k}xx_{j^{*}}$
+    $=\sum_{k\in\mathcal{I}}h_{k}x_{k}x(\sum_{j\in\mathcal{I}}N_{k^{*},i}^{j^{*}}x_{j^{*}})=\sum_{k\in\mathcal{I}}h_{k}x_{k}xx_{k^{*}}x_{i}=I(x)x_{i}.$
   2. (2)
-    Since $m{I(1)}^{n}={I(1)}^{n}m$, for all $m\in M$, this shows that both the left and the right $H_{ad}$-submodules have the same standard constituents.
+    Recall that $H_{ad}:=\langle I(1)\rangle$.
+By (1), $m{I(1)}^{n}={I(1)}^{n}m$, thus by RN-assumption, $mH_{ad}=H_{ad}m\subset M$, for all $m\in M$, meaning that $M$ is also a based right $H_{ad}$-submodule.
   3. (3)
     Clear. ∎
-Let $H=\oplus_{a\in A}H_{a}$ be a decomposition of $H$ as a $H_{ad}$-bimodule.
+Let $H=\oplus_{a\in A}H_{a}$ be a decomposition (into indecomposable) of $H$ as a $H_{ad}$-bimodule.
 This also corresponds to a partition $\mathcal{B}=\sqcup_{a\in A}\mathcal{B}_{a}$.
-One may suppose that there is an element $1\in A$ such that $H_{1}=H_{ad}$.
+Let 111 be an element of $A$ such that $H_{1}=H_{ad}$.
 7.2.1. Definition of $a^{*}$ for any $a\in A$
-Note that if $H_{a}$ is an indecomposable $H_{ad}$-bimodule then, $(H_{a})^{*}$ is also an indecomposable $H_{ad}$-bimodule.
-Therefore, we may suppose that $(H_{a})^{*}=H_{a^{*}}$, for some $a^{*}\in A$.
-Next lemma has the same proof as [GN08, Lemma 3.4].
+Since $H_{a}$ is an indecomposable $H_{ad}$-bimodule then $(H_{a})^{*}$ is also an indecomposable $H_{ad}$-bimodule.
+Let $a^{*}$ be an element of $A$ such that $(H_{a})^{*}=H_{a^{*}}$.
+Next lemma is as [GN08, Lemma 3.4].
 Lemma 7.4.
-If $x_{a},y_{a}\in\mathcal{B}_{a}$ then, $x_{a}y_{a}^{*}\in H_{ad}$.
+If $x_{a},y_{a}\in\mathcal{B}_{a}$ then $x_{a}y_{a}^{*}\in H_{ad}$.
 Proof.
-Note that
-$M:=\{y_{a}\in{\mathcal{B}}_{a}\;|\;m(y_{a},x_{a}{I(1)}^{n})>0,\;\;\text{for some}\;\;n\geq 0\}$
-is a $H_{ad}$-submodule of $H_{a}$.
-Since $H_{a}$ is indecomposable, it follows that $M=H_{a}$.
-It follows that $m(y_{a}^{*},x_{a}{I(1)}^{n})>0$, which implies that $x_{a}y_{a}^{*}$ has a constituent in $H_{1}$.
-This gives that $x_{a}y_{a}^{*}$ is contained in $x_{a}x_{a}^{*}{I(1)}^{n}\in R_{1}.$
+Note that $M:=H_{a}\cap(H_{ad}y_{a})$ is a $H_{ad}$-submodule of $H_{a}$ (containing $y_{a}$).
+But $H_{a}$ is indecomposable (so irreducible by complete reducibility).
+It follows that $M=H_{a}$, so $H_{a}\subset H_{ad}y_{a}$.
+Thus $x_{a}\in H_{ad}y_{a}$ and $x_{a}y_{a}^{*}\in H_{ad}y_{a}y_{a}^{*}\subset H_{ad}$.
+∎
 Theorem 7.5.
-There is a well-defined binary operation on $A$ such that $a\star b=c$ if and only if $x_{a}\star y_{b}\in H_{c}$, for all $(x_{a},y_{b})\in\mathcal{B}_{a}\times{\mathcal{B}}_{b}$.
-It defines a group structure on $A$, with $a^{*}$ the inverse of $a$, and $1$ the unit of $A$.
+For all $a,b\in A$, there is $c\in A$ such that $H_{a}H_{b}\subset H_{c}$. This induces a group structure on $A$ given by $ab:=c$. Moreover, $a^{*}$ is the inverse of $a$, and 111 the unit of $A$.
 Proof.
@@
 Every ARN-hypergroup ${(H,\mathcal{B})}$ has a canonical faithful grading by the group $U(H)$.
-Any other faithful grading of ${(H,\mathcal{B})}$ by a group $G$ is determined by a surjective homomorphism $\pi:U(H)\rightarrow G$.
+Any other faithful grading of ${(H,\mathcal{B})}$ by a group $G$ is determined by a surjective group morphism $\pi:U_{H}\rightarrow G$.
 Proof.
-Let ${(H,\mathcal{B})}=\oplus_{g\in G}(S_{g},\mathcal{D}_{g})$ be another grading.
-It follows that $xx^{*}\in\mathcal{D}_{1}$, for any $x\in\mathcal{D}_{g}$, and therefore, $\mathcal{B}_{ad}\subseteq\mathcal{D}_{1}$.
+Let ${(H,\mathcal{B})}=\oplus_{g\in G}(S_{g},\mathcal{D}_{g})$ be another faithful grading.
+It follows that $xx^{*}\in S_{1}$, for any $x\in\mathcal{D}_{g}$, but $xx^{*}\in H_{ad}$ therefore, $\mathcal{B}_{ad}\subseteq\mathcal{D}_{1}$.
 Moreover, each $S_{g}$ is a ${H_{ad}}$-subbimodule of $H$.
 Since $H_{a}$ is an indecomposable module, then $H_{a}\subseteq S_{\pi(a)}$, for some well-defined $\pi(a)\in G$.
-It follows that $\pi:U(H)\rightarrow G$ is an epimorphism of groups, since we started with a grading.
+It follows that $\pi:U_{H}\rightarrow G$ is a surjective group morphism.
+Indeed, for all $a,b\in A$, then $H_{a}\subset S_{\pi(a)}$ and $H_{b}\subset S_{\pi(b)}$, thus $H_{a}H_{b}\subset S_{\pi(a)}S_{\pi(b)}\subset S_{\pi(a)\pi(b)}$.
+But $H_{a}H_{b}\subset H_{ab}$, thus $H_{a}H_{b}\subset H_{ab}\cap S_{\pi(a)\pi(b)}\neq\emptyset$, so $H_{ab}\subset S_{\pi(a)\pi(b)}$, meaning that $\pi(ab)=\pi(a)\pi(b)$.
+Finally, the surjectivity comes from the faithfulness.
 ∎
 7.3. On the dual universal grading group and invertible
-In this subsection, we use the notation $H$ for a hypergroup ${(H,\mathcal{B})}$, and $\widehat{H}$ for its dual ${\widehat{(H,\mathcal{B})}}$. The goal of this subsection is to prove the following:
+In this subsection, we use the notation $H$ for a hypergroup ${(H,\mathcal{B})}$, and $\widehat{H}$ for its dual $^{\widehat{(H,\mathcal{B})}}$.
+The goal of this subsection is to prepare for the proof (§7.4) of the following:
 Theorem 7.8.
-Let $H$ be an abelian RN-hypergroup. Then,
-$U_{H}\simeq{G(\widehat{H})}$
-Dually one has the following:
+Let $H$ be an ARN-hypergroup. Then,
+$U(H)\simeq{G(\widehat{H})}$
+Dually, we have the following:
 Theorem 7.9.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
-$U_{\widehat{H}}\simeq G(H)$
-as abelian groups.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup satisfying any of the two hypothesis from Lemma 2.23. Recall that this means that al least one of the following holds:
-  1. (1)
-    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
-  2. (2)
-    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
-By Corollary 2.32, for any such hypergroup, if $\mu_{j}\in{G(\widehat{H})}$ then $\mu_{j}^{n}=\mu_{1}$, for some $n\geq 0$.
-Thus, for all $i$, $\mu_{j}(x_{i})=\omega_{ij}d_{i}$ for some complex number $|\omega_{ij}|=1$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then
+$U(\widehat{H})\simeq G(H).$
+Recall from Lemma 2.28 that for any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ then $\mu_{j}\in{G(\widehat{H})}$ if and only if $|\omega_{ij}|=1$, for all $i\in\mathcal{I}$, with $\omega_{ij}:=\mu_{j}(\frac{x_{i}}{d_{i}})$.
 Definition 7.10.
-Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23.
-For any standard element $x_{i}\in\mathcal{B}$, we define a morphism $\omega_{i}:{G(\widehat{H})}\rightarrow\mathbb{C},\mu_{j}\mapsto\omega_{ij}=\mu_{j}({\frac{x_{i}}{d_{i}}})$.
-Next lemma shows that $\omega_{i}$ is a character of ${G(\widehat{H})}$.
-Lemma 7.11.
-Let ${(H,\mathcal{B})}$ be a hypergroup satisfying any of the two hypothesis from Lemma 2.23. For any $x_{i}$, then $\omega_{i}$ is a linear character of ${G(\widehat{H})}$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup.
+For any $x_{i}\in\mathcal{B}$, we define the linear character $\omega_{i}:={\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}$, the restriction to ${G(\widehat{H})}$ of the linear character $^{\frac{x_{i}}{d_{i}}}:\widehat{H}\rightarrow\mathbb{C}$, i.e.
+$\omega_{i}(\mu_{j})=\omega_{ij}$.
+Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Given $\psi\in\widehat{{G(\widehat{H})}}$, let
+(7.3) $\mathcal{B}_{\psi}:=\{x_{i}\in\mathcal{B}\;\mid\;\omega_{i}=\psi\}=\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j}),\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}.$
+Proposition 7.11.
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then $(\mathcal{B}_{\psi})_{\psi\in\widehat{{G(\widehat{H})}}}$ defines a grading on ${(H,\mathcal{B})}$.
 Proof.
-Since ${\frac{x_{i}}{d_{i}}}:\widehat{H}\rightarrow\mathbb{C}$ are algebra characters, then their restrictions ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}$ are also algebra characters.
-Note that ${\frac{x_{i}}{d_{i}}}\big{|}_{{G(\widehat{H})}}=\omega_{i}$.
+For any $\mu_{j}\in{G(\widehat{H})}$, $\psi,\phi\in\widehat{{G(\widehat{H})}}$, $x_{i}\in B_{\psi}$ and $x_{k}\in B_{\phi}$, then
+$\psi\phi(\mu_{j})=\psi(\mu_{j})\phi(\mu_{j})=\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j}(\frac{x_{k}}{d_{k}})=\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}).$
+By the triangle inequality:
+$1=|\psi\phi(\mu_{j})|=|\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l\in\mathcal{I}}d_{l}N^{l}_{ik}|\mu_{j}(\frac{x_{l}}{d_{l}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l\in\mathcal{I}}d_{l}N^{l}_{ik}=1,$
+which forces $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$ if $N^{l}_{ik}$ is nonzero, meaning that if $x_{i}\in\mathcal{B}_{\psi}$ and $x_{k}\in\mathcal{B}_{\phi}$, then $x_{i}x_{k}$ has all constituents in $\mathcal{B}_{\psi\phi}$.
 ∎
-Let ${(H,\mathcal{B})}$ be a satisfying any of the two hypothesis from Lemma 2.23. Given $\psi\in\widehat{{G(\widehat{H})}}$, let
-(7.3) $\mathcal{B}_{\psi}:=\{x_{i}\in\mathcal{B}\;\mid\;\omega_{i}=\psi\}=\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j}),\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
-Note that the above definition can be rewritten as
-$x_{i}\in\mathcal{B}_{\psi}\iff{\frac{x_{i}}{d_{i}}}\big{|}_{G(\widehat{H})}=\psi$
-Proposition 7.12.
-Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Then, Equation (7.3) defines a grading on ${(H,\mathcal{B})}$.
-Proof.
-It reduces to show that if $x_{i}\in\mathcal{B}_{\psi}$ and $x_{k}\in\mathcal{B}_{\phi}$, then $x_{i}x_{k}$ has all constituents in $\mathcal{B}_{\psi\phi}$.
-Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})=\psi(\mu_{j})$ and $\mu_{j}(\frac{x_{k}}{d_{k}})=\phi(\mu_{j})$, for all $\mu_{j}\in{G(\widehat{H})}$.
-It follows that
-$\psi\phi(\mu_{j})=\psi(\mu_{j})\phi(\mu_{j})=\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j}(\frac{x_{k}}{d_{k}})=\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}).$
-Now, since $|\psi\phi(\mu_{j})|=1$ and ${\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}}=\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}\frac{x_{l}}{d_{l}}$, the triangle inequality implies that $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$, for all $l$ with $N^{l}_{ik}>0$.
-Indeed, note that
-$1=|\psi\phi(\mu_{j})|=|\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{k}}{d_{k}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}|\mu_{j}(\frac{x_{l}}{d_{l}})|\leq\frac{1}{d_{i}d_{k}}\sum_{l=1}^{m}d_{l}N^{l}_{ik}=1,$
-and this forces $\mu_{j}(\frac{x_{l}}{d_{l}})=\mu_{j}(\frac{x_{i}}{d_{i}})\mu_{j}(\frac{x_{k}}{d_{k}})$. ∎
-7.3. Computations of the regular components
-For any grading component $\mathcal{B}_{\psi}$, we define (as above) the regular part of the grading:
+7.4. Computation of the regular components
+For any grading component $\mathcal{B}_{\psi}$, we define (as above):
 $R_{\psi}=\sum_{x_{i}\in\mathcal{B}_{\psi}}h_{i}d_{i}x_{i}.$
 Lemma 7.12.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. With the above notations, for all $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$,
-(7.4) $F_{j}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. For all $\mu_{j}\in{G(\widehat{H})}$,
+(7.4) $F_{j}=\frac{1}{n(H)}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}.$
 Proof.
-For any $j$ such that $\mu_{j}\in{G(\widehat{H})}$,
-$F_{j}$ $=$ $\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\sum_{x_{i}\in\mathcal{B}_{\psi}}\big{(}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=$
-$=$ $\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi^{-1}(\mu_{j})R_{\psi}\big{)}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
-If $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ then $n_{j}=n(H)$ by Corrolary 2.22. ∎
+For all $\mu_{j}\in{G(\widehat{H})}$, by Equation (2.4):
+$F_{j}$ $=$ $\frac{1}{n_{j}}\sum_{i\in\mathcal{I}}h_{i}\mu_{j}(x_{i^{*}})x_{i}=\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\sum_{x_{i}\in\mathcal{B}_{\psi}}h_{i}\mu_{j}(x_{i^{*}})x_{i}=$
+$=$ $\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi^{-1}(\mu_{j})R_{\psi}=\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}.$
+Finally, if $\mu_{j}\in{G(\widehat{H})}$ then $n_{j}=n(H)$ by Corollary 2.29. ∎
 Corollary 7.13.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23.
-With above notations, for any linear character $\psi$ of ${G(\widehat{H})}$, the set $\mathcal{B}_{\psi}$ is not empty.
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. The set $\mathcal{B}_{\psi}$ is not empty.
 Proof.
-Previous corollary shows that the linear span of $<F_{j}\;|\;\mu_{j}\in{G(\widehat{H})}>$ lies inside the span of $<\mathcal{B}_{\psi}\;|\;\mathcal{B}_{\psi}\neq\emptyset>$.
-By counting the dimension of these vector spaces, it follows that $\mathcal{B}_{\psi}$ is not empty, for any $\psi$ in the dual.
+Lemma 7.12 shows that the linear span of $\{F_{j}\;|\;\mu_{j}\in{G(\widehat{H})}\}$ lies inside the span of $\{\mathcal{R}_{\psi}\;|\;\psi\in\widehat{{G(\widehat{H})}}\text{ with }\mathcal{B}_{\psi}\neq\emptyset\}$.
+But $\widehat{{G(\widehat{H})}}\simeq{G(\widehat{H})}$ as finite abelian group, so have same order.
+Thus, an argument involving the dimension of these vector spaces implies that $\mathcal{B}_{\psi}$ is not empty, for any $\psi\in\widehat{{G(\widehat{H})}}$.
 ∎
-From Equation (7.4), it follows immediately:
+By Equation (7.4) and the second orthogonality relation (2.6) applied on the (hyper)group ${G(\widehat{H})}$, it follows that:
 Lemma 7.14.
-Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. Then
-$R_{\psi}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}\psi(\mu_{j})F_{j}\big{)}.$
-Proof of Theorem 7.8
-By Lemma 7.15, if $\mu_{j}=\mu_{1}$, then
-$R_{1}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}F_{j}\big{)}=\lambda_{{H_{ad}}}.$
-Corollary 6.10 proves that $\mathcal{B}_{1}=\mathcal{B}_{ad}$. Therefore the above grading
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then
+$R_{\psi}=\frac{n(H)}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}\psi(\mu_{j})F_{j}\big{)}.$
+Proof of Theorem 7.8 and Theorem 7.9
+By Lemma 7.14, and Corollary 6.6, if $\psi=1$, then
+$R_{1}=\frac{n(H)}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}F_{j}\big{)}=\frac{n(H)}{|{G(\widehat{H})}|}\lambda_{{H_{ad}}}.$
+Proposition 6.12 proves that $\mathcal{B}_{1}=\mathcal{B}_{ad}$. Therefore the above grading
 $\mathcal{B}=\bigsqcup_{\psi\in\widehat{G(\widehat{H})}}{\mathcal{B}}_{\psi}$
-coincides with the universal grading of ${(H,\mathcal{B})}$. Thus $U_{H}={G(\widehat{H})}$. By duality, $U_{\widehat{H}}=G(H)$ if ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is abelian.
-7.4. On the perp of the adjoint subring
-Recall from Definition 1.6 that an abelian RN-hypergroup is called dualizable if ${\widehat{(H,\mathcal{B})}}$ is also RN.
-For any sub-hypergroup ${(L,\;\mathcal{S})}$ of an abelian dualizable hypergroup ${(H,\mathcal{B})}$, define
+coincides with the universal grading of ${(H,\mathcal{B})}$. Thus $U_{H}={G(\widehat{H})}$. By duality, $U_{\widehat{H}}=G(H)$ if $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
+7.5. On the perp of the adjoint sub-hypergroup
+Recall from Definition 1.6 that an ARN-hypergroup is called dualizable if $^{\widehat{(H,\mathcal{B})}}$ is also (A)RN.
+For any sub-hypergroup ${(L,\mathcal{S})}$ of a dualizable ARN-hypergroup ${(H,\mathcal{B})}$, define
 $\mathcal{S}^{\perp}:=\{\mu_{j}\;|\;\mu_{j}(s)=\mathrm{FPdim}(s)\;\text{for all}\;s\in\mathcal{S}\}.$
 With the above notations, $\mathcal{S}^{\perp}=\bigcap_{s\in\mathcal{S}}\ker_{(H,\mathcal{B})}(s)$.
-Moreover, [Har79, Proposition 2.11] implies that $\mathcal{S}^{\perp}$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$, and $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$.
+Moreover, [Har79, Proposition 2.11] implies that $\mathcal{S}^{\perp}$ is a sub-hypergroup of $^{\widehat{(H,\mathcal{B})}}$, and $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$.
 Corollary 7.15.
-Let $H$ be an abelian dualizable RN-hypergroup. With the above notations,
+Let $H$ be a dualizable ARN-hypergroup. With the above notations,
 ${G(\widehat{H})}^{\perp}=H_{ad}\text{ and }G(H)^{\perp}={(\widehat{H})_{ad}}.$
+Thus, by $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$,
+${G(\widehat{H})}=(H_{ad})^{\perp}\text{ and }G(H)={(\widehat{H})_{ad}}^{\perp}.$
 Proof.
-Consider the trivial character $\psi_{1}=\mu_{1}\big{|}_{G(\widehat{H})}$. It follows that
+Consider the trivial character $\psi_{1}=x_{1}\big{|}_{G(\widehat{H})}$. It follows that
 ${H_{ad}}=B_{\psi_{1}}$ $=$ $\{x_{i}\;|\;\mu_{j}({\frac{x_{i}}{d_{i}}})=1\;\text{for all}\;\mu_{j}\in{G(\widehat{H})}\}$
 $=$ $\bigcap_{\mu_{j}\in{G(\widehat{H})}}\ker_{\widehat{H}}(\mu_{j})={G(\widehat{H})}^{\perp}.$
-By duality, one obtains the other equality. ∎
+By duality, we obtain the second equality. ∎
 7.6. Quotients of hypergroups
@@
 This is the analogue of the quotient construction from [Har79], in the case of probability groups.
-Recall that a probability group is a symmetric normalized RN-hypergroup.
-One can define the following equivalence relation on $\mathcal{B}$.
+Recall that a probability group is a normalized RN-hypergroup.
+We can define the following equivalence relation on $\mathcal{B}$.
 For two elements $a,b\in\mathcal{B}$, say $a\sim_{\mathcal{S}}b$ if and only if there are $s_{1},s_{2}\in{\mathcal{S}},x\in{\mathcal{B}}$ such that $m(x,as_{1})>0$ and $m(x,s_{2}b)>0$.
@@
 Define $H//{\mathcal{S}}$ as the set of all these equivalence classes of $\sim_{\mathcal{S}}$.
-Recall that $\lambda_{\mathcal{S}}$ is the primitive central idempotent of ${\mathcal{S}}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
-For an element $a\in\mathcal{B}$, one can see that $[a]_{\mathcal{S}}=[b]_{\mathcal{S}}$ if and only if $\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}=\lambda_{\mathcal{S}}b\lambda_{\mathcal{S}}$.
+Recall that $\lambda_{\mathcal{S}}$ is the central idempotent of ${\mathcal{S}}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
+For two elements $a,b\in\mathcal{B}$, we can see that $[a]_{\mathcal{S}}=[b]_{\mathcal{S}}$ if and only if $\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}=\lambda_{\mathcal{S}}b\lambda_{\mathcal{S}}$.
 Therefore there is a set bijection
 $\begin{array}[]{cccc}\phi:&\mathbb{C}[H//{\mathcal{S}}]&\to&\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}\\ &[a]_{\mathcal{S}}&\mapsto&\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}\end{array}$
-Then, it easy to verify that $H//{\mathcal{S}}$ becomes a ARN-hypergroup, with the multiplication inherited from $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, via the above isomorphism.
+Then, it is easy to verify that $H//{\mathcal{S}}$ becomes a RN-hypergroup, with the multiplication inherited from $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, via the above isomorphism.
 We denote by $\overline{m}({[c]},\;[a][b])$ the multiplicity structure of $H//{\mathcal{S}}$.
@@
 $[a][b]=\sum_{[c]\in H//{\mathcal{S}}}\overline{m}({[c]},\;[a][b])[c].$
-We write shortly $[a]$ instead of $[a]_{\mathcal{S}}$ when no confusion is possible. Following [Har79], if ${(H,\mathcal{B})}$ is abelian, one can show that
+We write shortly $[a]$ instead of $[a]_{\mathcal{S}}$ when no confusion is possible. Following [Har79], if ${(H,\mathcal{B})}$ is abelian, we can show that
 (7.5) $\overline{m}({[c]},\;[a][b])=\sum_{w\in[c]}m(w,ab).$
@@
 7.7. Applying Harrison’s results
-Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup.
-In this subsection, we also use the short notation $H$ (or $\mathcal{B}$) for this hypergroup, if $\mathcal{B}$ (or $H$) are implicitly understood.
-For any sub-hypergroup ${(L,S)}$ of an ANR-hypergroup ${(H,\mathcal{B})}$, note that [Har79, Proposition 2.11] implies that there is an isomorphism of hypergroups:
+In this subsection, we also use the short notation $H$ (or $\mathcal{B}$) for a hypergroup $(H,\mathcal{B})$, if $\mathcal{B}$ (or $H$) are implicitly understood.
+For any sub-hypergroup ${(L,\mathcal{S})}$ of a dualizable ANR-hypergroup ${(H,\mathcal{B})}$, note that [Har79, Proposition 2.11] implies that there is an isomorphism of hypergroups:
 $\widehat{H}//\mathcal{S}^{\perp}\simeq\widehat{\mathcal{S}},\;\;[\mu]_{\mathcal{S}^{\perp}}\mapsto\mu\big{|}_{L}.$
@@
 (7.6) $\widehat{H}//G(\widehat{H})\simeq\widehat{H_{ad}}$
-Dually, for ${\mathcal{S}}={(\widehat{H})_{ad}}$, one obtains that
+Dually, for ${\mathcal{S}}={(\widehat{H})_{ad}}$, we obtain that
 (7.7) $H//G(H)\simeq\widehat{{(\widehat{H})_{ad}}}$
@@
 Definition 8.2.
-Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian RN-hypergroup ${(H,\mathcal{B})}$.
-We define the commutator ${(L,\;\mathcal{S})}^{{co}}$ of ${(L,\;\mathcal{S})}$ in ${(H,\mathcal{B})}$ as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by all standard elements $x\in\mathcal{B}$ such that $xx^{*}\in\mathcal{{\mathcal{S}}}$.
+Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an ARN-hypergroup ${(H,\mathcal{B})}$.
+Let ${\mathcal{S}}^{co}$ be the set of standard elements $x\in\mathcal{B}$ such that $xx^{*}\in\mathcal{{\mathcal{S}}}$.
+The commutator ${(L,\mathcal{S})}^{{co}}$ of ${(L,\mathcal{S})}$ in ${(H,\mathcal{B})}$ as the sub-hypergroup generated by ${\mathcal{S}}^{co}$.
 As in [GN08, Remark 4.9], it follows that the linear span of ${\mathcal{S}}^{co}$ is already a sub-hypergroup of $H$.
@@
 (8.1) $(\mathcal{S}^{{co}})_{ad}\subseteq\mathcal{S}\subseteq(\mathcal{S}_{ad})^{{co}}$
-for any sub-hypergroup $\mathcal{S}$ of $\mathcal{B}$. By applying $\perp$ to (8.1), one obtains
+for any sub-hypergroup $\mathcal{S}$ of $\mathcal{B}$. By applying $\perp$ to (8.1), we obtain
 (8.2) $\big{(}(\mathcal{S}^{{co}})_{ad}\big{)}^{\perp}\supseteq\mathcal{S}^{\perp}\supseteq\big{(}(\mathcal{S}_{ad})^{{co}}\big{)}^{\perp}.$
-For an ARN-hypergroup ${(H,\mathcal{B})}$, define $H_{(0)}=\mathbb{C}$ and $H_{(n)}=(H_{(n-1)})^{{co}}$, for all $n\geq 1$. Then, $H_{(1)}=\mathbb{C}^{{co}}=G({(H,\mathcal{B})})=H_{pt}$.
+For an ARN-hypergroup ${(H,\mathcal{B})}$, define $H_{(0)}=\mathbb{C}$ and $H_{(n)}=(H_{(n-1)})^{{co}}$, for all $n\geq 1$. Then,
+$H_{(1)}=\mathbb{C}^{{co}}=G{(H,\mathcal{B})}=H_{pt}.$
 Definition 8.3.
@@
 will be called the lower central series of $H$.
-Similarly to [GN08, Theorem 4.16], one can prove that $H^{(n)}=\mathbb{C}$ if and only if $H_{(n)}=H$, where $H$ is an ARN-hypergroup.
+Similarly to [GN08, Theorem 4.16], we can prove that $H^{(n)}=\mathbb{C}$ if and only if $H_{(n)}=H$, where $H$ is an ARN-hypergroup.
 Definition 8.4.
-An ARN-hypergroup is called nilpotent when $H^{(n)}=\mathbb{C}$, for some $n\in\mathbb{N}$.
+An ARN-hypergroup is termed nilpotent if there exists some $n\in\mathbb{N}$ such that $H^{(n)}=\mathbb{C}$.
+The smallest such $n$ for which this condition is satisfied is referred to as the nilpotency class of $H$.
 Lemma 8.5.
@@
 Proof.
-It reduces to show that $\mu_{j}\mu_{j^{\#}}\in({\mathcal{S}}^{co})^{\perp}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$.
+It reduces to show that $\mu_{j}\star\mu_{j^{\#}}\in({\mathcal{S}}^{co})^{\perp}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$.
 Let $x\in{\mathcal{S}}^{co}$ be a standard element.
@@
 Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. With the above notations, for all $n\geq 0$,
-(8.4) $\widehat{H}^{(n)}\subseteq H_{(n)}^{\perp}$
+(8.4) $\widehat{H}^{(n)}\subseteq(H_{(n)})^{\perp}$
 and
@@
 We will prove the first inclusion by induction on $n$. If $n=0$, then
-$\widehat{H}^{(0)}=H=\mathbb{C}^{\perp}=H_{(0)}^{\perp}.$
+$\widehat{H}^{(0)}=\widehat{H}=\mathbb{C}^{\perp}=H_{(0)}^{\perp}.$
 Now, suppose that ${\widehat{H}^{(n)}}\subseteq{H_{(n)}}^{\perp}$. Then
 $\widehat{H}^{(n+1)}=({\widehat{H}^{(n)}})_{ad}\subseteq({H_{(n)}}^{\perp})_{ad}\subseteq({H_{(n)}}^{co})^{\perp}=H_{(n+1)}^{\perp}.$
-By duality, one obtains the second inclusion. ∎
+By duality, we obtain the second inclusion. ∎
 Observe that the equality holds for $n=1$ also, because
@@
 If ${(H,\mathcal{B})}$ is nilpotent then $H_{(n)}=H$, for some $n$.
-Then, $\widehat{H}^{(n)}=\mathbb{C}$ by Equation (8.4), and therefore ${\widehat{(H,\mathcal{B})}}$ is nilpotent.
+Then, $\widehat{H}^{(n)}=\mathbb{C}$ by Equation (8.4), and therefore $^{\widehat{(H,\mathcal{B})}}$ is nilpotent.
 Idem for the converse with Equation (8.5).
 8.2. Burnside property for nilpotent ARN-hypergroups
-Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Recall that we call $x_{i}\in\mathcal{B}$ a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Recall that we call $x_{i}\in\mathcal{B}$ a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
 Lemma 8.7.
-Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup and ${(L,\;\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$.
+Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and ${(L,\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$.
 Consider a standard element $x_{i}\in{\mathcal{B}}$ such that $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$.
-Then, $x_{i}$ is also vanishing element in ${(H,\mathcal{B})}$.
+Then, $x_{i}$ is also a vanishing element in ${(H,\mathcal{B})}$.
 Proof.
 Recall that $H//{\mathcal{S}}\simeq\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ as hypergroups, with $[x]\mapsto\lambda_{\mathcal{S}}x\lambda_{\mathcal{S}}$.
-Moreover, the hypergroup $\widehat{H//{\mathcal{S}}}$ can be identified with a subring of $\widehat{H}$.
-If $\lambda_{\mathcal{S}}=\sum_{j\in\mathcal{J}_{\mathcal{S}}}F_{j}$, then $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}=\oplus_{j\in\mathcal{J}_{\mathcal{S}}}\mathbb{C}F_{j}$, as a subalgebra of $H$.
-Thus, the characters $\mu_{j}$, with $j\in\mathcal{J}_{\mathcal{S}}$, have distinct restriction to $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, and cover all the characters of $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$.
-So if $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$, then $\lambda_{\mathcal{S}}x_{i}\lambda_{\mathcal{S}}$ is a vanishing element in $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, which in turn implies that $x_{i}$ is a vanishing element in $H$.
+If $\lambda_{\mathcal{S}}=\sum_{j\in\mathcal{I}_{\mathcal{S}}}F_{j}$ (Definition 6.3), then $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}=\oplus_{j\in\mathcal{I}_{\mathcal{S}}}\mathbb{C}F_{j}$, as a subalgebra of $H$.
+Thus, the characters $\mu_{j}$, with $j\in\mathcal{I}_{\mathcal{S}}$, have distinct restriction to $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ (since $\mu_{i}(F_{j})=\delta_{i,j}$), and so cover all the characters of $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ (since $\dim_{\mathbb{C}}(\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}})=|I_{\mathcal{S}}|$).
+If $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$, then $\lambda_{\mathcal{S}}x_{i}\lambda_{\mathcal{S}}$ is a vanishing element in $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, thus by above, $x_{i}$ is a vanishing element in $H$.
 ∎
@@
 Let ${(H,\mathcal{B})}$ be an ARN-hypergroup.
-A grouplike element $g\in{G{(H,\mathcal{B})}}$ is a constituent of $x_{i}x_{i^{*}}$ if and only if $gx_{i}=\mathrm{FPdim}(g)x_{i}$.
-The set of all grouplike elements that are constituents of $x_{i}x_{i^{*}}$ form a subgroup of ${G{(H,\mathcal{B})}}$.
+A grouplike element $x_{i}\in{G{(H,\mathcal{B})}}$ is a constituent of $x_{j}x_{j^{*}}$ if and only if $x_{i}x_{j}=\mathrm{FPdim}(x_{i})x_{j}$.
+The set of all grouplike elements that are constituents of $x_{j}x_{j^{*}}$ form a subgroup of ${G{(H,\mathcal{B})}}$.
 Proof.
-Equation (5.1) gives that
-$m(g,x_{i}x_{i}^{*})=\tau(x_{i}x_{i^{*}}g^{*})=\tau(x_{i^{*}}g^{*}x_{i})=m(x_{i^{*}},x_{i^{*}}g^{*})=m(x_{i},gx_{i}).$
-Thus, by Lemma 2.27 one has $g^{-1}x_{i}=\mathrm{FPdim}(g^{-1})x_{i}$ if $m(g,x_{i}x_{i}^{*})>0$. The second assertion is a straightforward consequence of this. ∎
-Proof of Theorem 1.12 It is enough to prove the following:
+By Frobenius reciprocity (2.18), $N_{j,j^{*}}^{i}=\frac{h_{j}}{h_{i}}N_{i,j}^{j}$, thus $x_{i}$ is a constituent of $x_{j}x_{j^{*}}$ if and only if $x_{j}$ is a constitutent of $x_{i}x_{j}$, if and only if (by Lemma 2.34) $\frac{x_{i}}{d_{i}}\frac{x_{j}}{d_{j}}=\frac{x_{j}}{d_{j}}$, meaning that $x_{i}x_{j}=\mathrm{FPdim}(x_{i})x_{j}$.
+The second assertion is a straightforward consequence of this.
+∎
+Proof of Theorem 1.15 By Theorem 1.14, it is enough to prove the following:
 Proposition 8.9.
@@
 We will proceed by induction on the nilpotency class of ${(H,\mathcal{B})}$.
-If ${(H,\mathcal{B})}$ is of nilpotency class $1$, then ${(H,\mathcal{B})}$ is pointed and we are done.
+If ${(H,\mathcal{B})}$ is of nilpotency class 111, then ${(H,\mathcal{B})}$ is pointed and we are done.
 Now, suppose that ${(H,\mathcal{B})}$ is nilpotent of class $n$ and that the result holds for the nilpotency classes less than $n$.
@@
 Consider $[x_{i}]\in H//G(H)$.
-Since $H//G(H)\simeq{(\widehat{H})_{ad}}$, it is nilpotent, with nilpotency class strictly less that of $H$, by Theorem 1.11.
+Since $H//G(H)\simeq\widehat{{(\widehat{H})_{ad}}}$ by (7.7), it is nilpotent, with nilpotency class $n-1$, by Theorem 1.14.
 By induction hypothesis, we know that $[x_{i}]$ is either a grouplike element or a vanishing element in $H//G(H)$.
 If $[x_{i}]$ is a vanishing element, then Lemma 8.7 shows that $x_{i}$ is a vanishing element.
-Now, if $[x_{i}]$ is a grouplike element of $H//G(H)$, then
-$\lambda_{G(H)}h_{[x_{i}]}x_{i}x_{i}^{*}\lambda_{G(H)}=\lambda_{G(H)},$
+Now, if $[x_{i}]$ is a grouplike element of $H//G(H)$ and by abelian assumption,
+$\lambda_{G(H)}=h_{[x_{i}]}\lambda_{G(H)}x_{i}\lambda_{G(H)}x_{i}^{*}\lambda_{G(H)}=\lambda_{G(H)}h_{[x_{i}]}x_{i}x_{i}^{*}\lambda_{G(H)},$
 which implies that $h_{[x_{i}]}x_{i}x_{i}^{*}\in G(H)$.
-Thus, $h_{[x_{i}]}x_{i}x_{i}^{*}=\sum_{g\in G_{1}}g$, for a subgroup $G_{1}\subset G$.
+Thus, $h_{[x_{i}]}x_{i}x_{i}^{*}=\sum_{g\in G_{1}}g$, for a subgroup $G_{1}\subset G$ (by Lemma 8.8).
 Since $x_{i}$ is not a grouplike element in $H$, $G_{1}\neq\{1\}$.
@@
 Let $\mathcal{C}$ be a pivotal fusion category with a commutative Grothendieck ring.
-As in §1, we denote the set of isomorphism classes of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$ and let :=$\mathcal{I}:=\{1,\dots,m\}$.
-Let also $d_{i}:=\dim(X_{i})$ be the quantum dimension of $X_{i}$ for all $i\in\mathcal{I}$.
+As in §1, we denote the set of isomorphism class representatives of simple objects of $\mathcal{C}$ by $\mathrm{Irr}(\mathcal{C}):=\{X_{1},\dots,X_{m}\}$ and let :=$\mathcal{I}:=\{1,\dots,m\}$.
+Let also $d_{i}:=\dim(X_{i})$ be the categorical dimension of $X_{i}$ for all $i\in\mathcal{I}$.
 For any simple object $X_{i}\in\mathrm{Irr}(\mathcal{C})$ denote by $x_{i}:=[X_{i}]$ the class of $X_{i}$ in the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
@@
 The forgetful functor $F:{\mathcal{Z}(\mathcal{C})}\rightarrow\mathcal{C}$ admits a right adjoint functor $R:\mathcal{C}\rightarrow{\mathcal{Z}(\mathcal{C})}$.
-It is known that $A:=R(\mathbf{1})$ has the structure of a central commutative algebra in $\mathcal{Z}(\mathcal{C})$, where $\mathbf{1}$ is the unit object of $A$.
-The vector space $\mathrm{{CE}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(\mathbf{1},A)$ is called the set of central elements.
-The space ${\mathrm{CF}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(A,\mathbf{1})$ is called the space of class functions of $\mathcal{C}$.
-Recall [Shi17] that ${\mathrm{CF}(\mathcal{C})}\simeq K(\mathcal{C})$, the complexification of the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
-For any simple object $X_{i}$ of $\mathcal{C}$ we denote by $\chi_{i}:=\mathrm{ch}(X_{i})\in{\mathrm{CF}(\mathcal{C})}$ its associated character, [Shi17].
+It is well-known that $A:=FR(\mathbf{1})$ has the structure of a central commutative algebra in $\mathcal{C}$ (meaning that $R(\mathbf{1})$ is a commutative algebra in ${\mathcal{Z}(\mathcal{C})}$), where $\mathbf{1}$ is the unit object of $\mathcal{C}$.
+The vector spaces
+$\mathrm{{CE}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(\mathbf{1},A)\text{ and }{\mathrm{CF}(\mathcal{C})}:=\mathrm{{Hom}}_{\mathcal{C}}(A,\mathbf{1})$
+are respectively called the space of central elements and the space of class functions of $\mathcal{C}$.
+Recall [Shi17] that ${\mathrm{CF}(\mathcal{C})}\simeq K(\mathcal{C})_{\mathbb{C}}$, the complexification of the Grothendieck ring $K(\mathcal{C})$ of $\mathcal{C}$.
+For any simple object $X_{i}$ of $\mathcal{C}$ we denote by $\chi_{i}:=\mathrm{ch}(X_{i})\in{\mathrm{CF}(\mathcal{C})}$ its associated character [Shi17, Definition 3.9].
 The central element space $\mathrm{{CE}(\mathcal{C})}$ has a basis of primitive orthogonal idempotents $E_{i}$ such that $<\chi_{i},E_{j}>=\delta_{i,j}d_{i}$ for all $1\leq i,j\leq m$.
@@
 where ${\mathrm{CF}(\mathcal{C})}$ is the algebra of class functions on $\mathcal{C}$.
-Since ${\mathcal{Z}(\mathcal{C})}$ is also fusion category we can write $R(\mathbf{1})=\bigoplus_{j=1}^{m}\mathcal{C}^{j}$ as a direct sum of simple objects in ${\mathcal{Z}(\mathcal{C})}$.
+Since ${\mathcal{Z}(\mathcal{C})}$ is also a fusion category we can write $R(\mathbf{1})=\bigoplus_{j\in\mathcal{I}}\mathcal{C}^{j}$ as a direct sum of simple objects in ${\mathcal{Z}(\mathcal{C})}$.
 Recall that $\mathcal{C}^{j}$ are called conjugacy classes for $\mathcal{C}$.
-The above isomorphism also gives a canonical bijection between the set of central primitive idempotents $\{F_{j}\}_{j=1}^{m}$ and the set of conjugacy classes $\{\mathcal{C}^{j}\}$.
-We denote as above $\mathcal{J}:=\{1,\dots m\}$.
-For any fusion category $\mathcal{C}$, by abuse of notation, the symmetric RN-hypergroup ${(H,\mathcal{B},\mu_{1})}:=(K(\mathcal{C}),\mathrm{Irr}(\mathcal{C}),\mathrm{FPdim})$ is also denoted by $K(\mathcal{C})$.
-We denote by ${\widehat{{K(\mathcal{C})}}}$ its dual hypergroup.
+The above isomorphism also gives a canonical bijection between the set of primitive central idempotents $\{F_{j}\}_{j\in\mathcal{I}}$ and the set of conjugacy classes $\{\mathcal{C}^{j}\}$.
+For any fusion category $\mathcal{C}$, by abuse of notation, the RN-hypergroup ${(H,\mathcal{B},\mu_{1})}:=(K(\mathcal{C}),\mathrm{Irr}(\mathcal{C}),\mathrm{FPdim})$ is also denoted by $K(\mathcal{C})$.
+We denote by $^{\widehat{{K(\mathcal{C})}}}$ its dual hypergroup.
 Recall also from [GN08] the notion of a universal group grading of a fusion category $\mathcal{C}$ and its adjoint subcategory ${\mathcal{C}_{\mathrm{ad}}}$.
@@
 (9.2) $\widehat{h}_{j}=\dim(\mathcal{C}^{j})$
-in the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$.
-In particular, for a weakly-integral fusion category $\mathcal{C}$, by [ENO05, Proposition 8.27], $\dim(\mathcal{C}^{j})\in\mathbb{Z}$, and therefore ${\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
+in the dual hypergroup $^{\widehat{{K(\mathcal{C})}}}$.
+In particular, for a weakly-integral fusion category $\mathcal{C}$, by [ENO05, Proposition 8.27], $\dim(\mathcal{C}^{j})\in\mathbb{Z}$, and therefore $^{\widehat{{K(\mathcal{C})}}}$ is $h$-integral.
 Corollary 9.2.
-Let $\mathcal{C}$ be a a fusion category with a Burnside commutative Grothendieck ring and let ${\widehat{{K(\mathcal{C})}}}$ be its dual. Then
-(9.3) $\prod_{j=1}^{m}\mu_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
-where ${\mathrm{sgn}}(x_{i})$ is the determinant of the permutation matrix determined by $L_{{\frac{x_{i}}{d_{i}}}}$ on $K(\mathcal{C})$.
+Let $\mathcal{C}$ be a a fusion category with a Burnside commutative Grothendieck ring and let $^{\widehat{{K(\mathcal{C})}}}$ be its dual. Then
+(9.3) $\prod_{j=1}^{m}\mu_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})\widehat{F}_{i}$
+where ${\mathrm{sgn}}(x_{i})$ is the determinant of the permutation matrix $L_{{\frac{x_{i}}{d_{i}}}}$ on $K(\mathcal{C})$.
 Proof.
 Since $K(\mathcal{C})$ is a Burnside hypergroup the result follows from Corollary 4.4. ∎
-Note that the dual hypergroup ${\widehat{{K(\mathcal{C})}}}$ is denoted by $\widehat{{\mathrm{CF}(\mathcal{C})}}$ in [Bur21].
-Theorem 3.4 from the same paper implies that for any pseudo-unitary fusion category there is a canonical isomorphism of hypergroups
-(9.4) $\alpha:\widehat{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})},\;\mu_{j}\mapsto\frac{C_{j}}{\dim(\mathcal{C}^{j})}.$
-For a pivotal fusion category, Shimizu has defined $C_{j}:={\mathcal{F}_{\lambda}}^{-1}(F_{j})\in\mathrm{{CE}(\mathcal{C})}$ to be the conjugacy class sums corresponding to the conjugacy class $\mathcal{C}^{j}$.
+Note that the dual hypergroup $^{\widehat{{K(\mathcal{C})}}}$ is denoted by $\widehat{{\mathrm{CF}(\mathcal{C})}}$ in [Bur21].
+Theorem 3.4 from the same paper implies that for any pivotal fusion category there is a canonical isomorphism of $\mathbb{C}$-algebra
+(9.4) $\alpha:\widehat{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})},\;\mu_{j}\mapsto\frac{C_{j}}{\dim(\mathcal{C}^{j})},$
+where $C_{j}\in\mathrm{{CE}(\mathcal{C})}$ is the conjugacy class sum corresponding to the conjugacy class $\mathcal{C}^{j}$, defined for a pivotal fusion category by Shimizu as ${\mathcal{F}}^{-1}(F_{j})$.
 Here $\lambda\in{\mathrm{CF}(\mathcal{C})}$ is a cointegral of $\mathcal{C}$ such that $\langle\lambda,u\rangle=1$, see [Shi17, §5].
@@
 (9.5) $\mathcal{F}_{\lambda}:\mathrm{{CE}(\mathcal{C})}\rightarrow{\mathrm{CF}(\mathcal{C})}\;\;\text{given by}\;\;a\mapsto\lambda\leftharpoonup\mathcal{S}(a)$
-Since any weakly-integral fusion category is pseudo-unitary, by applying the canonical morphism $\alpha$ one obtains the following:
+By [Bur20, Lemma 4.1 and (4.7)], for any spherical fusion category $\mathcal{C}$ over $\mathbb{C}$ with a commutative Grothendieck ring, we have
+$\langle F_{i},C_{j}\rangle=\delta_{i,j}\dim(\mathcal{C}^{j}).$
+This relation implies that the set $(F_{i})$ forms a dual basis to $(C_{j}/\dim(\mathcal{C}^{j}))$ with respect to the bilinear form $\langle$,$\rangle$ , ⟩.
+Consequently, this provides an easier way for defining the conjugacy class sums $(C_{j})$ in this case (see [Bur20, (4.11)] for a formula).
 Corollary 9.3.
-For any weakly-internal fusion category $\mathcal{C}$ the following identity holds in $\mathrm{{CE}(\mathcal{C})}$:
-(9.6) $\prod_{j=1}^{m}C_{j}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})E_{i}$
+For any weakly-integral fusion category $\mathcal{C}$ with a commutative Grothendieck ring the following identity holds in $\mathrm{{CE}(\mathcal{C})}$:
+(9.6) $\prod_{j=1}^{m}\frac{C_{j}}{\dim(\mathcal{C}^{j})}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})E_{i}$
 where $E_{i}\in\mathrm{{CE}(\mathcal{C})}$ is the primitive central idempotent of $X_{i}\in{\mathcal{C}_{\mathrm{pt}}}$.
+Proof.
+By Corollary 4.21, we can apply Corollary 9.2. The result follows by applying the canonical isomorphism $\alpha$. ∎
 For a spherical fusion category note the following:
 Proposition 9.4.
-Let $\mathcal{C}$ be a spherical fusion category with the Grothendieck ring ${K(\mathcal{C})}$ commutative.
-Let also $\mu_{j}\in{\widehat{{K(\mathcal{C})}}}$ a character of ${K(\mathcal{C})}$.
-Suppose that ${\widehat{{K(\mathcal{C})}}}$ is a RN-hypergroup.
-Then $\mu_{j}$ is a grouplike element of ${\widehat{{K(\mathcal{C})}}}$ if and only if $\dim(\mathcal{C}^{j})=1$.
+Let $\mathcal{C}$ be a spherical fusion category with a commutative Grothendieck ring ${K(\mathcal{C})}$ and RN dual.
+Let $\mu_{j}\in{\widehat{{K(\mathcal{C})}}}$ be a linear character.
+Then $\mu_{j}$ is a grouplike element of $^{\widehat{{K(\mathcal{C})}}}$ if and only if $\dim(\mathcal{C}^{j})=1$.
 Proof.
-By Lemma 2.17, $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}d_{j}^{2}=1$.
-Since ${\widehat{{K(\mathcal{C})}}}$ is normalized it follows that $d_{j}=1$ for all $j\in\mathcal{J}$.
-Thus $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}=1$.
-On the other hand, it follows from Equation (9.2) that ${\widehat{h}}_{j}=1$ if and only if $\mathcal{C}^{j}$ is a one dimensional conjugacy class of $\mathcal{C}$.
+By Lemma 2.24 and the fact that $^{\widehat{{K(\mathcal{C})}}}$ is normalized, $\mu_{j}$ is a grouplike element if and only if ${\widehat{h}}_{j}=1$.
+This is also equivalent to $\dim(\mathcal{C}^{j})=1$, by Equation (9.2).
 ∎
-Suppose now that $\mathcal{C}$ is a pivotal fusion category and ${\mathcal{D}}\subseteq\mathcal{C}$ is a fusion subcategory. There is a subset $\mathcal{J}_{\mathcal{D}}\subseteq\mathcal{J}$ such that
-(9.7) $\lambda_{{\mathcal{D}}}=\sum_{j\in\mathcal{J}_{\mathcal{D}}}F_{j}$
-since $\lambda_{\mathcal{D}}$ is an idempotent element inside ${\mathrm{CF}(\mathcal{C})}$.
-Note that this set is denoted by $\mathcal{L}_{\mathcal{D}}$ in [Bur20, §4.2].
-Also, this is the corresponding notion of the support from Definition 6.3.
-Note that [Bur20, Lemma 4.6 ] implies
-(9.8) $J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}=\{j\in\mathcal{J}\;|\;\dim(\mathcal{C}^{j})=1\},$
+Now, suppose that $\mathcal{C}$ is a pivotal fusion category and ${\mathcal{D}}\subseteq\mathcal{C}$ is a fusion subcategory.
+Following the notion of support defined in Definition 6.3, there exists a subset $\mathcal{I}_{\mathcal{D}}\subseteq\mathcal{I}$ such that
+$\lambda_{{\mathcal{D}}}=\sum_{j\in\mathcal{I}_{\mathcal{D}}}F_{j},$
+Please note that this set is called $\mathcal{L}_{\mathcal{D}}$ in [Bur20, §4.2]. Additionally, [Bur20, Lemma 4.6] implies
+(9.7) ${\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}=\{j\in\mathcal{I}\;|\;\dim(\mathcal{C}^{j})=1\},$
 for any fusion category with a commutative Grothendieck ring ${K(\mathcal{C})}$.
 Proposition 9.5.
-Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then ${\widehat{{K(\mathcal{C})}}}$ is Burnside if and only if:
-(9.9) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{j\in J_{{\mathcal{C}_{\mathrm{ad}}}}}{\mathrm{sgn}}(\mu_{j})F_{j}$
+Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring. Then $^{\widehat{{K(\mathcal{C})}}}$ is Burnside if and only if:
+(9.8) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{j\in{{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}{\mathrm{sgn}}(\mu_{j})F_{j}$
 Proof.
@@
 Proof.
-Suppose that the dual ${\widehat{{K(\mathcal{C})}}}$ is Burnside. Squaring Equation (9.9) it follows that
-$\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\sum_{j\in J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)}.$
-The converse also follows immediately from Proposition 9.5. ∎
-Let us provide a dual statement in the group case:
+It follows from Corollary 4.9 and Theorem (6.5), together with
+$\sum_{j\in{{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\frac{1}{\dim({\mathcal{C}_{\mathrm{ad}}})}\big{(}\sum_{x_{i}\in{\mathcal{C}_{\mathrm{ad}}}}d_{i}x_{i}\big{)},$
+where the last equality comes from Equation (5.2). ∎
+Here is a reformulation of Theorem 1.12 for the case of a nilpotent finite group:
 Theorem 9.6.
-For any finite nilpotent group $G$ the dual $\widehat{K(G)}$ is Burnside and therefore
+For any finite nilpotent group $G$, the dual $\widehat{\text{ch}(G)}$ is Burnside and therefore
 (9.9) $\left(\prod_{x_{i}\in\mathrm{Irr}(G)}{\frac{x_{i}}{d_{i}}}\right)^{2}=\frac{|Z(G)|}{|G|}\left(\sum_{x_{i}\in\mathrm{Irr}(G/Z(G))}d_{i}x_{i}\right).$
 Proof.
-Let $\mathcal{C}=\mathrm{Rep}(G)$.
-It is easy to verify that $\widehat{K(G)}$ is a real-non negative hypergroup.
-For, example see [Bur21, Equation (3.12)].
-Then Theorem B of [INW99] says that if $G$ is a nilpotent group then ${K(G)}$ is dual-Burnside.
+Recall that $K(\text{Rep}(G))$ is the character ring $\text{ch}(G)$.
+The result that $\widehat{\text{ch}(G)}$ is a RN hypergroup is established in [Bur21, Equation (3.12)].
+Furthermore, [INW99, Theorem B] states that if $G$ is a nilpotent group, then $\text{ch}(G)$ is dual-Burnside.
 ∎
-Theorem 1.12 implies the following:
+Theorem 1.15 leads to the following conclusion:
 Corollary 9.7.
-Suppose that $\mathcal{C}$ is a nilpotent fusion category. Then ${K(\mathcal{C})}$ and ${\widehat{{K(\mathcal{C})}}}$ are both Burnside.
+If $\mathcal{C}$ is a nilpotent fusion category with a commutative Grothendieck ring and RN dual, then both ${K(\mathcal{C})}$ and $^{\widehat{{K(\mathcal{C})}}}$ are Burnside.
+Remark 9.8.
+A nilpotent fusion category is weakly-integral [GN08], which means it is also pseudo-unitary and spherical [EGNO15].
+It remains an open question whether every pseudo-unitary fusion category can be given a unitary structure [EGNO15, Remark 9.4.7].
+However, a commutative fusion ring with a unitary categorification does have a RN dual [LPW21].
+Therefore, we believe that the assumption of a RN dual in Corollary 9.7 could be omitted.
+Additionally, in the braided case, its Grothendieck ring possesses a RN dual according to [Bur21, Theorem 1.2].
 10. Premodular categories
-Recall that a premodular category is a braided and spherical fusion category.
-For the rest of this section suppose that $\mathcal{C}$ is a pseudo-unitary premodular category.
+Recall that a premodular category is defined as a braided spherical fusion category.
+In the rest of this section, we will assume that $\mathcal{C}$ is a premodular category.
 By [Shi17, Example 6.14] there is $\mathbb{C}$-algebra map $\mathrm{f}_{Q}:{{\mathrm{CF}(\mathcal{C})}}\rightarrow\mathrm{{CE}(\mathcal{C})}$ given by the following formula:
-(10.1) $\mathrm{f}_{Q}(\chi_{i})=\sum_{i^{\prime}=1}^{m}\frac{s_{ii^{\prime}}}{d_{i^{\prime}}}E_{i^{\prime}}.$
-where $S=(s_{ij})$ is the $S$-matrix of $\mathcal{C}$ and $E_{i^{\prime}}$ are the primitive central idempotents of $\mathrm{{CE}(\mathcal{C})}$ as defined in the previous section.
-As in [Bur22, §4] there is a braided function $M:\mathcal{I}\rightarrow\mathcal{J}$ such that if
-$\mathrm{f}_{Q}(F_{j})=\sum_{i\in\mathcal{B}_{j}}E_{i}$
-[…]
-Recall from [BB15] the notion of cosets of a fusion category with respect to a fusion subcategory.
-By [Bur22, Theorem 4.10] two simple objects $X_{i},X_{i^{\prime}}$ of a pseudo-unitary premodular fusion category $\mathcal{C}$ are in the same coset with respect to $\mathcal{Z}_{2}(\mathcal{C})$ if and only if $M(i)=M(i^{\prime})$.
-For any $j\in\mathcal{J}_{2}$ we denote by $\mathcal{R}_{j}:=\{[X_{i}]\;|\;M(i)=j\}$.
-The aforementioned theorem implies also that $\mathcal{R}_{j}$ with $j\in\mathcal{J}_{2}$ are exactly the right cosets of $\mathcal{C}$ with respect to $\mathcal{C}^{\prime}:=\mathcal{Z}_{2}(\mathcal{C})$.
+(10.1) $\mathrm{f}_{Q}(\chi_{i})=\sum_{j\in\mathcal{I}}\frac{s_{ij}}{d_{j}}E_{j}.$
+where $S=(s_{ij})$ is the $S$-matrix of $\mathcal{C}$ and $(E_{j})$ are the primitive central idempotents of $\mathrm{{CE}(\mathcal{C})}$ as defined in the previous section.
+As described in [Bur22, §4], there is a map $M:\mathcal{I}\to\mathcal{I}$ (referred to as the braided partition function) such that if
+$\mathrm{f}_{Q}(F_{j})=\sum_{i\in\mathcal{B}_{j}}E_{i},$
+then $M(i)=j$.
+Let $\mathcal{I}_{2}\subseteq\mathcal{I}$ denote the set of all indices $j$ for which $\mathrm{f}_{Q}(F_{j})\neq 0$, meaning that $\mathcal{B}_{j}$ is not an empty set.
+Given that $\mathrm{f}_{Q}(1)=1$, we can thus partition the set of all isomorphism classes of simple objects $\text{Irr}(\mathcal{C})=\bigsqcup_{j\in\mathcal{I}_{2}}\mathcal{R}_{j}$, where $\mathcal{R}_{j}=\{[X_{i}]\;|\;i\in\mathcal{B}_{j}\}$.
+In other words, $\mathcal{R}_{j}=\{[X_{i}]\;|\;M(i)=j\}$.
+Thus, we obtain a unique function $M:\mathcal{I}\to\mathcal{I}_{2}$ with the property that $E_{i}\mathrm{f}_{Q}(F_{M(i)})\neq 0$ for all $i\in\mathcal{I}$.
+The paper [BB15] introduces the notion of cosets of a fusion category $\mathcal{C}$ with respect to a fusion subcategory $\mathcal{D}$: two simple objects $X,Y$ in $\mathcal{C}$ are in the same (right) coset if and only if there is a simple object $S$ in $\mathcal{D}$ such that $X$ is a constituent of $Y\otimes S$.
+By [Bur22, Theorem 4.10] two simple objects $X_{i},X_{i^{\prime}}$ of a pseudo-unitary premodular fusion category $\mathcal{C}$ are in the same coset with respect to $\mathcal{C}^{\prime}:=\mathcal{Z}_{2}(\mathcal{C})$ if and only if $M(i)=M(i^{\prime})$.
+In other words, $(\mathcal{R}_{j})_{j\in\mathcal{I}_{2}}$ correspond exactly to these cosets.
 Denote also ${\mathrm{R}}_{j}:=\sum_{[X_{i}]\in\mathcal{R}_{j}}d_{i}\chi_{i}\in{\mathrm{CF}(\mathcal{C})}$ the regular part of their characters.
 By [Bur22, Equation (4.20)],
-(10.2) $\dim({\mathrm{R}}_{j})=\dim({\mathcal{C}^{\prime}})\dim(\mathcal{C}^{j}),\;\text{for all}\;j\in\mathcal{J}_{2}.$
+(10.2) $\dim({\mathrm{R}}_{j})=\dim({\mathcal{C}^{\prime}})\dim(\mathcal{C}^{j}),\;\text{for all}\;j\in\mathcal{I}_{2}.$
 Proposition 10.1.
-Suppose that $\mathcal{C}$ is a weakly-integral premodular category such that $\mathcal{Z}_{2}(\mathcal{C})\subseteq{\mathcal{C}_{\mathrm{pt}}}$ and it acts freely on the set $\mathrm{Irr}(\mathcal{C})$ of isomorphism classes of simple objects of $\mathcal{C}$.
-Then for any $j\in\mathcal{J}_{2}$ such that $\dim(\mathcal{C}^{j})>1$ there is an $x_{i}$ such that $\mu_{j}(x_{i})\neq 0$.
+Suppose that $\mathcal{C}$ is a weakly-integral premodular category such that $\mathcal{C}^{\prime}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ and acting freely on the set $\mathrm{Irr}(\mathcal{C})$.
+Then for any $j\in\mathcal{I}_{2}$ such that $\dim(\mathcal{C}^{j})>1$ there is an $x_{i}$ such that $\mu_{j}(x_{i})\neq 0$.
 Proof.
-Note that in ${K(\mathcal{C})}$ one has $h_{i}=1$ for all $i$, and therefore ${K(\mathcal{C})}$ is $h$-integral.
-If $\mathcal{C}$ is weakly-integral, the fact that the dual ${\widehat{{K(\mathcal{C})}}}$ is rational RN follows from [Bur21, Equation (4.5)].
-Thus one may try to apply Theorem 4.22 for ${(H,\mathcal{B},\mu_{1})}={K(\mathcal{C})}$.
-In order to do be able to do this, one needs to verify that the additional condition (4.18) is satisfied for all $1\leq i\leq m$, i.e.
-$\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}\in\mathbb{A}$ for all $i\in\mathcal{I}$.
-In [Bur22, Lemma 4.2] it was shown that
+First of all, ${K(\mathcal{C})}$ is $h$-integral because as a fusion ring, $h_{i}=1$ for all $i$.
+If $\mathcal{C}$ is weakly integral, then the dual $^{\widehat{{K(\mathcal{C})}}}$ is a rational RN hypergroup, as shown by [Bur21, Equation (4.3)].
+Therefore, we can attempt to apply Theorem 4.22 for ${(H,\mathcal{B},\mu_{1})}={K(\mathcal{C})}$.
+To do this, we need to ensure that the additional condition (4.18) holds for all $i\in\mathcal{I}$, specifically that $\dim(\mathcal{C}^{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}\in\mathbb{A}$ for all $i\in\mathcal{I}$, where ${\alpha_{{}_{ij}}}:=\mu_{j}(x_{i})$.
+It was demonstrated in [Bur22, Lemma 4.2] that
 (10.3) $\frac{\alpha_{{}_{iM(i^{\prime})}}}{d_{i}}=\frac{s_{{}_{ii^{\prime}}}}{d_{i}d_{i^{\prime}}}=\frac{\alpha_{{}_{i^{\prime}M(i)}}}{d_{i^{\prime}}}.$
-for all $1\leq i,i^{\prime}\leq m$. Fix an index $i_{j}\in\mathcal{I}$ such that $M(i_{j})=j$. It follows that
+for all $i,i^{\prime}\in\mathcal{I}$. Fix an index $i_{j}\in\mathcal{I}$ such that $M(i_{j})=j\in\mathcal{I}_{2}$, and take $i^{\prime}=i_{j}$. It follows that
 (10.4) ${\alpha_{{}_{ij}}}=\frac{d_{i}}{d_{i_{j}}}\alpha_{i_{j}M(i)}.$
@@
 […]
-Since $\mathcal{C}$ is spherical one has $d_{i}=\overline{d_{i}}$ and therefore $|d_{i}|^{2}=d_{i}^{2}$. Thus
-(10.6) $\dim(\mathcal{C}_{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}=\frac{\dim(\mathcal{C}_{j})}{d_{i_{j}}^{2}}\mid\alpha_{i_{j}M(i)}\mid^{2}$
-We will show that $\dim(\mathcal{C}^{j})=d_{i_{j}}^{2}$ and then the proof follows since $\alpha_{i_{j}M(i)}\in\mathbb{A}$.
-Indeed, applying Equation (10.2), then $\dim(\mathcal{C}^{j})=\frac{\dim(R_{j})}{\dim(\mathcal{C}^{\prime})}=d_{i_{j}}^{2}$ since the action of $\mathcal{C}^{\prime}$ is free on $\mathrm{Irr}(\mathcal{C})$.
+But $\mathcal{C}$ is spherical, so $d_{i}=\overline{d_{i}}$ and therefore $|d_{i}|^{2}=d_{i}^{2}$. Thus
+(10.6) $\dim(\mathcal{C}^{j})\frac{|{\alpha_{{}_{ij}}}|^{2}}{d_{i}^{2}}=\frac{\dim(\mathcal{C}^{j})}{d_{i_{j}}^{2}}\mid\alpha_{i_{j}M(i)}\mid^{2}$
+But $\mathcal{C}^{\prime}\subseteq{\mathcal{C}_{\mathrm{pt}}}$, acting freely on $\mathrm{Irr}(\mathcal{C})$, thus $|\mathcal{R}_{j}|=\dim(\mathcal{C}^{\prime})$, and so by Equation (10.2), $\dim(\mathcal{C}^{j})=\frac{\dim(R_{j})}{\dim(\mathcal{C}^{\prime})}=d_{i_{j}}^{2}$, and the proof follows since $\alpha_{i_{j}M(i)}\in\mathbb{A}$.
 ∎
+Roughly speaking, Proposition 10.1 can be intuitively understood as follows: a weakly integral premodular fusion category $\mathcal{C}$ is dual-Burnside with respect to $\mathcal{C}^{\prime}$.
 10.1. Modular category case
-In this section let $\mathcal{C}$ be a weakly-integral modular fusion category.
-We apply the results of previous section to $K(\mathcal{C})$ and prove Theorem 1.13 and Theorem 1.16.
-Recall the definition of $\mathcal{J}_{{}_{\mathcal{D}}}$ from Equation (9.7).
-Composing Drinfeld’s map $\mathrm{f}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\mathrm{{CE}(\mathcal{C})}$ with the inverse of the natural isomorphism $\alpha:\mathrm{{CE}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$ from Equation (9.4) one obtains an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{\mathrm{CF}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$.
-This can also be written as an algebra isomorphism ${\widehat{\mathrm{f}_{Q}}}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$.
-It was shown in [Bur21, Theorem 4.1] that in fact this is an isomorphism of normalized hypergroup ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}.$ Recall that $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$.
+In this section, let $\mathcal{C}$ denote a modular fusion category.
+We will utilize the results from the previous section to $K(\mathcal{C})$ and prove Theorems 1.16 and 1.19.
+Recall the definition of $\mathcal{I}_{{}_{\mathcal{D}}}$ as provided in Definition (6.3).
+By composing Drinfeld’s map $\mathrm{f}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\mathrm{{CE}(\mathcal{C})}$ with the inverse of the isomorphism $\alpha:\mathrm{{CE}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$ from Equation (9.4), we obtain an algebra isomorphism $\widehat{\mathrm{f}}_{Q}:{\mathrm{CF}(\mathcal{C})}\rightarrow\widehat{{\mathrm{CF}(\mathcal{C})}}$.
+This can also be expressed as an algebra isomorphism $\widehat{\mathrm{f}}_{Q}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$.
+It was shown in [Bur21, Remark 4.2] that this is, in fact, an isomorphism of normalized hypergroups: $\widehat{\mathrm{f}}_{Q}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$, where $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$, so that $\widehat{\mathrm{f}}_{Q}(\frac{x_{i}}{d_{i}})=\mu_{i}$.
 Remark 10.2.
-This isomorphism establishes a canonical bijection between the sets $\mathcal{I}$ and $\mathcal{J}$ which is given by $\mathrm{f}_{Q}(F_{i})=E_{i}$ for all $i\in\mathcal{I}$.
-Through the rest of the paper we identify $\mathcal{I}$ and $\mathcal{J}$ via the above bijection.
+Drinfeld’s map $\mathrm{f}_{Q}$ is an algebra isomorphism since the S-matrix is invertible.
+It establishes a canonical bijection between the index sets of $\{F_{i}\}$ and $\{E_{i}\}$ which is given by $\mathrm{f}_{Q}(F_{i})=E_{i}$, for all $i\in\mathcal{I}$.
 Corollary 10.3.
@@
 Proof.
-As mentioned above ${\widehat{\mathrm{f}_{Q}}}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$ is an isomorphism of normalized hypergroup, where $\overline{{K(\mathcal{C})}}$ denotes the normalized version of ${K(\mathcal{C})}$.
-Therefore $\overline{{K(\mathcal{C})}}$ (and thus ${K(\mathcal{C})}$) is Burnside if and only if ${\widehat{{K(\mathcal{C})}}}$ is also Burnside.∎
+As mentioned above $\widehat{\mathrm{f}}_{Q}:\overline{{K(\mathcal{C})}}\rightarrow{\widehat{{K(\mathcal{C})}}}$ is an isomorphism of normalized hypergroup.
+Therefore $\overline{{K(\mathcal{C})}}$ (and thus ${K(\mathcal{C})}$) is Burnside if and only if $^{\widehat{{K(\mathcal{C})}}}$ is also Burnside.∎
 Lemma 10.4.
 Let $\mathcal{C}$ be a modular fusion category. Then with the above notations,
-(10.7) $X_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{pt}}})\iff i\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$
+(10.7) $X_{i}\in\mathrm{Irr}({\mathcal{C}_{\mathrm{pt}}})\iff i\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}$
 Proof.
-The result follows since $({\mathcal{C}_{\mathrm{ad}}})^{\prime}={\mathcal{C}_{\mathrm{pt}}}$.
-Then $X_{i}\in\mathrm{Irr}(({\mathcal{C}_{\mathrm{ad}}})^{\prime})$ if and only if $i\in J_{\mathcal{C}_{\mathrm{ad}}}$ by the second item of [Bur20, Theorem 6.2].
-∎
+Apply Theorem 6.5 with the isomorphism of normalized hypergroups $\widehat{\mathrm{f}}_{Q}$. ∎
 Theorem 10.5.
 In any modular fusion category $\mathcal{C}$ with ${K(\mathcal{C})}$ Burnside the following identity holds:
-(10.8) $\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}{\mathrm{sgn}}(x_{j})F_{j}$
+(10.8) $\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}=\sum_{j\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}}{\mathrm{sgn}}(x_{j})F_{j}$
 Proof.
-Since $K(\mathcal{C})$ is Burnside we may apply ${\widehat{\mathrm{f}_{Q}}}^{-1}$ to Equation (9.3).
-Under the above bijection between $\mathcal{I}$ and $\mathcal{J}$ one has ${\widehat{\mathrm{f}_{Q}}}({\frac{x_{i}}{d_{i}}})=\mu_{i}$, see [Bur20, Theorem 6.1].
-Thus the above Equation (9.3) becomes
-$\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})F_{i}$
+Since $K(\mathcal{C})$ is Burnside, we can apply $\widehat{\mathrm{f}}_{Q}^{-1}$ to Equation (9.3), but $\widehat{\mathrm{f}}_{Q}({\frac{x_{i}}{d_{i}}})=\mu_{i}$, so we get
+$\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{X_{i}\in{\mathcal{C}_{\mathrm{pt}}}}{\mathrm{sgn}}(x_{i})F_{i}$
 and Lemma 10.4 finishes the proof. ∎
@@
 Proof.
-Squaring Equation (10.8) one obtains that $\big{(}\prod_{i=1}^{m}\frac{x_{i}}{d_{i}}\big{)}^{2}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$.
-Note that $\lambda_{{\mathcal{C}_{\mathrm{ad}}}}=\sum_{j\in{J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}}F_{j}$ by the definition of ${J_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$.∎
-Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside.
+Squaring Equation (10.8) we obtain that $\big{(}\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}\big{)}^{2}=\sum_{j\in{\mathcal{I}}_{\mathcal{C}_{\mathrm{ad}}}}F_{j}=\lambda_{{\mathcal{C}_{\mathrm{ad}}}}$ by the definition of ${{\mathcal{I}}_{{}_{{\mathcal{C}_{\mathrm{ad}}}}}}$.
+The result follows by Equation (5.2).∎
+Recall that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [GNN09, Appendix] or [Bur23, Theorem 2].
 Proof of Corollary 1.17
@@
 Proposition 10.6.
-Let $\mathcal{C}$ be a fusion category such that the dual ${\widehat{{K(\mathcal{C})}}}$ is a Burnside hypergroup. Then
-$\frac{({\prod_{i=1}^{m}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
+Let $\mathcal{C}$ be a fusion category such that ${K(\mathcal{C})}$ is commutative and the dual $^{\widehat{{K(\mathcal{C})}}}$ is Burnside. Then
+$\frac{(\prod_{i\in\mathcal{I}}d_{i})^{2}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}\in\mathbb{Z}.$
 Proof.
-For any category $\mathcal{C}$ define $P_{\mathcal{C}}:=\prod_{i=1}^{m}d_{i}$.
-Suppose that $\big{(}\prod_{i=1}^{m}x_{i}\big{)}^{2}=\sum_{i=1}^{m}N_{i}x_{i}$ with $N_{i}\in\mathbb{Z}_{\geq 0}$.
-Then equalizing coefficients in Equation (1.7) one obtains the following $\frac{N_{i}}{P_{\mathcal{C}}^{2}}=\frac{d_{i}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$.
-For $i=1$ it follows that $n_{1}\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})=P_{\mathcal{C}}^{2}$ which proves the divisibility.
+Let $P_{\mathcal{C}}$ be $\prod_{i\in\mathcal{I}}d_{i}$.
+Suppose that $\big{(}\prod_{i\in\mathcal{I}}x_{i}\big{)}^{2}=\sum_{i\in\mathcal{I}}N_{i}x_{i}$ with $N_{i}\in\mathbb{Z}_{\geq 0}$.
+Then equalizing coefficients in Equation (1.6), we obtain the following $\frac{N_{i}}{P_{\mathcal{C}}^{2}}=\frac{d_{i}}{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$.
+Take $i=1$ then $N_{1}\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})=P_{\mathcal{C}}^{2}$ which proves the divisibility.
 ∎
@@
 By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$ since $\mathcal{C}$ is weakly-integral.
-Note that since ${\widehat{{K(\mathcal{C})}}}$ is Burnside the first divisibility result follows from Proposition 10.6.
+Note that since $^{\widehat{{K(\mathcal{C})}}}$ is Burnside the first divisibility result follows from Proposition 10.6.
+Thus ${\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$ is an integer (as a rational algebraic integer).
 Moreover, if $\mathcal{C}$ is nilpotent then [GN08, Corollary 5.3] implies $d_{i}^{2}\mid{\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})}$.
-Thus $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})$.
+Thus $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{2})$.
 ∎
@@
 In the case of an integral fusion category Equation (1.10) can be written as
-$\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}).$
+$\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}).$
+Recall that a weakly-integral fusion category with a commutative Grothendieck ring is Burnside by [Bur23, Theorem 2].
+Moreover, a Burnside modular fusion category is dual-Burnside by Corollary 10.3.
+So a weakly-integral modular fusion category is dual-Burnside.
 Proof of Theorem 1.19
@@
 The first divisibility follows from Proposition (10.6).
-Suppose now that $\mathcal{C}$ is a weakly-integral modular fusion category.
-By [ENO05, Proposition 8.27] one has that $\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})$ is an integer.
+Now, let $\mathcal{C}$ be a weakly-integral modular fusion category.
+By [ENO05, Proposition 8.27] we have that $\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})$ is an integer.
 By [GN08, Theorem 3.10] $d_{i}^{2}\in\mathbb{Z}$ for all $i$.
-The second item follows since for a modular category one has $\mathrm{FPdim}(\mathcal{C})=\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$.
+The second item follows since for a modular category we have $\mathrm{FPdim}(\mathcal{C})=\mathrm{FPdim}({\mathcal{C}_{\mathrm{ad}}})\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$, because the group associated to ${\mathcal{C}_{\mathrm{pt}}}$ is the universal grading group by [EGNO15, Lemma 8.22.5].
 Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\cup\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})$.
-On the other hand, from the first item, $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\subseteq\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})$ which proves the inclusion $\mathcal{V}(\mathcal{C})\subseteq\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{{2}})\big{)}$.
-Conversely, $\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i=1}^{m}\mathcal{V}(d_{i}^{2})\big{)}\subseteq\mathcal{V}(\mathcal{C})$ since $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})\mid\mathrm{FPdim}(\mathcal{C})$ and $d_{i}^{{2}}\mid\mathrm{FPdim}(\mathcal{C})$ for all $1\leq i\leq m$.
+On the other hand, from the first item, $\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})\subseteq\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{{2}})$ which proves the inclusion $\mathcal{V}(\mathcal{C})\subseteq\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{{2}})\big{)}$.
+Conversely, $\mathcal{V}({\mathcal{C}_{\mathrm{pt}}})\cup\big{(}\bigcup_{i\in\mathcal{I}}\mathcal{V}(d_{i}^{2})\big{)}\subseteq\mathcal{V}(\mathcal{C})$ since $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})\mid\mathrm{FPdim}(\mathcal{C})$, and by [EGNO15, Proposition 8.14.6], $d_{i}^{{2}}\mid\mathrm{FPdim}(\mathcal{C})$ for all $i\in\mathcal{I}$.
 ∎
@@
 Proof.
-Since $g.c.d(d_{i}^{2},\;d)=1$ it follows from Equation (1.14) that $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$.
-Let now ${\mathcal{D}}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ a pointed fusion subcategory such that $\mathrm{FPdim}({\mathcal{D}})=d$.
-Since $\mathrm{FPdim}({\mathcal{D}}^{\prime})=N$ and $\mathrm{g.c.d}(d,N)=1$ it follows that ${\mathcal{D}}\cap{\mathcal{D}}^{\prime}=\mathrm{Vec}$ and therefore ${\mathcal{D}}$ is non-degenerate.
-Thus $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$ and ${\mathcal{D}}^{\prime}$ is non-degenerate as well.
+Since $\gcd(d_{i}^{2},d)=1$ and $d$ is square-free, it follows from Equation (1.13) that $d$ divides $\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$.
+Now, consider a pointed fusion subcategory ${\mathcal{D}}\subseteq{\mathcal{C}_{\mathrm{pt}}}$ such that $\mathrm{FPdim}({\mathcal{D}})=d$.
+The existence of such a subcategory is straightforward to demonstrate because the associated groups are abelian, given that the category is braided.
+According to [EGNO15, Theorem 8.21.5], with $\mathcal{C}^{\prime}=\mathrm{Vec}$, we have $\mathrm{FPdim}({\mathcal{D}}^{\prime})=m$, but $\gcd(d,m)=1$.
+Consequently, by Lagrange’s theorem [EGNO15, Theorem 7.17.6], ${\mathcal{D}}\cap{\mathcal{D}}^{\prime}=\mathrm{Vec}$, which implies that ${\mathcal{D}}$ is non-degenerate, as stated in [EGNO15, Corollary 8.20.10].
+Thus, by [EGNO15, Theorem 8.21.4], $\mathcal{C}\simeq{\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$, and ${\mathcal{D}}^{\prime}$ is also non-degenerate.
 ∎
 Remark 10.8.
-Note that the above Corollary improves [OY23, Theorem 4.5] since the authors also assume the existence of a Tannakian subcategory ${\mathcal{E}}=\mathrm{Rep}(G)\subseteq\mathcal{C}$ such that $\mathcal{C}^{0}_{G}\simeq\mathcal{C}(\mathbb{Z}_{d},q)\boxtimes\mathcal{A}$ for some non-degenerate braided fusion category $\mathcal{A}$.
+Note that Theorem 1.21 improves [OY23, Theorem 4.5] since the authors also assume the existence of a Tannakian subcategory ${\mathcal{E}}=\mathrm{Rep}(G)\subseteq\mathcal{C}$ such that $\mathcal{C}^{0}_{G}\simeq\mathcal{C}(\mathbb{Z}_{d},q)\boxtimes\mathcal{A}$ for some non-degenerate braided fusion category $\mathcal{A}$.
 Proof of Corollary 1.22:
 Proof.
-Immediate from Theorem 1.17, we just need to check that $g.c.d(d_{i}^{2},\;d)=1$, but it is clear since $d_{i}^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ and $d$ is a square-free factor. ∎
-Lemma 10.9.
-With the above notations, for an integral perfect modular category, $\mathcal{V}(\mathcal{C})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ and
-$d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})\mid P_{\mathcal{C}}^{2}$
-Proof.
-Since ${\mathcal{C}_{\mathrm{pt}}}$ is trivial then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$, see Remark 11.13.
-Thus $\mathcal{V}(\mathcal{C})=\mathcal{V}({\mathcal{C}_{\mathrm{ad}}})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i})$ by Theorem 1.16.
-Since $d_{i}^{2}\mid\mathrm{FPdim}(\mathcal{C})$ the result follows.
-∎
-Lemma 10.9 extends trivially to the weakly-integral case because:
+Immediate from Theorem 1.21, we just need to check that $gcd(d_{i}^{2},d)=1$, but it is clear since $d_{i}^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ and $d$ is a factor of the square-free part. ∎
+Corollary 10.9.
+An integral modular fusion category $\mathcal{C}$ can always be decomposed into ${\mathcal{D}}\boxtimes{\mathcal{D}}^{\prime}$, where both ${\mathcal{D}}$ and ${\mathcal{D}}^{\prime}$ are modular, ${\mathcal{D}}$ is pointed, and $\mathrm{FPdim}({\mathcal{D}})$ is the maximal square-free part of $\mathrm{FPdim}(\mathcal{C})$.
+This means that the classification of integral modular fusion categories $\mathcal{C}$ simplifies to those where $\mathrm{FPdim}(\mathcal{C})$ lacks a square-free part, up to a Deligne tensor product with a pointed modular fusion category.
+Corollary 10.10.
+For an integral perfect modular fusion category $\mathcal{C}$, $\mathrm{FPdim}(\mathcal{C})$ has no square-free part.
+In other words, if a prime $p$ divides $\mathrm{FPdim}(\mathcal{C})$, then $p^{2}$ also divides $\mathrm{FPdim}(\mathcal{C})$.
+Corollary 10.10 obviously extends to the weakly-integral case due to the following:
 Proposition 10.11.
@@
 Proof.
-By the proof of Lemma 10.9 then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$. But [EGNO15, Exercice 9.6.12] states that the adjoint subcategory of a weakly-integral fusion category is integral. ∎
+Since ${\mathcal{C}_{\mathrm{pt}}}$ is trivial then ${\mathcal{C}_{\mathrm{ad}}}=\mathcal{C}$, because $({\mathcal{C}_{\mathrm{pt}}})^{\prime}={\mathcal{C}_{\mathrm{ad}}}$ by [EGNO15, Corollary 8.22.8], see also Remark 11.13.
+Moreover, [EGNO15, Exercise 9.6.12] asserts that the adjoint subcategory of a weakly-integral fusion category is integral.
+∎
 Corollary 10.12.
-Let $\mathcal{C}$ be an integral perfect modular fusion category. If $p$ is a prime number dividing $\mathrm{FPdim}(\mathcal{C})$ then $p^{2}\mid\mathrm{FPdim}(\mathcal{C})$.
-Corollary 10.12.
-Any integral modular category $\mathcal{C}$ such that there is a prime $p$ whose exponent in $\mathrm{FPdim}(\mathcal{C})$ is $1$ (i.e $m\neq 1$) it is not perfect.
-Corollary 10.13.
-If $\mathcal{C}$ is a perfect, integral, even dimensional modular fusion category then $4\mid\mathrm{FPdim}(\mathcal{C})$.
-10.2. On the Statements (1.21) and (1.22) and Conjecture (1.20)
-In [LPR23, Theorem 5.6] it was shown that the statements (1.21) and (1.22) are equivalent. Next we will show that any of these statements implies Conjecture 1.20 formulated in [CP22].
+If $\mathcal{C}$ is a perfect, integral, even-dimensional modular fusion category, then 444 divides $\mathrm{FPdim}(\mathcal{C})$.
+10.2. On Statements (1.25) and (1.26) and Conjecture (1.24)
+In [LPR23, Theorem 5.6] it was shown that Statements (1.25) and (1.26) are equivalent. We will show that any of these statements implies Conjecture 1.24 formulated in [CP22].
 Proposition 10.13.
-Existence of a perfect odd dimensional modular fusion category implies the existence of a non-pointed simple integral modular fusion category.
+The existence of a perfect odd dimensional modular fusion category implies the existence of a non-pointed simple integral modular fusion category.
 Proof.
-Suppose that there exists a perfect modular odd-dimensional fusion category $\mathcal{C}$.
-Note first that $\mathcal{C}$ has no pointed subcategories since it is perfect.
-It also has no symmetric subcategories, since then these subcategories are of the type $\mathrm{Rep}(G)$ with $G$ a perfect group.
-But then $|G|$ is divisible by $4$ and this is a contradiction since the $\mathrm{FPdim}(\mathcal{C})$ is odd.
-On the other hand, by a dimension argument, $\mathcal{C}$ has at least one minimal subcategory under the inclusion (and hence a simple subcategory), say ${\mathcal{D}}$.
-If ${\mathcal{D}}$ is a modular fusion category then ${\mathcal{D}}$ is an example of a non-pointed simple integral modular fusion category.
-If ${\mathcal{D}}$ is not modular then ${\mathcal{D}}^{\prime}\cap{\mathcal{D}}$ is a non-trivial symmetric fusion subcategory which as explained above, it is not possible. ∎
+Let $\mathcal{C}$ be a perfect modular fusion category with an odd dimension. By Proposition 10.11, $\mathcal{C}$ must be integral.
+[…]
+However, due to the finite number of isomorphism classes of simple objects, $\mathcal{C}$ must contain at least one minimal subcategory.
+Let us call this minimal subcategory ${\mathcal{D}}$.
+If ${\mathcal{D}}$ is itself a modular fusion category, then ${\mathcal{D}}$ serves as an example of a non-pointed, simple, integral modular fusion category.
+If ${\mathcal{D}}$ is not modular, then ${\mathcal{D}}^{\prime}\cap{\mathcal{D}}$ is non-trivial according to [EGNO15, Corollary 8.20.10], yet it is symmetric by design.
+But as discussed earlier, such symmetric subcategories cannot exist, leading to a contradiction.
+∎
 Recall that in [CP22] the authors have also shown that Conjecture 1.24 is equivalent to either of the following:
@@
 11.1. Perfect Drinfeld center
-We will characterize the perfect fusion categories with a perfect center. That will provide a large class of perfect modular fusion categories.
+We will characterize the perfect fusion categories with a perfect Drinfeld center. That will provide a large class of perfect modular fusion categories.
 Lemma 11.1.
@@
 Proposition 11.2.
-Let $\mathcal{C}$ be a perfect fusion category. Then its universal grading group is perfect if and only if $\mathcal{Z}(\mathcal{C})$ is perfect.
+Let $\mathcal{C}$ be a perfect fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if the universal grading group of $\mathcal{C}$ is perfect.
 Proof.
@@
 Corollary 11.3.
-Let $\mathcal{C}$ be a braided fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if $\mathcal{C}$ is perfect with a perfect universal grading group.
+Let $\mathcal{C}$ be a braided fusion category. Then $\mathcal{Z}(\mathcal{C})$ is perfect if and only if $\mathcal{C}$ is perfect with a trivial universal grading group.
 Proof.
-Immediate by Proposition 11.2 because $\mathcal{C}$ is isomorphic to a fusion subcategory of $\mathcal{Z}(\mathcal{C})$. ∎
+If $\mathcal{Z}(\mathcal{C})$ is perfect, then $\mathcal{C}$ is also perfect because it can be embedded into $\mathcal{Z}(\mathcal{C})$ as a braided fusion category.
+According to Proposition 11.2, the universal grading group of $\mathcal{C}$ is perfect.
+However, since $\mathcal{C}$ is braided, its Grothendieck ring is commutative, which means its universal grading group must be abelian due to the faithful grading.
+But a perfect abelian group is trivial.
+Conversely, the result follows directly from Proposition 11.2, as the trivial group is perfect.
+∎
+Proposition 11.2 and Corollary 11.3 provide a complete characterization of when the Drinfeld center $\mathcal{Z}(\mathcal{C})$ is perfect assuming the fusion category $\mathcal{C}$ is perfect or braided.
+However, it is important to note that there are fusion categories that are neither perfect nor braided, yet their Drinfeld centers are perfect.
+An example of this is the fusion category $\mathcal{C}=\mathrm{Vec}(G)$, where $G$ is a non-abelian finite simple group.
+This fusion category is Morita equivalent to $\mathrm{Rep}(G)$, as illustrated in [EGNO15, Example 7.12.19].
+As a result, their Drinfeld centers are braided equivalent, as shown in [EGNO15, Theorem 8.12.3].
+Nonetheless, Corollary 11.7 below confirms that $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect in this case.
 Lemma 11.4.
@@
 Proof.
-Immediate by Theorem 7.8, because a grouplike element in the dual hypergroup of the Grothendieck ring of $\mathrm{Rep}(G)$ corresponds to a column of the character table of $G$ with squared norm $|G|$, so to a conjugacy class of size one, thus to a central element.
+Immediate by Theorem 7.8, because a grouplike element $\mu_{j}$ in the dual hypergroup of the Grothendieck ring of $\mathrm{Rep}(G)$ corresponds to a column of the character table of $G$ with squared norm $n_{j}=|G|$, by Lemma 2.31, so to a conjugacy class of size $|G|/n_{j}=1$, thus to a central element.
 ∎
 Corollary 11.5.
-Let $G$ be a finite group. Then $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if $G$ is perfect with a trivial center $Z(G)$.
+Let $G$ be a finite group. The Drinfeld center $\mathcal{Z}(\mathrm{Rep}(G))$ is perfect if and only if $G$ is perfect with a trivial center $Z(G)$.
 Proof.
-Immediate by Corollary 11.3 and Lemma 11.4, because $\mathrm{Rep}(G)$ is perfect if and only if $G$ is perfect, and (as an abelian group) $Z(G)$ is perfect if and only if it is trivial. ∎
+Immediate by Corollary 11.3 and Lemma 11.4, because $\mathrm{Rep}(G)$ is perfect if and only if $G$ is perfect. ∎
 Remark 11.6.
-Note that Corollary 11.5 can be proved alternatively using the fact (see [EGNO15, §8.5]) that the simple objects of $\mathcal{Z}(\mathrm{Rep}(G))$ are in bijection with the pairs $(C,V)$ where $C$ is a finite conjugacy class in $G$, and $V$ is an irreducible finite dimensional representation of the centralizer of $g$ in $C$, and the Frobenius-Perron dimension of the object corresponding to a pair $(C,V)$ is $|C|dim_{\mathbb{C}}(V)$.
+[…]
 Corollary 11.7.
@@
 Immediate by Corollary 11.5, because a non-abelian finite simple group is perfect with a trivial center. ∎
-By Corollary 11.7, $\mathcal{Z}(\mathrm{Rep}(A_{5}))$ is a perfect integral modular fusion category of $\mathrm{FPdim}$ $360060^{2}=$ 3600. Using Remark 11.6 and GAP, it is of rank $22$ and type
-$[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].$
+By Corollary 11.7, $\mathcal{Z}(\mathrm{Rep}(A_{5}))$ is a perfect integral modular fusion category of $\mathrm{FPdim}$ $60^{2}=3600$. Using Remark 11.6 and GAP, it is of rank 222222 and type
+[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].[[1,1],[3,2],[4,1],[5,1],[12,10],[15,4],[20,3]].[ [ 1 , 1 ] , [ 3 , 2 ] , [ 4 , 1 ] , [ 5 , 1 ] , [ 12 , 10 ] , [ 15 , 4 ] , [ 20 , 3 ] ] .
 Question 11.8.
-Is there a perfect integral modular fusion category of rank less than $22$? of $\mathrm{FPdim}$ less than $3600$?
+Is there a non-trivial perfect integral modular fusion category of rank less than 222222, or $\mathrm{FPdim}$ of less than 360036003600?
 Recall that a Hopf algebra $A$ is called perfect if $\mathrm{Rep}(A)$ is perfect, i.e.
-$G(A^{*})$ is a trivial group; that a finite dimensional semisimple Hopf algebra $A$ (over $\mathbb{C}$) is factorizable if and only if $\mathrm{Rep}(A)$ is modular; that $\mathcal{Z}(\mathrm{Rep}(G))=\mathrm{Rep}(D(G))$ where the Hopf algebra $D(G)$ is the Drinfeld double of $G$.
+$G(A^{*})$ is a trivial group; that a finite dimensional semisimple Hopf algebra $A$ (over $\mathbb{C}$) is factorizable if and only if $\mathrm{Rep}(A)$ is modular; that $\mathcal{Z}(\mathrm{Rep}(A))=\mathrm{Rep}(D(A))$ where the Hopf algebra $D(A)$ is the Drinfeld double of $A$.
 So, for any non-abelian finite simple group $G$ then $D(G)$ is a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension $|G|^{2}$.
+Let us clarify Question 11.8 for the case of Hopf algebras:
 Question 11.9.
-Is there a finite dimensional semisimple factorizable perfect Hopf algebra (over $\mathbb{C}$) of dimension less than $\dim(D(A_{5}))=3600$?
+Is there a non-trivial finite-dimensional semisimple factorizable perfect Hopf algebra over $\mathbb{C}$ that has less than 222222 irreducible representations (up to isomorphism), or a dimension of less than 360036003600?
 Lemma 11.10.
@@
 Remark 11.11.
-Let $\mathcal{C}$ be a fusion category with a commutative Grothendieck ring $K(\mathcal{C})$.
-By Theorem 7.8, the dual of $K(\mathcal{C})$ is perfect if and only if its universal grading group is trivial.
+According to Theorem 7.8, the dual hypergroup of a commutative fusion ring is perfect if and only if the universal grading group of the fusion ring is trivial.
 Remark 11.12.
-The Grothendieck ring of a modular fusion category is perfect if and only if its dual is perfect (as it is isomorphic to its dual).
+The Grothendieck ring of a modular fusion category is perfect if and only if its dual is perfect (since it is isomorphic to its dual as normalized hypergroup, see §10.1).
 Remark 11.13.
-By Remarks 11.11 and 11.12, a modular fusion category $\mathcal{C}$ is perfect if and only if its universal grading group is trivial (i.e. $\mathcal{C}_{ad}=\mathcal{C}$).
+By Remarks 11.11 and 11.12, a modular fusion category $\mathcal{C}$ is perfect if and only if its universal grading group is trivial (i.e.
+$\mathcal{C}_{ad}=\mathcal{C}$).
+Alternatively, recall that $({\mathcal{C}_{\mathrm{pt}}})^{\prime}={\mathcal{C}_{\mathrm{ad}}}$ by [EGNO15, Corollary 8.22.8].
 Proof of Theorem 1.13
 Proof.
-We use the Galois action on $\widehat{\mathcal{B}}$ from Lemma 3.2.
+We utilize the Galois action on $\widehat{\mathcal{B}}$ as described in Lemma 3.2.
 Let $\mu_{1}$ be the $\mathrm{FPdim}$ character.
-By Proposition 3.5, it follows that any Galois conjugate of $\mu_{1}$, say $\mu_{j}=\sigma.\mu_{1}$ has for formal codegree $n_{j}=\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))$.
-On the other hand, since $\mu_{j}=\sigma.\mu_{1}$ is also a non-vanishing character, it should be a grouplike element, by the definition of a dual-Burnside hypergroup.
-Thus, its formal codegree should also be $\mathrm{FPdim}(H,{\mathcal{B}})$, which implies that $\sigma(\mathrm{FPdim}(H,{\mathcal{B}}))=\mathrm{FPdim}(H,{\mathcal{B}})$, for all $\sigma$ in the Galois group.
-Then, we are done since this implies that $\mathrm{FPdim}(H,{\mathcal{B}})$ is a rational number.
+According to Proposition 3.5, any Galois conjugate of $\mu_{1}$, denoted by $\mu_{j}=\sigma\cdot\mu_{1}$, has a formal codegree given by $n_{j}=\sigma(\mathrm{FPdim}(H,\mathcal{B}))$.
+However, because $\mu_{j}=\sigma\cdot\mu_{1}$ is also a non-vanishing character, it must be a grouplike element due to the definition of a dual-Burnside hypergroup (Definition 4.7).
+Therefore, its formal codegree must also be $\mathrm{FPdim}(H,\mathcal{B})$, by Corollary 2.29.
+Thus, $\sigma(\mathrm{FPdim}(H,\mathcal{B}))=\mathrm{FPdim}(H,\mathcal{B})$ for every $\sigma$ in the Galois group.
+It follows that $\mathrm{FPdim}(H,\mathcal{B})$ is a rational number, completing the proof.
 ∎
-The following result is partially due to Andrew Schopieray, see [Sch23].
+As a consequence of Theorem 1.13, every commutative dual-Burnside fusion ring is weakly integral.
+However, the converse is not true, even for unitary integral fusion categories such as $\mathrm{Rep}(A_{7})$; refer to §12.2 for more examples.
+The following result is partially attributed to Andrew Schopieray; see [Sch23].
 Corollary 11.14.
 Let ${(H,\mathcal{B},\mu_{1})}$ be rational abelian normalizable hypergroup.
-If the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect and Burnside then ${(H,\mathcal{B},\mu_{1})}$ is integral, i.e $d_{i}\in\mathbb{Q}$.
+If the dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect and Burnside then ${(H,\mathcal{B},\mu_{1})}$ is integral, i.e $d_{i}\in\mathbb{Q}$.
 Proof.
-The dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect, so as for the above proof of Theorem 1.10, $\sigma.\mu_{1}=\mu_{1}$ for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$.
-Thus $\sigma(d_{i})=d_{i}$, which shows that $d_{i}\in\mathbb{Q}$.
+The dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is perfect, meaning that $\mu_{1}$ is the only grouplike linear character.
+As in the proof of Theorem 1.13, for any $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma.\mu_{1}$ is also non-vanishing and thus grouplike by the dual-Burnside assumption.
+Consequently, $\sigma.\mu_{1}=\mu_{1}$, and $\sigma(d_{i})=d_{i}$ for all $\sigma$, which implies that $d_{i}\in\mathbb{Q}$.
 ∎
@@
 Recall that a modular fusion category is Burnside if and only if it is dual-Burnside (Corollary 10.3).
-Now, by Remarks 11.11 and 11.12, the universal grading group is trivial, so by Corollary 11.15, it must be integral.
+Now, by Remark 11.13, the universal grading group is trivial, so by Corollary 11.16, it must be integral.
 The converse follows from [GNN09, Theorem 6.1].
@@
 Remark 11.17.
-By Theorem 1.23, a simple non-integral modular fusion category is not (dual-)Burnside.
-Plenty of modular fusion categories of Lie type (Verlinde categories) are simple non-integral, so is the one from [Sch22, Theorem 1], and the center of Extended-Haagerup, so none is (dual-)Burnside.
+According to Theorem 1.27, a simple non-integral modular fusion category cannot be (dual-)Burnside.
+Many modular fusion categories of Lie type, known as Verlinde categories, belong to this class, as do the one described in [Sch22, Theorem 1] and the Drinfeld center of the Extended-Haagerup fusion categories.
+Consequently, none of these are (dual-)Burnside.
+Therefore, for the Grothendieck rings in this class, we can still infer the existence of a non-group-like and non-vanishing linear character (and basic element).
 11.3. Near-group modular fusion categories
-Let $G$ be a finite abelian group and $m$ be a non-negative integer. Let $K(G,m)$ be the fusion ring with basis $G\cup\{\rho\}$ and fusion rules:
+Let $G$ be a finite abelian group and $m$ be a non-negative integer. Let $K(G,m)$ be the fusion ring with basis $\mathcal{B}=G\cup\{\rho\}$ and fusion rules:
 $\rho^{2}=\sum_{g\in G}g+m\rho\text{ and }g\rho=\rho g=\rho,$
@@
 Proposition 11.18.
-There is no modular fusion category whose Grothendieck ring is $K(G,m)$, with $G$ non-trivial and $m>0$.
+There is no modular fusion category of Grothendieck ring $K(G,m)$ if $G$ is non-trivial and $m>0$.
 Proof.
-Let $\mathcal{C}$ be such a category and $R=K(\mathcal{C})$ be its Grothendieck ring.
-Since $\rho^{2}=m\rho+\sum_{g\in G}g$, then $\mathrm{FPdim}(\rho)^{2}=m\mathrm{FPdim}(\rho)+|G|$.
-Let $x_{1}>0$ and $x_{2}<0$ be the two solutions of the equation $x^{2}-mx-|G|=0$, then $\mathrm{FPdim}(\rho)=x_{1}$.
-Any non-trivial character $\eta\in\widehat{G}$ can be extended to a character of $R$ (denoted also by $\eta$) by declaring $\eta(\rho)=0$.
-On the other hand, the trivial character $\eta=\epsilon$ ($\epsilon(g)=1$, for all $g\in G$) has two extensions on $R$: $\psi_{+}(\rho)=x_{1}$ and $\psi_{-}(\rho)=x_{2}$.
-It follows that $\psi_{+}=\mathrm{FPdim}$. Thus
-^=${\widehat{\mathcal{B}}}=\{\eta|\;\eta\in\widehat{G},\;\eta\neq\epsilon\}\cup\{\psi_{\pm}\}$
-The multiplication on $\widehat{R}$ is given by
-$\eta\star\psi_{-}=\eta,\;\eta\star\eta^{\prime}=\eta\eta^{\prime}\in\widehat{G},\;$
-Moreover, it is easy to compute that:
-$\psi_{-}^{2}=\frac{x_{2}^{2}+|G|}{x_{1}^{2}+|G|}\psi_{+}+\frac{x_{1}^{2}-x_{2}^{2}}{x_{1}^{2}+|G|}\psi_{-}.$
-Thus $R$ has RN-dual since $x_{1}^{2}=mx_{1}+|G|>mx_{2}+|G|=x_{2}^{2}$.
-Note that $R$ is a self dual hypergroup since $\mathcal{C}$ is modular.
-In particular, Drinfeld’s map $\mathrm{f}_{Q}:R\rightarrow\widehat{R}$, since an isomorphism of hypergroups, sends grouplike elements of $R$ to grouplike elements of $\widehat{R}$.
-Since $G(\widehat{R})=1$ and $G(R)=G$ is non-trivial, this shows that $\mathrm{f}_{Q}$ is not bijective, which is a contradiction.
-∎
+Let $\mathcal{C}$ be a modular fusion category of Grothendieck ring $K(G,m)$.
+Given that $\rho^{2}=m\rho+\sum_{g\in G}g$, it follows that $\mathrm{FPdim}(\rho)^{2}=m\mathrm{FPdim}(\rho)+|G|$.
+Let $x_{+}>0$ and $x_{-}<0$ be the solutions to the equation $x^{2}-mx-|G|=0$.
+Thus, $\mathrm{FPdim}(\rho)=x_{+}$.
+Now, consider $\eta$, a linear character of $R$.
+For all $g\in G$, we have $\eta(\rho)=\eta(\rho g)=\eta(\rho)\eta(g)$.
+Therefore, $\eta(\rho)$ is nonzero if and only if $\eta(g)=1$ for all $g\in G$.
+This condition is equivalent to solving $\eta(\rho)^{2}-m\eta(\rho)-|G|=0$, which implies $\eta(\rho)=x_{\pm}$.
+We denote such characters as $\psi_{\pm}$.
+It follows that $\psi_{+}=\mathrm{FPdim}$ is the unit of $^{\widehat{\mathcal{B}}}$.
+Let $\epsilon$ be the trivial character of $G$.
+We identify any $\eta$ in $\widehat{G}\setminus\{\epsilon\}$ (non-empty since $G$ is non-trivial) with the corresponding element in $^{\widehat{\mathcal{B}}}$ that vanishes on $\rho$.
+Let $q=-\frac{x_{-}}{x_{+}}$, which is positive.
+The multiplication on $\widehat{R}$ is defined by:
+$\eta\star\psi_{-}$ $=\eta,$
+$\eta\star\eta^{\prime}$ $=(1-\delta_{\eta^{\prime},\eta^{*}})\eta\eta^{\prime}+\delta_{\eta^{\prime},\eta^{*}}(\frac{q}{1+q}\psi_{+}+\frac{1}{1+q}\psi_{-}),$
+$\psi_{-}\star\psi_{-}$ $=q\psi_{+}+(1-q)\psi_{-}.$
+[…]
 Proposition 11.19.
-Let $G$ be a finite group and $\mathcal{C}$ be a modular category whose Grothendieck ring ${K(\mathcal{C})}$ is $K(G,0)$. Then $G\simeq C_{2}$.
+Let $\mathcal{C}$ be a modular fusion category of Grothendieck ring $K(G,0)$. Then $G\simeq C_{1}$ or $C_{2}$.
 Proof.
-As above, if $m=0$ then $\psi_{-}^{2}=\mathrm{FPdim}$ and therefore ${\widehat{{K(\mathcal{C})}}}$ has two grouplike elements $\psi_{+}=\mathrm{FPdim}$ and $\psi_{-}$.
-Note that Drinfeld’s map $\mathrm{f}_{Q}:{K(\mathcal{C})}\rightarrow{\widehat{{K(\mathcal{C})}}}$ send bijectively grouplike elements of ${K(\mathcal{C})}$ into grouplike elements of ${\widehat{{K(\mathcal{C})}}}$.
-Since all $g\in G$ are grouplike elements of ${K(\mathcal{C})}$ it follows $|G|=2$.
-∎
+[…]
 Remark 11.20.
-Ising’s category is the only example of a modular category whose Grothendieck ring is of type $K(C_{2},0)$, see [DGNO07, Appendix B].
+Ising categories are the only modular fusion categories whose Grothendieck ring are $K(C_{2},0)$, see [DGNO10, Appendix B].
 Corollary 11.21.
@@
 Proof.
-Immediate by Propositions 11.18, 11.19, Remark 11.20 and [Ost03b]. ∎
+According to Proposition 11.18, either $G$ is trivial or $m=0$.
+If $G$ is trivial, the category has rank 2, and the conclusion follows from [Ost03b].
+On the other hand, if $G$ is non-trivial and $m=0$, the result follows from Proposition 11.19 and Remark 11.20.
+∎
 Remark 11.22.
@@
 12. Concrete examples
-This section focus on providing some concrete examples of fusion rings/categories to which the main results of the paper apply as categorification criteria, and also some counter-examples showing that some results cannot be generalized.
+This section focuses on providing concrete examples of fusion rings and categories to which the main results of the paper apply as categorification criteria, along with counter-examples demonstrating the limits of these results.
 12.1. Burnside property
-Recall that a fusion ring is called Burnside if for every basis element then its fusion matrix has norm $1$ if and only if its determinant is nonzero (in other words, grouplike is equivalent to invertible, see Proposition 4.3).
-Recall that William Burnside proved that the Grothendieck ring of $\mathrm{Rep}(G)$ is Burnside, for every finite group $G$, and that first author’s [Bur23, Theorem 2] generalizes this result to every weakly-integral fusion category (over $\mathbb{C}$).
-A fusion ring $R$ is called $\alpha$-Frobenius if for all basis element $x$ then $\frac{\mathrm{FPdim}(R)^{\alpha}}{\mathrm{FPdim}(x)}$ is an algebraic integer.
-Kaplansky’s 6th conjecture states that the Grothendieck ring of a complex fusion category is $1$-Frobenius.
-Under the following bounds:
-$\begin{array}[]{c|c|c|c|c|c|c|c|c}\text{rank}(R)&\leq 5&6&7&8&9&10&11&12\\ \hline\cr\mathrm{FPdim}(R)<&1000000&150000&15000&10000&4000&2000&500&200\end{array}$
-there are exactly $192$ non-pointed simple integral $1$-Frobenius fusion rings with $\mathrm{FPdim}\neq p^{a}q^{b},pqr$ (the last restriction is motivated by [ENO11]).
-This is the current state of a classification initiated in [LPW21].
-Among them, we found only $4$ non-Burnside ones.
+Recall that a fusion ring is called Burnside if, for every basis element, its fusion matrix has a norm of 111 if and only if its determinant is nonzero.
+In other words, being grouplike is equivalent to being invertible (see Proposition 4.3).
+William Burnside originally proved that the Grothendieck ring of $\mathrm{Rep}(G)$ is Burnside for every finite group $G$.
+This result was expanded in [Bur23] to include all weakly integral fusion categories with a commutative Grothendieck ring, but first within the context of modular categories in [GNN09, Appendix].
+In Theorem 1.2, it extends to every commutative fusion ring with an $h$-integral dual, and further to hypergroups in Theorem 4.17.
+Remark 12.1.
+By Theorems 1.8 and 7.9, along with the fact that $\mathrm{Rep}(G)_{pt}\simeq\mathrm{Rep}(G/G^{\prime})$, Burnside’s result can be restated as follows: for any finite group $G$,
+$\left(\prod_{j\in\mathcal{I}}\frac{C_{j}}{|\mathcal{C}^{j}|}\right)^{2}=\frac{1}{|G^{\prime}|}\sum_{\mathcal{C}^{j}\subseteq G^{\prime}}C_{j},$
+where $G^{\prime}$ is the commutator subgroup of $G$, $(\mathcal{C}^{j})$ are the conjugacy classes of $G$, and $C_{j}:=\sum_{g\in\mathcal{C}^{j}}g$ is the class sum associated with $\mathcal{C}^{j}$.
+It was already observed in [Har08].
+A fusion ring $R$ is termed $\alpha$-Frobenius if, for every basis element $x$, the expression $\frac{\mathrm{FPdim}(R)^{\alpha}}{\mathrm{FPdim}(x)}$ is an algebraic integer [ENO25].
+Kaplansky’s 6th conjecture posits that the Grothendieck ring of a complex fusion category is 111-Frobenius [ENO11, Question 1].
+An exhaustive classification of all 111-Frobenius simple integral fusion rings, within certain specified limits, was provided in [LPW21].
+These limits have been updated in [BP24] as follows:
+$\begin{array}[]{c|cccccccc}\text{Rank}&\leq 5&6&7&8&9&10&11&12\\ \hline\cr\operatorname{\mathsf{FPdim}}\leq&10^{7}&10^{6}&10^{5}&20000&10000&5000&3000&1000\end{array}$
+There are exactly 505505505 non-pointed examples (including 888 that are character rings of groups).
+Among them, we found only 444 non-Burnside ones.
 They have the type of $\mathrm{Rep}(A_{7})$, but different fusion data.
-  • Rank: $9$,
+  • Rank: 999,
   • FPdim: $2520=2^{3}\times 3^{2}\times 5\times 7$,
-  • Type: $[[1,1],[6,1],[10,2],[14,2],[15,1],[21,1],[35,1]]$,
+  • Type: [[1,1],[6,1],[10,2],[14,2],[15,1],[21,1],[35,1]][[1,1],[6,1],[10,2],[14,2],[15,1],[21,1],[35,1]][ [ 1 , 1 ] , [ 6 , 1 ] , [ 10 , 2 ] , [ 14 , 2 ] , [ 15 , 1 ] , [ 21 , 1 ] , [ 35 , 1 ] ],
   • Fusion data 1:
@@
     $\begin{smallmatrix}0&0&0&0&0&1&0&0&0\\ 0&1&0&0&2&0&1&0&1\\ 0&0&1&1&0&1&1&1&2\\ 0&0&1&1&0&1&1&1&2\\ 0&2&0&0&3&0&2&2&2\\ 1&0&1&1&0&3&0&3&2\\ 0&1&1&1&2&0&2&1&3\\ 0&0&1&1&2&3&1&4&3\\ 0&1&2&2&2&2&3&3&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1&0&0\\ 0&0&1&1&0&1&0&1&1\\ 0&1&2&1&1&1&2&1&1\\ 0&1&1&2&1&1&2&1&1\\ 0&0&1&1&0&2&1&2&3\\ 0&1&1&1&2&0&2&1&3\\ 1&0&2&2&1&2&2&2&2\\ 0&1&1&1&2&1&2&2&5\\ 0&1&1&1&3&3&2&5&8\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&1&0\\ 0&1&0&0&2&0&1&2&1\\ 0&0&1&1&0&1&1&1&4\\ 0&0&1&1&0&1&1&1&4\\ 0&2&0&0&5&2&2&4&2\\ 0&0&1&1&2&3&1&4&3\\ 0&1&1&1&2&1&2&2&5\\ 1&2&1&1&4&4&2&6&4\\ 0&1&4&4&2&3&5&4&12\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&0&0&1\\ 0&0&1&1&0&1&1&1&4\\ 0&1&0&0&3&2&1&4&5\\ 0&1&0&0&3&2&1&4&5\\ 0&0&3&3&0&2&3&2&9\\ 0&1&2&2&2&2&3&3&8\\ 0&1&1&1&3&3&2&5&8\\ 0&1&4&4&2&3&5&4&12\\ 1&4&5&5&9&8&8&12&14\end{smallmatrix}$
-The determinant of the second fusion matrix (of each fusion data above) is ±$\pm 36$ whereas its norm is $6$, proving that these integral fusion ring are non-Burnside, and so cannot be categorified.
-Similar classification were done in the non-1-Frobenius case, and unexpectedly, most (not all) of the ones we found are non-Burnside (see one example below), so without categorification, which can be seen as an evidence for Kaplansky 6th conjecture.
-  • Rank: $7$,
-  • FPdim: $798=2\times 3\times 7\times 19$,
-  • Type: $[[1,1],[7,1],[8,1],[9,3],[21,1]]$,
+The determinant of the second fusion matrix for each fusion data set mentioned above is ±36$\pm$ 36± 36, while its norm is 666.
+This indicates that these integral fusion rings are non-Burnside and, consequently, cannot be categorified.
+Similar classifications have been conducted for the non-1-Frobenius case.
+Interestingly, most—but not all—of the fusion rings identified are non-Burnside, and thus, not amenable to categorification (see the example below).
+This serves as supporting evidence for Kaplansky’s 6th conjecture.
+  • Rank: 777,
+  • FPdim: 798=2×3×7×19798=2$\times$ 3$\times$ 7$\times$ 19798 = 2 × 3 × 7 × 19,
+  • Type: [[1,1],[7,1],[8,1],[9,3],[21,1]][[1,1],[7,1],[8,1],[9,3],[21,1]][ [ 1 , 1 ] , [ 7 , 1 ] , [ 8 , 1 ] , [ 9 , 3 ] , [ 21 , 1 ] ],
   • Fusion data:
     $\begin{smallmatrix}1&0&0&0&0&0&0\\ 0&1&0&0&0&0&0\\ 0&0&1&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&1&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0&0\\ 1&0&0&1&1&1&1\\ 0&0&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&1&5\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0&0\\ 0&0&1&1&1&1&1\\ 1&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&1&1&1&6\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&2&1&1&1\\ 1&1&2&1&1&2&1\\ 0&1&1&1&2&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&2&1&1\\ 0&1&1&1&2&2&1\\ 1&1&2&2&1&1&1\\ 0&1&1&2&1&2&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1&0\\ 0&1&1&1&1&1&1\\ 0&1&1&1&1&2&1\\ 0&1&1&2&2&1&1\\ 0&1&1&2&1&2&1\\ 1&1&2&1&2&1&1\\ 0&1&1&1&1&1&7\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&0&1\\ 0&1&1&1&1&1&5\\ 0&1&1&1&1&1&6\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 0&1&1&1&1&1&7\\ 1&5&6&7&7&7&8\end{smallmatrix}$
-The determinant of the third fusion matrix is $16$ whereas its norm is $8$, proving that this integral fusion ring is non-Burnside, and so cannot be categorified.
+The determinant of the third fusion matrix is 161616, while its norm is 888. This shows that the integral fusion ring is non-Burnside and therefore cannot be categorified.
 12.2. Dual-Burnside property
-We made the abbreviation (A)RN for (abelian) real non-negative.
-Note that [LPW21, Corollary 8.5] means that the Grothendieck ring, if commutative, of a unitary fusion category, has a RN dual (hypergroup).
-Question 12.1.
+We abbreviate “(A)RN” for “(abelian) real non-negative.” According to [LPW21, Corollary 8.5], the Grothendieck ring of a unitary fusion category, if commutative, has an RN dual (hypergroup).
+Question 12.2.
 Is the dual of the Grothendieck ring of a complex fusion category always RN?
-Recall that a commutative fusion ring $F$ with RN dual is called dual-Burnside when a column of its character table has a zero entry if and only if the squared norm of the column (i.e.
-the corresponding formal codegree) has value strictly less than $\mathrm{FPdim}(F)$, or alternatively, when $P^{2}$ is an idempotent, where
-$P=\prod_{i\in I}\frac{x_{i}}{\mathrm{FPdim}(x_{i})}$
-and $\{x_{i}\}_{i\in I}$ is the set of basis elements (Corollary 4.16).
-We warn the reader that for a non-abelian group $G$, the product of all its elements depends on the ordering, in fact, the set of all possible products forms a $G^{\prime}$-coset, see [DH82].
-Recall that [INW99, Theorem B] states that for all finite nilpotent group $G$ then the Grothendieck ring of $\mathrm{Rep}(G)$ is dual-Burnside.
-It is extended to every nilpotent dualizable ARN-hypergroup by Theorem 1.12.
-So every commutative nilpotent fusion ring with RN dual is both Burnside and dual-Burnside.
-Note that all the commutative nilpotent fusion rings we checked have RN dual.
-Question 12.2.
-Is there a commutative nilpotent fusion ring without RN dual?
-Let us call a finite group $G$ dual-Burnside if $K(\mathrm{Rep}(G))$ is dual-Burnside.
-So a nilpotent finite group is dual-Burnside.
-There are exactly $144$ finite groups of order less than $32$, and among them, exactly $30$ ones are non-nilpotent, and among them, only $\mathrm{SL}(2,3)$ is dual-Burnside (so the $29$ other ones, e.g.
-$S_{3}$ and $D_{5}$, are not dual-Burnside).
-Observe that a finite group $G$ is centerless if and only if $K(\mathrm{Rep}(G))$ has a perfect dual (see Definition 2.39, Theorem 7.8 and Lemma 11.4).
-So a centerless finite group is dual-Burnside if and only if every non-$\mathrm{FPdim}$ column of its character table has a zero entry.
-We checked that every centerless and dual-Burnside finite group $G$ of order $|G|\leq 1000$ is almost simple (i.e.
-$S\subseteq G\subseteq\mathrm{Aut}(S)$ with $S$ non-abelian simple); but $G=A_{5}\times A_{5}$, of order $3600$, is not almost simple.
-A finite non-abelian simple group is centerless, so by [Pal] and its answers, all the non-alternating finite simple groups are dual-Burnside, except the Mathieu groups $M_{22}$ and $M_{24}$.
-The alternating group $A_{n}$, $5\leq n\leq 19$, is dual-Burnside if and only if $n\not\in\{7,11,13,15,16,18,19\}$.
-Recall that a finite group $G$ is simple if and only if $K(\mathrm{Rep}(G))$ is simple.
+Recall that a commutative fusion ring $F$ is called dual-Burnside if a column of its character table has a zero entry if and only if the squared norm of the column (the corresponding formal codegree) is strictly less than $\mathrm{FPdim}(F)$.
+When $F$ has an RN dual, it is equivalent to $P^{2}$ being idempotent, where
+$P=\prod_{i\in\mathcal{I}}\frac{x_{i}}{\mathrm{FPdim}(x_{i})}$
+and $\{x_{i}\}_{i\in I}$ is the set of basis elements (see Corollary 4.16).
+Note that for a non-abelian group $G$, the product of all its elements depends on the ordering; in fact, the set of all possible products forms a $G^{\prime}$-coset (see [DH82]).
+According to [INW99, Theorem B], for any finite nilpotent group $G$, the Grothendieck ring of $\mathrm{Rep}(G)$ is dual-Burnside.
+This extends to every nilpotent dualizable ARN-hypergroup by Theorem 1.15.
+Thus, every commutative nilpotent fusion ring with an RN dual is both Burnside and dual-Burnside.
+All commutative nilpotent fusion rings we examined have RN duals.
+Question 12.3.
+Is there a commutative nilpotent fusion ring without an RN dual?
+We define a finite group $G$ as dual-Burnside if $K(\mathrm{Rep}(G))$ is dual-Burnside.
+Therefore, a nilpotent finite group is dual-Burnside.
+There are exactly 144 finite groups of order less than 32, among which 30 are non-nilpotent.
+Of these, only $\mathrm{SL}(2,3)$ is dual-Burnside, while the other 29, such as $S_{3}$ and $D_{5}$, are not.
+[…]
+Remark 12.4.
+By Theorem 1.8, a finite group $G$ is dual-Burnside if and only if
+$\left(\prod_{\chi\in\mathrm{Irr}(G)}\frac{\chi}{\chi(1)}\right)^{2}=\frac{|Z(G)|}{|G|}\left(\sum_{\chi\in\mathrm{Irr}(G/Z(G))}\chi(1)\chi\right).$
 12.3. Modular categories
-Let us provide some applications of Theorem 1.16.
-There are exactly $43$ different types for a $\frac{1}{2}$-Frobenius integral fusion ring up to rank $11$ (see [ABPP]), and among them, exactly $7$ ones (below) are excluded from modular categorification by Theorem 1.18.
-$[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],$
-$[1,1,1,1,1,1,1,1,1,3],[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4].$
+Let us present some applications of Theorem 1.19.
+There are exactly 71 distinct types of half-Frobenius integral fusion rings with ranks up to 12 (see [ABPP]).
+Of these, 15 types (listed below) are excluded from modular categorification by Theorems 1.19 or 1.22.
+[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],[1,1,1,1,1,1,1,1,1,3],[1,1,1,1,2,2],[1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2],[1,1,1,1,2,2,2,4,4],[1,1,1,1,1,1,1,1,1,3],[ 1 , 1 , 1 , 1 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 ] ,
+[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4],[1,1,1,1,1,1,1,1,1,3,3,3],[1,1,1,1,1,1,1,1,2,2,2,2],[1,1,1,1,2,2,2,2,2,2],[1,1,1,1,2,2,2,4,4,4,4],[1,1,1,1,1,1,1,1,1,3,3,3],[1,1,1,1,1,1,1,1,2,2,2,2],[ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 , 3 , 3 ] , [ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 ] ,
+[1,1,1,1,1,1,1,1,2,2,4,4],[1,1,1,1,2,2,2,2,2,2,2,2],[1,1,1,1,2,2,2,2,2,2,2,4],[1,1,1,1,2,2,2,4,4,4,4,4],[1,1,1,1,1,1,1,1,2,2,4,4],[1,1,1,1,2,2,2,2,2,2,2,2],[1,1,1,1,2,2,2,2,2,2,2,4],[1,1,1,1,2,2,2,4,4,4,4,4],[ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 2 , 2 , 4 , 4 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 2 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 4 ] , [ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 4 , 4 ] ,
+[1,1,1,1,2,2,2,4,4,4,8,8],[1,1,2,2,2,2,3,3,6,6,6,6].[1,1,1,1,2,2,2,4,4,4,8,8],[1,1,2,2,2,2,3,3,6,6,6,6].[ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 4 , 4 , 4 , 8 , 8 ] , [ 1 , 1 , 2 , 2 , 2 , 2 , 3 , 3 , 6 , 6 , 6 , 6 ] .
 For example, let us provide all the fusion data for the first type above:
-  • Rank: $6$,
+  • Rank: 666,
   • FPdim: $12=2^{2}\times 3$,
-  • Type: $[1,1,1,1,2,2]$,
+  • Type: [1,1,1,1,2,2][1,1,1,1,2,2][ 1 , 1 , 1 , 1 , 2 , 2 ],
   • Fusion data 1:
@@
     $\begin{smallmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&2\\ 1&1&1&1&0&0\end{smallmatrix},\ \begin{smallmatrix}0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 0&0&0&0&0&1\\ 1&1&1&1&0&0\\ 0&0&0&0&2&0\end{smallmatrix}$
-They are excluded by Theorem 1.18 from modular categorification, because $3$ is a powerless factor of the $\mathrm{FPdim}$, whereas $3$ does not divide $4$, the number of grouplike elements.
-Note that the above fusion data $3$ and $4$ can be generalized as follows.
-Let $G$ be a group of order $n^{2}$, and let $K$ be a group of order $m+1$.
-Consider the following $\frac{1}{2}$-Frobenius integral fusion ring:
+They are excluded from modular categorification by Theorem 1.19 (1.14) because $\mathcal{V}(\mathcal{C})=\mathcal{V}(12)=\{2,3\}$, while $3\not\in\mathcal{V}(d_{i})$ for any $i$, and $\mathcal{V}(\mathcal{C}_{pt})=\{2\}$.
+Additionally, Theorem 1.22 rules them out since 3 is a (powerless) factor of $\mathrm{FPdim}(\mathcal{C})$, but does not divide $\mathrm{FPdim}(\mathcal{C}_{pt})=4$.
+Note that the fusion data for cases 3 and 4 can be generalized.
+Consider a group $G$ with order $n^{2}$ and a group $K$ with order $m+1$.
+Then, analyze the following half-Frobenius integral fusion ring:
   • Rank: $n^{2}+m$,
@@
     • $x_{g}\rho_{k}=\rho_{k}x_{g}=\rho_{k}$, for all $(g,k)\in G\times K$,
-It is excluded by Theorem 1.16 (1.13) from modular categorification if $n>1$ and $m+1$ has a prime factor not dividing $n$.
-Availability of data and materials
-The datasets generated during and/or analyzed during the current study are available from the second author on reasonable request.
-Declaration
-Conflict of interests: The authors declare that they have no conflict of interest.
+According to Theorem 1.19 (1.14), it is excluded from modular categorification if $n>1$ and $m+1$ has a prime factor that does not divide $n$.
+Note that the types [1,1,1,1,1,1,1,1,1,3,3,3][1,1,1,1,1,1,1,1,1,3,3,3][ 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 3 , 3 , 3 ] and [1,1,1,1,2,2,2,2,2,2,2,2][1,1,1,1,2,2,2,2,2,2,2,2][ 1 , 1 , 1 , 1 , 2 , 2 , 2 , 2 , 2 , 2 , 2 , 2 ] are excluded from modular categorification by Theorem 1.19, but not by Theorem 1.22, since their $\mathrm{FPdim}=2^{2}3^{2}$ have no powerless prime factor.
+Availability of data and materials. Data for the computations in this paper are available on reasonable request from the authors.
+Conflict of interest statement. On behalf of all authors, the corresponding author declares that there are no conflicts of interest.